### v36

# Fast Polynomial Root-Finding and Approximation of the Eigenvalues of a Matrix

Abstract
Univariate polynomial root-finding has been studied for four millennia and very intensively in the last decades. Our new near-optimal root-finders approximate all zeros of a polynomial $p$ almost as fast as one accesses its coefficients with the precision required for the solution within a prescribed error bound. Furthermore, our root-finders can be applied to a black box polynomial, defined by an oracle (black box subroutine) for its evaluation rather than by its coefficients. Due to this feature our root-finders (i) support approximation of the eigenvalues of a matrix in a record Las Vegas expected bit operation time and (ii) are particularly fast for a polynomial that can be evaluated fast such as the sum of a few shifted monomials of the form $(x-c)^{d}$ or a Mandelbrot-like polynomial defined by a recurrence. Our divide and conquer algorithm of ACM STOC 1995 is the only other known near-optimal polynomial root-finder, but it extensively uses the coefficients, is quite involved, and has never been implemented, while according to extensive numerical experiments with standard test polynomials, already a slower initial implementation of our new root-finders competes with user’s choice package of root-finding subroutines MPSolve and supersedes it more and more significantly as the degree of a polynomial grows large. Our auxiliary algorithms and techniques for computations with black box polynomials can be of independent interest.
Keywords:
polynomial root-finding, matrix eigenvalues, computational complexity, black box polynomials, symbolic-numeric computing.
2020 Math. Subject Classification:
65H05, 26C10, 30C15
PART I: BACKGROUND AND OUR PROGRESS
1 State of the Art and our Main Results
In this part of the paper we state our main results and outline our algorithms. We present them in some detail in Parts II, III, and IV, which can be read independently of each other.
1.1 Polynomial Root-finding: Problems, Impact, and Earlier Studies
According to [7, 99, 100, 83], univariate polynomial root-finding, that is, the solution of a polynomial equation $p(x)=0$, had been the central problem of Mathematics and Computational Mathematics for about four millennia since Sumerian times; it was still a central problem well into the 19th century. In a long way from its rudimentary level to modern version, its study has been responsible for the introduction of mathematical rigor and notation as well as the concepts of irrational, negative, and complex numbers, algebraic groups, rings, and fields etc.
The existence of complex roots of the equation $p(x)=0$ is called the Fundamental Theorem of Algebra. The attempts of its proof by d’Alembert (1746), Euler (1749), de Foncenex (1759), Lagrange (1772), Laplace (1795), and Gauss (1799) had gaps (cf. [133]), and the first rigorous proof is due to an amateurish mathematician Argand (1806, revised in 1813).
Already about 500 BC it was proved in Pythagorean school that the roots of the equation $x^{2}=2$ were irrational radicals, while even with the help of radicals one cannot express the roots rationally if the degree $d$ of $p(x)$ exceeds 4, according to Ruffini (1799,1813) and Abel (1824); moreover, one cannot do this already for specific equations such as $x^{5}-x-1=0$ or $x^{5}-4x-2$, according to Galois theory of 1830, 1832.
Nevertheless, in 1924 in [149], Herman Weyl published his constructive proof of the Fundamental Theorem of Algebra where he rationally approximated within any fixed positive error tolerance all roots of $p(x)=0$ lying in a fixed square in the complex plane, that is, all roots if the square is large enough. This was the solution of the following Problem 0 of modern polynomial root-finding:
INPUT: $d+1$ complex coefficients $p_{0},p_{1},\dots,p_{d}$ of a polynomial
$p=p(x):=\sum_{i=0}^{d}p_{i}x^{i}:=p_{d}\prod_{j=1}^{d}(x-z_{j}),$ (1.1)
a positive error tolerance $\epsilon=1/2^{b}>0$, an integer $m$, $1\leq m\leq d$, and a bounded domain $\mathbb{D}$ in the complex plane, e.g., a disc or a rectangle, such that $z_{j}\in\mathbb{D}$ if and only if $j\leq m$.
OUTPUT: approximations $\tilde{z}_{j}\approx z_{j}$, $j=1,\dots,m$, to the $m$ zeros of $p(x)$ (aka roots of $p(x)=0$) [^1] Hereafter we freely say just “root” or “zero” for short, always count them with their multiplicities, and do not distinguish multiple roots from pairwise isolated root clusters of diameters less than $\epsilon$. lying in the domain $\mathbb{D}$ such that
$|\tilde{z}_{j}-z_{j}|\leq\epsilon~{}{\rm for}~{}j=1,\dots,m.$ (1.2)
Example 1.1.
[Approximation of all roots.] Fix $R\geq\max_{j}|z_{j}|$ (see [52] for simple expressions for such $R$) and let $\mathbb{D}$ denote the disc $D(0,R):=\{x:~{}|x\leq R|\}$, centered in the origin and having radius $R$. Then Problem 0 requires us to approximate all $d$ roots.
Example 1.2.
[Numerical approximation of real roots.] Root-finding in the real line or its segment (real root-finding) is highly important because in many applications, e.g., to optimization in algebraic geometry and geometric modeling, only real roots of a polynomial are of interest, and typically they are much less numerous than all $d$ complex roots (cf. Kac [66], Edelman and Kostlan [39]). More generally consider Problem 0 where $\mathbb{D}$ denotes the minimal rectangle in the complex plane that covers the $\epsilon/\sqrt{2}$-neighborhood of a fixed real line segment. To solve this problem we should approximate both real and near-real roots, that is, their perturbations due to small input and rounding errors. The known real root-finders (cf. [80, Chs. 2 and 3]) do not address this important issue.
Major motivation for modern study of Problem 0 comes from application to approximation of the eigenvalues of a matrix (see our Secs. 1.3 and 1.5) and the solution of a multivariate polynomial system of equations [29], both problems being highly important for the theory and practice of modern computations. [82, Introduction] specifies some applications of polynomial root-finding to computer algebra, signal processing, geometric modeling, control, and financial mathematics. Furthermore, Problem 0 for a domain $\mathbb{D}$ containing all roots is closely linked to polynomial factorization (see Secs. 24 and 25), which has applications to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [150, 18, 8, 35, 36, 147].
Hundreds of efficient root-finders have been proposed [80, 82] and keep appearing (see [122, 22, 40, 70, 76, 127, 108, 23, 24, 64, 110, 111, 75, 60, 62, 81, 63, 59, 115], and the bibliography therein). In [133] Smale proposed to study Problem 0 in terms of arithmetic complexity and computational precision and later, jointly with Blum, Cucker, and Shub (cf. [14, 30]), linked this study to rounding errors and conditioning of the roots. In [134] Schönhage proposed to estimate the complexity of polynomial root-finders in terms of the number of bit operations involved and presented an algorithm that solved Problem 0 for a domain that contains all roots by using $\tilde{O}(bd^{3})$ bit operations, [^2] Here and hereafter we assume that $b\geq d$, recalling that error amplification factors $2^{cd}$ are inevitable in polynomial root-finding, e.g., in routine shift of the variable. that is, $bd^{3}$ bit operations up to a polylogarithmic factor in $bd$.
His work prompted intensive study of root-finding by researchers in the Theory of Computing in the 1980s and 1990s, culminated at ACM STOC 1995 with our near-optimal divide and conquer root-finder of [96] (also see its journal versions [97, 103] and surveys [99, 100]); it supports bound $\tilde{O}(bd^{2})$ and showed the information lower bound $0.25(d+1)bd$ (see Cor. 24.2 in Sec. 24.2).
If we only consider polynomials whose zeros and tiny clusters of zeros are pairwise sufficiently isolated, as we specify in Def. 24.1 of Sec. 24, then the lower bounds on the computational precision and the bit operation complexity of root-finding decrease by a factor of $d$ (see Sec. 24), and so do the upper bounds supported by the algorithm of [96, 97, 103] as well; hence the algorithm remains near-optimal (see Thms. 24.1 and 24.4 and Cor. 24.3 in Sec. 24).
1.2 Reduction to the case of the unit disc
Hereafter T shows transposition of a vector or a matrix, $\mathbb{R}^{k}$ and $\mathbb{C}^{k}$ denote the spaces of the column vectors $(v_{1},\dots,v_{k})^{T}$ of length $k$ with real and complex coordinates, respectively, and $D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$ and $C(c,\rho):=\{x:~{}|x-c|=\rho\}$ denote the disc of radius $\rho>0$ centered at $c\in\mathbb{C}$ and its boundary circle, respectively. Furthermore, let Problem $\widehat{0}$ denote Problem 0 where the domain $\mathbb{D}$ is the unit disc $D(0,1)$; we can extend its solution to root-finding in the disc $D(c,\rho)$ for any pair of complex $c$ and positive $\rho$ by means of applying the linear map
$x\mapsto\frac{x-c}{\rho},~{}p(x)\mapsto t(x)=p\Big{(}\frac{x-c}{\rho}\Big{)},~{}D(c,\rho)\mapsto D(0,1),$ (1.3)
which shifts the variable $x$ by $c$ and then scales $x-c$ by $1/\rho$.
Next we reduce Problem 0 for any domain $\mathbb{D}$ to Problem $\widehat{0}$.
(i) Two ways from the roots in the unit disc to all roots. Given a value
$R\geq\max_{j=1}^{d}\{|x_{j}|\}$ (1.4)
(see [52] for simple expressions for such $R$), we can scale the variable $x\mapsto y=\frac{x}{R}$, $p(x)\mapsto p(y)=p(\frac{x}{R})$ and approximate every root $x_{y}$, for $j=1,\dots,d$, within $\epsilon R$ by solving Problem $\widehat{0}$ for $p(y)$.
Schönhage in [134] reduced approximation of all $d$ roots to Problem $\widehat{0}$ without assuming (1.4) or using (1.3) as follows. Given an algorithm $\mathcal{A}$ that solves Problem $\widehat{0}$, compute also approximations $\tilde{y}_{j}$ within $\epsilon$ to the reciprocals $y_{j}=1/z_{j}$ of all roots lying outside the disc $D(0,1)$ – simply by applying this algorithm to the reverse polynomial
$p_{\rm rev}(x):=x^{d}p\Big{(}\frac{1}{x}\Big{)}=\sum_{i=0}^{d}p_{i}x^{d-i},~{}p_{\rm rev}(x)=p_{0}\prod_{j=1}^{d}\Big{(}x-\frac{1}{z_{j}}\Big{)}~{}{\rm if}~{}p_{0}\neq 0.$ (1.5)
Write $\Delta_{j}:=\tilde{y}_{j}-y_{j}$ and deduce that $\frac{1}{y_{j}}-\frac{1}{\tilde{y}_{j}}=\frac{\Delta_{j}}{\tilde{y}_{j}~{}y_{j}}$, and so $\frac{\tilde{y}_{j}}{y_{j}}-1=\frac{\Delta_{j}}{y_{j}}=\Delta_{j}z_{j}$. Recall that $|\Delta_{j}|\leq\epsilon$ by assumption and obtain that $|\frac{\tilde{y}_{j}}{y_{j}}-1|\leq\epsilon|z_{j}|$, thus arriving at the following
Proposition 1.1.
The reciprocal of an approximation within $\epsilon$ to a zero of $p_{\rm rev}(x)$ in the disc $D(0,1)$ approximates the associated zero of $p(x)$ lying outside that disc within a relative error at most $\epsilon$.
(ii) From roots in the unit disc to roots in a bounded domain. Given a bounded domain $\mathbb{D}$ on the complex plane, we can successively (a) cover it with a minimal disc $D(c,\rho)$, (b) map this disc into the disc $D(0,1)$, (c) solve Problem $\widehat{0}$ for the resulting polynomial $t(x)$ of (1.3), and finally, (d) use the map (1.3) again – this time to approximate all roots lying in the intersection of the disc $D(c,\rho)$ with the domain $\mathbb{D}$. In this way we also approximate extraneous roots lying in $D(c,\rho)$ but not in $\mathbb{D}$; in Sec. 2.2 we control the computational cost of that additional work.
1.3 Black box root-finders; the isolation of roots, eigenvalues, and their clusters
A nontrivial variant of Newton’s iterations by Louis and Vempala [76], at FOCS 2016, approximates within a fixed positive $\epsilon\leq 1/2$ an absolutely largest zero of a real-rooted polynomial $p$ of a degree $d$ at the cost of its evaluation with a precision in $O((b+\log(R))\log(R~{}d)))$ at $O((b+\log(R))\log(d))$ points for $b:=\log_{2}(R/\epsilon)$ and a fixed $R$ satisfying (1.4). While the previous polynomial root-finders have been devised for polynomials given in the monomial basis, that is, with their coefficients, the algorithm of [76] never involves them and works for a black box polynomial $p$ – given with an oracle (black box subroutine) for its evaluation. [76] shows benefits of extension of such algorithms to approximation of the eigenvalues of a matrix. Next we state the relevant tasks.
Problem 1 (Black Box Polynomial Root Approximation) is Problem 0 for a black box polynomial $p$.
Problem 2: Approximation of the Eigenvalues of a Matrix. [^3] As in [76], the problem covers or can be extended to the case where an input is a polynomial matrix, aka matrix polynomial.
INPUT: $\epsilon=1/2^{b}>0$, integers $m$ and $d$ such that $0<m\leq d$, and a $d\times d$ matrix $M$, and a complex domain $\mathbb{D}$ that contains precisely $m$ eigenvalues of $M$ counted with their multiplicities.
OUTPUT: Approximations within $\epsilon$ to each of all $m$ eigenvalues of $M$ lying in the domain $\mathbb{D}$.
The output of Problems 0, 1, and 2 may include multiple approximations within $\epsilon$ to the same root or eigenvalue, but the next modifications exclude such a confusion.
Problems $0^{*}$, $1^{*}$, $2^{*}$: Isolation of Clusters of the Zeros of a polynomial and the Eigenvalues of a matrix. Under the assumptions of Problems 0, 1, and 2 compute at most $d$ disjoint discs of radii at most $\epsilon$ (hereafter $\epsilon$-discs) covering all roots or eigenvalues that lie in the domain $\mathbb{D}$; also compute the number of the roots or eigenvalues lying in each $\epsilon$-disc.
Remark 1.1.
If we collapse all zeros of $p$ lying in every $\epsilon$-disc into a single multiple zero placed in the center of the disc, then these zeros will lie at the distances greater than $2\epsilon$ from each other. If a polynomial $p$ is square-free (having only simple zeros) and has integer coefficients, then Problem 0 turns into the well-known Problem of Root-isolation (see Problem ISOL and Remark 25.4 in Sec. 25.5). These comments can be immediately extended to Problems 1 and 2.
1.4 More on Black Box Polynomial Root-finders
1.4.1 Why we should use them
One can perform black box evaluation at $q>d$ points $x_{1},\dots,x_{q}$ and then interpolate to $p$. The bit operation cost of interpolation, however, is unbounded unless the $q$ points are well isolated from the roots, but it is a challenge to yield such isolation. Furthermore, even if the isolation is ensured, the cost of interpolation exceeds the cost of the solution of Problems 0, 0, 1 ,1, 2, and 2 in the cases where domain $\mathbb{D}$ contains only a small number of roots.
Moreover, here are three examples of important benefits of using black box algorithms even where an input polynomial is given with its coefficients.
(i) Application of a black box root-finder to the characteristic polynomial $p(x)=\det(xI-M)$ of a $d\times d$ matrix $M$ only involves the values of this polynomial but not its coefficients, whose overall size exceeds the size (a norm) of the matrix $M$ by a factor of $d$. Consequently, extension of such a root-finder to Problem 2 saves a factor of $d$ in bit precision and bit operation count versus the known solutions of Problems 2 and 2, based on [134, Sec. 21].
(ii) Map (1.3) is a valuable root-finding tool, but its application is limited because it destroys sparsity and blows up computational precision and bit operation cost. Similar comments can be applied to maps (3.6) and (3.7) of Sec. 3.4. Black box polynomial root-finders keep all benefits of using those maps but avoid the latter notorious problems by involving no coefficients of $t(x)$ and $p_{h}(x)$ of these maps.
(iii) Black box polynomial root-finders are particularly efficient where a $d$th degree polynomial $p$ can be evaluated fast, say, in $O(\log(d))$ arithmetic operations (hereafter (ops) such as a Mandelbrot-like polynomial $p:=M_{k}(x)$, where $M_{0}:=x,~{}M_{i+1}(x):=xM_{i}^{2}(x)+1$, $i=0,1,\dots,k$, $d=2^{k}-1$, or the sum of a bounded number of shifted monomials, such as $p:=\alpha(x-a)^{d}+\beta(x-b)^{d}+\gamma(x-c)^{d}$ for six constants $a$, $b$, $c$, $\alpha$, $\beta$, and $\gamma$. For comparison, if we compute the coefficients of such a polynomial and then use them to evaluate it at a point $x$ by applying Horner’s algorithm, then we would need $2d$ ops, and this bound is optimal (see [89, 69] or [139, Section “Pan’s method"]).
1.4.2 What we adopt and what cannot adopt for black box polynomial root-finding
The known polynomial root-finders, except [76], extensively use the coefficients of $p$, and we must be careful in adopting their techniques for Problems 1 and 1.
On the other hand, the root-finder of [76] cannot help solve even Problem 1. Indeed, repeatedly deflating $p$, by peeling off large roots, can accumulate errors that grow too quickly, together with the coefficient length of the factors (notice, e.g., the swell of the coefficients of $p(x)=x^{d}+1$ in transition to its factors of degree $d/2$, say). This is true even if one performs deflation in the black box setting implicitly, based on Eqn. (3.3) of Sec. 3. Moreover, even in that setting the coefficient swell in deflation is translated into swelling of the values of the high order derivatives involved in the algorithm of [76].
Even if the algorithm of [76] worked for Problem 1 it would only solve the well-known and much easier problem of root-finding for a real-rooted polynomial (cf. [93, 25, 33]). Moreover, the resulting algorithm for that much easier problem would involve more than $b^{2}d^{2}$ bit operations, which would by a factor of $b$ exceed the upper bound of [96, 103] for general polynomial root-finding.
Recursive combination of [76] with the Inverse Power Iterations [48] could lead to approximation of all eigenvalues of a symmetric (although not unsymmetric) matrix, together with its eigenvectors. One can be challenged to estimate the increased complexity of this extended computations, but we study other approaches.
By applying a root-finder to Given a polynomials $p$ of (1.1) given with its coefficients We accelerate known algorithms for Problem 0 by a factor of $m/\log^{2}(m)$ by incorporating fast multipoint polynomial evaluation algorithms of [42, 84, 68, 81], but we could not use them if we successively approximated zeros of $p$ or matrix eigenvalues by using deflation. Thus, instead of applying high order Newton’s iterations of [76] and deflation, we recall and extend subdivision iterations of [149, 52, 128, 101].
1.5 Our complexity estimates
Definition 1.1.
$m=m(\mathbb{D},\beta)$ for $\beta>0$ is the number of roots of polynomial $p(x)$ in the $\beta$-neighborhood of a domain $\mathbb{D}$ on the complex plane.
Main Theorems 1-4 below hold for $m$ of Def. 1.1, any fixed constant $\beta>0$, and either $p(x)=\det(xI-M)$ in the case of Problems 2 and 2 or $p(x)$ of (1.1) otherwise.
Theorem 1.1.
(Main Theorem 1.) One can solve Problems 1 and 1 by querying the polynomial evaluation oracle at $\tilde{O}(m^{2})$ locations with a computational precision $O(b\log(d))$.
Theorem 1.2.
(Main Theorem 2.) One can solve Problems 2 and 2 at Las Vegas randomized bit operation cost
$\mathbb{B}_{\rm eigen}=\tilde{O}(\tilde{b}d^{\omega}m^{2})$ (1.6)
for $\tilde{b}:=\log_{2}(||M||_{F}/\epsilon)$, the Frobenius norm $||M||_{F}$ of a matrix $M$, and $\omega$ denoting an exponent of feasible or unfeasible matrix multiplication (MM). [^4] The current records are about 2.7734 for exponents of feasible MM [90, 107, 109, 72, 57], unbeaten since 1982, and about 2.372 for exponents of unfeasible MM [6, 37]. The exponents below 2.77 have only been reached at the expense of blowing up the size of MM by immense factors.
Theorem 1.3.
(Main Theorem 3.) One can solve Problems 0 and 0 at Las Vegas randomized bit operation cost $\tilde{O}(bdm))$ for a polynomial $p$ of (1.1) normalized such that $||p(x)||_{1}:=\sum_{i=1}^{d}|p_{i}|=1$.
Random Root Model: The zeros of $p$ (in particular for $p$ denoting the characteristic polynomial of a matrix $M$) are independent identically distributed (iid) random variables sampled under the uniform probability distribution on the disc $D(0,R)$ for a large $R\in\mathbb{R}$ such that the value $\frac{R^{2}}{d}-1$ exceeds a positive constant.555(i) The assumption of the model is a little stronger than that of our Def. 24.1, supporting such a decrease in [96, 103]. (ii) Random coefficient models may seem more natural for polynomial root-finding but are useless because no known probability distribution withstands shift of the variable $x$.
Theorem 1.4.
(Main Theorem 4) The upper bounds of Thms. 1.1-1.3 on computational precision and bit operation complexity can be decreased by a factor of $d$ under the Random Root Model.
In view of Remarks 1.1 and 25.4 we can readily extend our Main Theorems to Polynomial Root Isolation.
2 Accelerated subdivision polynomial root-finders: overview and reduction to soft exclusion/inclusion tests
For a polynomial $p$ given in monomial basis (with its coefficients) the deterministic algorithm of [96, 103] supports our Main Theorems 3 and 4, but not our Main Theorems 1 and 2. This is because the algorithm of [96, 103] cannot be applied to a black box polynomial, unlike our present Las Vegas randomized algorithms, which we outline next and elaborate upon in Parts II–IV. Namely, in the next two sections we reduce our root-finders to recursive application of exclusion/inclusion (e/i) test, which decides whether a fixed complex disc contains roots or not. In Secs. 3.3.2 and 3.4 we outline our two approaches to devising e/i tests for a black box polynomial and unit disc $\mathcal{D}=D(0,1)$. We cover these two approaches in some detail in Part II, made up of Secs. 612, and Part III, made up of Secs. 1318, respectively, and describe yet another fast e/i test in Sec. 23.4. Our root-finders of Parts II and III incorporate these e/i tests and are quite fast but still only support root approximation (rather than isolation) and the complexity estimates exceeding those of our Main Theorems by a factor of $b$. In Part IV, made up of Secs. 1923 and outlined in Sec. 4, we fix both deficiencies, thus proving our Main Theorems. In Sec. 5.4 we list some tentative directions to further progress.
2.1 Subdivision root-finders and reduction to e/i tests
We depart from the popular subdivision root-finders, traced back to [149, 52, 128, 101]. They extend root-finding in a line segment by means of bisection to polynomial root-finding in a square on the complex plane and have been extensively used in Computational Geometry under the name of Quad-tree Construction.
Next we outline them.
(A) First let a square on the complex plane contain precisely $m>0$ roots (with multiplicities) and let us approximate them. Call the square suspect, recursively divide every suspect square into four congruent sub-squares, and to each of them apply exclusion/inclusion (e/i) test in order to decide whether the square contains a root or does not. If it does, call the square suspect and process it in the next iteration; otherwise discard it (see Fig. 1a). In Sec. 9 and then in our Part IV we use a natural extension of e/i tests to $m$-tests, which decide whether a disc in the complex plane covering at most $m$ roots does or does not contain $m$ roots.
(a) Four roots marked by asterisks lie in suspect sub-squares; the other (empty) sub-squares are discarded.
(b) Both exclusion and $\sqrt{2}$-soft inclusion criteria hold for e/i test applied to the disc bounded by internal circle.
Figure 1:
Observation 2.1.
A subdivision iteration applied to a square with $m$ roots (i) decreases by twice the diameter of a suspect square and (ii) processes at most $4m$ suspect squares, (iii) whose centers approximate every root within the distance of one half of their diameter.
Proof.
Readily verify claim (i). Claim (ii) holds because a root can lie in at most four suspect squares, and even in at most two unless it is the shared vertex of the four. Claim (iii) holds because a subdivision iteration never discards a square containing a root. ∎
Corollary 2.1.
Let subdivision iterations be applied to a square $S$ with side length $\Delta$ and diameter $\Delta\sqrt{2}$ containing precisely $m$ roots. Then (i) the $k$-th iteration outputs at most $4m$ suspect squares whose centers approximate within $\epsilon:=\Delta/2^{k-0.5}$ all the $m$ roots lying in $S$, so that $k\leq\lceil\log_{2}(\frac{\Delta}{2\epsilon}\sqrt{2})\rceil$, and (ii) the first $k$ iterations together involve at most $4mk$ e/i tests.
(B) Seeking roots in any fixed domain $\mathbb{D}$ in the complex plane, we apply subdivision iterations to its minimal covering square $S$ and while performing them also discard all squares that do not intersect the domain $\mathbb{D}$, even if they contain roots.
Then the $(\Delta_{k}\sqrt{2})$-neighborhood of $\mathbb{D}$ covers the union $\mathbb{U}_{k}$ of all suspect squares defined after the $k$th subdivision iteration, while $\Delta_{k}=\Delta_{0}/2^{k}$ fast converges to 0 as $k$ increases. Hence the neighborhood fast converges to the domain $\mathbb{D}$: the width of the neighborhood decreases by a factor of $2^{k}$ in $k$ subdivision iterations, still at the price of performing at most $4km$ e/i tests.
Remark 2.1.
[Real and near-real root-finding.] In the case of Problem 0 of Example 1.2, generalize bisection as follows: (i) recursively partition (subdivide) every suspect rectangle into two congruent rectangles, to each apply e/i test, and then discards a rectangle if it contains no roots but call it suspect and subdivides it otherwise, (ii) stop where the rectangles have diameter at most $2\epsilon$, so that their centers approximate the roots within $\epsilon$.
2.2 Subdivision iterations with soft e/i tests
Instead of testing whether a suspect square $S$ contains some roots, Henrici in [52] tested whether the minimal disc $D(c,\rho)$ covering $S$ does this. Furthermore, in the presence of rounding errors no e/i test can decide whether a root lies in or outside a disc if it actually lies on or very close to its boundary circle. To address both problems we fix $\sigma>1$ and apply soft e/i tests.
Definition 2.1.
[See the end of Sec. 6 for a generalization, covering in particular soft $m$-tests.] Given a real $\sigma>1$, a polynomial $p$, a square $S$, and its minimal covering disc $D(c,\rho)$, the $\sigma$-soft e/i test reports either exclusion or $\sigma$-soft inclusion. Exclusion certifies that the square $S$ contains no roots and therefore should be discarded from subdivision process. $\sigma$-soft inclusion certifies that the disc $D(c,\sigma\rho)$, said to be the $\sigma$-cover of $S$, contains a root; then the square $S$ is said to be suspect and is subdivided into four congruent sub-squares in the next subdivision step. Criteria for both exclusion and $\sigma$-soft inclusion can hold simultaneously (see Fig. 1b), but the test ends as soon as it verifies any of them.
Observation 2.2.
Let $\gamma(\sigma)$ denote the maximal number of $\sigma$-covers at a fixed subdivision step that share a complex point, e.g., a root. Then $\gamma(\sigma)$ is in $O(1)$ if so is $\sigma$, and $\gamma(\sigma)\leq 4$ if $\sigma<\sqrt{2}$.
Corollary 2.2.
Cor. 2.1 still holds if e/i tests are replaced with $\sigma$-soft e/i tests for $\sigma<\sqrt{2}$, while a subdivision root-finder using $\sigma$-soft e/i tests involves $O(km)$ such tests if $\sigma$ is a constant.
3 Fast soft e/i tests
3.1 Background: Newton’s Inverse Ratio and its approximation
We reduce e/i tests of subdivision root-finders to the multipoint evaluation of the ratio
${\rm NIR}(x)=\frac{p^{\prime}(x)}{p(x)}=\sum_{j=1}^{d}\frac{1}{x-z_{j}},$ (3.1)
which we call Newton’s inverse ratio. To prove the latter well-known and much used equation, differentiate $p(x)=p_{d}\prod_{j=1}^{d}(x-z_{j})$.
The arithmetic operations (hereafter ops) involved in our evaluation of NIR$(x)$ are more numerous than the other ops involved in our root-finders, even in the case where $p(x)=x^{d}$.
By following [76] we fix a complex $\delta$, with $|\delta|$ small in context, and approximate $p^{\prime}(x)$ with the first order divided difference $\frac{p(x)-p(x-\delta)}{\delta}$, equal to $p^{\prime}(y)$ for some $y\in[x-\delta,x]$, by virtue of Taylor-Lagrange’s formula. Hence $y\mapsto x$ and $p^{\prime}(y)\mapsto p^{\prime}(x)$ as $\delta\mapsto 0$, and
${\rm NIR}(x)\approx\frac{p(x)-p(x-\delta)}{p(x)\delta}=\frac{p^{\prime}(y)}{p(x)}~{}{\rm for~{}some}~{}y\in[x-\delta,x].$ (3.2)
We deduce the same bound $O(b\log(d))$ of [76] on computational precision of our root-finders, but its derivation is greatly simplified versus [76] because we only use the first order derivatives and the first order divided differences.
Unlike [76] we estimate bit-operation cost of our root-finders and allow random choice of $x$ to control the value of the ratio $\frac{|p(x)-p(x-\delta)|}{p(x)\delta}$ with a high probability (whp). This is Las Vegas randomization, and so is any randomization in polynomial root-finding, as we show next.
3.2 Background: Error detection and correction
A randomized root-finder can lose some roots, although with a low probability, but we can detect such a loss at the end of root-finding process, simply by observing that among the $m\leq d$ roots lying in an input disc only $m-w$ tame roots, say, $z_{w+1},z_{w+1},\dots,z_{m}$ have been closely approximated, [^6] We call a complex $c$ a tame root for an error tolerance $\epsilon$ if $c$ lies in an isolated disc $D(c,\epsilon)$, whose index #$\#(D(c,\epsilon))$ we can readily compute (cf. Sec. 23.4). while $w>0$ wild roots, say, $z_{1},z_{2},\dots,z_{w}$ remain at large. We can compute the coefficients of the wild factor $q(x):=p(x)/\prod_{j=w+1}^{m}(x-z_{j})$, say, by applying evaluation-interpolation techniques, then apply to $q(x)$ the same or another root-finder, and repeat this computation recursively until we approximate all $m$ roots in the disc. [^7] By this recipe we detect output errors at the very end of computations in any root-finder that can only lose roots but cannot wrongly generate them. Our subdivision root-finders belong to this class, but with them we can detect the loss of a root even earlier – whenever in a subdivision process applied to a square with $m$ roots we detect that $m$ exceeds the sum of the indices of all suspect squares, that is, the overall number of roots contained in them.
Readily bounds the complexity and error probability of this recursive process:
Proposition 3.1.
Let the above recipe for error detection and correction be applied to the output of a randomized root-finder performing at the cost $\alpha$ (under a fixed model), let it be correct with a probability $1/\beta$ for $\beta>1$ and let the algorithm be applied recursively, each time with error detection at the cost $\delta$. Then in $v$ recursive applications, the output is certified to be correct at the cost of less than $(\alpha+\delta)v$ with a probability at least $1-1/\beta^{v}$.
Explicit deflation can blow up the coefficient length and destroy sparsity unless the degree $w$ is sufficiently small, but we avoid these mishaps by combining a fixed black box polynomial root-finder with implicit deflation, that is, by applying that root-finder to the quotient polynomial $q(x)$ without computing its coefficients – based on the following expression, implied by (3.1):
$\frac{q^{\prime}(x)}{q(x)}=\frac{p^{\prime}(x)}{p(x)}-\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}.$ (3.3)
One can accelerate low precision numerical approximation of the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ simultaneously at many points $x$ by applying Fast Multipole Method (FMM) of [50, 28, 32].
3.3 Cauchy-based root-counting and e/i tests
3.3.1 Introductory comments
Let #$\#(\mathcal{D})$ denote the number of roots lying in a complex domain $\mathcal{D}$ with the boundary $\Gamma$ and recall from [1] that
[…]
that is, #$\#(\mathcal{D})$ is the average of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ on the boundary $\Gamma$.
Having an approximation of the integral of (3.4) within less than 1/2, we can round it to the closest integer and obtain $s_{0}=\#(D(0,1))$. This is more than enough for our e/i tests, for which we only need to know whether $s_{0}=0$ or not.
3.3.2 Cauchy-based soft e/i tests: outline
In Part II we approximate the integral of (3.4) with finite sums made up of the scaled values of NIR$(x)$ at the scaled $q$th roots of unity. Namely, for a fixed $q>1$ let $\zeta:=\zeta_{q}$ denote a primitive $q$th root of unity:
$\zeta:=\zeta_{q}:=\exp\Big{(}\frac{2\pi{\bf i}}{q}\Big{)},~{}{\bf i}:=\sqrt{-1},$ (3.5)
and devise randomized $\sigma$-soft e/i tests as follows:
(i) Choose a real $\rho$ at random in a fixed range covered by $[1,\sigma]$.
(ii) Fix a sufficiently large integer $q$ of order $m\log(d)$.
[…]
(iv) Conclude that exclusion holds whp if $|s_{0,q}|<1/2$.
(v) Repeat the test $v$ times to decrease the probability of its output error exponentially in $v$, from $\frac{1}{\gamma}$ to $\frac{1}{\gamma^{v}}$ for $\gamma>1$ of our choice, say, for $\gamma=2$.
The $v$ tests are reduced to the evaluation of $p(x)$ at $O(mv\log(d))$ points.
3.4 Soft e/i test with root-lifting: generic algorithm
In Part III we decrease the complexity estimates of Part II for $q$ in e/i tests by a factor of $m$ by avoiding approximation of the integer #$\#(D(0,1))$ and only deciding whether this integer is 0 or not, which is sufficient for e/i tests. In these e/i tests we elaborate upon the generic one (see Alg. 3.1 below), where we use $h$-lifting of roots:
$p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{h}(x):=\prod_{g=0}^{h-1}p_{0}(\zeta_{h}^{g}x)=(-1)^{d}\prod_{j=1}^{d}(x-z_{j}^{h}),~{}\zeta_{h}:=\exp\Big{(}\frac{2\pi{\bf i}}{h}\Big{)},$ (3.6)
for $h=1,2,\dots$ and $t_{2^{k}}(x)=p_{k}(x)$ for all $k$. $h$-lifting maps roots to their $h$th powers; this keeps the unit disc $D(0,1)$ and its boundary circle $C(0,1)$ intact but moves the external roots away from the circle, thus strengthening its isolation from those roots.
Remark 3.1.
The $h$-lifting generalizes the DLG classical root-squaring iterations of Dandelin 1826, Lobachevsky 1834, and Gräffe 1837 (cf. [51]):
$p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{2^{i}}(x^{2}):=(-1)^{d}p_{2^{i-1}}(x)p_{2^{i-1}}(-x)=\prod_{j=1}^{d}(x-z_{j}^{2^{i}}),~{}i=1,2,\dots.$ (3.7)
Polynomial root-finding in the 19-th and well into the 20-th century relied on iterations (3.7) combined with Vieta’s formulas, which link the polynomial coefficients $p_{0},p_{1},\dots,p_{d}$ to the elementary symmetric polynomials of the roots. The combination enabled efficient approximation of the minimal and maximal distance of the roots from the origin and due to the map (1.3) from any complex point (see [86, 113, 115] for details). The approach became obsolete because of severe problems of numerical stability of the DLG iterations, but they disappeared in the case of root-finders that work for black box polynomials. Extensive tests with standard test polynomials have consistently demonstrated efficiency of the particular implementation of this approach in [115], but even the straightforward implementation of iterations (3.7) promises to be highly efficient as well.
We will elaborate upon the following generic algorithm in Sec. 14.
Algorithm 3.1.
A generic soft e/i test.
INPUT: $\sigma>1$, an integer $d>0$, and a $d$-th degree black box polynomial $p(x)$ given by an oracle for its evaluation.
INITIALIZATION: Fix a positive $q\in\mathbb{N}$ and a positive integer $h$.
COMPUTATIONS:
(1) Select $q$ points $x_{1},\dots,x_{q}$ on the circle $C(0,1)$.
(2) By extending (3.2) to $p_{h}(x)$ of (3.7) (we specify this extension in (13.3)) approximate the values
$v_{g}:=|{\rm NIR}_{p_{h}}(x_{g})|=\Big{|}\frac{p^{\prime}_{h}(x_{g})}{p_{h}(x_{g})}\Big{|},~{}g=1,\dots,q.$ (3.8)
(3) If $1+d/v_{g}\leq\sigma^{h}$ for some $g\leq q$, then certify that #$(D(0,\sigma))>0$, that is, claim $\sigma$-soft inclusion. Otherwise, claim that #$(D(0,1))=0$, that is, claim exclusion.
3.5 Soft e/i test with root-lifting: correctness of inclusion claim
Eqn. (3.1) implies two useful corollaries:
Corollary 3.1.
Map (1.3) scales NIR by $\rho$.
Proof.
Eqn. (3.1) implies that NIR$(x-c)=$NIR$(x)$ and that NIR$(x/a)=a$NIR$(x)$. ∎
Corollary 3.2.
It holds that $r_{d}(c)\leq d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}$, that is, #$\#\Big{(}D(c,d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}\Big{)}\Big{)}>0$, for any complex $c$.
Proof.
Let $x=c$ in Eqn. (3.1). If $|c-z_{j}|<\rho$ for all $j$, then
$\Big{|}\frac{p^{\prime}(c)}{p(c)}\Big{|}=\Big{|}\sum_{j=1}^{d}\frac{1}{c-z_{j}}\Big{|}\leq\sum_{j=1}^{d}\frac{1}{|c-z_{j}|}<\frac{d}{\rho},~{}{\rm and~{}so}~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}>\frac{\rho}{d}.$
We use Cor. 3.1 later but next apply Cor. 3.2 to prove correctness of the inclusion claim of Alg. 3.1. The bound of Cor. 3.1 on $r_{d}(c)$ is too large for that, but a similar bound on $r_{d}(c,p_{h})=(r_{d}(c))^{h}$, implies that $r_{d}(c)$ is close to 1 already for $h$ of order $\log(d)$.
Lemma 3.1.
Let $\rho:=d\Big{|}\frac{p_{h}(x)}{p^{\prime}_{h}(x)}\Big{|}$ and $\sigma:=(|x|+\rho)^{1/h}$ for a positive integer $h$. Then #$(D(0,\sigma))>0$.
Proof.
Apply Cor. 3.2 to the polynomial $p_{h}(x)$ and obtain that the equation $\rho=d\Big{|}\frac{p_{h}(x_{)}}{p^{\prime}_{h}(x)}\Big{|}$ implies that the disc $D(x,\rho)$ contains a zero of $p_{h}(x)$. Hence the disc $D(0,|x|+\rho)$ contains a zero of $p_{h}(x)$ by virtue of triangle inequality. Therefore, the disc $D(0,\sigma)$ contains a zero of $p(x)$ for $\sigma=(|x|+\rho)^{1/h}$. ∎
Apply this lemma for $x=x_{i}$, $|x_{i}|=1$, $\rho=d/v_{i}$, and $v_{i}=|{\rm NIR}_{p_{h}}(x_{i})|$ (cf. Alg. 3.1). Deduce that #$(D(0,\sigma))>0$ for $\sigma^{h}=1+d/v_{i}$, that is, Alg. 3.1 is correct when it claims $\sigma$-soft inclusion.
Observation 3.1.
$\sigma$-soft inclusion test by means of Alg. 3.1 is correct where
$d\geq 2,~{}x_{1}=1,~{}\frac{1}{|{\rm NIR}(1)|}=4\sqrt{d}-\frac{1}{d},~{}h=40\log_{2}(d),~{}{\rm and}~{}\sigma<1.1.$
Proof.
Indeed, in this case
$1+\rho=1+\frac{d}{|{\rm NIR}(1)|}=4d\sqrt{d}=2^{2+1.5\log_{2}(d)}.$
Apply Lemma 3.1 and obtain that #$(D(0,\sigma))>0$ for $\sigma=(1+\rho)^{1/h}=2^{\nu}~{}{\rm and}~{}\nu=\frac{2+1.5\log_{2}(d)}{40\log_{2}(d)}$.
Deduce that $\nu\leq\frac{7}{80}$ because $\frac{2}{40\log_{2}(d)}\leq 0.05$ for $d\geq 2$ and because $\frac{2+1.5\log_{2}(d)}{40\log_{2}(d)}<\frac{3}{80}$.
Hence $\sigma\leq 2^{\frac{7}{80}}=1.062527\dots<1.1$. ∎
3.6 Soft e/i test with root-lifting: further analysis (outline)
(i) Correctness of the exclusion claim of Alg. 3.1. We prove it separately for its two specifications, both of which strengthen isolation of the unit circle $C(0,1)$ by means of root-squaring (3.7) or root-lifting (3.6).
(a) In Sec. 14.3 we choose any integer $q$ exceeding $d$ and then prove deterministic correctness of exclusion claim, provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(2\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity. This bound is compatible with the bound of Example 3.1 because $4\sqrt{d}-\frac{1}{d}>3\sqrt{d}$ for $d\geq 2$. We also extend this deterministic e/i test by allowing any $q>m$ (rather than $q>d$) provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(3\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity.
(b) In Sec. 14.4 we sample $q=O(\log(d))$ iid random values of $x_{i}$ under the uniform probability distribution on the circle $C(0,1)$ and then prove that the exclusion claim is correct whp under the Random Root Model if $|{{\rm NIR}_{p_{h}}(x_{i})}|\leq 1/d^{a}$ for a sufficiently large constant $a$.
(ii) Counting query locations. Eqn. (3.7) reduces approximation of $p_{h}(x)$ at $x=c$ to approximation of $p(x)$ at all the $h$th roots of $c$. We can fix $h>1$ being a power of 2 and apply (3.7) or fix any integer $h>1$ and apply (3.6). In both cases Alg. 3.1 evaluates polynomial $p(x)$ either at $2jh$ points for $j\leq q$, when it claims inclusion, or at $2qh$ points, when it claims exclusion. We always choose $h=O(\log(d))$ and then can ignore a factor of $h$ in our $\tilde{O}$ estimates.
In our deterministic soft e/i test based on Alg. 3.1 we evaluate polynomial $p(x)$ at $\tilde{O}(m)$ points, combine these estimates with Cor. 2.2, and bound by $\tilde{O}(bm^{2})$ the overall number of points $x$ at which we evaluate $p(x)$ in subdivision root-finders incorporating our e/i tests.
Under the Random Root Model of Sec. 1.5 the roots are reasonably well isolated whp from the unit circle $C(0,1)$. We strengthen this isolation by applying $O(\log(\log(d)))$ root-squaring steps (3.7) or $O(\log(d))$ steps of root-lifting of (3.6) and then whp decrease the overall bit operation count from $\tilde{O}(bm^{2})$ to $\tilde{O}(bm)$.
(iii) Estimation of computational precision and bit-operation complexity. Louis and Vempala in [76] bound by $O(b\log(d))$ the computational precision of their novel high order Newton’s iterations. We combine a small part of their advanced analysis with Las Vegas randomization and some customary techniques of error analysis to extend this estimate to our root-finders.
For the solution of Problems 0, 1, and 3, where the coefficients are given, we apply fast algorithm of [81] for multipoint polynomial evaluation to decrease the estimated number of queries and the bit operation cost bound.
We devise two groups of efficient e/i tests in Parts II and III of this paper, and in both cases arrive at the estimates for the numbers of (i) evaluation points, that is, of the oracle queries, and (ii) bit-operations involved in our algorithms. Both of the estimates exceed those of our Main Theorems by factors of $b$ for the roots and $\tilde{b}$ for the eigenvalues.
4 Acceleration of subdivision root-finding by means of compression: outline
In Part IV we accelerate the algorithms of both Parts II and III by these factors (of $b$ for roots and $\tilde{b}$ for eigenvalues) to prove the Main Theorems in two ways. Next we outline that acceleration.
(i) Component tree and compact components. Apply subdivision process for a fixed $\sigma$, $1<\sigma<\sqrt{2}$, to Problem 1 and partition the union of all suspect squares tested at every subdivision step into connected components.
Represent all partitions of components with a component tree, $\mathbb{T}$, as follows. At most $m$ components, said to be the leaves, are covered by isolated $\epsilon$-discs for a fixed error tolerance $\epsilon\geq 0$. The other (at most $m-1$) vertices of the tree are components-parents, each broken by a subdivision step into at least two components-children. The initial suspect square is the root of the tree. The tree has at most $22m-$ 2 edges, connecting parents with their children and representing partition of the components but not showing modification of components between their partitions.
Recall that every component contains $O(m)$ suspect squares whose diameter is halved at every subdivision step. Hence in $O(\log(b))$ subdivision iterations, for $b:=\log_{2}(1/\epsilon)$, every component has diameter below $\epsilon/2$, and then its $\sigma$-cover lies in an $\epsilon$-disc.
Call a component compact if it consists of at most four squares sharing a vertex (see Fig. 2). Recall that a subdivision step never includes new roots into a component and deduce that unless a component stays compact it is soon broken into more than one component or shrinks enough to be covered by an $\epsilon$-disc.
[…]
Figure 2: Roots (asterisks) define compact components. A subdivision step halves their diameters and accordingly strengthens isolation of their minimal covering squares and discs.
(ii) Acceleration of subdivision process with compression of components. Subdivision process becomes quite inefficient where a component stays compact and not broken into more than one component in many consecutive subdivision steps (see Fig. 3). To accelerate it, we replace these steps with compression of such a component into an $\epsilon$-disc or a near-minimal disc $D$ that still keeps all roots lying in the component. In Sec. 19 we comment on two customary but inferior alternatives to our compression, relying on deflation or solely on functional iterations.
Figure 3: Five blue marked centers of suspect squares converge to a green-marked root cluster with linear rate; two red marked Schröder’s iterates of (23.1) converge with quadratic rate.
If the compression outputs $\epsilon$-disc, we just output it and do not subdivide it anymore. Otherwise we apply subdivision iterations to the minimal superscribing square $S$ of the disc $D$. Since the disc is near-minimal, the iterations very soon output more than one component or a sub-domain of an $\epsilon$-disc. From this we deduce that our subdivision process with compression reduces the solution of Problem 1 to at most $22m-$ 2 compression steps and $O(m)$ soft e/i steps overall.
(iii) Compression of components. We reduce compression of a compact component $C$ to the following two main stages (see more details in Sec. 23.5):
(a) compute a point $c^{\prime}$ lying in or near the minimal disc that covers all roots lying in the component $C$ and let $c^{\prime}$ denote the center of the compressed disc that we seek;
(b) compute a close upper bound $\rho^{\prime}$ on the maximal distance from $c^{\prime}$ to the roots in the component $C$ and output a compressed disc $D(c^{\prime},\rho^{\prime})$.
We prove that Schröder’s single iteration (23.1) following $O(\log(d))$ auxiliary subdivision steps performs step (a) above, that is, computes the desired center $c^{\prime}$.
At the beginning of step (b) above we have a range $[\rho_{-},\rho_{+}]$ for the bound $\rho^{\prime}$, $0<\epsilon\leq\rho_{-}<\rho_{+}\leq 1$. Hence we can approximate it in $\lceil b\rceil$ bisection steps, for $b=\log_{2}(1/\epsilon)$ but we greatly accelerate convergence to $\rho$ by applying bisection to the range of the logarithms, $[\log_{2}(\rho_{-}),\log_{2}(\rho_{+})]=[b,\log_{2}(\rho_{+})]$, for $\log_{2}(\rho_{+})\leq 0$.
We extend our fast soft e/i tests of Part II to fast soft $m$-tests (see part (A) of Sec. 2.1) and use them to define the pivot step of such a modified bisection.
Summarizing we reduce the solution of Problems 0 and 1 to $\tilde{O}(m)$ soft e/i tests and $m$ tests and $\tilde{O}(m)$ Schröder’s iterations. In this way we accelerate the root-finders of both Parts II and III by a factor of $b/\log(b)$ to complete the proof of our Main Theorems.
Remark 4.1.
Many “natural" directions to proving our Main Theorems have led us into dead end. E.g., we defined a tree whose vertices represented non-adjacent suspect squares; then we readily bounded by $12m-$ 1 their overall number at all partition stages but could not count them so readily between these stages until we decided to count all suspect squares. See more examples of this kind in the next remark and Remark 14.1.
Remark 4.2.
By using some of the advanced techniques of Louis and Vempala in [76], we readily bound the precision of computing in our root-finders. This enables us to deduce a priori estimates for the bit operation cost of our subdivision root-finders versus inferior known a posteriori estimates (see, e.g., [122, 24]). In particular, the bit operation cost estimates of the Main Theorem of [24] contain the term
$\log(1/{\rm Discr}(p(x))),$
with “Discr" standing for “Discriminant", such that
$\log(1/{\rm Discr}(p(x/2^{b})))=(d+1)db\log(1/{\rm Discr}(p(x))).$
Subdivision iterations of the known root-finders, in particular, of those of [24], scale the variable $x$ order of $m$ times by factors of $2^{b}$. Hence the bit-complexity specified in the Main Theorem of [24] has order of $mbd^{2}$, which exceeds our bound and that of [96, 103] by a factor of $d$.
5 Our novelties, related works, implementation, and further progress
5.1 Our novelties (a short list)
Root-finding for polynomial given in the monomial basis, with their coefficients, has been studied for four millennia and very intensively in the last decades, but we had to introduce a number of nontrivial novel techniques (i) to handle efficiently black box polynomials and (ii) to accelerate the known subdivision algorithms to a near optimal level.
Some of our techniques such as randomization, numerical stabilization of root-squaring, compression of a disc without losing roots, bisection of logarithms, approximation of root radii and near-real roots, and even our simple error detection recipe can be of independent interest.
5.2 Related works
The pioneering paper [76] proposed a novel algorithm for approximation of an absolutely largest zero of a real-rooted black box polynomial, estimated the computational precision of that algorithm, and extended it to approximation of an absolutely largest eigenvalue of a symmetric matrix. From that paper we adopt the model of computation, extension to approximation of matrix eigenvalues, and some patterns for estimation of the computational precision of root-finders, but otherwise use distinct approach and techniques.
Namely, instead of Newton’s iterations in [76], we accelerate the classical subdivision root-finders, based on e/i tests. Its closest variations in [128, 101, 23, 24] still use inferior e/i tests, for which we proposed two fast novel approaches – based on approximation of Cauchy integrals and on root-lifting; these tests run much faster than the tests in the cited papers and, unlike the latter tests, can be applied to a black box polynomial. The papers [128, 101, 23, 24] apply functional iterations (such as Newton’s and Schröders) to accelerate convergence of subdivision process but we apply our novel techniques of disc compression instead.
Estimating the bit operation complexity of our root-finders we use [134, Eqn. (9.4)] and [81, Thm. 2] for multipoint polynomial evaluation. [^8] The ingenious algorithm of Moroz supporting his [81, Thm. 2] departs from the simple observation that a polynomial $p(x-c)\mod(x-c)^{\ell-1}$ approximates $p(x)$ in the disc $D(c,\rho)$ within $O(\rho^{\ell+1})$ for small $\rho$. To approximate $p(c)$ for any $c\in D(0,1)$ Moroz used $\tilde{O}((\ell+\tau)d)$ bit operations to compute a set of complex $c_{i}$ and positive $\rho_{i}$ for $i=1,\dots,K$ and $K=O(d/\ell)$ such that $D(0,1)\subseteq\cup_{i=1}^{K}D(c_{i},\rho_{i})$; then he ensured approximation to $p(c)$ for every point $c\in D(0,1)$ within $\epsilon:=3\sum_{{}_{i}=0}^{d}|p_{i}|/2^{\ell}$ by means of approximation within $\epsilon/2$ of the lower degree polynomial $p(x-c_{i})\mod(x-c)^{\ell+1}$ for an appropriate $i$. At this point, application of [68, Alg. 5.3 and Thm. 3.9] yields new record fast algorithm for multipoint approximation of a polynomial given with its coefficients. The root-finders of Moroz in [81] and Imbach and Moroz in [59] use $\tilde{O}(bd)$ bit operations for polynomials given with their coefficients and having only well-conditioned zeros. In spite of the appeal to the concept of condition number, used for the design and analysis of polynomial root-finding in [14, 30, 22], the upper bound on the condition numbers of all zeros of $p(x)$ in [81, 59] is a little stronger restriction than the bound on the cluster sizes of Def. 24.1, under which [96, 103] proved the same complexity bound. Unlike our present root-finders, the root-finders of [81, 59] are not efficient for the worst case inputs and extensively use the coefficients.
5.3 Implementation: State of the Art
The near-optimal but quite involved algorithm of [96, 103] has never been implemented, while the slower algorithm of [134] has been implemented by Xavier Gourdon in [46] for the Magma and PARI/GP computer algebra systems. By now it is superseded by implementation of various non-optimal complex polynomial root-finders such as MATLAB “roots”, Maple “solve”, CCluster, and MPSolve; the latter one relies on Ehrlich’s iterations and is user’s current choice for approximation of all $d$ complex roots (in Sec. 26.4 we list some competitive implementations of Newton’s iterations).
Current user’s choice for real root-finding is the algorithm of [70], although it has already been superseded in [62] according to extensive tests for standard test polynomials.
We present algorithms that are rather simple for implementation and promise to compete for user’s choice. Already incorporation in 2020 of some of our techniques into the previous best implementation of subdivision root-finder [64] has enabled its noticeable acceleration (see [60]). The more advanced publicly available implementation in [63] “relies on novel ideas and techniques" from [112] (see [63, Sec. 1.2]). According to the results of extensive tests with standard test polynomials, the algorithm has performed dramatically faster than [64], has at least competed with user’s choice package MPSolve even for approximation of all $d$ zeros of a polynomial $p$ given with its coefficients, [^9] The computational cost of subdivision root-finders in a domain $\mathbb{D}$ as well as the root-finder of [96, 103] decreases at least proportionally to the number of roots in $\mathbb{D}$, while Ehrlich’s, Weierstrass’s, and a number of other functional iterations approximate the zeros of $p$ in a domain $\mathbb{D}$ almost as slow as all its $d$ complex roots. Thus already implementation in [64] of subdivision iterations, slightly outperformed MPSolve for root-finding in a domain $\mathbb{D}$ containing a small number of roots although is by far inferior for the approximation of all $d$ roots. and tended to supersede MPSolve more and more significantly as the degree $d$ of $p$ grew large.
There is still a large room for further improvement because the root-finder of [63] inherits some earlier sub-algorithms from [23, 64], which are slower by order of magnitude than our current ones.
Next we comment on some promising directions to further improvement of [63], but also notice that adjustment of our near-optimal subdivision root-finder in Remark 2.1 to approximation of real and near-real roots and consequently of matrix eigenvalues should be at least competitive with the current user’s choice [70] for real root-finding as well as with its significant improvement in [62]; moreover, it more efficiently counters the impact of the input and rounding errors.
5.4 Further progress and promise
(A) Acceleration of e/i tests. E/i tests are basic ingredient of subdivision root-finders. [60, 63] used the Cauchy-based e/i tests of our Part II and their heuristic acceleration. It is still a challenge to compare their empirical performance with that of the formally supported e/i tests of our Part III and Sec. 23.4, based on root-lifting. Furthermore, we can try to exploit potential benefits of fast numerical approximation of Cauchy integrals with the Exponentially Convergent Adaptive Trapezoidal Quadrature (see [146]), efficiently implemented in Boost C++ libraries.
(B) Acceleration of the compression of a disc. We assume that the unit disc $D(0,1)$ contains $m$ roots and is isolated from the external $d-m$ roots, and we compress that disc towards the minimal disc containing the same $m$ roots. We reduce this task essentially to (i) computing a point in or near the convex hull of the root set in $D(0,1)$ and (ii) approximation of the $m$th smallest root radius from a given point in $D(0,1)$. For task (i) we dramatically accelerate our algorithm of [60, 63] because our algorithms of Secs. 23.2 and 23.3 for task (i), using Schröder’s iterations, deterministic, and formally supported, run dramatically faster than those of [60, 63], based on approximation of the average of all roots in $D(0,1)$. For task (ii) we can apply our Las Vegas randomized algorithms of Secs. 8 and 9 combined or the alternative of [60, 63], based on the algorithm of our Sec. 15.1, which reduces this task to performing a constant number of e/i tests. Both approaches support the same asymptotic cost bound up to polylog factors, but fro $m\gg\log^{2}(d)$ the second approach is faster by a factor of $m/\log^{2}(d)$ where it uses fast heuristic e/i tests. Such an improvement, however, may be lost in practice because of the impact of polylog factors. Implementation of compression in [60, 63] is slowed down with verification based on Pellet’s theorem, but we can drop this verification because we readily detect highly unlikely output errors of our Las Vegas root-finders.
(C) Fast approximation of the eigenvalues of a matrix. [20] explored reduction of approximation of the eigenvalues of a matrix to root-finding by means of Ehrlich’s iterations, reduced to recursive computations of NIR$(x)$, and showed benefits of this approach. We can readily extend this reduction and its benefits by using subdivision iterations instead of Ehrlich’s. Furthermore, win the case of subdivision we obtain additional benefits, e.g., our algorithms are much faster for approximation of a small number of eigenvalues that lie in a fixed domain and unlike Ehrlich’s iterations isolate eigenvalues, rather than just approximate them.
(D) Acceleration of real and near real root-finding. In Remark 2.1 we adjust our near-optimal subdivision iterations to real and near-real root-finding, which counters the impact of input and rounding errors and can be readily extended to approximation of the eigenvalues of a matrix.
(E) Acceleration by using fast multipoint polynomial evaluation. Mere incorporation of fast algorithms of [81, 59] should greatly accelerate the performance of [60, 63] in the case of root-finding for polynomials given with their coefficients.
5.5 Organization of the rest of the paper
In Parts II and III we devise and analyze our two alternative e/i tests. Deterministic e/i test of Part III is as fast as the randomized one of Part II. We use Las Vegas randomization in both parts to bound computational precision and bit-operation cost, and in Part III we yield acceleration by a factor of $d$ under the Random Root Model. The bit-operation complexity bounds of Parts II and III are inferior by a factor of $b$ to that of [96, 97, 103], but we match it in Part IV by means of our acceleration of the algorithms of both Parts II and III based on compression of a disc without losing its roots. In Part V we cover some important root-finding subjects that involve no substitution iterations, namely, factorization, root isolation, and functional iterations. In Secs. 24 and 25 we outline the fundamental results of Schönhage from [134] on polynomial factorization, its links to root-finding and root-isolation, and some relevant bit-complexity estimates, including lower bounds. In Sec. 26 we discuss root-finding by means of functional iterations, namely, Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s), and their acceleration by means of incorporation of Fast Multipole Algorithm (FMM) and fast multipoint polynomial evaluation as well as their initialization both where the coefficients of $p$ are unknown or given to us.
5.6 Independence of Part III from Part II
In Parts II an III we solve Problems 0, 1, 0, and 1 in two different ways. In Part II we rely on approximation of a Cauchy integral associated with root-counting in a disc on the complex plane. In Part III we rely on lifting the roots to a fixed power $h=O(\log(d))$. In addition, in Part III, unlike Part II, we devise a faster root-finder under a random root model. We make our presentation in Part III independent of Part II; in particular we include again some definitions, figures, and theorems.
PART II: Classical subdivision root-finder with novel exclusion/inclusion (e/i) tests based on approximation of Cauchy integrals
Organization of the presentation in Part II: We devote the next section to some basic definitions. In Sec. 7 we study root-counting in a disc on the complex plane by means of approximation of a Cauchy integral over its boundary circle by finite sums. [^10] We do not use the results of Sec. 7.5, but they can be of independent interest. In Sec. 8 we apply this study to devise and analyze root-counters and e/i tests in a disc on the complex plane. In Sec. 9 we extend our e/i tests to estimating root radii, that is, the distances from a fixed complex point to the roots. In Secs. 10 we estimate precision of computing of our root-finders. We devote short Sec. 11 to their extension to approximation of the eigenvalues of a matrix or a matrix polynomial. In Sec. 12 we estimate the bit operation complexity of our root-finders assuming that $p(x)$ is a general polynomial of (1.1) represented with its coefficients.
6 Background: basic definitions
  • “Wlog" stands for “without loss of generality".
  • $S(c,\rho):=\{x:~{}|\Re(c-x)|\leq\rho$, $|\Im(c-x)|\leq\rho\},~{}D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$,
    $C(c,\rho):=\{x:~{}|x-c|=\rho\}$, and $A(c,\rho,\rho^{\prime}):=\{x:~{}\rho\leq|x-c|\leq\rho^{\prime}\}$
    denote a square, a disc, a circle (circumference), and an annulus (ring) on the complex plane, respectively.
  • $\Delta(\mathbb{S})$, $X(\mathbb{S})$, $CH(\mathbb{S})$, and #$\#(\mathbb{S})$ denote the diameter, the root set, the convex hull, and the index (root set’s cardinality) of a set $\mathbb{S}$ in the complex plane, respectively.
  • A disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is said to be $\theta$-isolated (from external roots) for $\theta>1$ if $X(D(c,\rho))=X(D(c,\theta\rho))$, $X(C(c,\rho))=X(A(c,\rho/\theta,\rho/\theta))$, or $X(S(c,\rho))=X(S(c,\theta\rho))$, respectively.
  • The largest upper bound on such a value $\theta$ is said to be the isolation of the disc $D(c,\rho)$, the circle $C(c,\rho)$, or the square $S(c,\rho)$, respectively (see Fig. 4), and is denoted $i(D(c,\rho))$, $i(C(c,\rho))$, and $i(S(c,\rho))$, respectively.
  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex point $c$ to the zeros $y_{1}\dots,y_{d}$ of a $d$th degree polynomial $t(x)$. $r_{j}(c):=r_{j}(c,p)$, $r_{j}:=r_{j}(0)$ for $j=1,\dots,d$.
  • Generalize Def. 2.1 of $\sigma$-soft e/i tests as follows: for $p(x)$ of (1.1), $\sigma>1$, and integers $\ell$ and $m$ such that $1\leq\ell\leq m\leq d$ and #$\#(D(0,1))\leq m$, a $\sigma$-soft $\ell$-test, or just $\ell$-test for short (so that e/i test turns into 1-test), either outputs 1 and stops if it detects that $r_{d-\ell+1}\leq\sigma$, that is, #$\#(D(0,\sigma))\geq\ell$, or outputs 0 and stops if it detects that $r_{d-\ell+1}>1$, that is, [^11] Bounds $r_{d-\ell+1}\leq\sigma$ and $r_{d-\ell+1}>1$ can hold simultaneously (cf. Fig. 1b), but as soon as an $\ell$-test verifies any of them, it stops without checking if the other bound also holds. #$\#(D(0,1))<\ell$. $\ell$-testc,ρ, aka $\ell$-test for the disc $D(c,\rho)$, is an $\ell$-test applied to the polynomial $t(y)$ of (1.3).
Observation 6.1.
Eqn. (1.3) maps the zeros $z_{j}$ of $p(x)$ into the zeros $y_{j}=\frac{z_{j}-c}{\rho}$ of $t(y)$, for $j=1,\dots,d$, and preserves the index #$\#(D(c,\rho))$ and the isolation $i(D(c,\rho))$.
$\rho$$\rho\theta$$\rho\theta^{2}$c********
Figure 4: The roots are marked by asterisks. The red circle $C(c,\rho\theta)$ and the disc $D(c,\rho\theta)$ have isolation $\theta$. The disc $D(c,\rho)$ has isolation $\theta^{2}$.
7 Cauchy root-counting in a disc
7.1 The power sums of the roots in a disc
Recall the following equations generalizing (3.4) (cf. [1]):
[…]
For $\mathcal{D}=D(0,1)$ and $\zeta$ of (3.5) we will approximate $s_{h}$ with finite sums
$s_{h,q}:=\frac{1}{q}\sum_{g=0}^{q-1}\zeta^{(h+1)g}~{}\frac{p^{\prime}(\zeta^{g})}{p(\zeta^{g})}~{}{\rm for}~{}h=0,1,\dots,q-1$ (7.2)
and recall that $s_{0}(D(0,1))=\#(D(0,1))$ (cf. (3.4)), while an e/i test is just the decision whether #$\#(D(0,1))>0$.
7.2 Cauchy sums in the unit disc as weighted power sums of the roots
Cor. 7.1 of the following lemma expresses the $h$th Cauchy sum $s_{h,q}$ as the sum of the $h$th powers of the roots $z_{j}$ with the weights $\frac{1}{1-x^{q}}$.
Lemma 7.1.
For a complex $z$, two integers $h\geq 0$ and $q>1$, and $\zeta$ of (3.5), it holds that
$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=\frac{z^{h}}{1-z^{q}}.$ (7.3)
Proof.
First let $|z|<1$ and obtain
$\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=\frac{\zeta^{hg}}{1-\frac{z}{\zeta^{g}}}=\zeta^{hg}\sum_{u=0}^{\infty}\Big{(}\frac{z}{\zeta^{g}}\Big{)}^{u}=\sum_{u=0}^{\infty}\frac{z^{u}}{\zeta^{(u-h)g}}.$
The equation in the middle follows from Newman’s expansion
$\frac{1}{1-y}=\sum_{u=0}^{\infty}y^{u}~{}{\rm for}~{}y=\frac{z}{\zeta^{g}}.$
We can apply it because $|y|=|z|$, while $|z|<1$ by assumption.
Sum the above expressions in $g$, recall that
$\zeta^{u}\neq 1~{}{\rm and}~{}\sum_{s=0}^{q-1}\zeta^{u}=0~{}{\rm for}~{}0<u<q,~{}\zeta^{q}=1,$ (7.4)
and deduce that
$\frac{1}{q}\sum_{g=0}^{q-1}\frac{z^{u}}{\zeta^{(u-h)g}}=z^{u}$
for $u=h+ql$ and an integer $l$, while $\sum_{g=0}^{q-1}\frac{z^{u}}{\zeta^{(u-h)g}}=0$ for all other values $u$. Therefore
$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=z^{h}\sum_{l=0}^{\infty}z^{ql}.$
Apply Newman’s expansion for $y=z^{q}$ and deduce (7.3) for $|z|<1$. By continuity extend (7.3) to the case of $|z|=1.$
Now let $|z|>1$. Then
$\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=-\frac{\zeta^{(h+1)g}}{z}~{}~{}\frac{1}{1-\frac{\zeta^{g}}{z}}=-\frac{\zeta^{(h+1)g}}{z}\sum_{u=0}^{\infty}\Big{(}\frac{\zeta^{g}}{z}\Big{)}^{u}=-\sum_{u=0}^{\infty}\frac{\zeta^{(u+h+1)g}}{z^{u+1}}.$
Sum these expressions in $g$, write $u:=ql-h-1$, apply (7.4), and obtain
$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=-\sum_{l=1}^{\infty}\frac{1}{z^{ql-h}}=-z^{h-q}\sum_{l=0}^{\infty}\frac{1}{z^{ql}}.$
Apply Newman’s expansion for $y=z^{q}$ again and obtain
$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=-\frac{z^{h-q}}{1-\frac{1}{z^{q}}}=\frac{z^{h}}{1-z^{q}}.$
Hence (7.3) holds in the case where $|z|>1$ as well. ∎
Sum equations (7.3) for $z=z_{j}$ and $j=1,\dots,d$, combine Eqns. 3.1 and (7.2), and obtain
Corollary 7.1.
For the roots $z_{j}$ and all $h$, the Cauchy sums $s_{h,q}$ of (7.2) satisfy $s_{h,q}=\sum_{j=1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}$ unless $z_{j}^{q}=1~{}{\rm for~{}some}~{}j.$
7.3 Approximation of the power sums in the unit disc
Cor. 7.1 implies that the values $|s_{h,q-}s_{h}|$ decrease exponentially in $q-h$ if the unit disc $D(0,1)$ is isolated.
Theorem 7.1.
[134, Eqn. (12.10)]. [^12] Unlike the proof in [134], we rely on Cor. 7.1 of independent interest. Let the unit disc $D(0,1)$ be $\theta$-isolated and let $d_{\rm in}$ and $d_{\rm out}=d-d_{\rm in}$ denote the numbers of the roots lying in and outside the disc $D(0,1)$, respectively. Then
$|s_{h,q}-s_{h}|\leq\frac{d_{\rm in}\theta^{-h}+d_{\rm out}\theta^{h}}{\theta^{q}-1}\leq\frac{d\theta^{h}}{\theta^{q}-1}~{}{\rm for}~{}h=0,1,\dots,q-1.$ (7.5)
Proof.
Enumerate the roots $z_{j}$ so that $|z_{j}|<1$ if and only if $j\leq d_{\rm in}$; then Cor. 7.1 implies that
$s_{h,q}-s_{h}=s_{h,q,{\rm in}}+s_{h,q,{\rm out}}$ (7.6)
where
$s_{h,q,{\rm in}}=\sum_{j=1}^{d_{\rm in}}z_{j}^{h}\Big{(}\frac{1}{1-z_{j}^{q}}-1\Big{)}~{}{\rm and}~{}s_{h,q,{\rm out}}=\sum_{j=d_{\rm in}+1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}=\sum_{j=d-d_{\rm out}+1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}.$ (7.7)
Notice that $\frac{1}{1-z_{j}^{q}}-1=\frac{z^{q}}{1-z_{j}^{q}}$, and so $|z_{j}^{h}(\frac{1}{1-z_{j}^{q}}-1)|\leq\frac{\theta^{-h}}{\theta^{q}-1}$ for $z_{j}\in D(0,1)$, while $\Big{|}\frac{z_{j}^{h}}{1-z_{j}^{q}}\Big{|}\leq\frac{\theta^{h}}{\theta^{q}-1}$ for $z_{j}\notin D(0,1)$. Combine these bounds with Eqns. (7.6) and (7.7) and obtain (7.5). ∎
By choosing $q$ that exceeds $\log_{\theta}(2d+1)$ we ensure that $|s_{0}-s_{0,q}|<1/2$.
Example 7.1.
Let the unit circle $C(0,1)$ be 2-isolated. Then Thm. 7.1 implies that $|s_{0,q}-s_{0}|<1/2$ for $d\leq 1,000$ if $q=10$ and for $d\leq 1,000,000$ if $q=20$.
Algorithm 7.1.
INPUT: a black box polynomial $p$ of a degree $d$ and $\theta>1$.
INITIALIZATION: Compute the integer $q=\lfloor\log_{\theta}(4d+2)\rfloor>\log_{\theta}(2d+1)$.
COMPUTATIONS: Compute Cauchy sum $s_{0,q}$ and output an integer $\bar{s}_{0}$ closest to it.
Observation 7.1.
(i) Alg. 7.1 evaluates $p(x)$ at $q=\lfloor\log_{\theta}(4d+2)\rfloor$ points. (ii) It outputs $\bar{s}_{0}=\#(D(0,1))$ if the circle $C(0,1)$ is $\theta$-isolated for $\theta>1$. (iii) If the algorithm outputs $\bar{s}_{0}>0$, then #$\#(D(0,\theta))>0$.
Proof.
Thm. 7.1 immediately implies claim (ii) but also implies that #$\#(D(0,1))>0$ if $\bar{s}_{0}>0$ unless #$\#(A(0,1/\theta,\theta))>0$. In both cases #$\#(D(0,\theta))>0$. ∎
7.4 Extension to any disc
Map (1.3), preserving #$\#(D(0,1))$ (cf. Observation 6.1), enables us to extend definition (7.2) of Cauchy sum to any disc $D(c,\rho)$ as follows:
$s_{0,q}(p,c,\rho):=s_{0,q}(t,0,1):=\frac{\rho}{q}\sum_{g=0}^{q-1}\zeta^{g}~{}\frac{p^{\prime}(c+\rho\zeta^{g})}{p(c+\rho\zeta^{g})},$ (7.8)
for $\zeta$ of (3.5), that is, $s_{h,q}(p,c,\rho)$ is the Cauchy sums $s_{h,q}(t,0,1)$ of the zeros of the polynomial $t(y)=p(\frac{y-c}{\rho})$ in the unit disc $D(0,1)$.
Observation 7.2.
Given a polynomial $p(x)$, a complex $c$, a positive $\rho$, a positive integer $q$, and the $q$-th roots of unity, we can compute $s_{0,q}(p,c,\rho)$ by means of the evaluation of NIR$(\zeta^{g})$ for $g=0,1,\dots,q-1$, reduced to evaluation of $p$ at $2q$ points, and in addition performing $q+1$ divisions, $12q-$ 1 multiplications, and $12q-$ 1 additions.
Refer to Alg. 7.1 applied to the polynomial $t(y)$ of (1.3) for an isolation parameter $\theta>1$ and $q=\lfloor\log_{\theta}(4d+2)\rfloor$ as Alg. 7.1c,ρ,q and also as Alg. 7.1 applied to $\theta$-isolated circle $C(c,\rho)$ for $2q$ evaluation points. Combine Observations 6.1 and 7.2 to extend Observation 7.1 as follows.
Proposition 7.1.
(i) Alg. 7.1c,ρ,θ evaluates polynomial $p(x)$ at $q=\lfloor\log_{\theta}(4d+2)\rfloor$ points $x$ and in addition performs $25q-$ 2 ops. (ii) It outputs $\bar{s}_{0}=\#(D(c,\rho))$ if the circle $C(c,\rho)$ is $\theta$-isolated. (iii) If the algorithm outputs $\bar{s}_{0}>0$, then #$\#(D(c,\theta\rho))>0$.
By applying Thm. 7.1 we can estimate the errors $|s_{h,q}(t,0,1)-s_{h}(t,0,1)|$ of the approximation of the $h$th power sums $s_{h}(t,0,1)$ of the roots of $t(x)$ in the unit disc $D(0,1)$. Next we specify this for $h=1$.
Proposition 7.2.
(Cf. [63].) Let for $\theta>1$ a $\theta$-isolated disc $D(c,\rho)$ contain exactly $m$ roots. Fix a positive integer $q$ and define the Cauchy sum $s_{1,q}(p,c,\rho)$ of (7.8). Then
$|mc+\rho s_{1,q}(p,c,\rho)-s_{1}(p,c,\rho)|\leq\frac{\rho d\theta}{\theta^{q}-1}.$
7.5 The real part of the 0-th Cauchy sum
Clearly, $|\Re(s_{0,q})-s_{0}|\leq|s_{0,q}-s_{0}|$ because $s_{0}$ is an integer. Next we estimate the bound $|\Re(s_{0,q})-s_{0}|$ for the unit disc $D(0,1)$, point out some tentative directions to algorithmic gain for root-finding, but leave further elaboration as a challenge for formal and empirical investigation. [^13] We will not use the results of this subsection, but they can be of independent interest.
Recall that $s_{0,q}=\sum_{j=1}^{d}\frac{1}{1-z_{j}^{q}}$ unless $z_{j}^{q}=1$ for some $q$ (cf. Cor. 7.1), write $z:=z_{j}^{q}$ and estimate the contribution of a roots $z_{j}$ to $\Re(s_{0,q})$.
Proposition 7.3.
Let $z=\rho\cdot(\cos(\phi)+\sin(\phi)\bf i)$ for a non-negative $\rho$ and a real $\phi\in[0,2\pi)$. Then
$\Re\Big{(}\frac{1}{1-z}\Big{)}=\frac{1-\rho\cos(\phi)}{1+\rho^{2}-2\rho\cos(\phi)},$ (7.9)
$1-\Re\Big{(}\frac{1}{1-z}\Big{)}=\rho\frac{\rho-\cos(\phi)}{1+\rho^{2}-2\rho\cos(\phi)}.$ (7.10)
Proof.
Multiply both numerator and denominator of the fraction
$\frac{1}{1-z}=\frac{1}{1-\rho\cdot(\cos(\phi)+\sin(\phi)\bf i)}$
by $1-\rho\cdot(\cos(\phi)-\sin(\phi)\bf i)$, substitute $\cos^{2}(\phi)+\sin^{2}(\phi)=1$, and obtain (7.9). Then readily verify (7.10). ∎
Corollary 7.2.
Under the assumptions of Prop. 7.3 it holds that
$\Re\Big{(}\frac{1}{1-z}\Big{)}\geq 0~{}{\rm if}~{}\rho\leq 1,$ (7.11)
$\Re\Big{(}\frac{1}{1-z}\Big{)}=\frac{1}{2}~{}{\rm if}~{}\rho=1~{}{\rm and~{}if}~{}\phi\neq 0,$ (7.12)
$\Re\Big{(}\frac{1}{1-z}\Big{)}\geq\frac{1-0.8\rho}{1+\rho^{2}-1.6\rho}~{}{\rm if}~{}\rho\leq 5/4~{}{\rm and~{}if}~{}\cos(\phi)\leq 0.8,$ (7.13)
$\Re\Big{(}\frac{1}{1-z}\Big{)}=\frac{1}{1-\rho}~{}{\rm if}~{}\phi=0~{}{\rm and~{}if}~{}\rho\neq 1,$ (7.14)
$\frac{1}{1+\rho}\leq\Big{|}\Re\Big{(}\frac{1}{1-z}\Big{)}\Big{|}\leq\frac{1}{|1-\rho|},$ (7.15)
$\frac{\rho}{1+\rho}\leq\Big{|}1-\Re\Big{(}\frac{1}{1-z}\Big{)}\Big{|}\leq\frac{\rho}{|1-\rho|}.$ (7.16)
Proof.
Readily deduce (7.11), (7.12), and (7.14) from (7.9).
Write
$y:=\cos(\phi),~{}f(y):=\Re\Big{(}\frac{1}{1-z}\Big{)},~{}{\rm and}~{}g(y):=\frac{1}{f(y)}.$
Then deduce from (7.9) that $g(y)=2+\frac{\rho^{2}-1}{1-\rho y}$ Hence $g^{\prime}(y)=\frac{(\rho^{2}-1)\rho}{(1-\rho y)^{2}}$ is a monotone increasing or monotone decreasing function, singular if and only if $y=\frac{1}{\rho}$. Extend these properties to $f(y)$ and $1-f(y)$ if $\rho<1$. This implies (7.13), (7.15), and (7.16) provided that $\rho<1$ because
$f(1)=\frac{1}{1-\rho},~{}1-f(1)=\frac{\rho}{1-\rho},~{}f(-1)=\frac{1}{|1+\rho|},~{}1-f(-1)=\frac{\rho}{|1+\rho|},$
and $f(0.8)=\frac{1-0.8\rho}{1+\rho^{2}-1.6\rho}$.
The proof can be extended to the case where $\rho\geq 1$ except that the zero $y=\frac{1}{\rho}$ of $f(y)$ turns into a pole of $\frac{1}{f(y)}$. We complete the proof of the bounds (7.15) for all positive $\rho$ by observing that $f(\frac{1}{\rho})$ is undefined for $\rho=1$ and that otherwise $f(\frac{1}{\rho})=0$ and $11-f(\frac{1}{\rho})=$ 1. ∎
Next we outline but do not elaborate upon some potential links of Cor. 7.2 to root-counting. Let the unit disc $D(0,1)$ be $\theta$-isolated, combine bounds (7.15) and (7.16), and obtain
$|\Re(s_{0,q})-s_{0}|\leq\frac{d}{\theta^{q}-1}.$
The bound is similar to (7.5) for $h=0$ but is overly pessimistic: it is only reached where $\cos(\phi)=1$ for $z=z_{j}$ and all $j$.
Equation (7.14) implies numerically unstable dependence of the fraction $\Re(\frac{1}{1-z})$ on $\rho$ where $\cos(\phi)\approx 1\approx\rho$, but bound (7.13) shows that outside a rather small neighborhood of the extreme values of $\cos(\phi)$ and $\rho$ the fraction is smooth and approaches 1/2 as $\rho\rightarrow 1$ and/or $\cos(\phi)\rightarrow-1$.
Remark 7.1.
In view of bound (7.13) random rotation of the circle $C(0,1)$ is a heuristic recipe for countering instability in $z$ of the fraction $\Re(\frac{1}{1-z})$ because $\cos(\phi)>0.8$ for only one of four equally spaced points $\phi$ on the unit circle $C(0,1)$ and hence occurs with a probability less than 1/4 for a random choice of $\phi$ under the uniform probability distribution in the range $[0,2\pi)$. Moreover, random rotation of the circle can limit the adverse and numerically unstable impact of boundary roots, lying on or near the boundary circle $C(0,1)$, on an e/i test.
We conclude with a simple but surprising result on root-counting on an isolated circle. We do not use it in this paper.
Corollary 7.3.
Suppose that $p(1)\neq 0$ and that $p(x)$ has precisely $m$ roots in an annulus $A(0,\frac{1}{\theta},\theta)$ for a fixed $\theta>1$ and that all of them lie on the unit circle $C(0,1)$. Let $s^{\prime}$ and $s^{\prime\prime}$ denote the 00th Cauchy sums for the discs $D(0,\frac{1}{\theta})$ and $D(0,\theta)$, respectively, and write $s_{0,q}=s^{\prime\prime}-s^{\prime}$. Then #$(C(0,1))=2\Re(s_{0,q})$.
Proof.
Combine Cor. 7.1 for $h=0$ with equation (7.12). ∎
8 Root-counting and $\ell$-tests
8.1 Overview
We seek fast root-counters and $\ell$-tests based on computing Cauchy sum $s_{0,q}$ for a smaller integer $q$, but the output can deceive us if we decrease $q$ too much: the polynomial $p(x):=x^{d}-dx+w$ has zero near the origin if $w\approx 0$, while NIR$(x)=p^{\prime}(x)/p(x)$ vanishes for $x$ equal to any $(d-1)$st root of unity.
Such deceptive e/i tests have been very rare guests in the extensive experiments in [60, 63] for $q$ of order $\log^{2}(d))$ and were completely excluded where the values $|s_{h,q}|$ were required to be small for $h=0,1,2$, but formal support for this empirical behavior is still a challenge.
In the next two subsections we formally support Las Vegas randomized root-counters and $\ell$-tests that evaluate polynomial $p$ at $O(m\log(d))$ points.
Combine this estimate with Cor. 2.2 for $k$ of order $b$ and obtain Las Vegas randomized solution of Problems 0, 1, 0, and 1 at the cost of evaluation of the polynomial $p$ at $O(m^{2}b)$ points.
8.2 A basic randomized root-counter
Next we fix a parameter $\gamma>1$ and devise a root-counter for a disc $D(c,\rho)$ with a fixed center $c$ (wlog we let $c=0$) and a radius $\rho$ sampled at random in a fixed small range. Then we prove that under such a Las Vegas randomization the test fails with a probability at most $1/\gamma$.
Algorithm 8.1.
Basic randomized root-counter.
INPUT: $\gamma\geq 1$ and a $d$-th degree black box polynomial $p$ having at most $m$ zeros in the disc $D(0,\sqrt{2})$.
INITIALIZATION: Sample a random value $w$ in the range $[0.2,0.4]$ under the uniform probability distribution in that range and write $\rho:=2^{w}$ and $q=\lfloor 10m\gamma\log_{2}(4d+2)\rfloor$.
COMPUTATIONS: Apply Alg. 7.10,ρ,q.
Proposition 8.1.
Alg. 8.1 queries evaluation oracle at $q=\lfloor 10m\gamma\log_{2}(4d+2)\rfloor$ locations and outputs the value $\bar{s}_{0}=\#(D(0,\rho))$ with a probability at least $1-1/\gamma$.
Proof.
Readily verify the bound on the number of query locations claimed in the proposition.
By assumption, the annulus $A(0,1,\sqrt{2})$ contains at most $m$ roots.
Hence at most $m$ root radii $r_{j}=2^{e_{j}}$, for $j=1,\dots,m^{\prime}\leq m$, lie in the range $[1,\sqrt{2}]$, that is, $0\leq e_{j}\leq 0.5$ for at most $m^{\prime}\leq m$ integers $j$.
Fix $m^{\prime}$ intervals, centered at $e_{j}$, for $j=1,\dots,m^{\prime}$, each of a length at most $1/(5m^{\prime}\gamma)$. Their overall length is at most $1/(5\gamma)$. Let $\mathbb{U}$ denote their union.
Sample a random $u$ in the range $[0.2,0.4]$ under the uniform probability distribution in that range and notice that Probability$(u\in\mathbb{U})\leq 1/\gamma$.
Hence with a probability at least $1-1/\gamma$ the circle $C(0,\rho)$ is $\theta$-isolated for $\theta=2^{\frac{1}{10m\gamma}}$, in which case Alg. 8.1 outputs $\bar{s}_{0}=s_{0}=\#(D(0,\rho))$ by virtue of claim (ii) of Prop. 7.1. ∎
8.3 Refining the bound on the error probability
When we apply Alg. 8.1 $v$ times, the number of query locations of the evaluation oracle grows linearly – by a factor of $v$, while the bound on Las Vegas error probability of outputting errors decreases exponentially – from $1/\gamma$ to $1/\gamma^{v}$.
First let $\ell=1$ or $\ell=m$ and then again reduce the root radius approximation problem to the decision problem of $\ell$-test. Namely, narrow Alg. 8.1 to an $\ell$-test for a fixed $\ell$ in the range $1\leq\ell\leq m$, so that with a probability at least $1-1/\gamma$ this $\ell$-test outputs 0 if #$\#(D(0,1))<\ell$ and outputs 1 if #$\#(D(0,2^{0.5}))\geq\ell$. Moreover, output 1 is certified by virtue of claim (iii) of Prop. 7.1 if $\ell=1$; similarly the output 0 is certified if $\ell=m$.
Next extend the $\ell$-test by means of applying Alg. 8.1 for $v$ iid random variables $w$ in the range $0.2\leq w\leq 0.4$. Call this $\ell$-test Alg. $\ell$.
[…]
Recall Prop. 8.1 and then readily verify the following proposition.
Proposition 8.2.
(i) For two integers $1\leq\ell\leq m$ and $v\geq 1$ and a real $\gamma\geq 1$, Alg. $\ell$ evaluates $p(x)$ at $\lfloor 10m\gamma\log_{2}(4d+2)\rfloor v$ points; this evaluation involves more ops than the other stages of the algorithm, even in the cases of the Mandelbrot and sparse polynomials $p$ (cf. Observation 7.2). (ii) Output 1 of Alg. 8.1v,1 and the output 0 of Alg. 8.1v,$m$ are certified. (iii) Output 0 of Alg. 8.1v,1 and output 1 of Alg. 8.1v,$m$ are correct with a probability at least $1-1/\gamma^{v}$.
The proposition bounds the cost of $\ell$-tests for $\ell=1$ (e/i tests) and $\ell=m$. This is sufficient for our root-finding, but our next natural extension covers $\ell$-tests for any integer $\ell$ in the range $1<\ell<m$.
Assign the values 0 or 1 to the output integers $g$ as follows: compute the average of the $v$ outputs in $v$ independent applications of the test and output $g$ denoting the integer closest to that average; if the average is 1/2, write $g=0$ to break ties.
This extension of Alg. $\ell$ is only interesting for $1<\ell<m$; otherwise its cost/error estimates are inferior to those of Prop. 8.2.
Proposition 8.3.
For an integer $\ell$ such that $1<\ell<m\leq d$ and real $v\geq 1$ and $\gamma>1$, Algs. $\ell$ queries evaluation oracle at $\lfloor 10m\gamma\log_{2}(4d+2)\rfloor v$ locations, and its output is correct with a probability at least $1-(4/\gamma)^{v/2}$.
Proof.
Since $D(0,1)\subset D(0,\sqrt{2})$, we only consider the following three cases:
(i) #$\#(D(0,1))<\ell$ and #$\#(D(0,2^{0.5}))\geq\ell$,
(ii) #$\#(D(0,2^{0.5}))<\ell$, and then also #$\#(D(0,1))<\ell$,
(iii) #$\#(D(0,1))\geq\ell$, and then also #$\#(D(0,2^{0.5}))\geq\ell$.
In case (i) the output $g$ is always correct.
In case (ii) the output 0 is correct, while the output 1 is wrong and occurs if and only if more than $v/2$ tests go wrong, by outputting 1.
Likewise, in case (iii) the output 1 is correct, while the output 0 occurs if and only if more than $v/2$ tests go wrong, by outputting 0.
In both cases (ii) and (iii) we deal with the binomial distribution having failure probability $1/\gamma$, and so (cf. [26]) the probability $P$ of at least $v/2$ failures among $v$ tests is at most $\sum_{j=\bar{v}}^{v}\begin{pmatrix}v\\ j\end{pmatrix}T_{j}$ for $T_{j}=\gamma^{-j}(1-1/\gamma)^{v-j}\leq\gamma^{-j}$.
Therefore, $P\leq\sum_{j=0}^{v}\begin{pmatrix}v\\ j\end{pmatrix}=T_{\bar{v}}2^{v}\max_{j}T_{j}$. Substitute $\sum_{j=0}^{v}\begin{pmatrix}v\\ j\end{pmatrix}=2^{v}$ and $\bar{v}=\lceil v/2\rceil$ and obtain $P<2^{v}T_{\bar{v}}\leq 2^{v}/\gamma^{v/2}=(4/\gamma)^{v/2}$. ∎
9 Root radius estimation based on $\ell$-test
In this section we assume that we are given an algorithm for $\ell$-test and seek approximation of $r_{d-\ell+1}(c^{\prime})$, the $\ell$th smallest root radius from a fixed complex point $c^{\prime}$, known to lie in a fixed range $[a_{-},a_{+}]$. [^14] For our root-finders we only need to approximate $r_{d-\ell+1}(c^{\prime})$ for $\ell=1$ and $\ell=m$, but we estimate $r_{d-\ell+1}(c^{\prime})$ for any $\ell$, $1\leq\ell\leq m$, which is of independent interest for root-finding.
Wlog let this be the range $[\epsilon,1]$ for a fixed small positive $\epsilon=1/2^{b}$ and seek a range $[\rho_{-},\rho_{+}]$ such that $0\leq\rho_{-}\leq r_{d-\ell+1}(c^{\prime})\leq\rho_{+}$ and either $\rho_{+}\leq\epsilon$ or else $0<\rho_{+}<\beta\rho_{-}$ for two fixed tolerance values $\epsilon=1/2^{b}$ and $\beta=2^{\phi}>1$. We can obtain the desired approximation by applying bisection algorithm to the range $[\epsilon,1]$, but this is too slow for us, and we apply bisection of logarithms or BoL to approximate the logarithm of the target value in the range $[-b,0]=[\log_{2}(\epsilon),\log_{2}(1)]$.
Assume that besides the above tolerance values the input of BoL includes a softness bound $\sigma=1+\alpha$, for a fixed small positive parameter $\alpha<1/2$, and a $\sigma$-soft $\ell$-testc,ρ, to be used as the pivot of BoL.
Since the test is soft, so is BoL as well: for a range $[b_{-},b_{+}]$ with midpoint $\widehat{b}=0.5(b_{+}+b_{-})$ regular rigid bisection continues the search in the range $[b_{-},\widehat{b}]$ or $[\widehat{b},b_{+}]$, but our soft BoL continues it in the range $[b_{-},(1+\alpha)\widehat{b}]$ or $[(1-\alpha)\widehat{b},b_{+}]$ for a fixed small positive $\alpha$. Accordingly we slightly increase range of search for $r_{d-\ell+1}(c^{\prime})$ to $[-\bar{b},0]$ for $\bar{b}:=\log_{2}(\frac{\sigma}{\epsilon})$. BoL still converges linearly and just a little slower than the regular rigid bisection: the approximation error decreases by a factor of $\frac{2}{1+\alpha}$ rather than by twice. [^15] If we applied BoL to the range $[0,1]$, we would have to bisect the infinite range $[-\infty,0]$, but we deal with the narrower range $[-b,0]$ since we have narrowed the range $[0,1]$ to $[\epsilon,1]$.
Algorithm 9.1.
Bracketing the $\ell$-th smallest root radius based on $\ell$-test.
INPUT: Two integers $\ell$ and $m$ such that $1\leq\ell\leq m$ and #$\#(D(0,1))\leq m$, $\sigma=1+\alpha$ for $20<\alpha<1/$ 2, a $\sigma$-soft $\ell$-test0,ρ, and two tolerance values $\beta:=2^{\phi}>1$ and $\epsilon=1/2^{b}$, $b\geq 1$.
OUTPUT: A range $[\rho_{-},\rho_{+}]$ such that $[\rho_{-},\rho_{+}]=[0,\epsilon]$ or $\rho_{-}\leq r_{d-\ell+1}\leq\rho_{+}\leq\beta\rho_{-}$.
INITIALIZATION: Write $\bar{b}:=\log_{2}(\frac{\sigma}{\epsilon})$, $b_{-}:=-\bar{b}$, and $b_{+}:=0$.
[…]
$\rho_{-}=2^{b_{-}},~{}\rho_{+}=2^{b_{+}},~{}\rho:=2^{\widehat{b}},~{}{\rm and}~{}\widehat{b}:=\lceil(b_{-}+b_{+})/2\rceil.$ (9.1)
Write $b_{-}:=(1-\alpha)\widehat{b}$ if the $\ell$-test outputs 0; write $b_{+}:=(1+\alpha)\widehat{b}$ if the test outputs 1. Then update $\widehat{b}$ and $\rho$ according to (9.1).
Continue recursively: stop and output the current range $[\rho_{-},\rho_{+}]=[2^{b_{-}},2^{b_{+}}]$ if $b_{+}-b_{-}\leq\phi$, that is, if $\rho_{+}\leq\beta\rho_{-}$; otherwise update $\widehat{b}$ and $\rho$ and reapply the computations.
The disc $D(0,\epsilon)$ contains at least $\ell$ roots, that is, the range $[0,\epsilon]$ brackets the $\ell$th smallest root radius $r_{d-\ell+1}$ if the $\sigma$-soft $\ell$-test applied to the disc $D(0,\frac{\epsilon}{\sigma})$ outputs 1. Otherwise, the disc $D(0,\frac{\epsilon}{\sigma})$ contains less than $\ell$ roots, and so the root radius $r_{d-\ell+1}$ lies in the range $[\frac{\epsilon}{\sigma},1]$, in which we shall continue our search.
Readily verify that our policy of recursively updating the range $[b_{-},b_{+}]$ maintains this property throughout and that updating stops where $b_{+}-b_{-}\leq\phi$, that is, where $\rho_{+}\leq\beta\rho_{-}$. This implies correctness of Alg. 9.1.
Let us estimate the number $\mathbb{I}$ of its invocations of an $\ell$-test.
Proposition 9.1.
Alg. 9.1 invokes an $\ell$-test $\mathbb{I}$ times where $\mathbb{I}\leq\lceil\log_{\nu}(\frac{\bar{b}}{\phi})\rceil+1$ for $\nu=\frac{2}{1+\alpha}$, $\mathbb{I}=O(\log(b/\phi))$.
Proof.
The computations in Alg. 9.1 stop after the first application of an $\ell$-test if it outputs 1. It remains to count the subsequent $\ell$-tests if it outputs 0.
Each test decreases the length $\bar{b}=b_{+}-b_{-}$ of its input range $[b_{-},b_{+}]$ for the exponent to $0.5(1+\alpha)\bar{b}$. Hence $i$ applications change it into $\frac{(1+\alpha)^{i}}{2^{i}}\bar{b}$ for $i=1,2,\dots$. These values linearly converge to 0 as $i$ increases to the infinity, and the proposition follows. ∎
Remark 9.1.
In the case where we know that #$\#(D(0,\rho))\geq\ell$ for a fixed $\rho$ in the range $[\epsilon,1]$, we only need to refine the initial range $[\rho,1]$ for the root radius $r_{d-\ell+1}$.
Clearly, $\mathbb{I}$ applications of $\ell$-test increase the number of evaluation points for the ratio $\frac{p^{\prime}(x)}{p(x)}$ by at most a factor of $\mathbb{I}$, and if the $\ell$-test is randomized, then the error probability also grows by at most a factor of $\mathbb{I}$. Let us specify these estimates for Alg. 9.1.
Corollary 9.1.
Let precisely $m$ zeros of a black box polynomial $p(x)$ of a degree $d$ lie in the unit disc, that is, let #$\#(D(0,1))=m$, and let $i(D(0,1))\geq\theta$, for $\theta-1$ exceeding a positive constant. Given two positive tolerance values $\epsilon=1/2^{b}$ and $\phi$, an integer $\ell$, $1\leq\ell\leq m$, and a black box $\ell$-test, which evaluates polynomial $p(x)$ at $\mathcal{NR}$ points, Alg. 9.1 approximates the $\ell$-th smallest root radius $|z_{j}|$ within the relative error $2^{\phi}$ or determines that $|z_{j}|\leq\epsilon$, at the cost of evaluation of polynomial $p(x)$ at $\mathbb{I}\cdot\mathcal{NR}$ points for $\mathbb{I}=O(\log(\frac{b}{\phi}))$ estimated in Prop. 9.1.
10 Precision of computing in our root-finders
10.1 Doubling computational precision towards its a posteriori estimation
Given an algorithm $\mathbb{A}$ for black box polynomial computations and target tolerance TOL to output errors, how should one choose computational precision? Here is a simple practical recipe from [10, 15, 22] and MPSolve for obtaining a posteriori estimate within a factor of two from optimal: first apply algorithm $\mathbb{A}$ with a lower precision, then recursively double it and stop where an output error decreases below TOL.
In the rest of this section we obtain a priori estimate for the precision of $O(b\log(d))$ bits.
10.2 Reduction to approximation of $x$NIR$(x)$ within $\frac{3}{8}$
In our e/i tests we approximate NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ with the first order divided difference of (3.2); we only apply the tests to $\theta$-isolated discs $D(c,\rho)$ for $\theta-1>0$ of order $\frac{1}{m}$ and $\rho\geq\epsilon=1/2^{b}$.
Map (1.3) reduces such a test for $p(x)$ to a test in the disc $D(0,1)$ for the polynomial $t(x)=p(\frac{x-c}{\rho})$ replacing $p(x)$ and decreases the error estimate by a factor of $\rho$. Hence map (1.3) increases the precision bound by at most a factor of $b$. Thus we obtain the claimed precision bound in $O(b\log(d))$ in case of any disc $D(c,\rho)$ if we prove bound $O(\log(d))$ in the case of the unit disc $D(0,1)$.
The output of an e/i test is an integer, and we only need to certify that the output error bound of Alg. 9.1 is less than $\frac{1}{2}$. The error can be magnified by a factor of $v$ in $v$ applications of Alg. 9.1 but this factor is immaterial in our notation $\tilde{O}$.
Our e/i tests compute the $q$ values $\rho x\frac{p^{\prime}(c+\rho x)}{p(c+\rho x)}=x\frac{t^{\prime}(x)}{t(x)}$, for $x=\zeta^{g}$, $\zeta$ of (7.2), and $g=0,1,\dots,q-1$, sum these values, and divide the sum by $q$.
By slightly abusing notation write ${\rm NIR}(x):=\frac{t^{\prime}(x)}{t(x)}$ rather than ${\rm NIR}(x)=\frac{p^{\prime}(x)}{p(x)}$. According to straightforward error analysis (cf. [21, Part A of Ch.3] and the references therein) rounding errors introduced by these additions and division are dominated, and we can ignore them; then it is sufficient to compute $\zeta^{g}{\rm NIR}(\zeta^{g})$ within, say, $\frac{3}{8}$ for every $g$. Indeed, in that case the sum of $q$ error bounds is at most $\frac{3q}{8}$ and decreases below $\frac{1}{2}$ in division by $q$.
10.3 Precision of representation of NIR$(x)$ within $\frac{1}{8}$
Recall Eqn. (3.1) and represent $x{\rm NIR}(x)$ for $|x|=1$ within $\frac{1}{8}$ by using a precision of $O(\log(d))$ bits.
Proposition 10.1.
Write $C:=C(c,\rho)$ and assume that $|x|=1$ and the circle $C$ is $\theta$-isolated for $\theta>1$. Then $|x{\rm NIR}(x)|\leq\frac{d}{1-1/\theta}=\frac{\theta d}{\theta-1}$.
Proof.
Eqn. (3.1) implies that $|{\rm NIR}(x)|=|\sum_{j=1}^{d}\frac{1}{x-z_{j}}|\leq\sum_{j=1}^{d}\frac{1}{|x-z_{j}|}$ where $|x-z_{j}|\geq(1-1/\theta)\rho$ for $|z_{j}|<1$, while $|x-z_{j}|\geq(\theta-1)\rho$ for $|z_{j}|>1$ since $i(C)\geq\theta$. Combine these bounds, write $m:=\#(D(c,\rho))$, and obtain $|x{\rm NIR}(x)|\leq\frac{m}{1-1/\theta}+\frac{d-m}{\theta-1}\leq\frac{d}{1-1/\theta}$ for $|x|=1$. ∎
Corollary 10.1.
One can represent $x{\rm NIR}(x)$ within $\frac{1}{8}$ for $|x|=1$ by using a precision of $3+\lceil\log_{2}(\frac{d}{1-1/\theta})\rceil$ bits, which is in $O(\log(d))$ provided that $\frac{1}{\theta-1}=d^{O(1)}$.
10.4 What is left to estimate?
Instead of $x{\rm NIR}(x)=\frac{xt^{\prime}(x)}{t(x)}$ we actually approximate $\frac{xt^{\prime}(y)}{t(x)}$ where $|y-x|\leq\delta$ and $t^{\prime}(y)$ is equal to the divided difference of (3.2). Thus, we shall increase the above error bound $\frac{1}{8}$ by adding to it the upper bounds $\frac{1}{8}$ on $\alpha:=|\frac{xt^{\prime}(x)}{t(x)}-\frac{xt^{\prime}(y)}{t(x)}|$ for $|x|=1$ and on the rounding error $\beta$ of computing $x\frac{t(x)-t(x-\delta)}{\delta t(x)}$. Next we estimate $\alpha$ in terms of $d$, $\theta$, and $\delta$ – by extending the proof of [76, Lemma 3.6], then readily ensure that $\alpha<\frac{1}{8}$ by choosing $\delta$ with a precision $\log_{2}(1/|\delta|)$ of order $\log(d)$, and finally estimate $\beta$. We simplify our estimates by dropping the factor $x$ where $|x|=1$.
10.5 Precision of approximation with divided difference
Recall the following lemma, which is [76, Fact 3.5].
Lemma 10.1.
For $t(x)=\prod_{j=1}^{d}(x-y_{j})$ and a non-negative integer $j\leq d$ it holds that
$t^{(j)}(x)=j!~{}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}$
where the summation $\sum_{S_{j,d}}$ is over all subsets $S_{j,d}$ of the set $\{1,\dots,d\}$ having cardinality $j$.
Proposition 10.2.
Under the assumptions of Prop. 10.1 let $w=w(x):=\max_{j=1}\frac{1}{|x-y_{j}|}$. Then
$|t(x)|\alpha=|\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)|\leq|t(x)\frac{(d~{}w)^{2}\delta}{1-\delta~{}d~{}w}|$.
Proof.
The claimed bound on $|\xi^{\prime}|$ follows from (3.1). It remains to apply the first five lines of the proof of [76, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, s respectively. Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
Substitute the expressions of Lemma 10.1 and obtain
$t(x)-t(x-\delta)=\delta t^{\prime}(x)+\sum_{j=2}^{\infty}\delta^{j}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}.$
Combine this equation with the assumed bounds on $\frac{1}{|x-y_{j}|}$ and deduce that
$\Big{|}\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)\Big{|}\leq\Big{|}\frac{f(x)}{\delta}\sum_{j=2}^{\infty}\delta^{j}w^{j}d^{j}\Big{|}\leq\Big{|}f(x)\frac{(d~{}w)^{2}~{}\delta}{1-\delta~{}d~{}w}\Big{|}.$
In our estimates for $\alpha$ and $\delta$ we apply Prop. 10.1 for $w\leq\frac{\theta}{\theta-1}$ and obtain
Corollary 10.2.
Under the assumptions of Prop. 10.1, let $w\leq\frac{\theta}{\theta-1}$. Then
$\alpha<1/8~{}{\rm for}~{}|\delta|\leq\frac{(\theta-1)^{2}}{10d^{2}\theta^{2}}.$
10.6 The impact of rounding errors
To approximate $|x{\rm NIR}_{c,\rho}(x)|$ within $\frac{3}{8}$ it remains to approximate $\beta$ within $\frac{1}{8}$.
Proposition 10.3.
Suppose that a black box oracle has computed the values $t(y)$ within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor. 10.2. Then we can ensure that $|\beta|<\frac{1}{8}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $\log(\frac{1}{\nabla(y)})=O(\log(d))$ bits.
Proof.
Write $\beta\delta=\frac{t(x-\delta)(1+\nabla(x-\delta))}{t(x)(1+\nabla(x))}-\frac{t(x-\delta)}{t(x)}=\frac{t(x-\delta)}{t(x)}~{}\frac{\nabla(x-\delta)-\nabla(x)}{1+\nabla(x)}$.
Taylor-Lagrange’s formula implies that $\frac{t(x-\delta)}{t(x)}=1+\frac{t(x-\delta)-t(x)}{t(x)}=1+\delta\frac{t^{\prime}(u)}{t(x)}$, for $u\in[x-\delta,x]$. Hence $|\frac{t(x-\delta)}{t(x)}|\leq 1+|(1+\alpha^{\prime})\delta|$ for $\alpha^{\prime}=\frac{t^{\prime}(u)-t(x)}{t(x)}$.
By extending Cor. 10.2 obtain $|\alpha^{\prime}|\leq\frac{1}{8}$.
Hence $|\frac{t(x-\delta)}{t(x)}|\leq w:=1+\frac{9|\delta|}{8}$, while $|\beta\delta|\leq w\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}$.
Now choose $\nabla(y)$ such that $|\nabla(y)|\leq\frac{|\delta|}{17w}$ for $y=x$ and $y=x-\delta$, and so $w(|\nabla(x)|+|\nabla(x-\delta)|)\leq\frac{2}{17\delta}$. Then verify that $\log(\frac{1}{|\nabla(y)|})=O(\log(d))$, $1-|\nabla(x)|\geq 1-\frac{|\delta|}{17w}>\frac{16}{17}$, and $|\delta\beta|<\frac{|\delta|}{8}$. ∎
By combining Cors. 10.1 and 10.2 and Prop. 10.3 obtain
Corollary 10.3.
Suppose that a black box oracle can output approximations of a $d$-th degree polynomial $t(x)$ with any relative error in $1/d^{O(1)}$ for $|x|\leq 1$ and let the unit disc be $\theta$-isolated for $t(x)$ and for $\frac{1}{\theta-1}=d^{O(1)}$. Then we can perform e/i test for $t(x)$ on the unit disc by using a precision of $O(b\log(d))$ bits.
11 Approximation of matrix eigenvalues
Given a $d\times d$ matrix $M$ that has characteristic polynomial $p(x)=\det(xI-M)$, our root-finders approximate all $m$ its zeros (that is, the eigenvalues of $M$) lying in a fixed disc $D$ on the complex plane isolated from external eigenvalues.
Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [138, Thm. 1.3.2]: $|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|$ for some $i$, $1\leq i\leq d$.
Hence, clearly, $|\lambda|\leq||M||_{F}$.
Combine this bound with the precision estimates of the previous section for our root-finders [^16] Practically, one can combine our root-finders with the recipe of doubling the precision of computing of [10, 15, 22]. and with the following result of Storjohann [137] (cf. Remark 11.1):
Theorem 11.1.
One can evaluate the determinant of a $d\times d$ integer matrix $A$ by using a Las Vegas expected number of $O(d^{\omega}\log^{2}(d)\log(|A||_{F}))$ bit operations.
This implies a record bit operation complexity bound $\tilde{O}(m^{2}\tilde{b}^{2}d^{\omega})$ (cf. (1.6)) for Las Vegas approximation within $\epsilon$ of all $m$ eigenvalues of $M$ lying in the disc $D$ provided that $\tilde{b}=\log_{2}(||M||_{F}/\epsilon)$.
Remark 11.1.
Our root-finders are essentially reduced to the evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$, and we can apply the following expressions:
$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}((xI-M)^{-1})}$
where $M$ is a matrix and $p(x)=\det(xI-M)$ (cf. (3.1)) and
$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}(M^{-1}(x)M^{\prime}(x))}$
where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [20, Eqn. (5)]. [^17] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and its determinant. Based on the above expressions one can reduce approximation of eigenvalues of a matrix or a matrix polynomial to its inversion. This computation tends to be more numerically stable than computing determinants and is significantly accelerated in the highly important case where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
12 Bit operation cost of the classical subdivision root-finding with e/i tests based on approximation of a Cauchy integral
12.1 Our goal and a basic theorem
Next we estimate the bit operation complexity of our root-finders by combining them with fast multipoint evaluation of a polynomial by means of the algorithm of Fiduccia [42] and Moenck and Borodin [84] (see [19, Sec. 4.5] for historical account), non-trivially extended in [81] to support the following theorem:
Theorem 12.1.
Given a positive $b$, the coefficients of a polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}$ such that $||t(x)||_{1}=\sum_{i=0}^{d}|t_{i}|\leq 2^{\tau}$, and $q$ complex points $x_{1},\dots,x_{q}$ in the unit disc $D(0,1)$, one can approximate the values $t(x_{1}),\dots,t(x_{q})$ within $1/2^{b}$ by using $\tilde{O}((d+q)(b+\tau))$ bit operations.
Proof.
For $q=d$ this is [81, Thm. 2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Moroz in [81]. Trivially extend this theorem to the cases where $q>d$ (by considering $t(x)$ a polynomial of degree $q$ with $q-d$ leading coefficients 0) and where $q<d$ (by removing $d-q$ points $x_{i}$, with no increase of the complexity bound). ∎
We only apply this theorem for $\tau=0$, in which case its bit operation bound turns into
$\mathbb{B}_{\rm Moroz}=\tilde{O}((d+q)b)$ (12.1)
For comparison, the previous State of the Art was represented by the bound
$\mathbb{B}_{\rm Kirr}=\tilde{O}((d+q)(b+q)),$ (12.2)
due to [68, Thm. 3.9, Alg. 5.3, and Appendix A.3] by Peter Kirrinnis, which relied on combining his and Schönhage’s study of the bit operation cost of basic polynomial computations (cf. [134]) with the algorithm of [42, 84]. [^18] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by applying the Fast Multipole Method of Greengard and Rokhlin [50, 28, 105, 106] to the equivalent task of multiplication of Vandermonde matrix by a vector. One can, however, support bounds of Eqns. (12.1) and (12.2) on the overall bit operation complexity by computing with multiple precision for larger values $b$. Moroz obtained [81, Thm. 2] by applying bound (12.2) to his lower degree polynomials approximating $p(x)$.
The algorithms of [42, 84, 68, 81] extensively involve the coefficients of $p(x)$.
12.2 Reduction of our root-finders to multipoint polynomial approximation
Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ roots, we approximate all these roots within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with e/i tests of Sec. 8. Then every subdivision step is reduced to application of e/i tests to $\bar{m}=O(m)$ discs $D(c_{\lambda},\rho_{\lambda})$ for $\rho_{\lambda}\geq R/2^{b}$ and all $\lambda$.
Such a test, based on our e/i test of Sec. 8, amounts to approximation of $p(x)$ at $O(m\log(d))$ equally spaced points $x$ on each of $v$ circles sharing the center $c_{\lambda}$ and having radii in the range $[\rho_{\lambda}/\sigma,\rho_{\lambda}\sigma]$ for $1<\sigma<\sqrt{2}$.
We only use Thm. 12.1 for $\tau=0$ to cover the case where a polynomial $t(x)$ with $||t(x)||_{1}\leq 1$ is evaluated at points $x_{j}\in D(0,1)$. To ensure these assumptions we first scale the variable $x\mapsto Rx$ to map the discs $D(c_{\lambda},\rho_{\lambda}\sigma)\mapsto D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})\subseteq D(0,1)$, for $\bar{c}_{\lambda}=c_{\lambda}/R$, $\bar{\rho}_{\lambda}=\rho_{\lambda}/R\geq 1/2^{b}$, and all $\lambda$.
Then we write $t(x):=p(Rx)/\psi$ for $\psi=O(R^{d}||p(x)||_{1})$ such that $||t(x)||_{1}=1$ and approximate $t(x)$ within $1/2^{\bar{b}}$ or equivalently within the relative error bound $1/(|t(x)|2^{\bar{b}})$ at $\bar{q}:=qv$ points $x$ in $\bar{m}$ e/i tests. [^19] Scaling $p$ and $x$ above is done once and for all e/i tests; to compensate for this we only need to change our bound on the output precision $b$ by $\lceil|\log_{2}R|\rceil$ bits.
Due to Cor. 10.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we achieve this by choosing any $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$. Map the disc $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ into the unit disc $D(0,1)$. This does not change isolation of the boundary circle and may increase the norm $||t(x)||_{1}$ by at most factor of $2^{b+d}$ because $|\bar{c}_{\lambda}|\leq 1$ and $\bar{\rho}_{\lambda}\geq 1/2^{b}$. Now substitute $||t(x)||_{1}\leq 2^{b+d}$ and $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d\log(m))$ (cf. [134, Eqn.(9.4)]) [^20] This bound is readily deduced from the following non-trivial result [134, Thm. 4.2]: Let $t(x)=f\prod_{j=1}^{m}(x-z_{j})\prod_{j=m+1}^{d}(\frac{x}{z_{j}}-1)$, $||t(x)||_{1}=1$, $|z_{j}|<1$ for $j\leq m$, and $|z_{j}|>1$ for $j>m$. Then $|f|\geq 1/2^{d}$. and obtain that $\bar{b}=\log(\frac{1}{|t(x)|})+O(\log(d))=O(d\log(m)+b)$.
Remark 12.1.
For $m=d$ the above bound on $\log(\frac{||t(x)||_{1}}{|t(x)|})$ of [134, Eqn.(9.4)]) is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$. For $m<d$ [134, Thm. 4.5] implies a little stronger bounds, namely, $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.
12.3 The bit operation complexity of our e/i tests and root-finders
Apply Thm. 12.1 with $q$ and $b$ replaced by $\bar{q}=O(m^{2})$ and $\bar{b}=O(b+d+m\log(m))$, respectively, [^21] Under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $m$ and then extend this decrease to the related subsequent estimates. and obtain that the bit operation cost of our $O(m)$ e/i tests of Sec. 8 at any fixed subdivision step is in $\tilde{O}((b+d+m)m^{2})$ for $1\leq m\leq d$. This implies the upper estimate
$\mathbb{B}_{\rm roots}=\tilde{O}((b+d)m^{2}b)$ (12.3)
for the overall bit operation cost of our $O(b)$ subdivision steps, which includes the time-complexity of the evaluation oracle. Based on [68, Alg. 5.3] and Eqn. (12.2) obtain a little larger upper bound on the bit operation cost of our subdivision root-finders:
$\mathbb{B}_{\rm roots}^{\prime}=\tilde{O}((m^{2}+d+b)(m^{2}+d)b).$ (12.4)
12.4 Higher bit operation complexity of general polynomial root-finding based on map (1.3) and root-squaring
Our bit operation complexity estimates for root-finding rely on restriction of computations to approximation of NIR without involving coefficients of $p$. Knowing the coefficients of the polynomial $t(x)=p(\frac{x-c}{\rho})$ one can devise very fast e/i test for the disc $D(c,\rho)$ by applying Pellet’s theorem to $t_{h}(x)$ for $h$ of order $\log(d)$, whose zero are the $h$th powers of the zeros of $t(x)$ [23, 24].
The paper [24] applies subdivision iterations of [149, 52, 128, 101] with e/i tests performed by means of Pellet’s theorem, which involves all coefficients of an input polynomial, thus losing the benefits where, say, $p(x)$ is the Mandelbrot polynomial. The tests must be applied to the discs covering all suspect squares, that is, up to $4m$ squares per subdivision iteration.
A subdivision algorithm of [24] first maps the covering disc into the unit disc $D(0,1)$ (cf. (1.3)), thus losing the benefits where $p(x)$ is the sum of a small number of shifted monomials. Then the algorithm performs order of $\log(\log(d))$ root-squaring steps (one can apply order of $\log(d)$ root-lifting steps instead but within the same bit operation complexity estimates), and finally applies Pellet’s theorem to the resulting polynomial $t_{c,\rho}(x)=\sum_{i=0}^{d}t_{i}x^{i}$.
Transition from $p(x)$ to $t_{c,\rho}(x)$, however, involves scaling of the variable $x$ by factors ranging from $R/2^{b}$ to $R=\max_{j-1}^{d}|z_{j}|$ (besides root-squaring and shift of the variable). Hence for every $i$, $i=0,1,\dots,d$, the coefficient $t_{i}$ must be computed within $1/2^{b-i}$ to approximate the zeros of $p$ within $1/2^{b}$ (see Observation 24.2).
This involves at least $bg$ bits for $g=1,\dots,d$, that is, at least $0.5(d+1)db$ bits and $0.25(d+1)db$ bit-operations overall for a single e/i test.
A single subdivision iteration for Problem 1m can involve up to $4m$ e/i tests and hence up to at least $0.25(d+1)dbm$ bit-operations. This is by a factor of $m$ greater than for the root-finder of [96, 103] overall.
The bit-count above and the bit operation cost bounds for the root-finders of [96, 103, 24] decrease by a factor of $d$ where the minimal pairwise distance between the zeros of $p$ exceeds a positive constant, but the ratio of the bounds of [24] and [96, 103] remains at least $m$, where $m$ can be as large as $d$.
Similar argument shows that even representation of the leading coefficients $t_{d}$ of $t(x)$ alone after scaling for a singe e/i test involves odder of $bd$ bits, which is amplified to $d^{2}b$ for $t_{d}^{d}$ in every e/i test (see the end of Sec. 24.4) and up to $d^{2}bm$ per a subdivision iteration with $m$ e/i tests.
It may be interesting that a shift by $c$ such that $|c|\leq 2$ is much less costly than scaling:
Theorem 12.2.
Bit operation cost of the shift of the variable, [136, Lemma 2.3] (cf. [68, Lemma 3.6]). Given a positive $b$, the coefficients of a polynomial $p(x):=\sum_{i=0}^{d}p_{i}x^{i}$ such that $||p(x)||_{1}\leq 1$, and a complex $c$ such that $|c|\leq 2$, one can approximate within $1/2^{b}$ the coefficients $t_{0},t_{1},\dots,t_{d}$ of the polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}:=p(x-c)$ by performing $O(\mu((b+d)d))$ bit operations for
$\mu(s):=O(s\log(s)=\tilde{O}(s)$ (12.5)
bounding the number of bit operations for integer multiplication modulo $2^{s}$ [53] (cf. [142, 69, 134, 135] ).
We can relax the restriction $|c|\leq 2$ above by means of scaling the variable $x\mapsto xc$, but as we showed above, this would greatly increase the bound on the overall number of bit operations involved.
PART III: Classical subdivision root-finder with novel e/i tests based on root-lifting
Organization of Part III: We devote the next section to background. In Secs. 14 and 15 we cover our e/i tests and their extensions to $m$-tests, respectively. In Secs. 16 and 18 we estimate computational precision of our root-finders and their bit operation cost, respectively. In short Sec. 17 we extend our study of root-finding to approximation of matrix eigenvalues.
13 Background
13.1 Definitions
We recall or redefine some concepts from the Introduction and Part II.
  • ${\rm NR}_{t}(x):=\frac{t(x)}{t^{\prime}(x)}$ is Newton’s ratio, ${\rm NIR}_{t}(x):=\frac{t^{\prime}(x)}{t(x)}$ is Newton’s inverse ratio for a polynomial $t(x)$. We write NR$(x):={\rm NR}_{p}(x)$ and NIR$(x):={\rm NIR}_{p}(x)$ (cf. (3.2)).
  • Define squares, discs, circles (circumferences), and annuli on the complex plane: $S(c,\rho):=\{x:~{}|\Re(c-x)|\leq\rho$, $|\Im(c-x)|\leq\rho\}$$D(c,\rho):=\{x:~{}|x-c|\leq\rho\},$ $C(c,\rho):=\{x:~{}|x-c|=\rho\}$,  $A(c,\rho,\rho^{\prime}):=\{x:~{}\rho\leq|x-c|\leq\rho^{\prime}\}.$
  • $\Delta(\mathbb{S})$, $CH(\mathbb{S})$, $X(\mathbb{S})$ and #$\#(\mathbb{S})$ denote the diameter, root set, convex hull, and index (root set’s cardinality) of a set $\mathbb{S}$ on the complex plane, respectively.
  • A disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is $\theta$-isolated for $\theta\geq 1$ if $X(D(c,\rho))=X(D(c,\theta\rho))$, $X(C(c,\rho))=X(A(c,\rho/\theta,\rho\theta)$ or $X(S(c,\rho))=X(S(c,\theta\rho)),$ respectively.
    The isolation of a disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is the largest upper bound on such a value $\theta$ (see Fig. 5) and is denoted $i(D(c,\rho))$, $i(C(c,\rho))$, or $i(S(c,\rho))$, respectively.
    We say isolated for “$\theta$-isolated" if $\theta-1$ exceeds a positive constant.
    $\rho$$\rho\theta$$\rho\theta^{2}$c********
    Figure 5: The roots are marked by asterisks. The red circle $C(c,\rho\theta)$ and the disc $D(c,\rho\theta)$ have isolation $\theta$. The disc $D(c,\rho)$ has isolation $\theta^{2}$.
  • A disc $D$ has rigidity $\eta(\mathbb{S})=\Delta(X(D))/\Delta(D)\leq 1$. A disc is $\eta$-rigid for any $\eta\leq\eta(D)$.
    We say rigid for “$\eta$-rigid" if $\eta$ exceeds a positive constant.
  • $|u|:=||{\bf u}||_{1}:=\sum_{i=0}^{q-1}|u_{i}|$, $||{\bf u}||:=||{\bf u}||_{2}:=(\sum_{i=0}^{q-1}|u_{i}|^{2})^{\frac{1}{2}}$, and $|u|_{\rm max}:=||{\bf u}||_{\infty}:=\max_{i=0}^{q-1}|u_{i}|$ denote the 1-, 2-, and $\infty$-norm, respectively, of a polynomial $u(x)=\sum_{i=0}^{q-1}u_{i}x^{i}$ and its coefficient vector [^22] Here and hereafter ${\bf v}$ denotes a column vector, while ${\bf v}^{T}$ denotes its transpose. ${\bf u}=(u_{i})_{i=0}^{d-1}$, and then (see [48, Eqs. (2.2.5)–(2.2.7)])
    $||{\bf u}||_{\infty}\leq||{\bf u}||_{2}\leq||{\bf u}||_{1}\leq{\sqrt{q}}~{}||{\bf u}||_{2}\leq q||{\bf u}||_{\infty}.$ (13.1)
  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex point $c$ to the roots $y_{1},\dots,y_{d}$ of a $d$th degree polynomial $t(x)$. Write $r^{h}_{j}(c,t):=(r_{j}(c,t))^{h}$, $r_{j}(c):=r_{j}(c,p)$.
  • Generalize Def. 2.1 of $\sigma$-soft e/i tests as follows: For $p(x)$ of (1.1), $\sigma>1$, and integers $\ell$ and $m$ such that $1\leq\ell\leq m\leq d$ and #$\#(D(0,1))\leq m$, a $\sigma$-soft $\ell$-test, or just $\ell$-test for short (in particular e/i test turns into 1-test), either outputs 1 and stops if it detects that $r_{d-\ell+1}\leq\sigma$, that is, #$\#(D(0,\sigma))\geq\ell$, or outputs 0 and stops if it detects that $r_{d-\ell+1}>1$, that is, [^23] Bounds $r_{d-\ell+1}\leq\sigma$ and $r_{d-\ell+1}>1$ can hold simultaneously (cf. Fig. 1b), but as soon as an $\ell$-test verifies any of them, it stops without checking if the other bound also holds. #$\#(D(0,1))<\ell$. An $\ell$-test applied to the polynomial $t(y)$ of (1.3) is said to be an $\ell$-test for the disc $D(c,\rho)$ and/or the circle $C(c,\rho)$ and also to be an $\ell$-testc,ρ. [^24] We will only use $\ell$-tests for $\ell=1$ (which are e/i tests) and $\ell=m$.
Observation 13.1.
Taylor’s shift, or translation, of the variable $x\mapsto y:=x-c$ and its scaling combined (see (1.3)) map the zeros $z_{j}$ of $p(x)$ into the zeros $y_{j}=\frac{z_{j}-c}{\rho}$ of $t(y)$, for $j=1,\dots,d$, and preserve the index and the isolation of the disc $D(c,\rho)$.
13.2 Reverse polynomial
Immediately observe the following properties of the reverse polynomial of (1.5).
Proposition 13.1.
(i) The roots of $p_{\rm rev}(x)$ are the reciprocals of the roots of $p(x)$, and hence
$r_{j}(0,p)r_{d+1-j}(0,p_{\rm rev})=1~{}{\rm for}~{}j=1,\dots,d.$ (13.2)
(ii) The unit circle $C(0,1)$ has the same isolation for $p(x)$ and $p_{\rm rev}(x).~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
${\rm(iii)}~{}p^{\prime}_{\rm rev}(x)=dx^{d-1}p\Big{(}\frac{1}{x}\Big{)}-x^{d-2}p^{\prime}\Big{(}\frac{1}{x}\Big{)}~{}{\rm for}~{}x\neq 0,~{}p_{d}=p_{\rm rev}(0),~{}p_{d-1}=p^{\prime}_{\rm rev}(0).$
[…]
13.3 Root-squaring and root-lifting
The DLG root-squaring iterations (3.7) by Dandelin 1826, Lobachevsky 1834, and Gräffe 1837 compute the polynomials $p_{2^{i}}(x):=\prod_{j=1}^{d}(x-z_{j}^{2^{i}})$, for $i=0,1,\dots$. By applying $h$-lifting of (3.6) we extend the definition of $p_{h}(x)$ to all $h\in\mathbb{N}$.
As we noticed in the introduction, the known studies of DLG iterations always focus and stumble on numerically unstable computation of the coefficients of the polynomials $p_{i}(x)$ [86, 113], but we only compute their values bypassing those problems.
[…]
${\rm NIR}_{p_{h}}(x)=\frac{1}{{\rm NR}_{p_{h}}(x)}\approx\frac{p_{h}(x)-p_{h}(x-\delta)}{p_{h}(x)\delta}=\frac{p_{h}^{\prime}(y)}{p_{h}(x)}.$ (13.3)
Remark 13.1.
[…]
Observation 13.2.
(i) Root-lifting preserves the index #$\#(D(0,1))$ of the unit disc. (ii) $h$-lifting (achieved with root-squaring for $h=2^{k}$ and a positive integer $k$) maps the values $i(D(0,1))$, $\sigma(D(0,1))$, $\eta(D(0,1))$, and $r_{j}(0,p)$ for all $j$, that is, the isolation of the unit disc $D(0,1)$, its softness, its rigidity and root radii, to their $h$-th powers.
13.4 The distance to a closest root and fast inclusion tests
Next we estimate the distance from the origin to the closest root. The distance to the farthest root $z_{1}$ is given by the same estimates applied to the reverse polynomial $p_{\rm rev}(x)$.
For bounds $r_{d}$ and $r_{1}$ in terms of the coefficients of $p$, see [80, Sec. 1.10]; in particular according to Fujihara’a bounds of [41]:
$r_{1}\leq 2\max_{i=1}^{d}|p_{d-j}/p_{d}|^{1/j}$ unless $p_{d}=0$ and $r_{d}\geq 2\min_{i=1}^{d}|p_{j}/p_{0}|^{1/j}$ unless $p_{0}=0$.
For a complex $c$ and positive integers $j$ and $h$, Eqn. (3.1) and Observation 13.2 combined imply the following relationships independent of $p_{d}$:
$r_{d}(c)\leq d~{}|{\rm NR}(c)|,$ (13.4)
$r^{h}_{j}(0,p)=r_{j}(0,p_{h}).$ (13.5)
The roots lying far from a point $c$ little affect NIR$(c)$, and we extend bounds (13.4) as follows.
Proposition 13.2.
If #$\#(D(c,1))=\#(D(c,\theta))=m$ for $\theta>1$, then $|{\rm NIR}(c)|-\frac{d-m}{\theta-1}<\frac{m}{r_{d}(c)}.$
Proof.
We can assume that $|z_{j}|>\theta$ if and only if $j>m$ and then deduce from Eqn. (3.1) that $|{\rm NIR}(c)|\leq\sum_{j=1}^{m}\frac{1}{|c-z_{j}|}+\sum_{j=m+1}^{d}\frac{1}{|c-z_{j}|}<\frac{m}{r_{d}(c)}+\frac{d-m}{\theta-1}$. ∎
Certify inclusion into the disc $D(0,\sigma)$ if $\rho_{d}(0,p_{h})\leq\sigma^{h}$, $\sigma>1$ and $h>0$, although this bound is very poor for worst case polynomials $p$:
Proposition 13.3.
The ratio $|{\rm NR}(0)|$ is infinite for $p(x)=x^{d}-u^{d}$ and $u\neq 0$, while
$r_{d}(c,p)=r_{1}(c,p)=|u|.$
Proof.
Notice that $p^{\prime}(0)=0\neq p(0)=-u^{d}$, $p(x)=\prod_{j=0}^{d-1}(x-u\zeta_{d}^{j})$ for $\zeta_{d}$ of (3.5). ∎
Next we argue that such deficiency only occurs for a narrow class of polynomials $p$ and comment on some heuristic recipes to counter it.
Notice that $\frac{1}{r_{d}(c)}\leq|{\rm NIR}(c)|=\frac{1}{d}\Big{|}\sum_{j=1}^{d}\frac{1}{c-z_{j}}\Big{|}$ by virtue of (3.1), and so approximation to the root radius $r_{d}(c,p)$ is poor if and only if severe cancellation occurs in the summation of the $d$ roots.
Formal probabilistic support for this observation is hard, but intuitively such a cancellation only occurs for a very narrow class of polynomials $p$. If so, the above non-costly test certifies inclusion for a very large class of polynomials $p$ [^26] One can try some empirical recipes. Clearly, the problem persists for the root radius $r_{d}(w)$ where $p^{\prime}(w)$ vanishes; rotation of the variable $p(x)\leftarrow t(x)=p(ax)$ for $|a|=1$ does not fix it, but the bound $r_{d}\leq|p_{0}/p_{d}|^{1/d}$ and shifts $p(x)\leftarrow t(x)=p(x-c)$ for $c\neq 0$ combined should fix it. We must also study the case where e/i tests output exclusion, however, because this occurs in more than 25% of all e/i tests in a subdivision process:
Proposition 13.4.
Let $\sigma_{i}$ suspect squares enter the $i$-th subdivision step for $i=1,\dots,t+1$. Let it stop at $f_{i}$ of them, discard $d_{i}$ and further subdivide $\sigma_{i}-d_{i}-f_{i}$ of them. Write $\sigma_{1}:=4$, $\sigma_{t+1}:=0$, $\Sigma:=\sum_{i=1}^{t}\sigma_{i}$, $D:=\sum_{i=1}^{t}d_{i}$, and $F:=\sum_{i=1}^{t}f_{i}$. Then $3\Sigma=4(D+F-1)$ and $D>\Sigma/4$.
Proof.
Notice that $\sigma_{i+1}=4(\sigma_{i}-d_{i}-f_{i})~{}{\rm for}~{}i=1,2,\dots,t+1$, sum these equations for all $i$, substitute $\sigma_{1}=4$ and $\sigma_{t+1}=0$, and obtain $3\Sigma+4\geq 4D+4F$. Represent subdivision process by a tree with $F\leq m$ leaves whose other nodes are components made up of suspect squares, to obtain pessimistic lower bounds $\Sigma+1\geq 2m\geq 2F$ and hence $D\geq\Sigma/4+1/2$. ∎
14 Our e/i tests
14.1 Overview
Map (1.3) and Cor. 3.1 together reduce an e/i test for any disc to the test for the unit disc $D=D(0,1)$, and we devise soft e/i tests for that disc assuming that #$\#(D(0,2))\leq m$ for a fixed $m$, $1\leq m\leq d$. Our Alg. 14.1 of Sec. 14.2 and Alg. 14.2 of Sec. 14.3 are deterministic specifications of Alg. 3.1, which is always correct when it claims inclusion, by virtue of Lemma 3.1. We prove that both of these specifications are also correct when they output exclusion. Namely, Algs. 14.1 and 14.2 evaluate ${\rm NIR}(x)$ at $hq$ equally spaced points of the unit circle $C(0,1)$ for $h=O(\log(d))$ and correctly claim exclusion if at all these points $|{\rm NIR}(x)|\leq\frac{1}{2\sqrt{q}}$ for $q>d$ or if $|{\rm NIR}(x)|\leq\frac{1}{3\sqrt{q}}$ for $q>m$, respectively. Our Example 14.1 shows that we cannot devise such deterministic algorithms for $q$ decreased below $m$, but in Sec. 14.4 we devise Las Vegas randomized e/i tests that evaluate polynomial $p(x)$ at $O(\log(d))$ points $x$ overall under the Random Root Model.
14.2 Basic deterministic e/i test
[…]
Algorithm 14.1.
Basic deterministic e/i test.
INPUT: a black box polynomial $p(x)$ of a degree $d$ given by an oracle for its evaluation and $\sigma$, $1<\sigma<\sqrt{2}$.
INITIALIZATION: Fix two integers $q>d$ and $h:=\lceil\log_{\sigma}(1+2d\sqrt{q})\rceil$, such that $\sigma^{h}\leq 1+2d\sqrt{q}$.
COMPUTATIONS: For $g=0,1,\dots,q-1$ recursively compute
$v_{g}:=\Big{|}\frac{p^{\prime}_{h}(\zeta^{g})}{p_{h}(\zeta^{g})}\Big{|}.$ (14.1)
Stop and certify $\sigma$-soft inclusion #$(D(0,\sigma))>0$ if $1+d/v_{g}\leq\sigma^{h}$ for an integer $g<q$. Otherwise stop and certify exclusion, that is, #$(D(0,1))=0$.
Alg. 14.1 outputs $\sigma$-soft inclusion correctly by virtue of Lemma 3.1. Next we prove that it also correctly claims exclusion.
Lemma 14.1.
A polynomial $p$ has no roots in the unit disc $D(0,1)$ if $|p(\gamma)|>2\max_{w\in D(0,1)}|p^{\prime}(w)|$ for a complex $\gamma\in C(0,1)$.
Proof.
By virtue of Taylor-Lagrange’s theorem $p(x)=p(\gamma)+(x-\gamma)p^{\prime}(w)$, for any pair of $x\in D(0,1)$ and $\gamma\in C(0,1)$ and for some $w\in D(0,1)$. Hence $|x-\gamma|\leq 2$, and so $|p(x)\geq|p(\gamma)|-2~{}|p^{\prime}(w)|.$ Therefore, $|p(x)|>0$ for $\gamma$ satisfying the assumptions of the lemma and for any $x\in D(0,1)$. ∎
Proposition 14.1.
#$\#(D(0,1))=0$ for a black box polynomial $p(x)$ of a degree $d$ if
[…]
Proof.
Bounds (14.1) and (14.2) together imply that $|p^{\prime}(\zeta^{g})|/v\leq|p(\zeta^{g})|$ for $g=0,\dots,q-1$ and hence $2\sqrt{q}||(p^{\prime}(\zeta^{g}))_{g=0}^{q-1}||_{\infty}<\widehat{P}~{}{\rm for}~{}\widehat{P}:=||(p(\zeta^{g}))_{g=0}^{q-1}||_{\infty}.$ Combine this bound with (13.1) to obtain
$2||(p^{\prime}(\zeta^{g}))_{g=0}^{q-1}||_{2}<\widehat{P}.$ (14.3)
Pad the coefficient vector of the polynomial $p^{\prime}(x)$ with $q-d-1$ initial coordinates 0 to arrive at $q$-dimensional vector ${\bf p^{\prime}}$. Then $(p^{\prime}(\zeta^{g}))_{g=0}^{q-1}=F{\bf p^{\prime}}$ for the $q\times q$ matrix $F:=(\zeta^{ig})_{i,g=0}^{q-1}$ of discrete Fourier transform. Substitute $F{\bf p^{\prime}}$ for $(p^{\prime}(\zeta^{g}))_{g=0}^{q-1}$ in (14.3) and obtain $2||F{\bf p^{\prime}}||_{2}<\widehat{P}$. Hence $2\sqrt{q}||{\bf p^{\prime}}||_{2}<\widehat{P}$ because the matrix $\frac{1}{\sqrt{q}}F$ is unitary (cf. [48, 102]). It follows (cf. (13.1)) that $2||{\bf p^{\prime}}||_{1}<\widehat{P}.$ Since $\max_{x\in D(0,1)}|p^{\prime}(x)|\leq||{\bf p^{\prime}}||_{1}$, deduce that $2\max_{x\in D(0,1)}|p^{\prime}(x)|<\widehat{P}=||(p(\zeta^{g}))_{g=0}^{q-1}||_{\infty}$. Complete the proof by combining this bound with Lemma 14.1. ∎
Now, if Alg. 14.1 outputs exclusion, then $1+\frac{d}{v_{g}}>\sigma^{h}$ for $g=0,1,\dots,q-1$, where $\sigma^{h}>1+2d\sqrt{q}$ according to initialization of Alg. 14.1. Hence $\frac{1}{v_{g}}>2\sqrt{q}$, and so the assumptions of Prop. 14.1 hold for $p_{h}(x)$ replacing $p(x)$. Therefore, the theorem implies that the disc $D(0,1)$ contains no zeros of $p_{h}(x)$, but then it contains no zeros of $p(x)$ as well (see Observation 13.2 (iii)), and so exclusion is correct.
Readily verify the following estimates.
Proposition 14.2.
Alg. 14.1 is a $\sigma$-soft e/i test, certifying inclusion at the cost of evaluation of polynomial $p(x)$ at $2hj$ points $x$ or exclusion at the cost of evaluation at $2hq$ points $x$ for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m$, and $h:=\lceil\log_{\sigma}(1+2d\sqrt{q})\rceil$, $h=O(\log(d))$ for $\sigma-1$ exceeding a positive constant.
14.3 A faster deterministic e/i test
[…]
Next we supply some relevant estimates.
Algorithm 14.2.
A faster deterministic e/i test.
INPUT: a black box polynomial $p(x)$ of a degree $d$, two integers $m$ and $q>m$, and real $\sigma$ and $\theta>1$ such that $1<\sigma<\sqrt{2}$,
$|z_{j}|\leq 1~{}{\rm for}~{}j\leq m,~{}|z_{j}|>\theta~{}{\rm for}~{}j>m,$
$\frac{d-m}{\theta^{h}-1}\leq\frac{1}{6\sqrt{q}}~{}{\rm for}~{}h:=\lceil\log_{\sigma}(1+6m\sqrt{q})\rceil~{}{\rm such~{}that}~{}\sigma^{h}\geq 1+6m\sqrt{q}.$
COMPUTATIONS: Recursively compute the values $v_{g}$ of (14.1) for $g=0,1,\dots,q-1$. Certify $\sigma$-soft inclusion, that is, #$(D(0,\sigma))>0$, if $v_{g}\geq\frac{1}{3\sqrt{q}}$ for some integer $g<q$. Otherwise certify exclusion, #$(D(0,1))=0$.
Prop. 13.2, for $p_{h}$ replacing $p$, and the assumed bound $\frac{d-m}{\theta^{h}-1}\leq\frac{1}{6\sqrt{q}}$ together imply
$\frac{m}{r_{d}(\zeta^{g},p_{h})}\geq v_{g}-\frac{d-m}{\theta^{h}-1}\geq v_{g}-\frac{1}{6\sqrt{q}}~{}{\rm for~{}all}~{}g.$
Alg. 14.2 claims inclusion if $v_{g}\geq\frac{1}{3\sqrt{q}}$ for some $g$, but then
$\frac{m}{r_{d}(\zeta^{g},p_{h})}\geq\frac{1}{6\sqrt{q}},~{}r_{d}(\zeta^{g},p_{h})\leq 6m\sqrt{q},$
and hence
$r_{d}(0,p_{h})\leq 1+6m\sqrt{q}~{}\leq\sigma^{h}~{}{\rm for}~{}|\zeta|=1.$
Now Observation 13.2 (iii) implies that Alg. 14.2 claims inclusion correctly.
Next prove that it also claims exclusion correctly.
Proposition 14.3.
Let a polynomial $p(x)$ of a degree $d$ have exactly $m$ roots in the disc $D(0,\theta)$ for
$\theta^{h}\geq 1+6(d-m)\sqrt{q},~{}h:=\lceil\log_{\sigma}(1+6m\sqrt{q})\rceil,~{}{\rm and~{}an~{}integer}~{}q>m$
and let $v<\frac{1}{3\sqrt{q}}$ for $v$ of (14.2). Then #$\#(D(0,1))=0$.
Proof.
Let $p(x)$ and its factor $f(x)$ of degree $m$ share the sets of their zeros in $D(0,\theta)$. Then (3.1) implies that $\Big{|}\frac{f^{\prime}(x)}{f(x)}\Big{|}<\Big{|}\frac{p^{\prime}(x)}{p(x)}\Big{|}+\frac{d-m}{\theta-1}~{}{\rm for}~{}|x|=1,$ and so $\max_{g=0}^{q-1}\Big{|}\frac{f^{\prime}(\zeta_{q}^{g})}{f(\zeta_{q}^{g})}\Big{|}<v+\frac{d-m}{\theta-1}$. Recall that $\frac{d-m}{\theta-1}\leq\frac{1}{6\sqrt{q}}$ and $v<\frac{1}{3\sqrt{q}}$ by assumptions.
Combine these bounds to obtain $\max_{g=0}^{q-1}\Big{|}\frac{f^{\prime}(\zeta_{q}^{g})}{f(\zeta_{q}^{g})}\Big{|}<\frac{1}{2\sqrt{q}}$. To complete the proof, apply Prop. 14.1 for $f(x)$ replacing $p(x)$ and for $m$ replacing $d$. ∎
Readily extend Prop. 14.2 to Alg. 14.2 by decreasing its bounds on $q$ and $h$ provided that #$\#(D(0,\theta))=m$ for $\theta\geq 1+6(d-m)\sqrt{q}$.
Proposition 14.4.
Alg. 14.2 is a $\sigma$-soft e/i test: it certifies either inclusion or exclusion at the cost of evaluation of $p(x)$ at $2hj$ or $2hq$ points $x$, respectively, for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m:=\#(D(0,\theta)),~{}\sigma^{h}>1+6m\sqrt{q}$, $\theta\geq 1+6(d-m)\sqrt{q}$, and $h:=\lceil\log_{\sigma}(1+6m\sqrt{q})\rceil$, so that $h=O(\log(m))$ if $\sigma-1$ exceeds a positive constant.
14.4 A randomized e/i test
14.4.1. Introductory comments. Would Algs. 14.1 and 14.2 stay correct if we apply them at fewer evaluation points? They have consistently stayed correct for $q$ of order $\log^{2}(d)$ in extensive numerical tests of [60, 63] applied to $p(x)$ rather than to $p_{i}(x)$ for $i>0$. The tests reported exclusion where the values $|s_{h,q}|$ were reasonably small for $s_{h,q}:=\frac{1}{q}\sum_{g=0}^{q-1}\zeta^{(h+1)g}~{}\frac{p^{\prime}(\zeta^{g})}{p(\zeta^{g})},~{}{\rm for}~{}h=0,1,2$ (cf. Eqn. 7.2 in Part II), and reported inclusion otherwise. Clearly, $|s_{h,q}|\leq\delta$ for all $h$ if $|p(\zeta^{g})|\leq\delta/q$ for $g=0,1,\dots,q-1$, while inclusion is certified unless $|p(\zeta^{g})|$ is small for all $g$.
The test results of [60, 63], however, cannot be extended to worst case input polynomials $p(x)$ already for $q=m-1$.
Example 14.1.
Let $q=m-1$, $p(x)=(1-ux^{d-m})(x^{m}-mx+u)$, $u\approx 0$ but $u\neq 0$. (i) First let $m=d$. Then $p(x)=(1-u)(x^{d}-dx+u)$, $p^{\prime}(x)=(1-u)d(x^{d-1}-1)$ and NIR$(x)=0$ for $x$ being any $(d-1)$-st roots of unity. Then our e/i test for $q=d-1$ would imply that #$\#(D(0,1))=0$, while $p(y)=0$ for $u\approx 0$ and some $y\approx 0$. (ii) For $0<m<d$ notice that $p^{\prime}(x)=uf(x)+m(1-ux^{d-m})(x^{m-1}-1)$ where $f(x)=(d-m)x^{d-m-1}(x^{m}-mx+u)$ and so $uf(x)\approx 0$ while $1-ux^{d-m}\approx 1$ for $|x|\leq 1$ and $u\approx 0$. Therefore, $p^{\prime}(x)\approx 0$ where $x$ is any $(m-1)$-st root of unity, and then again extension of our e/i test to $q=m-1$ would imply that #$\#(D(0,1))=0$, while $p(y)=0$ for $u\approx 0$ and some $y\approx 0$.
Clearly, [60, 63] have never encountered worst case inputs in their extensive tests. Next we formally support this observation by introducing probability structure in the space of polynomials $p(x)$ and $t(x)=p(\frac{x-c}{\rho})$. Shifts destroy all customary probability distributions in the space of the coefficients of $p(x)$, but we prove that our e/i tests that evaluate $p(x)$ at $O(\log(d)$ points are correct whp under the Random Root Model.
[…]
Lemma 14.2.
Fix any constant $\beta>1$ and a sufficiently large integer $h$ of order $\log(d)$. Then under the Random Root Model, Alg. 3.1(rrm) is correct with a probability at least $1-P$, where
$P\leq\frac{\gamma d}{\pi R^{2}}~{}{\rm for}~{}\gamma:=\pi\max\Big{\{}\frac{\beta^{2}}{\sigma^{2}},O\Big{(}\frac{1}{\beta^{2h}}\Big{)}\Big{\}}.$ (14.4)
Proof.
By virtue of Lemma 3.1 Alg. 3.1(rrm) outputs $\sigma$-soft inclusion correctly and, therefore, can be incorrect only where it outputs exclusion for a disc $D(0,1)$ containing a root.
Let this root be $z_{1}$, say. Then fix the values of the random variables $z_{2},\dots,z_{d}$ and $x_{1}\in C(0,1)$ and let $\gamma$ denote the area (Lebesgue’s measure) of the set $\mathbb{S}$ of the values of the variable $z_{1}$ for which
$v_{1}=|{\rm NIR}_{t_{h}}(x_{1})|<\frac{d}{\sigma^{h}-1}.$ (14.5)
Then Alg. 3.1(rrm) is incorrect with a probability at most
$P=\frac{\gamma d}{\pi R^{2}}$ (14.6)
where the factor of $d$ enables us to cover the cases where the value of any of the $d$ iid random variables $z_{1},\dots,z_{d}$, and not necessarily $z_{1}$, lies in the disc $D(0,1)$.
Fix two values $\bar{z},z\in\mathbb{S}$ such that
$\bar{z}={\rm argmax}_{y\in\mathbb{S}}~{}|y|~{}{\rm and}~{}z={\rm argmax}_{y\in\mathbb{S}}~{}|y-\bar{z}|$
and notice that
$\gamma\leq\pi\min\{|\bar{z}|^{2}|,|\bar{z}-z|^{2}\}.$ (14.7)
To estimate $|\bar{z}-z|$, first write $y:=z^{h}\in D(0,1),~{}\bar{y}:=\bar{z}^{h}\in D(0,1)$ and prove that
$|\Delta|=|\bar{y}-y|<\nu_{h}~{}{\rm for}~{}\nu_{h}:=\frac{8d}{\sigma^{h}-1}.$ (14.8)
[…]
$|\Delta_{t_{h}}|<\frac{2d}{\sigma^{h}-1}~{}{\rm for}~{}\Delta_{t_{h}}:={\rm NIR}^{\prime}-{\rm NIR}.$ (14.9)
Now let $y_{1}:=z_{1}^{h},\dots,y_{d}=z_{d}^{h}$ denote the $d$ zeros of $t_{h}(x)$ and deduce from (3.1) that
$|{\rm NIR}|=\Big{|}\frac{1}{x_{1}-y}+S\Big{|}~{}{\rm and}~{}|{\rm NIR}^{\prime}|=\Big{|}\frac{1}{x_{1}-\bar{y}}+S\Big{|}~{}{\rm for}~{}S:=\Big{|}\sum_{j=2}^{d}\frac{1}{x_{1}-y_{j}}\Big{|}.$
Hence
$|\Delta_{t_{h}}|=|{\rm NIR}-{\rm NIR}^{\prime}|=\Big{|}\frac{1}{x_{1}-y}-\frac{1}{x_{1}-\bar{y}}\Big{|}=\frac{|\Delta|}{|(x_{1}-y)(x_{1}-\bar{y})|}.$ (14.10)
Recall that $x_{1}\in C(0,1)$, while $y,\bar{y}\in D(0,1)$, and so $\max\{|x_{1}-y|,|x_{1}-\bar{y}|\}\leq 2.$
Combine this inequality with (14.10) and deduce that $|\Delta_{t_{h}}|\geq|\Delta|/4.$
Combine the latter bound with (14.10) and obtain (14.8).
Now fix a large integer $g$, substitute $\bar{y}=\bar{z}^{h}$ and $y=z^{h}$ for $h=g$ and $h=g+1$, and obtain
$\lambda_{g+1}:=z^{g+1}-\bar{z}^{g+1},~{}\lambda_{g}:=z^{g}-\bar{z}^{g},~{}{\rm and~{}so}~{}z^{g+1}=z\bar{z}^{g}+z\lambda_{g},~{}|\lambda_{h}|\leq\nu_{h}~{}{\rm for}~{}h=g,g+1.$
Therefore, $\lambda_{g+1}+\bar{z}^{g+1}=\lambda_{g}z+\bar{z}^{g}z$, and so
$z-\bar{z}=\frac{\lambda_{g+1}-z\lambda_{g}}{\bar{z}^{g}},~{}|\bar{z}-z|\leq\frac{|\lambda_{g+1}|+|z|\cdot|\lambda_{g}|}{|\bar{z}|^{g}}<\frac{\nu_{g+1}+|z|\nu_{g}}{|\bar{z}|^{g}}=O\Big{(}\frac{d}{(\sigma|\bar{z}|)^{g}}\Big{)}.$
Hence $|z-\bar{z}|=O(1/\beta^{g})$ for a fixed $\beta>1$. Combine this bound with (14.6) and (14.7), choosing sufficiently large $g$ of order $\log d$ unless $\sigma|\bar{z}|<\beta$.
Combine this bound with (14.6) and (14.7), recall that $h=g$ or $h=g+1$, and obtain the lemma. ∎
For a fixed $\sigma>1$, e.g., for $\sigma=1.2$, choose a constant $\beta>1$ and sufficiently large $g=O(\log(d))$ such that $\gamma<\pi$ in (14.4). Then under the Random Root Model the probability $P$ of the output error for Alg. 3.1(rrm) is less than a constant $\nabla<1$. We can decrease the error probability below $\nabla^{v}$ by applying the algorithm $v$ times, in particular below $d^{a}$ for any fixed positive constant $a$ and for $v=O(\log(d))$.
Based on Cor. 2.2 we reduce classical subdivision root-finder to performing $O(mb)$ soft e/i tests, but in Part IV we decrease that bound to $w=O(m\log(b))$ tests. These tests magnify the error bounds by at most a factor of $w$. Since $m\leq d$, we can ensure correctness of the resulting e/i test whp provided that $\log(b)=o(d^{a-1})$. We arrive at the following
Corollary 14.1.
Under Random Root Model one only need to apply a subdivision root-finder that involves $O(m\log(b)))$ soft Las Vegas randomized e/i tests to ensure the output error bound $\epsilon=1/2^{b}$ whp provided that $\log(b)=d^{O(1)}$ and that every test is reduced to $v=O(\log(d))$ applications of Alg. 3.1(rrm).
Remark 14.1.
It is tempting to combine root-lifting or root-squaring with algorithms of [115], which approximate Cauchy integral (3.4), to obtain alternative randomized e/i test supporting Cor. 14.1, but for that we would have to raise disc isolation from order of $1+1/m$ to a constant, and this would require too costly lifting of order $m$.
14.5 Counting the evaluation queries of our root-finders
Combine Cors. 2.2 and 14.1 and Prop. 14.4 and obtain
Corollary 14.2.
We can reduce root-finding to deterministic evaluation of $p(x)$ at $O(m^{2}b\log(d))$ points as well as to evaluation at $O(mb\log^{2}(d))$ points under the Random Root Model.
15 Reduction of $m$-test to e/i tests
Our two algorithms of Secs. 15.1 and 15.2 reduce $m$-test for a disc $D$ to bounded number of e/i tests. Namely, the algorithm of Sec. 15.2 uses just a single e/i test where $m=d$ but up to order of $\log(d)$ e/i tests otherwise, while the algorithm of Sec. 15.1 uses $O(1)$ e/i tests for any $m$.
15.1 $m$-test by means of e/i tests for any disc
Our alternative reduction of $m$-test for $1\leq m\leq d$ to a bounded number of e/i tests can be applied in the case of any disc, not necessarily isolated. If at least one of the e/i tests fails, then definitely there is a root in a narrow annulus about the boundary circle of an input disc, and hence the internal disc of the annulus would not contain $m$ roots. Otherwise we would obtain a root-free concentric annulus about the boundary circle of an input disc and then reduce $m$-test to root-counting in an isolated disc, which we can perform at a low cost (see, e.g., claim (ii) of Prop. 7.1).
Algorithm 15.1.
Reduction of $m$-test to e/i tests.
INPUT: a polynomial $p(x)$ of a degree $d$, two values $\bar{\theta}>1$ and $\bar{\sigma}>1$, an integer $m$ such that $1\leq m\leq d$ and #$\#(D(0,\bar{\sigma}))\leq m$, and two algorithms $\mathbb{ALG}~{}0$ and $\mathbb{ALG}~{}1$, which for a black box polynomial perform $\bar{\sigma}$-soft e/i test for in any disc and root-counting in an isolated disc, respectively.
INITIALIZATION (see Fig. 6): Fix a positive $\rho$ such that $\bar{\sigma}=O(\rho)$, say, $\rho=\bar{\sigma}/5$, and compute the integer
$v=\lceil 2\pi/\phi\rceil$ (15.1)
where $\phi$ is the angle defined by an arc of $C(0,1)$ with end points at the distance $\rho$, and so $v=O(1)$ for $\bar{\sigma}=O(\rho)$. Then compute the values
$c_{j}=\exp(\phi_{j}{\bf i}),~{}j=0,1,\dots,v,$ (15.2)
such that $|c_{j}-c_{j-1}|=\rho$ for $j=1,2,\dots,v$ and $\phi_{j}-\phi_{j-1}=\phi$.
$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c_{5}$$c_{6}$$c_{7}$$c_{8}$$c_{9}$$c_{0}$
Figure 6: $v=9$, $|c_{j}-c_{j-1}|=\rho>|c_{9}-c_{0}|$ for $j=1,2,\dots,9$.
OUTPUT: Either an upper bound $r_{d-m+1}<1-\rho\frac{\sqrt{3}}{2}$ or one of the two lower bounds $r_{d-m+1}\geq 1-\bar{\sigma}\rho$ or $r_{d-m+1}>1+(\sqrt{3}-1)\rho$.
COMPUTATIONS: Apply algorithm $\mathbb{ALG}~{}0$ (soft e/i test) to the discs $D(c_{j},\rho)$, $j=0,1,\dots,v$. (a) Stop and certify that $r_{d-m+1}\geq 1-\bar{\sigma}\rho$, that is, #$\#(D(0,1-\bar{\sigma}\rho))<m$ for the open disc $(D(0,1-\bar{\sigma}\rho)$, unless algorithm $\mathbb{ALG}~{}0$ outputs exclusion in all its $v+1$ applications.
(b) If it outputs exclusion $v+1$ times, then apply algorithm $\mathbb{ALG}~{}1$ (root-counter) to the $\theta$-isolated disc $D(0,1)$ for
${}\theta:=\min\Big{\{}1+(\sqrt{3}-1)\rho,\frac{1}{1-0.5\rho\sqrt{3}}\Big{\}}.$ (15.3)
Let $\bar{s}_{0}=\#(D(0,1))$ denotes its output integer. Then conclude that
(i) $r_{d-m+1}\leq 1-\rho\frac{\sqrt{3}}{2}$, and so #$\#(D(0,1-\rho\frac{\sqrt{3}}{2}))=m$ if $\bar{s}_{0}=m$, while
(ii) $r_{d-m+1}\geq 1+(\sqrt{3}-1)\rho$, and so #$\#(D(0,1+(\sqrt{3}-1)\rho))<m$ for the open disc $D(0,1+(\sqrt{3}-1)\rho)$ if $\bar{s}_{0}<m$.
To prove correctness of Alg. 15.1 we need the following result.
Theorem 15.1.
For the above values $c_{0},\dots,c_{v}$, and $\rho$, the domain $\cup_{j=0}^{v-1}D(c_{j},\rho)$ covers the annulus $A(0,1-0.5\rho\sqrt{3},1+(\sqrt{3}-1)\rho)$.
Proof.
Let $c_{j-1}$ and $c_{j}$ be the centers of two neighboring circles $C(c_{j-1},\rho)$ and $C(c_{j},\rho)$, let $z_{j}$ and $z^{\prime}_{j}$ be their intersection points such that $|z_{j}|=|z_{j}^{\prime}|-|z_{j}-z^{\prime}_{j}|$ for $j=1,2,\dots,v$, observe that the values $|z_{j}-z_{j}^{\prime}|$, $|z_{j}|$, $|z^{\prime}_{j}|$, and $|c_{j-1}-c_{j}|$ are invariant in $j$, and hence so is the annulus $A(0,|z_{j}|,|z_{j}^{\prime}|)$ of width $|z_{j}-z_{j}^{\prime}|$ as well; notice that it lies in the domain $\cup_{j=0}^{v-1}D(c_{j},\rho)$.
$c_{0}$$c_{1}$$z$$z^{\prime}$
Figure 7: $|c_{1}-c_{0}|=|c_{0}-z|=|c_{1}-z|=|c_{0}-z^{\prime}|=|c_{1}-z^{\prime}|=\rho$.
Let us prove that $|z_{j}-z_{j}^{\prime}|=\rho\sqrt{3}$ for $j=1$, and hence also for $j=2,\dots,v$.
Write $z:=z_{1}$ and $z^{\prime}:=z_{1}^{\prime}$ and observe (see Fig. 7) that $|c_{1}-c_{0}|=|c_{0}-z|=|c_{1}-z|=|c_{0}-z^{\prime}|=|c_{1}-z^{\prime}|=\rho$, and so the points $c_{0}$, $c_{1}$, and $z$ are the vertices of an equilateral triangle with a side length $\rho$. Hence $|z-z^{\prime}|=\rho\sqrt{3}$. The circle $C(c_{j},\rho)$ intersects the line segment $[z_{j},z_{j}^{\prime}]$ at a point $y_{j}$ such that $|y_{j}-z_{j}|>\rho\sqrt{3}/2$ and $|y_{j}-z_{j}^{\prime}|>(\sqrt{3}-1)\rho$ for $\rho<1$, and this implies the theorem if $|c_{v}-c_{0}|=\rho$. If, however, $|c_{v}-c_{0}|<\rho$ and if $C(c_{0},\rho)\cap C(c_{v},\rho)=\{\bar{z},\bar{z}\}$, then $|\bar{z}-\bar{z}^{\prime}|>|z-z^{\prime}|$, and the theorem follows in this case as well. ∎
Corollary 15.1.
The disc $D(0,1)$ is $\theta$-isolated for $\theta$ of (15.3) provided that the discs $D(c_{j},\rho)$ contain no roots for $j=0,1,\dots,v$.
Correctness of Alg. 15.1. (i) If algorithm $\mathbb{ALG}~{}1$ outputs inclusion in at least one of its applications, then one of the discs $D(c_{j},\bar{\sigma}\rho)$ contains a root, and then Alg. 15.1 correctly reports that $r_{d-m+1}\geq 1-\sigma\bar{\rho}$.
(ii) Otherwise Cor. 15.1 holds, and then correctness of Alg. 15.1 follows from correctness of $\mathbb{ALG}~{}1$.
Softness is at most $\frac{1}{1-\bar{\sigma}\rho}$, $\frac{1}{1-0.5\rho\sqrt{3}}$, and $1+(\sqrt{3}-1)\rho$ in cases (a), (b,i), and (b,ii), respectively.
Computational complexity. Alg. 15.1 invokes algorithm $\mathbb{ALG}~{}0$ at most $v+1$ times for $v=O(1)$ of (15.1). If $\mathbb{ALG}~{}0$ outputs exclusion $v+1$ times, then algorithm $\mathbb{ALG}~{}1$ is invoked and runs at a dominated cost.
Remark 15.1.
One can modify Alg. 15.1 as follows (see [63]): instead of fixing a positive $\rho$ and then computing $v$ of (15.1) fix an integer $v>0$, apply algorithm $\mathbb{ALG}~{}0$ at $v$ equally spaced points $c_{j}=\exp(\frac{2j\pi{\bf i}}{v}),~{}j=0,1,\dots,v-1$, let $\phi=\frac{2\pi}{v}$, and then define $\rho$. We also arrive at this algorithm if we compute or guess this value $\rho$ and then apply Alg. 15.1, but notice that $c_{v}=c_{0}$ for $c_{j}$ of (15.2) and only invoke algorithm $\mathbb{ALG}~{}0$ for $c_{0},\dots,c_{v-1}$.
15.2 Simple reduction to a single e/i test in the case of an isolated disc
Clearly, we can immediately reduce the $d$-test in the unit disc $D(0,1)$ for $p(x)$ to performing 1-test (that is, e/i test) in that disc for the reverse polynomial $p_{\rm rev}(x)$ of (1.5).
Let us next reduce an $m$-test, for $0<m<d$, to a single application of 1-test provided that an input disc $D$ with $m$ roots is $\theta$-isolated for $\theta$ of order $\log(d)$; we can ensure that provision by applying root-lifting or root-squaring, in both cases at the cost of evaluation of $p(x)$ at $O(\log(d))$ points if an initial disc is just isolated.
Proposition 15.1.
Let the root sets of two polynomials $f(y)$ and $g(y)$ of degrees $m$ and $d-m$, respectively, lie outside the disc $D(0,1)$ and in the disc $D(0,1/\theta)$, respectively, for a sufficiently large real $\theta$. Then
${\rm NIR}_{f}(z)={\rm NIR}_{fg}(z)-\frac{d-m}{z}+\Delta~{}{\rm for}~{}|z|=1~{}{\rm and}~{}|\Delta|\leq\frac{d-m}{\theta-1}.$
Proof.
Notice that $\frac{1}{z-y_{j}}-\frac{1}{z}=\frac{y_{j}}{(z-y_{j})z}$, recall that $|y_{j}|\leq\frac{1}{\theta}$ for $\theta>1$ and all $d-m$ roots $y_{1},\dots,y_{d-m}$ of $g(y)$ by assumption, and deduce that
$\Big{|}\frac{1}{z-y_{j}}-\frac{1}{z}\Big{|}\leq\frac{1}{\theta-1}~{}{\rm for}~{}|z|=1~{}{\rm and}~{}j=1,\dots,d-m.$ (15.4)
[…]
Next fix an integer $q$ such that $m<q\leq 2m$, e.g., $q:=m+1$ or $q:=2^{h}$ for $h=\lceil\log_{2}(m+1)$. Assume that $\theta-1$ exceeds a positive constant and ensure that
$\theta>1+4(d-m)\sqrt{q}$ (15.5)
by applying root-lifting or root-squaring at the cost of evaluation of $p(x)$ at $O(\log(d))$ points.
[…]
$V_{f}:=\max_{g=0}^{q-1}|{\rm NIR}_{fg}(\zeta^{g})-(d-m)\zeta^{-g}|.$ (15.6)
Now ensure that $V_{f}<\frac{1}{2\sqrt{q}}$, say, by requiring that $\max_{g=0}^{q-1}|{\rm NIR}_{fg}(\zeta^{g})|<\frac{1}{4\sqrt{q}}$ because bound (15.5) implies that $\frac{d-m}{\theta-1}<\frac{1}{4\sqrt{q}}$.
16 Precision of computing
16.1 Doubling computational precision towards its a posteriori estimation
Given an algorithm $\mathbb{A}$ for black box polynomial computations and target tolerance TOL to the output errors, how should one choose computational precision? Here is a simple practical recipe from [10, 15, 22] and MPSolve for obtaining a posteriori estimate within a factor of two from optimal: first apply algorithm $\mathbb{A}$ with a lower precision, then recursively double it and stop where an output error decreases below TOL.
16.2 A priori estimates for computational precision
In the rest of this section we deduce a priori estimate of $O(b\log(d))$ bits for computational precision, proving that this precision supports correctness of our subdivision root-finders. Such an estimate is deduced in [76] based on advanced non-trivial techniques, of which we only use a very small part.
Estimating computational precision of our soft e/i tests we assume that they are applied to the unit disc $D(0,1)$ because we can extend the results to any disc $D(c,\rho)$ by applying map (1.3) and increasing the upper bounds on the precision by a factor of $\log_{2}(1/\rho)$. This factor is at most $b$ because we only handle discs of radii $\rho>\epsilon=1/2^{b}$ until we stop reaching $\epsilon$-discs. Therefore, our bound $O(\log(d))$ on the computational precision of our e/i tests for the disc $D(0,1)$ will imply the bound $O(b\log(d))$ for any disc.
We also make Assumption 0 that we can request and obtain from evaluation oracle (black box subroutine) the values of $p$ with a relative error bound in $1/d^{\nu}$ for any fixed constant $\nu>0$ of our choice; we can represent these values with the bit-precision $\log_{2}(d)+\log_{2}(\nu)$.
In our e/i tests we only need to compare the values $|{\rm NIR}_{t_{h}}(x)|$ at some complex points $x$ with some fixed values $v_{d}\geq 1/d^{\beta}$ for a constant $\beta>0$, and next we estimate that under Assumption 0 we can indeed do this by performing our computations with a precision of $O(\log(d))$ bits.
Multiplication with a precision $s$ contributes a relative error at most $2^{s}$ (cf. [21, Eqn. A.3 of Ch. 3]) and otherwise just sums the relative error bounds of the factors ignoring the dominated impact of higher order errors. Hence we deduce that $h=O(\log(d))$ multiplications for root-lifting to their $h$th powers increase these bounds by at most a factor of $h$, which is immaterial since we allow relative errors of order $1/d^{\nu}$ for a constant $\nu>0$ of our choice.
We only apply this analysis to approximation of NIR$(x)$ for $p_{h}(x)$ based on (3.2) and also use the following
Assumption 1: Whp all $q$ evaluation points $x\in C(0,1)$ of an e/i test for the unit disc $D(0,1)$ lie at the distance at least $\phi:=\frac{2\pi}{mqw}$ from all roots for any fixed $w>1$ of our choice.
This assumption holds because we apply our e/i tests under random rotation $x\mapsto vx$ for the random variable $v$ specified in our next proposition.
Proposition 16.1.
Let $\phi_{g}:=\min_{j=1}^{d}|v\zeta^{g}-z_{j}|$ and let $\phi:=\min_{g=0}^{q-1}\phi_{g}$ for the roots $z_{j}$, $\zeta$ of (3.5) and random variable $v$ sampled from a fixed arc of the circle $C(0,1)$ of length $2\pi/q$ under the uniform probability distribution on that arc. Let $i(D(0,1))=m$. Then $\phi\leq\frac{2\pi}{mqw}$ for $w>1$ with a probability at most $\frac{\pi}{w}$.
Proof.
Let $\psi_{j}$ denote the length of the arc $C(0,1)\cap D(z_{j},\frac{2\pi}{mqw})$, $j=1,2,\dots,d$. Then $\psi_{j}=0$ for at least $d-m$ integers $j$, and $\max_{j=1}^{d}\psi_{j}\leq\frac{\pi^{2}}{mqw}$ for sure. Hence $\sum_{j=1}^{d}\psi_{j}<\frac{\pi^{2}}{qw}$, and so $\phi_{g}\leq\frac{2\pi}{mqw}$ for a fixed $g$ with a probability at most $\frac{\pi}{qw}$. Therefore, $\phi\leq\frac{2\pi}{mqw}$ with a probability at most $\frac{\pi}{w}$. ∎
Next we prove that a precision of $O(\log(d))$ bits supports our e/i tests under Assumptions 0 and 1 by applying Prop. 16.1 for sufficiently large $w=d^{O(1)}$. We will (i) prove that $|{\rm NIR}(z)|=d^{O(1)}$, (ii) approximate this value within $d^{\nu}\delta$ for any fixed constant $\nu$ and for $\delta$ of (3.2) chosen such that $\log(\frac{1}{\delta})=O(\log(d))$, and (iii) estimate that such a bound withstands the impact of rounding errors of computations for (3.2).
Eqn. (3.1) immediately implies
Proposition 16.2.
Under Assumption 1, $\max_{g=1}^{q}|{\rm NIR}(\zeta^{g})|\leq{d^{\eta}}$ for $\eta=O(1)$.
Corollary 16.1.
A precision of $\lceil(\eta+\beta)\log_{2}(d)\rceil$ bits is sufficient to represent ${\rm NIR}(x)$ within $\frac{1}{d^{\beta}}$ for $g=1,2,\dots,q$.
Instead of ${\rm NIR}(x)=\frac{t^{\prime}(x)}{t(x)}$ for $x\in C(0,1)$ we actually approximate $\frac{t^{\prime}(y)}{t(x)}$ where $|y-x|\leq\delta$ and $t^{\prime}(y)$ is equal to the divided difference of (3.2). Thus we shall increase the above error bound by adding upper bounds $\alpha$ on $|\frac{t^{\prime}(x)}{t(x)}-\frac{t^{\prime}(y)}{t(x)}|$ and $\beta$ on the rounding error of computing $\frac{t(x)-t(x-\delta)}{\delta t(x)}$, for $x=\zeta^{g}$ and $g=1,2,\dots,q$.
In the proof of the next theorem we use the following lemma, which is [76, Fact 3.5].
Lemma 16.1.
For $t(x)=\prod_{j=1}^{d}(x-y_{j})$ and a non-negative integer $j\leq d$ it holds that
$t^{(j)}(x)=j!~{}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}$
where the summation $\sum_{S_{j,d}}$ is over all subsets $S_{j,d}$ of the set $\{1,\dots,d\}$ having cardinality $j$.
Proposition 16.3.
Under the assumptions of Prop. 16.2 it holds that
$|t(x)\alpha|=|\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)|\leq|t(x)\frac{(d\xi^{\prime})^{2}\delta}{1-d\xi^{\prime}\delta}|$ for $\xi^{\prime}=\frac{1}{\rho}$.
Proof.
The claimed bound on $|\xi^{\prime}|$ follows from (3.1). It remains to apply the first five lines of the proof of [76, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, respectively. Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
Substitute the expressions of Lemma 16.1 and obtain
$t(x)-t(x-\delta)=\delta t^{\prime}(x)+\sum_{j=2}^{\infty}\delta^{j}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}.$
Combine this equation with the assumed bounds on $\frac{1}{|x-y_{j}|}$ and deduce that
$\Big{|}\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)\Big{|}\leq\Big{|}\frac{f(x)}{\delta}\sum_{j=2}^{\infty}\delta^{j}w^{j}d^{j}\Big{|}\leq\Big{|}f(x)\frac{(d~{}w)^{2}~{}\delta}{1-\delta~{}d~{}w}\Big{|}.$
Corollary 16.2.
We can ensure that $\alpha<1/d^{\bar{\eta}}$ for any fixed constant $\bar{\eta}$ by choosing $\delta$ of order $d^{O(1)}$, represented with a precision of $O(\log(d))$ bits.
Proposition 16.4.
Suppose that the values $t(y)$ have been computed by a black box oracle within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor. 16.2. Then one can ensure that $\beta<\frac{1}{d^{\bar{\eta}}}$ for any fixed constant $\bar{\eta}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $O(\log(d))$ bits, which we can do under Assumption 0.
Proof.
Write
$\beta\delta=\frac{t(x-\delta)(1+\nabla(x-\delta))}{t(x)(1+\nabla(x))}-\frac{t(x-\delta)}{t(x)}=\frac{t(x-\delta)}{t(x)}~{}\frac{\nabla(x)-\nabla(x-\delta)}{1+\nabla(x)}.$
Taylor-Lagrange’s formula implies that
$\frac{t(x-\delta)}{t(x)}=1+\frac{t(x-\delta)-t(x)}{t(x)}=1+\delta\frac{t^{\prime}(u)}{t(x)}$, for $u\in[x-\delta,x]$.
Hence $|\frac{t(x-\delta)}{t(x)}|\leq 1+|\delta\alpha^{\prime}|+|\frac{t^{\prime}(x)}{t(x)}|$ for $\alpha^{\prime}=\frac{t^{\prime}(u)-t(x)}{t(x)}$.
By extending Cor. 16.2 obtain $|\alpha^{\prime}|\leq\frac{1}{d^{O(1)}}$ and deduce from (3.1) and Assumptions 0 and 1 that $|\frac{t^{\prime}(x)}{t(x)}|\leq\frac{1}{d^{O(1)}}$ (cf. the proof of Prop. 10.1). Hence $w:=|\frac{t(x-\delta)}{t(x)}|\leq d^{O(1)}$ and $|\beta\delta|\leq|w|\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}$.
Now choose $\nabla(y)$ such that $|\nabla(y)|\leq\frac{|\delta|}{20|w|}$ for $y=x$ and $y=x-\delta$. Then verify that $\log(\frac{1}{|\nabla(y)|})=O(\log(d))$, $1-|\nabla(x)|\geq 1-\frac{|\delta|}{20|w|}>\frac{19}{20}$, and $|\delta\beta|<\frac{|\delta|}{8}$. ∎
Combine Cors. 16.1 and 16.2 and Prop. 16.4 to obtain
Corollary 16.3.
Given a black box (oracle) for the evaluation of a polynomial $p(x)$ of a degree $d$ with a relative error of any order in $1/d^{O(1)}$ (this holds under Assumption 0), we can perform our black box subdivision root-finders with a precision of $O(\log(d))$ bits.
Remark 16.1.
We derived our precision bound $O(\log(d))$ (in the case of root-finding in the unit disc $D(0,1)$) based on operating entirely with NIR$(x)$. For comparison, if we relied on computing $p(x)$ and on the bound $\min_{x:~{}|x|=1}(|p(x)|/\sum_{i=0}^{d}|p_{i}|)\geq(\frac{\theta-1}{2\theta})^{d}$ of [134, Eqn. (9.4)], we would have only obtained a bound of order $m^{2}d$ on the computational precision of our root-finders.
17 Approximation of matrix eigenvalues
Given a $d\times d$ matrix $M$ that has characteristic polynomial $p(x)=\det(xI-M)$, our root-finders can approximate all $m$ its zeros (that is, the eigenvalues of $M$) lying in a fixed disc $D$ on the complex plane isolated from external eigenvalues.
Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [138, Thm. 1.3.2]: $|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|$ for some $i$, $1\leq i\leq d$.
Hence, clearly, $|\lambda|\leq||M||_{F}$.
Combine this bound with the precision estimates of the previous section for our root-finders [^27] For a posteriori estimates, one can combine our root-finders with the recipe of doubling the precision of computing of (see Sec. 16.1). and with the following result of Storjohann [137] (cf. Remark 17.1):
Theorem 17.1.
One can evaluate the determinant of a $d\times d$ integer matrix $A$ by using an expected number of $O(d^{\omega}\log^{2}(d)\log(|A||_{F}))$ bit operations.
This implies a record bit operation complexity bound $\tilde{O}(m^{2}\tilde{b}^{2}d^{\omega})$ for any matrix $M$ which decreases whp by a factor of $m$ for matrices $M$ whose eigenvalues are iid random variables sampled under the uniform probability distribution from a large disc on the complex plane (cf. (1.6)) for approximation within $\epsilon$ of all $m$ eigenvalues of $M$ lying in the disc $D$ provided that $\tilde{b}=\log_{2}(||M||_{F}/\epsilon)$. This supports a bound exceeding (1.6) by a factor of $b$, to be removed with acceleration of our root-finding in Part III.
Remark 17.1.
Our root-finders are essentially reduced to the evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$, and we can rely on the following expressions for approximation of the eigenvalues of a matrix or a matrix polynomial:
$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}((xI-M)^{-1})}$
where $M$ is a matrix and $p(x)=\det(xI-M)$ (cf. (3.1)) and
$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}(M^{-1}(x)M^{\prime}(x))}$
where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [20, Eqn. (5)]. [^28] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and of its determinant. Based on the above expressions one can reduce approximation of eigenvalues of a matrix or a matrix polynomial to its inversion. This computation tends to be more numerically stable than computing determinants and is significantly accelerated the highly important case where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
18 Bit operation complexity of general polynomial root-finding
Next we estimate the bit operation complexity of our root-finders, which accelerate classical subdivision root-finding by incorporating our e/i tests and are applied to a general polynomial $p$ of (1.1) whose coefficients are given with a precision sufficient to support our root-finding.
We could simply combine the precision bound $O(b\log(d))$ of Sec. 16 for our e/i tests with the estimate $O(bm)$ of Cor. 14.2 for the number of evaluation points for our root-finders, the bound of $2d$ ops for the evaluation of $p(x)$, and (12.5) for bit-operation cost of an op and then would arrive at the bit-operation complexity bound $\tilde{O}(b^{2}dm)$ for our root-finders. [^29] Recall that under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $m$.
Next, however, we decrease this bound roughly by a factor of $d$ by using benefits of fast multipoint evaluation of a polynomial by means of the algorithm of Fiduccia [42] and Moenck and Borodin [84] (see [19, Sec. 4.5] for historical account), non-trivially extended in [81] to support the following theorem:
Theorem 18.1.
Given a positive $b$, the coefficients of a polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}$ such that $||t(x)||_{1}=\sum_{i=0}^{d}|t_{i}|\leq 2^{\tau}$, and $q$ complex points $x_{1},\dots,x_{q}$ in the unit disc $D(0,1)$, one can approximate the values $p(x_{1}),\dots,p(x_{q})$ within $1/2^{b}$ by using $\tilde{O}((d+q)(b+\tau))$ bit operations.
Proof.
For $q=d$ this is [81, Thm. 2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Guillaume Moroz in [81]. Trivially extend this theorem to the cases where $q>d$ (by considering $t(x)$ a polynomial of degree $q$ with $q-d$ leading coefficients 0) and where $q<d$ (by removing $d-q$ points $x_{i}$, with no increase of the complexity bound). ∎
We only apply this theorem for $\tau=0$, in which case its bit operation cost bound turns into
$\mathbb{B}_{\rm Moroz}=\tilde{O}((d+q)b)$ (18.1)
For comparison, the previous State of the Art was represented by the bound
$\mathbb{B}_{\rm Kirr}=\tilde{O}((d+q)(b+q)),$ (18.2)
due to [68, Thm. 3.9, Alg. 5.3, and Appendix A.3] by Peter Kirrinnis, which relied on combining his and Schönhage’s study of bit operation complexity of basic polynomial computations (cf. [134]) with the algorithm of [42, 84]. [^30] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by applying the Fast Multipole Method of Greengard and Rokhlin [50, 28, 105, 106] to the equivalent task of multiplication of Vandermonde matrix by a vector. One can, however, support bounds of Eqns. (18.1) and (18.2) on the overall bit operation complexity by computing with multiple precision for larger values $b$. Moroz obtained [81, Thm. 2] by applying bound (18.2) to his lower degree polynomials approximating $p(x)$.
The algorithms of [42, 84, 68, 81] extensively involve the coefficients of $p(x)$.
Bound (18.1) combined with the precision bound $O(\log(d))$ of Sec. 16 implies the bit operation cost bound $\tilde{O}((d+q)b)$ dominated by an estimated bit operation cost $\mathbb{B}$ for the evaluation of $p(x)$ with relative error $1/d^{O(1)}$ at $\tilde{O}(db)$ points. It remains to estimate $\mathbb{B}$, which is our next goal.
18.1 Reduction of our root-finders to multipoint polynomial approximation
Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ roots, we approximate all these roots within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with e/i tests of Sec. 14. Then every subdivision step is reduced to application of our e/i tests to $\bar{m}=O(m)$ discs $D(c_{\lambda},\rho_{\lambda})$, where $\rho_{\lambda}\geq R/2^{b}$ for all $\lambda$. Such a test amounts to approximation of $p(x)$ at $q$ equally spaced points $x$ on each circle $C(c_{\lambda},\rho_{\lambda})$.
In Thm. 18.1 a polynomial $t(x)$ with $||t(x)||_{1}\leq 1$ is approximated at some points $x\in D(0,1)$. To ensure the assumption about $x$, we scale the variable $x\mapsto Rx$ to map the discs $D(c_{\lambda},\rho_{\lambda}\sigma)\mapsto D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})\subseteq D(0,1)$, for $\bar{c}_{\lambda}=c_{\lambda}/R$, $\bar{\rho}_{\lambda}=\rho_{\lambda}/R\geq 1/2^{b}$, and all $\lambda$.
Then we write $t(x):=p(Rx)/\psi$ for $\psi=O(R^{d})$ such that $||t(x)||_{1}=1$ and approximate $t(x)$ within $1/2^{\bar{b}}$ or equivalently within the relative error bound $1/(|t(x)|2^{\bar{b}})$ at $\bar{m}$ points $x$ in $\bar{m}$ e/i tests. [^31] Under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $\bar{m}/\log(\bar{m})$ and then extend this decrease to the related subsequent estimates. Due to Cor. 16.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we achieve this by choosing $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$ (cf. Assumption 1 of Sec. 16.2). Then $\log(\frac{1}{|t(x)|})=O(d\log(m))$ by virtue of the bound of [134, Eqn. (9.4)]). [^32] That bound follows from the estimate $|f|\geq 1/2^{d}$ of [134, Thm. 4.2], which holds provided that $t(x)=f\prod_{j=1}^{m}(x-x_{j})\prod_{j=m+1}^{d}(\frac{x}{x_{j}}-1)$, $||t(x)||_{1}=1$, $|x_{j}|<1$ for $j\leq m$, and $|x_{j}|>1$ for $j>m$.
Substitute $\bar{\rho}_{\lambda}\geq 1/2^{b}$ and obtain $\bar{b}=\log(\frac{1}{|t(x)|})+O(\log(d))=O(d\log(m)+b)$.
Remark 18.1.
For $m=d$, the above bound on $\log(\frac{1}{|t(x)|})$ is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$. For $m<d$, [134, Thm. 4.5] implies a little stronger bounds, namely, $\log(\frac{1}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.
18.2 Bit operation cost of our e/i tests and root-finders
Apply Thm. 18.1 with $q$ and $b$ replaced by $\bar{q}=O(m^{2})$ and $\bar{b}=O(b+d+m\log(m))$, respectively, and obtain that the bit operation cost of our $O(m)$ e/i tests of Sec. 14 at any fixed subdivision step is in $\tilde{O}((b+d+m)m^{2})$ for $1\leq m\leq d$. This implies the upper estimate
$\mathbb{B}_{\rm roots}=\tilde{O}((b+d)(m^{2}+d)b)$ (18.3)
for the overall bit operation cost of our $O(b)$ subdivision steps. [^33] Under the Random Root Model of Sec. 1.5 we can decrease this bound by a factor of $m$ whp and can extend this decrease to the related subsequent estimates.
Based on [68, Alg. 5.3] and Eqn. (18.2) we obtain a little larger but still reasonably small upper bound
$\mathbb{B}_{\rm roots}^{\prime}=\tilde{O}(((m^{2}+d)(m^{2}+d+b)b).$ (18.4)
In Part IV we decrease the number of subdivision steps of our root-finders from order of $b$ to $O(\log(b))$, thus decreasing bounds (18.3) and (18.4) by a factor of $b$, and arrive at our Main Theorems 2, 3, and 4.
PART IV: Compression of a Disc and Acceleration of Subdivision Root-Finders
19 Overview and two alternative approaches
Classical subdivision iterations converge to the roots with linear convergence rate (see Sec. 2), which is slow when many consecutive iterations process a component $C$ made up of a single suspect square or up to four of them sharing a vertex. In this case the algorithms of [128, 101] dramatically accelerate convergence by means of applying Newton’s or Schröder’s iterations (see Fig. 8), which converge with superlinear rate to the convex hull of all roots lying in $C$. The diameter of the convex hull, however, can be too large, and then [128, 101] apply subdivision iterations again.
The resulting root-finder recursively intertwines subdivision iterations with Newton’s or Schröder’s, and this complicates the complexity estimates for a root-finder. [128, 101] have only estimated the number of ops but not of the bit operations involved. [23, 24] estimated the bit operation complexity, but only under a very strong assumption about the output (see Remark 4.2 and Sec. 12.4).
Figure 8: Five blue marked centers of suspect squares converge to a green-marked root cluster with linear rate; two red marked Schröder’s iterates converge with quadratic rate.
Instead of Newton’s or Schröder’s iterations, one can apply the algorithm of [134], which splits $p(x)$ into the product of two polynomials $f(x)$ and $g(x)$ whose zero sets lie in and outside the minimal disc covering the component $C$, respectively. Then one can apply a root-finder to the polynomials $f=f(x)$ and $g=g(x)$. Splitting of $p(x)$ into factors does not preserve sparsity of $p(x)$, and the factors may have significantly larger coefficient length than $p(x)$. (E.g., split $p(x)=x^{1000}+1$ into two factors of degrees 300 and 700.)
Overall asymptotic bit-complexity for high precision root-finding can still be near-optimal if a root-finder recursively splits $p(x)$ into factors of comparable degrees. This is achieved in [96, 103], but the resulting root-finder is much involved and extensively uses the coefficients of $p$.
The problems caused by splitting disappear if one of the factors, say, $g$ has a small degree; then we can deflate it, approximate its zeros readily, and approximate the zeros of $f(x)=p(x)/g(x)$ by applying to it a black box root-finder and using implicit deflation based on (3.3) or the equation NIR$(p)=$NIR(f)+limit-$(f)+$NIR$(g)$ for $p=fg$.
One would lose optimality, however, by always deflating low-degree factors: to deflate $k$ times one would need at least order of $bkmd$ bit operations overall, that is, would exceed the lower bound $0.25(d+1)db$ of Cor. 24.2 by at least a factor of $k$.
We stay with subdivision framework but instead of Newton’s or Schröder’s iterations, used in [128, 101], apply our novel algorithm that compresses a component isolated from its external roots into a near-minimal disc that keeps all internal roots. This enables us to decrease the estimates of Secs. 12.3 and 18.2 for the overall complexity of root-finding by a factor of $b$, thus completing proof of our Main Theorems 2-4.
We organize our presentation in this part as follows. In the next section we study a tree representing recursive partition of the union of suspect squares in a subdivision process into connected components. In Sec. 21 we devise and analyze our first compression algorithm. In particular we observe in Sec. 21.4 that the whole subdivision process involves at most $12m-$ 1 compression steps, reduced to $O(m)$ e/i tests. In Sec. 22 we prove that in addition it also involves $O(m)$ e/i steps. We reach a near-optimal bound on the bit operation complexity of our root-finders in Sec. 23 by devising and analyzing our compression algorithm and recalling our estimates for the complexity of an e/i step. In that section we also present a fast root-counter in an isolated disc.
We use some results of Part I, Alg. 9.1 of Part II, and Sec. 13 (Background) of Part III. Otherwise our presentation in this part of the paper is self-contained.
20 Component tree and compact components
20.1 Component tree, edge length, non-adjacent suspect squares, and compact components
Recall the definition of the component tree $\mathbb{T}$ of Sec. 4, which represent partition of the union of suspect squares at every subdivision step and representation of such partition with the component tree $\mathbb{T}$ of Sec. 4. The tree has at most $m$ components-leaves that represent the output $\epsilon$-discs and at most $m-1$ other vertices, which represent components-parents, each broken by a subdivision step into at least two components-children. The initial suspect square is the root of the tree $\mathbb{T}$. The edges of the tree connect parents and children, representing subdivision steps between partitions of components. Cor. 20.1 below bounds the number of such steps; one can consider this number the edge length.
Call a pair of suspect squares non-adjacent if they share no vertices.
Call a component compact if it contains just a single non-adjacent suspect square, that is, consists of at most four suspect squares sharing a vertex (see Fig. 10).
Figure 9: Two discs superscribing two non-adjacent suspect squares are not minimal but do not overlap.
Observation 20.1.
(See Fig. 9.) The $\sigma$-covers of two non-adjacent suspect squares are separated by at least the distance $a\Delta$ for the side length $\Delta$ and separation coefficient
$a:=2-\sigma\sqrt{2}>0~{}{\rm for}~{}1<\sigma<\sqrt{2}.$
Lemma 20.1.
Let $|z_{g}-z_{h}|\geq\sigma\Delta$, for a pair of roots $z_{g}$ and $z_{h}$, the softness $\sigma$ of e/i tests, and the side length $\Delta$ of a suspect square, and let $i>0.5+\log_{2}(m)$. Then in at most $i$ subdivision steps the roots $x$ and $y$ lie in distinct components.
Proof.
Two roots $z_{g}$ and $z_{h}$ lying in the same component at the $i$th subdivision step are connected by a chain of $\sigma$-covers of at most $m$ suspect squares, each having diameter $\sigma\Delta/2^{i-0.5}$. Hence
$|z_{g}-z_{h}|\leq m\sigma\Delta/2^{i-0.5},~{}i\leq\log_{2}\Big{(}\frac{m\sigma\Delta}{|z_{g}-z_{h}|}\Big{)}+0.5.$
Substitute the bound $|z_{g}-z_{h}|\geq\sigma\Delta$ and deduce that $m\leq 2^{i-0.5}$ or equivalently $i\leq 0.5+\log_{2}(m)$ if the roots $z_{g}$ and $z_{h}$ are connected by a chain of suspect squares, that is, lie in the same component. This implies the lemma. ∎
Corollary 20.1.
A component is broken in at most $1+\lfloor 0.5+\log_{2}(m)\rfloor$ successive subdivision steps unless it is compact.
20.2 Isolation of compact components
Observation 20.2.
(i) A compact component is surrounded by a frame of discarded squares (see Fig. 10) and hence is covered by a 2-isolated square. (ii) If it stays compact in $h$ successive subdivision steps, then its minimal covering square and disc become $2^{h+1}$-isolated and $2^{h}\sqrt{2}$-isolated, respectively.
Proof.
Readily verify claim (i). Claim (ii) follows because a subdivision step can only thicken the frame of discarded squares while decreasing at least by twice the diameter of any component that stays compact. ∎
21 $\epsilon$-compression: algorithm and analysis
21.1 Compression algorithm: high level description
Recall that the classical subdivision iterations converge slowly where they successively process only compact components (cf. Fig. 8), but then we accelerate this process by interrupting it with compression steps in at most three iterations.
Next we specify compression for a 3.5-isolated unit disc $D=D(0,1)$ covering a component; map (1.3) extends compression to any 3.5-isolated disc $D(c,\rho)$; the map $\rho\epsilon\mapsto\epsilon$ adjusts the error tolerance $\epsilon$ by scaling by $\rho$. Observation 20.2 implies that we can readily ensure $\theta$-isolation at the cost of performing $O(\log(\theta))$ additional subdivision steps.
Algorithm 21.1.
$\epsilon$-compression (high level description).
INPUT: a positive $\epsilon\leq 1$ and a black box polynomial $p(x)$ of a degree $d$ such that [^34] Newton’s and Schröder’s iterations of [128, 101] accelerate convergence of subdivision iterations to a root cluster in a disc if the disc is $5d^{2}$-isolated, while Alg. 21.1 only requires 3.5-isolation (see (21.1)).
$i(D(0,1))\geq 3.5~{}{\rm and}~{}\#(D(0,1))\geq 1.$ (21.1)
OUTPUT: an integer $m$, a complex $c^{\prime}$, and a positive $\rho^{\prime}$ such that $m=\#(D(0,1))$, (i) $X(D(c^{\prime},\rho^{\prime}))=X(D(0,1))$, and (ii) $\rho^{\prime}\leq\epsilon$ or $8|z_{g}-z_{h}|>5\rho^{\prime}$ for a pair of roots $z_{g},z_{h}\in D(c^{\prime},\rho^{\prime})$. [Hence the disc $D(c^{\prime},\rho^{\prime})$ is 5/8-rigid.]
INITIALIZATION: Fix $\delta\geq 0$ such that $8\delta\leq\epsilon\leq 1$.
COMPUTATIONS: Compute $m=\#(D(0,1))$, $c^{\prime}$, and $\rho^{\prime}$ such that
$|c^{\prime}-c|\leq\delta~{}{\rm and}~{}\rho\leq\rho^{\prime}\leq\max\{\epsilon,\rho+2\delta\}$ (21.2)
for $\rho:=r_{d-m+1}(c^{\prime})$ and some $c\in CH(X(D(0,1)))$.
21.2 Correctness of compression algorithm
First prove correctness in a simple case where $\epsilon=\delta=0$, $c^{\prime}=c$, $\rho^{\prime}=\rho$.
In this case $c^{\prime}=c\in CH(X(D(0,1)))\subset D(0,1)$, and so $\rho^{\prime}=\rho=r_{d+1-m}(c^{\prime})\leq 2$. Hence $D(c^{\prime},\rho^{\prime})\subseteq D(0,3)$.
Furthermore, #$\#(D(c^{\prime},\rho^{\prime}))=m$ since $\rho^{\prime}=\rho:=r_{d+1-m}(c^{\prime})$, while #$\#(D(0,1))=m$ by assumption.
Now (21.1) implies that #$\#(D(0,3))=m$, which supports claim (i) for the output of Alg. 21.1.
Next let $|c^{\prime}-z_{g}|=\rho:=r_{d+1-m}(c^{\prime})$ for a root $z_{g}$. Then all $m$ roots in $D(0,1)$ also lie in the disc $D(c^{\prime},\rho)=D(c^{\prime},\rho^{\prime})$ by virtue of claim (i), just proved.
Since $c^{\prime}=c\in CH(X(D(0,1))$, however, there is a root, $z_{h}$, lying in $D(c^{\prime},\rho^{\prime})$ but not in $D(z_{g},\rho)$, such that $|z_{g}-z_{h}|>\rho\geq\rho^{\prime}-2\delta$ (cf. (21.2)).
Hence $|z_{g}-z_{h}|>3\rho^{\prime}/4$ (implying claim (ii) for the output of Alg. 21.1) because $42\delta\leq\epsilon/4<\rho^{\prime}/$ 4.
Next extend this correctness proof to the general case based on bounds (21.2).
Figure 10: Roots (asterisks) define compact components. A subdivision step halves their diameters and at least doubles isolation of their minimal covering squares and discs.
Notice that $c^{\prime}\in D(0,1+\delta)$ because $c\in CH(X(D(0,1)))\subset D(0,1)$ and $|c^{\prime}-c|\leq\delta$.
Hence $\rho=r_{d-m+1}(c^{\prime})\leq 2+\delta$, and so $\rho^{\prime}\leq\rho+2\delta\leq 2+3\delta$ (cf. (21.2)) and $D(c^{\prime},\rho^{\prime})\subseteq D(0,3+4\delta)$.
Therefore $D(c^{\prime},\rho^{\prime})\subseteq D(0,3.5)$ because $8\delta\leq\epsilon\leq 1$ by assumption.
Now recall that #$\#(D(0,3.5))=\#(D(0,1))=m$, while #$\#(D(c^{\prime},\rho^{\prime}))\geq\#(D(c^{\prime},\rho))=m$ for $\rho^{\prime}\geq\rho$ and deduce claim (i) for the output of Alg. 21.1.
Next assume wlog that $c=0$, $c^{\prime}=\delta$, and $z_{g}=c^{\prime}+\rho$ for a root $z_{g}$.
Then $\Re(z_{g})\geq\rho-\delta\geq\rho^{\prime}-3\delta$ because $|c^{\prime}|=|c^{\prime}-c|\leq\delta$ and $\rho^{\prime}\leq\rho+2\delta$ by assumptions.
Furthermore, $\Re(z_{h})<0$ for some root $z_{h}$ because $c=0\in CH(X(D(0,1)))$. (Actually, $\Re(z_{h})$ reaches its maximum $-\frac{z_{h}}{m-1}$ where $z_{h}$ is a simple root and $-\frac{z_{h}}{m-1}$ is a root of multiplicity $m-1$.)
Therefore, $|z_{g}-z_{h}|\geq\Re(z_{g})-\Re(z_{h})>\rho^{\prime}-3\delta\geq\rho^{\prime}-\frac{3}{8}\epsilon$.
This implies claim (ii) for the output of Alg. 21.1 since $\rho^{\prime}>\epsilon$.
21.3 Elaboration upon the compression algorithm
We readily (i) count roots in isolated input disc $D(0,1)$ of Alg. 21.1 by applying the fast root-counter of Sec. 23.4 or alternatively Alg. 7.1 and Prop. 7.1 and (ii) compute a complex point $c^{\prime}$ lying in or near the convex hull of the roots in $C$ by applying fast algorithm of Sec. 23. Finally, we estimate the root radius $\rho^{\prime}:=r_{d-m+1}(c^{\prime})$ by applying Alg. 9.1, which incorporates the $m$-test of of Sec. 23.4, or alternatively Alg. 7.1 and Prop. 7.1.
21.4 Counting $\epsilon$-compression steps
Recall from Sec. 1.3 that the output disc $D(c^{\prime},\rho^{\prime})$ of Alg. 21.1 is said to be an $\epsilon$-disc if $\rho^{\prime}\leq\epsilon$; in this case call the compression by Alg. 21.1 final and output that disc.
Otherwise call the square $S=S(c^{\prime},\rho^{\prime})$ suspect and apply to it subdivision iterations.
Extend Lemma 20.1 to the case where $|z_{g}-z_{h}|\geq\frac{5}{8}\rho^{\prime}$ for a pair of roots $z_{g},z_{h}\in D(c^{\prime},\rho^{\prime})$, combine it with claim (ii) for the output of Alg. 21.1, and obtain that at most $2.5+\log_{2}(8\sigma m/5)$ subdivision steps break the square $S$ into more than one component.
From part (ii) of Sec. 4 recall the bound $22m-$ 2 on the number of compression steps and obtain
Theorem 21.1.
For a fixed $\epsilon\geq 0$ let subdivision and compression steps be applied to a component $C$ of suspect squares containing at most $m$ roots. Then the computations involve at most $66m-$ 6 compact components, made up of at most $2424m-$ 24 suspect squares; that is, these computations involve at most $2424m-$ 24 e/i tests.
Proof.
A compact component stays unbroken and uncompressed in at most $3$ consecutive subdivision steps, which end with a final compression step or follow a non-final compression step and end with breaking a component. Hence all such components are processed in at most $66m-$ 6 subdivision step, each involving at most 4 e/i tests. ∎
22 Counting e/i tests
22.1 An upper estimate
The upper estimate $22m-$ 2 of part (ii) of Sec. 4 on the number of compression steps is sharp – reached in the case of the complete binary tree $\mathbb{T}$. Each compression step is preceded by an e/i test, and so we perform at least $22m-$ 2 e/i steps. According to the following result, which is the main theorem of this section, this lower bound on the number of e/i tests involved is sharp up to a constant factor:
Theorem 22.1.
Let subdivision iterations with compression be applied to a suspect square containing at most $m$ roots. Then they involve $O(m)$ e/i tests overall.
22.2 The first lemma
We begin our proof of Thm. 22.1 with
Lemma 22.1.
In every component $C$ made up of $\sigma(C)$ suspect squares we can choose $n(C)\geq\sigma(C)/5$ non-adjacent suspect squares.
Proof.
[…]
22.3 The second lemma
Lemma 22.2.
Fix two integers $i\geq 0$ and $g\geq 3$, fix a value $\sigma$ of softness, $1<\sigma<\sqrt{2}$, and define the separation fraction $a:=2-\sigma\sqrt{2}$ of Observation 20.1. For all $j$ let $j$-th subdivision step process $v_{j}$ components made up of $w_{j}$ suspect squares, $n_{j}$ of which are non-adjacent, let $\Delta_{j}$ denote their side length, and let the $j$-th steps perform no compression for $j=i,i+1,\dots,i+g$. Then
$w_{i+g}\geq(n_{i}-v_{i})a2^{g-0.5}~{}{\rm if}~{}\Delta_{i}\geq 2^{g}\Delta_{i+g}.$ (22.1)
Proof.
At the $i$th step a pair of non-adjacent suspect squares of the same component is separated by at least the gap $a\Delta_{i}$, and there are at least $n_{i}-v_{i}$ such gaps with overall length $(n_{i}-v_{i})a\Delta_{i}$. Neither of these gaps disappears at the $(i+g)$th step, and so they must be covered by $w_{i+g}$ suspect squares, each of diameter $\sqrt{2}~{}\Delta_{i+g}=\Delta_{i}/2^{g-0.5}$ under (22.1). Hence $(n_{i}-v_{i})a\Delta_{i}\leq w_{i+g}\Delta_{i}/2^{g-0.5}$. ∎
Next we extend Lemma 22.2 by allowing compression steps. We only need to extend the equation $2\Delta_{j+1}=\Delta_{j}$ to the case where the $j$th subdivision step for $i\leq j\leq i+g$ performs non-final compression, each followed with subdivision steps applied to a single suspect square with side length at most $4\Delta_{j}$.
The next two subdivision steps perform only five e/i tests, making up at most $105m-$ 10 e/i tests performed at all pairs of subdivision steps following all (at most $m-2$) non-final compression steps.
It remains to bound by $O(m)$ the number of e/i tests in the remaining subdivision steps, not interrupted with compression steps.
Furthermore, after a compression step we resume subdivision process with a suspect square whose side length bound decreases in two next subdivision steps from at most $4\Delta_{j}$ to at most $\Delta_{j}$.
Hence we can apply to it Lemma 22.2, because now it accesses the same number $n_{j}$ of non-adjacent suspect squares and an increased number $\sigma_{j}$ of all suspect squares (cf. (22.1)).
22.4 Counting e/i tests at partition steps
Let $h$ denote the height of our partition tree $\mathbb{T}$ and count suspect squares processed in subdivision steps that partition components, temporarily ignoring the other steps.
Clearly, bounds (22.1) still hold and can only be strengthened. Notice that $v_{i}\geq 1$ for all $i$, sum $v_{i+g}$ as well as bounds (22.1) for a fixed $g>0$ and $i=0,1,\dots,h-g$, and obtain
$\sum_{i=0}^{h-g}n_{i}-\lambda\sum_{i=0}^{h-g}\sigma_{i+g}\leq\sum_{i=0}^{h-g}v_{i+g}~{}{\rm for}~{}\lambda:=\frac{\sqrt{2}}{a2^{g}}.$ (22.2)
Notice that
$\sum_{i=0}^{h-g}v_{i+g}\leq\sum_{i=0}^{h}v_{i}-\sum_{i=0}^{g-1}v_{i}<2m-g,$
Combine the above estimate and the bounds $n_{i}\leq m$ and (cf. Lemma 22.1) $w_{i}\leq 5n_{i},$ for all $i$, write $n^{\prime}:=\sum_{i=0}^{h}n_{i}$, obtain
$\sum_{i=0}^{h-g}n_{i}\leq n^{\prime}-\sum_{i=h-g+1}^{h}n_{i}\leq n^{\prime}-mg,~{}\sum_{i=0}^{h-g}w_{i+g}\leq\sum_{i=0}^{h}w_{i}\leq 5n^{\prime},$
and so
[…]
Obtain that
$0.5n^{\prime}\leq 2m-1-g+mg~{}{\rm for}~{}g=1+\Big{\lceil}\log_{2}\Big{(}\frac{5\sqrt{2}}{a}\Big{)}\Big{\rceil}.$
Hence
$n^{\prime}\leq 4m-2-2g+2mg,~{}{\rm and~{}so}~{}n^{\prime}=O(m)~{}{\rm for}~{}g=O(1).$ (22.3)
22.5 Extension to counting all e/i steps
Next consider a non-compact component, count all suspect squares processed in all subdivision steps applied until it is partitioned, and then prove in Lemma 22.3 that up to a constant factor the overall number of these suspect squares is bounded by their overall number at all partition steps. Such a result extends to all subdivision steps the bound $n^{\prime}=O(m)$ of (22.3), so far restricted to partition steps. Together with Lemma 22.1 this will complete our proof of Thm. 22.1.
Lemma 22.3.
Let $k+1$ successive subdivision steps with $\sigma$-soft e/i tests be applied to a non-compact component for $\sigma$ of our choice in the range $1<\sigma<\sqrt{2}$. Let the $j$-th step output $n_{j}$ non-adjacent suspect squares for $j=0,1,\dots,k$. Then $\sum_{j=0}^{k-1}n_{j}\leq n_{k}\frac{10\sqrt{2}}{a}$ for the separation fraction $a:=2-\sigma\sqrt{2}$ (cf. Observation 20.1.)
Proof.
Substitute $g=k-i$ into (22.3) and obtain $w_{k}\geq(n_{i}-v_{i})a2^{k-i-0.5}$, for $i=0,1,\dots,k-1$.
Recall that $w_{j}\leq 5n_{j}$ (cf. Lemma 22.1) and that in our case $v_{i}=1$ and $n_{i}\geq 2$ and obtain that $n_{i}-v_{i}\geq n_{i}/2$ for all $i$.
Hence $5n_{k}\geq n_{i}a2^{k-i-1.5}$ for all $i$, and so
$\sum_{i=0}^{k-1}n_{i}\leq n_{k}\frac{10\sqrt{2}}{a}\sum_{j=1}^{k}2^{-j}<n_{k}\frac{10\sqrt{2}}{a}.$
Remark 22.1.
The overhead constants hidden in the "O" notation of our estimates are fairly large but overly pessimistic. E.g., (i) the roots $x$ and $y$ involved in the proof of Lemma 22.2 are linked by a path of suspect squares, but we deduce the estimates of the lemma by using only a part of this path, namely, the part that covers the gaps between non-adjacent suspect squares. (ii) Our proof of that lemma covers the worst case where a long chains between the roots $x$ and $y$ can consist of up to $4m$ suspect squares, but for typical inputs such chains are much shorter and can consist of just a few suspect squares. (iii) In the proof of Lemma 22.3 we assume that $n_{i}-1\geq n_{i}/2$; this holds for any $n_{i}\geq 2$, but the bounds are strengthened as $n_{i}$ increases. (iv) In a component made up of five suspect squares at least two are non-adjacent, and so the bound of Lemma 22.1 can be strengthened at least in this case.
Remark 22.2.
Alg. 21.1 compresses an input disc $D(0,1)$ into $\eta$-rigid disc for $\eta=5/8$, but our proof of Thm. 22.1 only involves $\eta$ in the argument in Sec. 21.4, and this argument and therefore the whole proof of the theorem can be immediately extended if it relies on compression into any rigid disc.
23 A point near root set and root-count in a disc
We prove that the algorithms of this section work where an input disc is $d^{\nu}$-isolated for a fixed positive constant $\nu$. In application to compression we begin with an isolated disc, say, 3.5-isolated, but can turn it into $d^{\nu}$-isolated for any fixed positive constant $\nu$ by applying $h$ additional subdivision steps or $h$-lifting for $h=O(\log(d))$. This comment can be extended to application of e/i test of Sec. 23.4 instead of Cauchy-based e/i tests of Part II.
23.1 Computation of a point near root set: introductory comments
Recall Schröder’s iterations [132],
$z=y-m~{}{\rm NR}(y),$ (23.1)
which fast converge to a root of multiplicity $m$ [80, Sec. 5.4]. Due to map (1.3) we only need to consider the case where the input disc $D$ is the unit disc $D(0,1)$. We first assume that it contains all $d$ roots and readily verify
Observation 23.1.
Schröder’s single iteration (23.1) for $m=d$ and $p(x)=p_{d}(x-z_{1})^{d}$ outputs the root $z_{1}$ unless $y=z_{1}$.
Next we extend this observation by proving that the output $z$ of Schröder’s iteration (23.1) lies near the convex hull $CH(X(D))$ provided that $m=d$, a disc $D$ contains all $d$ roots, and point $y$ lies far enough from the disc $D$. Here are our specific estimates.
23.2 The case where a disc contains all $d$ roots
Theorem 23.1.
Write $D:=D(0,1)$, $CH:=CH(D)$, and $\Delta:=\Delta(D)$, fix $y$ such that $|y|>2$, let $z:=y-d~{}{\rm NR(y)}$, which is (23.1) for $m=d$. Let $\bar{x}\in CH$. Then $|z-\bar{x}|\leq\alpha~{}\Delta$ for $\alpha:=\frac{(|y|-1)^{2}}{(y|-2)^{2}}=(1+\frac{1}{(y|-2})^{2}$, and so $\alpha=2.25$ for $u|=4$ and $\alpha=25/16$ for $u|=6$, $\alpha\mapsto 1$ $|y|\mapsto\infty$.
Proof.
Recall that $|\bar{x},z_{1},\dots,z_{d}\in D(0,1)$, $|z_{j}|\leq 1$ and $|\bar{x}-z_{j}|\leq\Delta$ for $j=1,\dots,d$. Hence (cf. (3.1))
$\Big{|}{\rm NIR}(y)-\frac{d}{y-\bar{x}}\Big{|}\leq\sum_{j=1}^{d}\Big{|}\frac{1}{y-z_{j}}-\frac{1}{y-\bar{x}}\Big{|}\leq\sum_{j=1}^{d}\frac{|\bar{x}-z_{j}|}{|y-\bar{x}|\cdot|y-z_{j}|}\leq\frac{\Delta d}{(|y|-1)^{2}}.$
Therefore,
$|u-w|\leq\beta~{}{\rm for}~{}u:=\frac{{\rm NIR}(y)}{d},~{}w:=\frac{1}{y-\bar{x}},~{}{\rm and}~{}\beta:=\frac{\Delta}{(|y|-1)^{2}},$ (23.2)
and so
$|z-\bar{x}|=\Big{|}\bar{x}-y+\frac{d}{{\rm NIR}(y)}\Big{|}=\Big{|}\frac{1}{u}-\frac{1}{w}\Big{|}=\frac{|u-w|}{|uw|}\leq\frac{\beta d}{|{\rm NIR}(y)|}|y-\bar{x}|.$ (23.3)
Deduce from (3.1) and the bounds $\max_{j=1}^{d}|z_{j}|\leq 1<|y|$ that
$|{\rm NIR}(y)|=\Big{|}\sum_{j=1}^{d}\frac{1}{y-z_{j}}\Big{|}\geq d\max_{j=1}^{d}\frac{1}{|y-z_{j}|}\geq\frac{d}{|y|-1}.$
Substitute this bound into Eqn. (23.3) and obtain that
$|z-\bar{x}|\leq\beta~{}||y|-1|\cdot|y-\bar{x}|$
Substitute $\beta:=\frac{\Delta}{(|y|-\Delta)^{2}}$ of (23.2) and obtain
$|\bar{x}-z|\leq\alpha^{\prime}\Delta~{}{\rm for}~{}\alpha^{\prime}:=\frac{(|y|-1)|y-\bar{x}|}{(|y|-\Delta)^{2}}.$
Substitute the bounds $\Delta\leq 2$ and $|\bar{x}|\leq 1$ and obtain the theorem. ∎
Corollary 23.1.
Under the assumptions of Thm. 23.2 the disc $D(z,r_{1}(z))$ is $\frac{1}{2+2\alpha}$-rigid.
Proof.
Verify the claim in the case where $x_{d}=0$, $x_{1}=\Delta=1$, e.g., $0\leq x_{j}\leq 1$ for all $j$, and $z=1+\alpha$ and notice that this is a worst case input. ∎
Now we devise our fast compression algorithm inthe case where $m=d$.
Algorithm 23.1.
$\epsilon$-compression 1.
INPUT: a positive tolerance $\epsilon=1/2^{b}\leq 1$ and a black box polynomial $p(x)$ of a degree $d$ such that
#$\#(D(0,1))=d.$ (23.4)
OUTPUT: a complex $z$, and a positive $r$ such that #$\#(D(z,r))=d$ and either $r\leq\epsilon$ or otherwise the disc $D(z,r)$ is $\frac{2}{13}$-rigid.
INITIALIZATION: Fix $y\geq 4$.
COMPUTATIONS:
(i) Compute and output $z=y-d~{}{\rm NR}~{}(y)$.
(ii) Compute an approximation $\tilde{r}\geq 0$ to $r_{1}(z)$ by applying Alg. 9.1 for $m=d$, a constant $\beta>1$, and $\epsilon$ replaced by $\epsilon/4$.
If $\tilde{r}\leq 3\epsilon/4$, output $z$ and $r=\epsilon$, defining $\epsilon$-compression into the disc $D(z,r)$ of radius $\epsilon$.
Otherwise output $r=\tilde{r}+0.25\epsilon$, defining compression into rigid disc $D(z,r)$.
Correctness of the algorithm is readily verified if Alg. 9.1 outputs a range $[\rho_{-},\rho_{+}]$ of length at most $\epsilon/4$, for in this case $r_{1}\leq r=\tilde{r}+\epsilon/4\leq\epsilon$, and #$\#(D(z,\epsilon))=d$.
Otherwise $1/\beta\leq\tilde{r}/r_{1}\leq\beta$ for a constant $\beta\geq 1$ of Alg. 9.1, while the disc $D(x,r_{1})$ is $\frac{1}{2\alpha+2}$-rigid by virtue of Cor. 23.1 where $\alpha\leq 2.25$ by virtue of THm. 23.1 because $y\geq 4$ in our case. It follows that the disc $\bar{D}:=D(z,r_{1}(z))$ is $\frac{2}{13}$-rigid.
Remark 23.1.
By the same argument we can ensure that unless $\tilde{r}\leq 3\epsilon/4$ the disc $D(z,r_{1}(z))$ is $\eta$-isolated for any $\eta>4$ if we initialize Alg. 23.1 at a sufficiently large $y$.
This compression queries evaluation oracle twice while performing Schröder’s single iteration for computing $z$ and otherwise runs at the cost of performing Alg. 9.1, estimated in Cor. 9.1. We can decrease the latter cost bound for a large class of polynomials $p$ by choosing larger tolerance $\epsilon$. Then we will yield desired compression if the algorithm outputs $r$ exceeding $\epsilon/4$. Otherwise we can recursively repeat the computations for tolerance recursively decreased to $\epsilon/4$.
23.3 Extension to any well-isolated disc
Now suppose that #$\#(D(0,1))=m<d$ and $i(D(0,1))\geq\theta$ for a sufficiently large $\theta$ of order $d^{O(1)}$. Then verify that the $d-m$ external roots, lying outside the disc $D(0,1)$, little affect the proof of Thm. 23.1 and our whole study in the previous subsection.
Thus we first perform $O(\log(d))$ subdivision steps (see Observation 20.2). We do not need compression if these steps break the component into more than one component or compress it into an $\epsilon$-disc. Otherwise the component remains compact and hence in $O(\log(d))$ subdivision steps shrinks enough to become $d^{\nu}$-isolated for any fixed constant $\nu$, in which case we apply our compression algorithm.
23.4 A fast root-counter and a fast e/i test based on Schröder’s iterations
Let the unit disc be sufficiently strongly isolated and let a single Schröder’s iteration (23.1) be applied at a point $y$ lying sufficiently far from the unit disc $D(0,1)$.
Then our study in this subsection implies that the iteration outputs a point $z=z(m)$ lying in or near the disc $D(0,1)$, say, in the disc $D(0,2)$, if and only if #$\#(D(0,1))=m$, that is, we count roots in $D(0,1)$ at the cost of evaluation of $p(x)$ at two points.
Indeed, if #$\#(D(0,1))=m$, then, as we have showed in this section, the iteration outputs a point $z=z(m)$ lying near the convex hull of the root set in $D(0,1)$ and hence lying in or near the disc $D(0,1)$, say, in the disc $D(0,2)$ if the value $|y|$ is large enough.
In this case $|m~{}{\rm NR}(y)-y|\leq 2$, and so $|{\rm NR}(y)-y|/m\leq 2/m$.
If #$\#(D(0,1))=\bar{m}\neq m$, then the iteration outputs a point $z=z(m)$ that lies at the distance at least $|\bar{m}-m|\cdot|{\rm NR}(y)|-2$ from the disc $D(0,2)$, which must be large if $|y|$ is sufficiently (although reasonably) large.
Extension to a fast e/i test. Now suppose that the disc $D(0,1)$ is sufficiently well isolated and apply this root-counter to the polynomial $xp(x)$ for a large $y$ and $m=1$, in which case Schröder’s iterations turn into Newton’s. In this case root-counter will show that the disc $D(0,1)$ contains no roots of $p(x)$ if and only if its output lies in that disc.
23.5 Overall complexity of our root-finders
We have reduced our root-finding in an isolated disc to performing at most $12m-$ 1 compression steps and in addition $O(m)$ e/i tests. We have also reduced a compression step to (i) performing $O(\log(d))$ additional subdivision steps (included to strengthen isolation of input discs for compression), (ii) application of Schröder’s single iteration, which evaluates polynomial $p(x)$ at two points, and (iii) performing $O(\log(b))$ times $m$-tests for approximation of the root radius $r_{d-m+1}(z,p)$.
Overall our root-finders apply e/i tests $O(m\log(d))$ times, $m$-tests $O(m\log(b))$ times, and at most $22m-$ 2 Schröder’s iterations. Clearly, the computational cost of Schröder’s is dominated, while in Sec. 15 we reduced an $m$-test is to $O(1)$ e/i tests, specified in Sec. 14. Based on the latter results we can strengthen our estimates of Sec. 18 and arrive at the complexity bounds of our Main Theorems. Alternatively, instead of estimating the cost of e/i tests and reduction to them of $m$- tests, we can use our estimates for the complexity of $\ell$- tests of Part II for $\ell=1$ and $\ell=m$ and would still arrive at our Main Theorems 2 – 4.
Part V: Factorization, Root Isolation, and Functional Iterations
24 Polynomial Factorization, its links to Root-finding, and some bit-complexity estimates
24.1 Polynomial Factorization and its links to Root-finding (the results)
Hereafter write $|u|=\sum_{i=0}^{d}|u_{i}|$ for a polynomial $u(x)=\sum_{i=0}^{d}u_{i}x^{i}$.
The root-finders of [134, 96, 103] solve Problems 0 and 0 by extending the the solution of the following problem of independent importance:
Problem F. Approximate Polynomial Factorization: [^35] The solution of Problem F can be also extended to isolation of the zeros of a square-free polynomial with integer coefficients and approximation of matrix eigenvalues (see our Remark 25.4 and Sec. 11) and has various other applications to modern computations, e.g., to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [150, 18, 8, 35, 36, 147]. For a polynomial of degree $d$ the known sharp upper and lower bounds on the input precision and number of bit operations for root-finding exceed those for numerical factorization by a factor of $d$. Given a real $b$ and $d+1$ coefficients $p_{0},p_{1},\dots,p_{d}$ of (1.1), compute $2d$ complex numbers $u_{j},v_{j}$ for $j=1,\dots,d$ such that
$|p-\prod_{j=1}^{d}(u_{j}x-v_{j})|\leq\epsilon|p|~{}{\rm for}~{}\epsilon=1/2^{b}.$ (24.1)
Hereafter $\mathbb{B}_{0}(b,d)$ and $\mathbb{B}_{F}(b,d)$ denote the optimal number of bit operations required for the solution of Problems 0 and F, respectively, for an error bound $\epsilon=1/2^{b}$, and next we specify a class of polynomials whose zeros make up no large clusters (cf. [14, 30]).
Definition 24.1.
Polynomials with bounded root cluster sizes. The zero set of $p$ consists of clusters, each made up of a constant number of zeros and having a diameter at most $1/2^{3-b/d}$; furthermore, these clusters are pairwise separated by the distances at least $1/d^{\tilde{O}(1)}$.
[96, 97, 99, 100, 103] proved the following near optimal bound for Problem F and readily extended it to near optimal solution of Problem 0.
Theorem 24.1.
$\mathbb{B}_{F}(b,d)=\tilde{O}((b+d)d)$.
In Sec. 24.3 we reduce Problem 0 to Problem F and vice versa by applying Schönhage’s results of [134, Cor. 4.3 and Sec. 19]:
Theorem 24.2.
(i) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}(bd,d))$ for any polynomial $p(x)$ of degree $d$, (ii) $\mathbb{B}_{F}(b,d)=\tilde{O}(\mathbb{B}_{0}(b+d+\log(d),d))$ for any polynomial $p(x)$ of degree $d$, and (iii) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}2(b,d))$ for polynomials of Def. 24.1, with bounded cluster sizes.
By combining Thms. 24.1 and 24.2 obtain the following estimates.
Theorem 24.3.
One can solve Problem 0 by using (i) $\tilde{O}((b+d)d^{2})$ bit operations for any polynomial $p$ of (1.1) and (ii) $\tilde{O}((b+d)d)$ bit operations for a polynomial $p$ with bounded cluster sizes.
Corollary 24.1.
One can solve Problem 0 by using $\tilde{O}((b+d)d+(b+m)m^{2})$ bit operation for $m$ of Def. 1.1 of Sec. 1.5. The bound turns into $\tilde{O}((b+d)d$ if $m^{2}=\tilde{O}(d)$.
Proof.
Proceed in two steps:
(i) by using $\tilde{O}((b+d)d)$ bit operations solve Problem F and then
(ii) apply the algorithm that supports Thm. 24.3 to the $m$th degree factor of the polynomial $\prod_{j=1}^{d}(u_{j}x-v_{j})$ of (24.1) that shares with this polynomial all its zeros lying in the disc $D(1+\beta)$ for positive $\beta$ of Def. 1.1.363636We can slightly speed up stage (i) because at that stage we actually only need to approximate factorization of the factor $f$ of degree $m$ of $p$ that shares with $p$ its zeros in $D$.
24.2 Lower bounds on the bit operation complexity
Readily verify the following information lower bounds.
Observation 24.1.
To solve Problem 0 or F one must access $(d+1)b$ bits of input coefficients and perform at least $0.5(d+1)b$ bit operations – at most one operation for two input bits.
Next strengthen the lower bound for Problem 0.
Observation 24.2.
Let $p(x)=(x-v_{1})^{j}f(x)$ for an integer $j>0$ and a monic polynomial $f(x)$ of degree $d-j$. Fix a real $b$. Then the polynomial $p_{j}(x)=p(x)-2^{-jb}f(x)$ has the $j$ zeros $v_{1}+\zeta^{g}2^{-b}$ for $g=0,1,\dots,j-1$.
Observation 24.2 implies that one must access the coefficient $p_{d-j}$ within $2^{-jb}$ to approximate within $2^{-b}$ a $j$-multiple zero of $p$ or its zero in a cluster of $j$ zeros, and so a root-finder must process at least $jb$ bits of the coefficient $p_{d-j}$. Sum these bounds for $j=1,\dots,d$ and obtain
Corollary 24.2.
One must access at least $0.5(d+1)db$ bits of the coefficients of a monic polynomial $p$ and perform at least $0.25(d+1)db$ bit operations to approximate even its single within $2^{-b}$.
The lower bounds of this subsection imply that the algorithm of [96, 103] and its extension supporting Thm. 24.1 solve Problems 0 and F of large size by using optimal number of bit operations up to polylogarithmic factors.
24.3 Linking Factorization and Root-finding (with more details)
Next we prove Thm. 24.2, by applying the results of [134, Cor. 4.3 and Sec. 19].
Recall that $\mathbb{B}_{0}(b,d)$ and $\mathbb{B}_{F}(b,d)$ denote the optimal number of bit operations for the solution of Problems 0 and F, respectively, for an error bound $\epsilon=1/2^{b}$, and furthermore, recall the lower bounds
$\mathbb{B}_{0}(b,d)\geq 0.5(d+1)b~{}{\rm and}~{}\mathbb{B}_{F}(b,d)\geq 0.25(d+1)db$ (24.2)
of [96, 103] and the following result (cf. [134, Thm. 19.1]).
Theorem 24.4.
Let $r_{1}=\max_{j=1}^{d}|z_{j}|\leq 1$. Then a solution of Problem 0 within an error bound $1/2^{b}$ can be recovered at a dominated bit operation cost from the solution of Problem F within an error bound $1/2^{bd+2}$, and so $\mathbb{B}_{0}(b,d)\leq\mathbb{B}_{F}(bd+2,d)$.
Based on [134, Cor. 4.3] we can also readily relax the assumption that $p(x)$ is monic.
Furthermore, in view of (24.2), relaxing the bound $r_{1}\leq 1$ would little affect the latter estimate because we can apply map (1.3), which changes the error bound of Problem 0 by at most $\log_{2}(\max\{\rho,1/\rho\})$, for $\rho\geq 1/2^{b}$.
[136, Thm. 2.7], which we reproduce in Sec. 25.4 as Thm. 25.3, shows an alternative way to relaxing both assumptions – that $r_{1}\leq 1$ and $p(x)$ is monic.
Schönhage in [134, Sec. 19] has greatly strengthened the estimate of Thm. 24.4 unless “the zeros of $p$ are clustered too much". Namely, he proved (see [134, Eqn. (19.8)]) the following result.
Lemma 24.1.
Assume that the polynomial $p(x)$ is monic, the values $u_{j}=1$ for all $j$ and complex $v_{1},\dots,v_{d}$ satisfy bound (24.1), and $|z_{j}-v_{j}|\leq r:=2^{2-b/d}$ for all $j$. Furthermore, write $\Delta_{j,1}:=|v_{j}-v_{1}|$ for $j=2,3,\dots,d$ and suppose that $1.25\Delta_{j,1}\leq r$ for $j=2,3,\dots,d$ and $\Delta_{1}:=\min_{j}\Delta_{j,1}>r$. Then $|z_{1}-v_{1}|\leq 2^{2-b/m}/|\Delta_{1}-r|^{\frac{d-m}{m}}$.
If $1/|\Delta-r|^{\frac{d-m}{m}}=2^{\tilde{O}(d)}$, then Lemma 24.1 implies that $|z_{1}-v_{1}|\leq 1/2^{b/m-\tilde{O}(d)}$, and we can similarly estimate $|v_{j}-z_{j}|$ for all $j=1,2,\dots,d$.
By extending these estimates we strengthen Thm. 24.4 as follows.
Corollary 24.3.
Under bounding cluster sizes of Def. 24.1, it holds that $|z_{j}-v_{j}|\leq 1/2^{b/m-\tilde{O}(d)},~{}j=1,2,\dots,d$. Hence if $m=O(1)$ is a constant, then a solution of Problem 0 within an error bound $1/2^{b}$ can be recovered at a dominated bit operation cost from the solution of Problem F with a precision $1/2^{\bar{b}}$ for a sufficiently large $\bar{b}=\tilde{O}(b+d)$, and so $\mathbb{B}_{1}(b,d)\leq\mathbb{B}_{2}(\tilde{O}(b+d),d)$.
For converse reduction of Problem F to Problem 0 recall [134, Cor. 4.3] in the case of $k=2$, covering the product of two polynomials:
Theorem 24.5.
Let $p=gh$ for three polynomials $p$ of (1.1), $g$, and $h$. Then $||g||_{1}||h||_{1}\leq 2^{d-1}||p||_{1}$.
Now consider the products $p=gh$ for $g=x-z_{j}$, for $j=1,\dots,d$. For a fixed $j$, an error of at most $\epsilon/2^{d}$ for the root $z_{j}$ contributes at most $\epsilon||h||_{1}$ to overall factorization error of Problem F. Therefore, $||h||_{1}\leq 2^{d-1}||p||_{1}/||x-z_{j}||_{1}\leq 2^{d-1}||p||_{1}$ by virtue of Thm. 24.5. Hence the approximation errors of all the $d$ roots together contribute at most $2^{d-1}d\epsilon$ to the relative error of the solution of Problem F (ignoring dominated impact of higher order terms), and we arrive at
Corollary 24.4.
If $b=O(b-d-\log(d))$, then Problem F for an error bound $1/2^{b}$ can be reduced at a dominated bit operation cost to Problem 0 for an error bound $1/2^{\widehat{b}}$ and a sufficiently large $\widehat{b}=O(b+d+\log(d))$, and so $\mathbb{B}_{F}(b,d)=O(\mathbb{B}_{0}(b,d))$.
It follows that the complexity estimate of Thm. 1.3 can be also applied to the solution of Problem F of factorization provided that $b=O(b-d-\log(d))$. Then again, the bit operation complexity of this solution of Problem F is optimal (up to polylogarithmic factors) under the random root model provided that $d=\tilde{O}(b)$ (cf. Observation 24.1).
24.4 Two root-finders of 2016 and 2018
Becker et al in [23, 24] claimed that their extension of subdivision root-finders of [128, 101] (i) runs within the same near optimal bit operation bounds of [96, 103] but (ii) unlike [96, 103] is easy to implement. This seemed to diminish incentive for further research in this area, but were the claims valid?
Implementation of the root-finder of [23] by Imbach et al in [64] provided no empirical support to optimality claimed by Becker et al because the root-finder of [63], using some novel techniques from [112] (see [63, Sec. 1.2]), runs much faster. Unfortunately, formal support for optimality claimed by Becker et al has been missing as well.
Indeed, for all proofs the paper [23] refers to a preprint of [24], with a distinct list of the authors and with the main algorithms and the Main Theorem significantly different from those of [23].
Furthermore, as we pointed out in the introduction, the the bit-complexity specified in the Main Theorem of [24] has order of $mbd^{2}$, that is, exceeds the near optimal bound of [96, 103] by a factor of $d$.
Can one salvage the claim by Becker et al in another way? Next we argue that this can be problematic because the algorithms of [23, 24] for Problem 1m are some variants of subdivision iterations. The basic tool and the main algorithmic novelty of the variant of subdivision in [96, 103] are e/i tests based on the following theorem by Pellet [88], which extended the previous results by Cauchy and Rouché:
Theorem 24.6.
Given a polynomial $(x)$ of (1.1) with complex coefficients and $p_{g}=1$. Let the polynomial $s(x)=x^{d}+p_{d-1}x^{d-1}+\cdots+|p_{i+1}|x^{i+1}-|p_{i}|x^{i}|+p_{i-1}|x^{i-1}+\cdots+|p_{0}|$ have two distinct positive roots $r$ and $R$, $r<R$. Then the polynomial $p$ has exactly $i$ zeros in or on the circle $C(0,r)=\{z:~{}|z|=r\}$ and no zeros in the annulus $A(0,r,R)=\{z:~{}r<|z|<R\}$.
[23, 24] apply Pellet’s theorem on each subdivision step to more than $4m$ polynomials $t(x)=\sum_{i=0}^{d}t_{i}x^{i}$ obtained from the polynomial $p$ by means of shifting the variable $x$, its scaling by up to $2^{b}$, and root-squaring.
To apply Pellet’s theorem to $t(x)$ one must compute the values $|\frac{t_{i}}{t_{d}}|$ for $i=0,1,\dots,d-1$. Hence, by virtue of Observation 24.2, one must process at least $0.5(d+1)db$ bits of the coefficients of $t(x)$ to support approximation of its zeros within $1/2^{b}$.
The algorithms of [24] cannot avoid this because they apply e/i tests to the polynomials obtained in $\log(\log(d))$ steps of root-squaring (cf. Sec. 13.3) from polynomials $t_{c,\rho}(x)=p(\frac{x-c}{\rho})$ where $\rho>0$ can be as small as $1/2^{b}$.
Already a single subdivision iteration can perform more than $4m$ soft e/i tests; this means that one must process more than $2(d+1)dmb$ bits of memory and $(d+1)dmb$ bit-operations to solve Problem 1m by applying such a Pellet-based root-finder.
25 Recursive Factorization; extensions to Root-finding and Root Isolation
In this section we briefly recall some algorithms and complexity estimates from [134, 136]. We first recursively decompose a polynomial into the product of linear factors and then extend the factorization to approximation and isolation of the roots.
25.1 Auxiliary norm bounds
Recall two simple lemmas.
Lemma 25.1.
It holds that $|u+v|\leq|u|+|v|$ and $|uv|\leq|u|~{}|v|$ for any pair of polynomials $u=u(x)$ and $v=v(x)$.
Lemma 25.2.
Let $u=\sum_{i=0}^{d}u_{i}x^{i}$ and $|u|_{2}=(\sum_{i=0}^{d}|u_{i}|^{2})^{\frac{1}{2}}$. Then $|u|_{2}\leq|u|$.
The following lemma relates the norms of a polynomial and its factors.
Lemma 25.3.
If $p=p(x)=\prod_{j=1}^{k}f_{j}$ for polynomials $f_{1},\dots,f_{k}$ and $\deg p\leq d$, then $\prod_{j=1}^{k}|f_{j}|\leq 2^{d}|p|_{2}\leq 2^{d}|p|.$
Proof.
The leftmost bound was proved by Mignotte in [78]. The rightmost bound follows from Lemma 25.2. ∎
Remark 25.1.
[136, Lemma 2.6] states with no proof a little stronger bound $\prod_{j=1}^{k}|f_{j}|\leq 2^{d-1}|p|$ under the assumptions of Lemma 25.3. From various factors of the polynomial $p(x)=x^{d}-1$ such as $\prod_{j=1}^{\frac{d}{2}}(x-\zeta_{d}^{j})$ for even $d$ and $\zeta:=\zeta_{q}:=\exp(\frac{2\pi{\bf i}}{q})$ denoting a primitive $q$-th root of unity, one can see some limitations on strengthening this bound further.
25.2 Errors and complexity
Suppose that we split a polynomial $p$ into a pair of factors over an isolated circle and recursively apply this algorithm to the factors until they become linear of the form $ux+v$; some or all of them can be repeated. Finally, we arrive at complete approximate factorization
$p\approx p^{\ast}=p^{\ast}(x)=\prod_{j=1}^{d}(u_{j}x+v_{j}).$ (25.1)
Next, by following [134, Sec. 5], we estimate the norm of the residual polynomial
$\Delta^{\ast}=p^{\ast}-p.$ (25.2)
We begin with an auxiliary result.
Theorem 25.1.
Let
$\Delta_{k}=|p-f_{1}\cdots f_{k}|\leq\frac{k\epsilon|p|}{d},$ (25.3)
$\Delta=|f_{1}-fg|\leq\epsilon_{k}|f_{1}|,$ (25.4)
for some non-constant polynomials $f_{1},\ldots,f_{k},f$ and $g$ such that
$\epsilon_{k}d2^{d}\prod_{j=1}^{k}|f_{j}|\leq\epsilon.$ (25.5)
Then
$|\Delta_{k+1}|=|p-fgf_{2}\cdots f_{k}|\leq\frac{(k+1)\epsilon|p|}{d}.$ (25.6)
Proof.
$\Delta_{k+1}=|p-f_{1}\cdots f_{k}+(f_{1}-fg)f_{2}\cdots f_{k}|.$ Apply Lemma 25.1 and deduce that $\Delta_{k+1}\leq\Delta_{k}+|(f_{1}-fg)f_{2}\cdots f_{k}|$ and, furthermore, that
$|(f_{1}-fg)f_{2}\cdots f_{k}|\leq|f_{1}-fg|~{}|f_{2}\cdots f_{k}|=\Delta|f_{2}\cdots f_{k}|.$
Combine the latter inequalities and obtain $\Delta_{k+1}\leq\Delta_{k}+\Delta|f_{2}\cdots f_{k}|$. Combine this bound with (25.3)-(25.5) and Lemmas 25.1 and 25.3 and obtain (25.6). ∎
Write $f_{1}:=f$ and $f_{k+1}=g.$ Then (25.6) turns into (25.3) for $k$ replaced by $k+1.$ Now compute one of the factors $f_{j}$ as in (25.4), apply Thm. 25.1, then recursively continue splitting the polynomial $p$ into factors of smaller degrees, and finally arrive at factorization (25.1) with
$|\Delta^{\ast}|\leq\epsilon|p|$
for $\Delta^{*}$ of (25.2). Let us call this computation Recursive Splitting Process provided that it begins with $k=1$ and $f_{1}=p$ and ends with $k=d$.
Theorem 25.2.
In order to support (25.3) for $0<\epsilon\leq 1$ and $j=1,2,\ldots,d$ in the Recursive Splitting Process it is sufficient to choose $\epsilon_{k}$ in (25.4) satisfying
$\epsilon_{k}~{}d~{}2^{2d+1}\leq\epsilon~{}\textrm{for~{}all}~{}k.$ (25.7)
Proof.
Prove bound (25.3) by induction on $j$. Clearly, the bound holds for $k=1$. It remains to deduce (25.6) from (25.3) and (25.7) for any $k$. By first applying Lemma 25.3 and then bound (25.3), obtain
$\prod_{i=1}^{k}|f_{i}|\leq 2^{d}|\prod_{i=1}^{k}f_{i}|\leq 2^{d}~{}\Big{(}1+\frac{k\epsilon}{d}\Big{)}~{}|p|.$
The latter bound cannot exceed $2^{d+1}|{p}|$ for $k\leq d$, $\epsilon\leq 1$. Consequently (25.7) ensures (25.5), and then (25.6) follows by virtue of Thm. 25.1.
Remark 25.2.
The theorem implies that we can ensure the output precision bound $b^{\prime}$ by working with the precision of
$\bar{b}\geq\bar{b}_{\rm inf}=2d+1+\log_{2}d+b^{\prime}$
bits throughout Recursive Splitting Process.
25.3 Overall complexity of Recursive Factorization and extension to Root Isolation
The overall complexity of recursive factorization is dominated by the sum of the bounds on the complexity of all splittings into pairs of factors. The algorithm of [103] adopts recursive factorization of [134] but decreases the overall complexity bound by roughly a factor of $d$. This is due to additional techniques, which ensure that in every recursive splitting of a polynomial into the product of two factors the ratio of their degrees is at most 11. Consequently, $O(\log(d)$ recursive splitting steps are sufficient in [103], and so the overall complexity of root-finding is proportional to that of the first splitting. In contrast, the degrees of a factor in any splitting
in [134] can have low, e.g., 1, and so there can be allows up to $d-1$ splittings overall, with each of the first $0.5d$ of them almost as expensive as the very first one.
25.4 From factors to roots
Theorem 25.3.
[136, Thm. 2.7]. Suppose that
$p=p_{d}\prod_{j=1}^{d}(x-z_{j}),~{}\tilde{p}=\tilde{p}_{d}\prod_{j=1}^{d}(x-y_{j}),~{}|\tilde{p}-p|\leq\epsilon|p|,~{}\epsilon\leq 2^{-7d}$
and
$|z_{j}|\leq 1~{}{\rm for}~{}j=1,\dots,d^{\prime}\leq d,~{}|z_{j}|\geq 1~{}{\rm for}~{}j=d^{\prime}+1,\dots,d.$
Then up to reordering the roots it holds that
$|y_{j}-z_{j}|<9\epsilon^{\frac{1}{d}}~{}{\rm for}~{}j=1,\dots,d^{\prime};~{}\Big{|}\frac{1}{y_{j}}-\frac{1}{z_{j}}\Big{|}<9\epsilon^{\frac{1}{d}}~{}{\rm for}~{}j=d^{\prime}+1,\dots,d.$
By virtue of Thm. 25.3 for $b^{\prime}=O(bd)$ we can bound the bit operation complexity of the solution of Problem 0 by increasing the estimate for the complexity of factorization in Sec. 25.3 by a factor of $d$.
Corollary 25.1.
The bit operation complexity of the solution of Problem 0. Given a polynomial $p$ of degree $d$ and a real $b^{\prime}\geq 7d$, one can approximate all its $d$ zeros within the error bound $2^{-b^{\prime}}$ by using $O(\mu(b^{\prime}+d)d^{2}\log(d)(\log^{2}d+\log(b^{\prime})))=\tilde{O}((b^{\prime}+d)d^{2})$ bit operations.
Remark 25.3.
The study in this section covers the solution of Problem 1 for the worst case input polynomial $p$. If its roots are isolated from each other, then the upper estimates of Thm. 25.3 and Cor. 25.2 can be decreased by a factor of $d$ (cf. [103]).
25.5 Polynomial Root Isolation
Based on his results recalled in this section, Schönhage in [134, Sec. 20] estimates the bit operation complexity of the following well-known problem by reducing it to Problem F.
Problem ISOL. Polynomial Root Isolation. Given a polynomial $p$ of (1.1) that has integer coefficients and is square-free, that is, has only simple roots, compute $d$ disjoint discs on the complex plane, each containing exactly one root.
Remark 25.4.
Our Problem 0 can be viewed as the Problem of Polynomial Root Cluster Isolation. It turns into Problem ISOL if all the coefficients of $p$ are integers and the roots are simple, and so our algorithms for Problem 0 as well as the Algorithms of [96, 103] solve Problem ISOL within a record complexity bounds.
In [134, Sec. 20] Schönhage reduces to factorization Problem F the root-isolation problem of computing $d$ disjoint discs each covering a single simple root of a square-free polynomial with integer coefficients. Problem ISOL is readily solved if the zeros are approximated within $\epsilon=1/2^{b}<0.5\delta$ for $\delta$ denoting the minimal pairwise distance between the zeros of $p$ (see the known lower bounds on $\delta$ in Kurt Mahler and Maurice Mignotte [77, 78, 79]), and Schönhage has estimated that solution of Problem F (e.g., by means of the algorithms of [134, 94, 103]) implies isolation provided that
$b=O((\tau+\log(d))+d)~{}{\rm and}~{}\tau=\log_{2}\max_{i=0}^{d}|p_{i}|.$ (25.8)
This implied the following result.
Corollary 25.2.
The bit operation complexity of polynomial root isolation. Suppose that a polynomial $p$ of (1.1) has integer coefficients and only simple roots. Let $\sigma_{p}$ denotes its root separation, that is, the minimal distance between a pair of its roots. Write $\epsilon:=0.4\sigma_{p}$ and $b^{\prime}:=\log_{2}(\frac{1}{\epsilon})$. Let $\epsilon<2^{-d}$ and let $m=m_{p,\epsilon}$ denote the maximal number of the roots of the polynomial $p(x)$ in $\epsilon$-clusters of its roots. Then one can solve Problem ISOL by using $\tilde{O}(bdm)$ bit operations for $b=\frac{b^{\prime}}{m}$.
26 Newton’s, Weierstrass’s, and Ehrlich’s iterations
26.1 Contents
In the next subsection we recall Newton’s, Weierstrass’s, and Ehrlich’s iterations, comment on their convergence, and acceleration by means of applying fast multipoint evaluation, which is dramatic but involves the coefficients of $p$. In Sec. 26.3 we cover mre limited acceleration of these iterations by means of incorporating FMM, which can be applied to a black box polynomial. In Secs. 26.5 and 27) we cover initialization of Newton’s, Weierstrass’s, and Ehrlich’s iterations, required for their fast global empirical convergence. For Newton’s iterations we only need to compute a disc that contains all zeros of a polynomial $p$ (cf. [71, 58, 143, 130, 141]). This is given by the well-known estimates via the coefficients, but we present simple but novel recipes in the case of black box polynomial $p$. The customary initialization is more demanding for Weierstrass’s and Ehrlich’s iterations but is still fast for polynomials represented with their coefficients, as we recall in Sec. 27), where we also point out how this initialization can be extended to fast real and nearly real root-finding. It is still a challenge to extend such an initialization to the case of a black box polynomial.
26.2 Definitions
1. Definitions. Consider Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s) iterations [71, 58, 130, 143, 148, 31, 67, 38, 3, 22]):
$z\leftarrow z-N(z),~{}z_{j}\leftarrow z_{j}-W(z_{j}),~{}z_{j}\leftarrow z_{j}-E_{j}(z_{j}),~{}j=1,\dots,d,~{}{\rm where}$ (26.1)
$N(x)=W(x)=E_{j}(x)=0~{}{\rm for~{}all}~{}j~{}{\rm if}~{}p(x)=0$ (26.2)
and otherwise
$N(z):=\frac{p(z)}{p^{\prime}(z)},~{}W(x):=\frac{p(x)}{l^{\prime}(x)},~{}\frac{1}{E_{j}(x)}:=\frac{1}{N(x)}-\sum_{i=1,i\neq j}^{d}\frac{1}{x-z_{i}}$ (26.3)
for $l(x):=\prod_{j=1}^{d}(x-z_{j})$.
Newton’s iterations can be initiated at a single point $z$ or at many points and can approximate a single root, many or all $d$ roots at the cost decreasing as the number of roots decreases [71, 58, 143, 141, 85, 130].
In contrast, Weierstrass’s and Ehrlich’s iterations begin at $d$ points for a $d$th degree polynomial and approximate a subset of its $d$ roots about as fast and almost as slow as they approximate all its $d$ roots.
Near the roots Newton’s and Weierstrass’s iterations converge with quadratic rate, Ehrlich’s iterations with cubic rate [145], but empirically all of them consistently converge very fast globally, right from the start, assuming a proper initialization, which involves the coefficients of $p$ (see [15, 22]). [^37] Formal support for such global convergence is much weaker for Newton’s iterations (see [71, 58, 143], and the references therein) and is completely missing for Weierstrass’s and Ehrlich’s iterations.
All expressions in (26.3) for the functions $N(x)$, $W(x)$, and $E_{j}(x)$ are already presented or can be rewritten as ratios $\frac{u(x)}{v(x)}$ for pairs of polynomials $u(x)$ and $v(x)$, whose coefficients are given or readily available at low cost if the polynomial $p(x)$ is gien with its coeffcients (see [102, Problems 3.6.1 and 3.6.3]). Then multipoint evaluation of such a function amounts to multipoint evaluation of $u(x)$ and $v(x)$ and division of the computed values, and one can accelerate Newton’s, Weierstrass’s, and Ehrlich’s iterations dramatically by incorporating the algorithms of [81, 59].
26.3 Acceleration with Fast Multipole Method (FMM)
The algorithms of [81, 59] deeply involve the coefficients of $p(x)$. Next we cover acceleration of Newton’s, Weierstrass’s, and Ehrlich’s iterations applied to black box polynomial root-finding. Given $3d$ complex values $w_{j}$, $z_{i}$, $v_{j}$ such that $z_{i}\neq v_{j}$ for all $i,j=1,\dots,d$, Fast Multipole Method (FMM) by Greengard and Rokhlin [50] and by Carrier, Greengard, and Rokhlin [28] computes the sums $\sum_{j=1}^{d}\frac{w_{j}}{x_{i}-z_{j}}$ as well as $\sum_{j=1,j\neq i}^{d}\frac{w_{j}}{z_{i}-z_{j}}$ for all $i$ by using $O(db)$ arithmetic operations, with the precision of order $b$, versus $2d^{2}$ arithmetic operations in the straightforward algorithm. The paper [121] proposes and extensively tests its acceleration to Superfast Multipole Method. The papers [105, 107] list the recent bibliography on FMM and related computations with Hierarchical Semiseparable (HSS) matrices. Relevant software libraries developed at the Max Planck Institute for Mathematics in the Sciences can be found in HLib,
The acceleration based on FMM is practically valuable in spite of the limitation of the working and output precision to the order of $\log(d)$ and the resulting limitation on the decrease of the bit operation complexity.
Next we incorporate FMM into Newton’s, Weierstrass’s, and Ehrlich’s iterations. First relate the roots of a polynomial $p(x)$ to the eigenvalues of a diagonal-plus-rank-one matrix
$A=D+{\bf v}{\bf e}^{T}~{}{\rm for}~{}D=\operatorname{diag}(z_{j})_{j=1}^{d},~{}{\bf e}=(1)_{j=1}^{d},~{}~{}{\bf w}=(w_{j})_{j=1}^{d},~{}w_{j}=W(z_{j}),$
$W(x)$ of (26.3) and all $j$ (cf. [16]). Let $I_{d}$ denote the $d\times d$ identity matrix, apply the Sherman-Morrison-Woodbury formula, and obtain
$xI_{d}-A=(xI_{d}-D)(I_{d}-(xI_{d}-D)^{-1}{\bf w}{\bf e}^{T}).$
Hence
$\frac{1}{p_{d}}p(x)=\det(xI_{d}-A)=-\sum_{i=1}^{d}w_{i}\prod_{j\neq i}(x-z_{j})+\prod_{j=1}^{d}(x-z_{j}).$ (26.4)
Indeed, the monic polynomials of degree $d$ on both sides have the same values for $x=z_{j}$, $j=1,\dots,d$ (cf. [16])). [^38] Gerschgorin’s theorem (cf. [48]) for the matrix $xI_{d}-A$ defines neighborhoods of its eigenvalues (the roots of $p$), which helps to analyze and to modify the iterations (cf. [22]). Equations (26.4) imply that
$p(z_{i})=-w_{i}\prod_{j\neq i}(z_{i}-z_{j}),~{}p^{\prime}(z_{i})=\prod_{j\neq i}(z_{i}-z_{j})\Big{(}1-\sum_{j\neq i}\frac{w_{i}+w_{j}}{z_{i}-z_{j}}\Big{)},$
and therefore
$\frac{w_{i}}{N(z_{i})}=\sum_{j\neq i}\frac{w_{i}+w_{j}}{z_{i}-z_{j}}-1.$ (26.5)
Hence, given the $2d$ values of $z_{j}$ and $w_{j}=W(z_{j})$ for all $j$, one can fast approximate $N(z_{j})$ and consequently $E_{j}(z_{j})$ for all $j$ (cf. (26.3)) by applying FMM.
Every iteration but the first one begins with some values $z_{j}$, $w_{j}$, and updated values $z_{j}^{\rm(new)}$ of (26.1), for $j=1,\dots,d$, and outputs updated values $w_{j}^{\rm(new)}=W(z_{j}^{\rm(new)})$ for all $j$. Then equations (26.3) and (26.5) define $N(z_{j}^{\rm(new)})$ and $E_{j}(z_{j}^{\rm(new)})$ for all $j$. To update the values $w_{j}$, apply [22, equation (17)]:
$w_{i}^{\rm new}=(z_{i}^{\rm new}-z_{i})\prod_{j\neq i}\frac{z_{i}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}\Big{(}\sum_{k=1}^{d}\frac{w_{k}}{z_{i}^{\rm new}-z_{k}}-1\Big{)},~{}i=1,\dots,d.$
In the case where $z_{i}^{\rm new}=z_{i}$ if and only if $i>k$ this updating is still used for $i=1,\dots,k$ but is simplified to
$w_{i}^{\rm new}=w_{i}\prod_{j=1}^{k}\frac{z_{i}-z_{j}}{z_{i}^{\rm new}-z_{j}}~{}{\rm for}~{}i=k+1,\dots,d.$
Now fast compute the values $\sum_{k=1}^{d}\frac{w_{k}}{z_{i}^{\rm new}-z_{k}}$ for all $i$ by applying FMM.
Also compute the values
$R_{i}=(z_{i}^{\rm new}-z_{i})\prod_{j\neq i}\frac{z_{i}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}$
fast in two ways.
(i) Use the following customary trick (see [32, Remark 3]): approximate the values $\ln(R_{i})$ for $i=1,\dots,d$ fast by applying FMM and then readily approximate $R_{i}=\exp(\ln(R_{i}))$ for all $i$. [32] proves numerical stability of the reduction to FMM of a similar albeit a little more involved rational expressions.
(ii) Observe that
$\ln\Big{(}\frac{R_{i}}{z_{i}^{\rm new}-z_{j}}\Big{)}=\sum_{j\neq i}\ln\Big{(}1-\frac{z_{j}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}\Big{)}~{}{\rm for}~{}i=1,\dots,d$
and that
$\ln\Big{(}1-\frac{z_{j}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}\Big{)}\approx\frac{z_{j}-z_{j}^{\rm new}}{z_{i}^{\rm new}-z_{j}^{\rm new}}~{}{\rm if}~{}|z_{j}^{\rm new}-z_{j}|\ll|z_{i}^{\rm new}-z_{j}^{\rm new}|.$
Let the latter bound hold for $j\in\mathbb{J}_{i}$ and let $\mathbb{J}^{\prime}_{i}$ denote the set of the remaining indices $j\neq i$. Then
$R_{i}\approx\exp(\sigma_{i})\prod_{j\in\mathbb{J}^{\prime}_{i}}\Big{(}1-\frac{z_{j}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}\Big{)},~{}\sigma_{i}=\sum_{j\in\mathbb{J}_{i}}\frac{z_{j}-z_{j}^{\rm new}}{z_{i}^{\rm new}-z_{j}^{\rm new}},~{}i=1,\dots,d.$
By applying FMM fast approximate $\sigma_{i}$ for all $i$ and compute the values $\prod_{j\in\mathbb{J}^{\prime}}\frac{z_{i}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}$ for all $i$ in at most $4\sum_{i=1}^{d}~{}|\mathbb{J}_{i}^{\prime}|-d$ arithmetic operations for $|\mathbb{J}^{\prime}_{i}|$ denoting the cardinality of the set $\mathbb{J}^{\prime}_{i}$. Compare the latter cost bound with order of $d^{2}$ arithmetic operations in the straightforward algorithm.
Clearly the above updating computation is simpler and more stable numerically for $W(z_{j})$ versus $N(z_{j})$ and even more so versus $E_{j}(z_{j})$, for $j=1,\dots,d$.
Let us comment on the changes where one incorporates FMM into Newton’s, Weierstrass’s, and Ehrlich’s iterations implicitly applied to the wild factor $f(x)$ of $p(x)$. The same expression (3.3) defines Ehrlich’s iterations for both $f(x)$ and $p(x)$ by virtue of Thm. 13.1. Newton’s iterations to a wild factor one additionally subtract the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ (cf. (3.3)) and can apply FMM to accelerate numerical computation of this sum. To apply Weierstrass’s iterations to a wild factor one performs additional division by $t(x)$ and decreases the degree of the divisor $l^{\prime}(x)$ from $d-1$ to $w-1$.
26.4 MPSolve and its competition with Newton’s iterations
Since its creation in 2000 the package MPSolve by Bini and Fiorentino [15], revised in 2014 by Bini and Robol [22], has been remaining user’s choice for approximation of all $d$ complex roots of a polynomial $p$. It implements Ehrlich’s iterations of 1967 [38], aka Aberth’s of 1973 [3], which empirically converge very fast globally – right from the start, under proper initialization, involving the coefficients of $p$. Such convergence has been consistently observed in decades-long worldwide applications of these iterations as well as in the case of some other root-finders by means of functional iterations, notably Weierstrass’s iterations [148], aka Durand and Kerner’s [31, 67]).
Unsuccessful attempts of supporting this empirical behavior formally (e.g., see [4, 54]), however, have recently ended with surprising proofs that Weierstrass’s and Ehrlich’s iterations diverge in the case of inputs that lie on some complex manifolds, being an open domain in the case of Weierstrass’s iterations (see Reinke, Schleicher, and M. Stoll [131, 129]).
Serious competition to MPSolve came in 2001 from the package EigenSolve by Fortune [43], but the later version of MPSolve [22] combines the benefits of both packages. An implementations of Newton’s iterations for approximation of all roots by Schleicher and R. Stoll in [143] competed with MPSolve in 2017, but then MPSolve has regained upper hand by means of incorporation of Fast Multipole Method (FMM), even though such implementation is simpler and can be more efficient for Newton’s than Ehrlich’s iterations (see the previous subsection). The competition seems to resume in 2024 (see [12, 130, 85]).
26.5 Computation of a disc covering all zeros of a black box polynomial
26.5.1 Introductory comments
Given coefficients of a polynomial $p$ of a degree $d$, we can readily compute its largest and smallest root radii $r_{1}$ and $r_{d}$, respectively (cf. [41, 52]). Next we compute a range $[\rho_{-},\rho_{+}]$ that brackets all root radii or, equivalently, compute an upper estimate for the largest root radius $r_{1}$ and a low estimate for the smallest root radius $r_{d}$ of $p(x)$.
We can apply the estimates of Sec. 13.4 to the reverse polynomial $p_{\rm rev}(x)$ if its coefficients are available, but next we study this problem for a black box polynomial. Its solution is required, e.g., in order to initialize algorithms for all $d$ roots by means of functional iterations.
Suppose that we have an algorithm that computes a disc $D(0,\rho_{-})$ containing no zeros of a black box polynomial $p(x)$; clearly, $\rho_{-}\leq r_{d}$. By applying this algorithm to the reverse polynomial $p_{\rm rev}(x)$ we will obtain the upper bound $1/\rho_{-,{\rm rev}}$ on $r_{1}$.
It remains to describe computation of the disc $D(0,\rho_{-})$. Empirically we can do this by applying the algorithms of [113, 49], but next we study formally some alternative algorithms.
To obtain a range $[\rho_{-},\rho_{+}]$ that brackets all root radii, we need upper estimate for the root radii $r_{1}$ and $r_{d}$, but next we only estimate $r_{d}$ because we can extend them to $r_{1}$ by estimating $r_{d}$ for the reverse polynomial $p_{\rm rev}(x)$. We obtain a low bound on $r_{d}$ by computing a disc that contains no zeros of a black box polynomial $p(x)$.
Hereafter let $p(x)$ denote a black box polynomial.
26.5.2 Probabilistic solution
Seek a disc $D(c,\rho)$ containing no zeros of $p(x)$ in its interior, where we can choose any positive $\rho$ Sample a random $c$ in the unit disc $D(0,1)$ under the uniform probability distribution in $D(0,1)$.
The disc has area $\pi$, while the union of the $d$ discs $D(x_{j},\rho)$, for $j=1,2,\dots,d$, has area at most $\pi d\rho^{2}$. Therefore, $c$ lies in that union with a probability at most $d\rho^{2}$, which we can decrease at will by decreasing $\rho$. Unless $c$ lies in such a union, we obtain a desired disc $D(c,\rho)$.
By applying our e/i tests of Secs. 14.414.3 to this disc, we can verify correctness of its choice.
26.5.3 Deterministic solution
We can deterministically compute a desired disc $D(0,\rho_{-})$ at a higher cost, based on a soft e/i test.
[…]
27 Fast approximation of the $d$ root radii with applications
The customary initialization recipes of Weierstrass"s and Ehrlich’s iterations involve the $d$ root radii at the origin and consequently require to use the coefficients of $p$. To initialize Ehrlich’s iterations MPSolve first applies Bini’s heuristic recipe of [10], reused in [15, 22], for the approximation of the $d$ root radii. Given such approximations MPSolve defines $k$ concentric annuli $A(c,\rho_{i},\rho_{i}^{\prime})$ and $k$ integers $m_{i}$ such that #$\#(A(c,\rho_{i},\rho_{i}^{\prime}))=m_{i}$ for $i=1,\dots,k$, $k\leq d$, and $\sum_{i=1}^{k}m_{i}=d$. Then MPSolve initializes Ehrlich’s iterations at $d$ points $z_{j}$, $j=1,2,\dots,d$, exactly $m_{i}$ of them being equally spaced on the circle $C(c,(\rho_{i}+\rho_{i}^{\prime})/2)$. This heuristic initialization rec;ipe turns out to be highly efficient.
The algorithm supporting the next theorem also approximates the $d$ root radii but is certified, deterministic, and faster than Bini’s. It was first outlined and briefly analyzed in [134] for fast approximation of a single root radius and then extended in [92, Sec. 4] (cf. also [93, Sec. 5]) to approximation of all root radii of a polynomial. As in [10, 15, 22] it uses all $d$ coefficients of $p$. A variation of this algorithm in [62] supports the following theorem.
Theorem 27.1.
[62, Prop. 3]. Given the coefficients of a polynomial $p=p(x)$, one can approximate all the $d$ root radii $|x_{1}|,\dots,|x_{d}|$ within the relative error bound $4d$ at a Boolean cost in $O(d\log(||p||_{\infty}))$.
[62] extends the supporting algorithm to approximation of all root radii within a positive relative error bound $\Delta$ by performing $\lceil\log_{2}\log_{1+\Delta}(2d)\rceil$ DLG root-squaring iterations (3.7).
Combine Thm. 27.1 with estimates for the Boolean cost of performing these iterations implicit in the proof of [134, Cor. 14.3] and obtain
Corollary 27.1.
Given the coefficients of a polynomial $p=p(x)$ and a positive $\Delta$, one can approximate all the $d$ root radii $|x_{1}|,\dots,|x_{d}|$ within a relative error bound $\Delta$ in $\frac{1}{d^{O(1)}}$ at a Boolean cost in $O(d\log(||p||_{\infty})+d^{2}\log^{2}(d))$.
The algorithm supporting the corollary and said to be Alg. 27.1a computes at a low cost $d$ suspect annuli $A(0,r_{j,-},r_{j,+})$ for $j=1,\dots,d$ (some of them may pairwise overlap), whose union $\mathbb{U}$ contains all $d$ roots. The algorithm was the point of departure for amazing progress in [59], which extends the novel ingenious algorithm of [81] for multipoint polynomial evaluation.
We conclude this subsection with a list of some other benefits of fast approximation of the $d$ root radii.
(i) In a subdivision step one can skip testing exclusion and discard a square unless it intersects the set $\mathbb{U}$.
(ii) The search for real roots can be narrowed to the intersection $\mathbb{I}$ of $\mathbb{U}$ with the real axis, which consists of $2m$ segments of small length (see Fig. 11). [126] has tested a simple real root-finder that approximates the roots in $\mathbb{I}$ by initiating Newton’s iterations at the centers of the segments. According to these tests the algorithm is highly efficient for the approximation of all isolated real roots. One can extend Newton’s iterations at the centers of the segments by applying also our root-radii algorithm of Sec. 9. By complementing the root radii approach with some advanced novel techniques, the paper [62] has noticeably accelerated the real root-finder [70], which is currently user’s choice.
(iii) Initialization based on Alg. 27.1a should improve global convergence of Ehrlich’s, Weierstrass’s, and other functional iterations for root-finding as well as of Kirrinnis’s factorization algorithm of [68].
[…]
We can apply this recipe to every annulus containing a small number of zeros of $p$ and isolated from its external zeros and finally approximate the remaining wild zeros of $p$ by applying implicit deflation (cf. (3.3)).
$x$$y$may contain rootsmay contain roots
Figure 11: Two approximate root radii define two narrow suspect annuli and four suspect segments of the real axis.