### v37

# Polynomial Root-Finding and Algebraic Eigenvalue Problem

Abstract
Univariate polynomial root-finding has been studied for four millennia and very intensively in the last decades. Our new near-optimal root-finders approximate all complex zeros of a polynomial almost as fast as one accesses its coefficients with the precision required for the solution within a prescribed error bound. The root-finders work even for a black box polynomial, defined by an oracle (black box subroutine) for its evaluation rather than by its coefficients. Hence they (i) support approximation of the eigenvalues of a matrix in a record Las Vegas expected bit operation time and (ii) are particularly fast for polynomials that can be evaluated fast such as the sum of a few shifted monomials of the form $(x-c)^{d}$ or a Mandelbrot-like polynomial defined by a recurrence. Our divide and conquer algorithm of ACM STOC 1995 is the only other known near-optimal polynomial root-finder, but it extensively uses the coefficients, is quite involved, and has never been implemented, while according to extensive numerical experiments with standard test polynomials already a slower initial implementation of our new root-finders competes with user’s choice package of root-finding subroutines MPSolve and supersedes it more and more significantly as the degree of a polynomial grows large. We elaborate upon design and analysis of our algorithms, comment on their potential heuristic acceleration, and briefly cover polynomial root-finding by means of functional iterations. Our techniques can be of independent interest.
Keywords:
polynomial root-finding, matrix eigenvalues, computational complexity, black box polynomials, symbolic-numeric computing.
2020 Math. Subject Classification:
65H05, 26C10, 30C15
PART I: State of the Art and Our Progress (Outline)
1 State of the Art and our Main Results
1.1 Polynomial Root-finding: Problems, Impact, and Earlier Studies
According to [4, 80, 81, 66], univariate polynomial root-finding, that is, the solution of a polynomial equation $p(x)=0$, had been the central problem of Mathematics and Computational Mathematics for about four millennia since Sumerian times; it was still a central problem well into the 19th century. In a long way from its rudimentary level to modern version, its study has been responsible for the introduction of mathematical rigor and notation as well as the concepts of irrational, negative, and complex numbers, algebraic groups, rings, and fields etc.
The existence of complex roots of the equation $p(x)=0$ is called the Fundamental Theorem of Algebra. The attempts of its proof by d’Alembert (1746), Euler (1749), de Foncenex (1759), Lagrange (1772), Laplace (1795), and Gauss (1799) had gaps (cf. [105]), and the first rigorous proof is due to an amateurish mathematician Argand (1806, revised in 1813).
Already about 500 BC it was proved in Pythagorean school that the roots of the equation $x^{2}=2$ were irrational radicals, while even with the help of radicals one cannot express the roots rationally if the degree $d$ of $p(x)$ exceeds 4, according to Ruffini (1799,1813) and Abel (1824); moreover, one cannot do this already for specific equations such as $x^{5}-x-1=0$ or $x^{5}-4x-2=0$, according to Galois theory of 1830, 1832.
Nevertheless, in his constructive proof of the Fundamental Theorem of Algebra in [120], Herman Weyl has rationally approximated within any fixed positive error tolerance all roots of $p(x)=0$ lying in a fixed square in the complex plane, that is, all roots if the square is large enough.
This solved the following Problem 0 of modern polynomial root-finding:
INPUT: $d+1$ complex coefficients $p_{0},p_{1},\dots,p_{d}$ of a polynomial
$p=p(x):=\sum_{i=0}^{d}p_{i}x^{i}:=p_{d}\prod_{j=1}^{d}(x-z_{j}),$ (1.1)
a positive error tolerance $\epsilon=1/2^{b}>0$, and a bounded domain $\mathbb{D}$ in the complex plane, e.g., a disc or a rectangle.
OUTPUT: approximations within $\epsilon$ to every zero of $p(x)$ (aka every root of $p(x)=0$) [^1] Hereafter we freely say just “root” or “zero” for short, always count them with their multiplicities, and do not distinguish multiple roots from pairwise isolated root clusters of diameters less than $\epsilon$. lying in the domain $\mathbb{D}$ such that
$|\tilde{z}_{j}-z_{j}|\leq\epsilon~{}{\rm if}~{}z_{j}\in\mathbb{D}.$ (1.2)
Example 1.1.
[Approximation of all $d$ roots.] For a real
$R\geq\max_{j=1}^{d}\{|x_{j}|\}$ (1.3)
(see [39] for simple expressions for such bound $R$) and $\mathbb{D}$ denoting the disc $D(0,R):=\{x:~{}|x|\leq R\}$, Problem 0 amounts to approximation of all $d$ roots $z_{j}$ within $\epsilon$.
Example 1.2.
[Numerical real root-finding.] Root-finding in the real line or its segment (real root-finding) is highly important because in many applications, e.g., to optimization in algebraic geometry and geometric modeling, only real roots of a polynomial are of interest, and typically they are much less numerous than all $d$ complex roots [50, 29]. Now consider Problem 0 where $\mathbb{D}$ denotes the minimal rectangle in the complex plane that covers the $\epsilon$-neighborhood of a fixed real line segment. This task addresses approximation of both real and near-real roots, that is, their perturbations due to small input and rounding errors. The known real polynomial root-finders (cf. [63, Chs. 2 and 3]) ignore this important issue.
Solution of Problem 0 can be extended to the eigenvalue problem for matrices and matrix polynomials (see our Secs. 1.3 and 1.5 and Part (C) of Sec. 5.4) and is involved into the solution of a multivariate polynomial system of equations [21]; both of them are much studied areas of modern computations. [65, Introduction] specifies some applications of polynomial root-finding to computer algebra, signal processing, geometric modeling, control, and financial mathematics. Furthermore, Problem 0 for a domain $\mathbb{D}$ containing all roots is closely linked to polynomial factorization (see Secs. 22 and 23), which has applications to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [121, 11, 5, 26, 27, 118].
Hundreds of efficient polynomial root-finders have been proposed [63, 65] and keep appearing (see [97, 15, 30, 54, 59, 99, 89, 16, 17, 49, 91, 92, 58, 45, 47, 64, 48, 44, 96], and the bibliography therein). In [105] Smale proposed to study Problem 0 in terms of its arithmetic complexity and computational precision. Together they would define its bit operation complexity, but Smale deduced no complexity estimates. [106] Schönhage did this in [106]. His much involved root-finder, presented with terse complexity analysis on 46 pages, approximates all $d$ roots by using $\tilde{O}(bd^{3})$ bit operations, that is, $bd^{3}$ bit operations up to a factor in $bd$ [^2] Error amplification factors $2^{cd}$ are inevitable in polynomial root-finding, e.g., in routine shift of the variable, and assume that $b\geq d$ throughout our paper; alternatively one can replace $b$ with $b+d$ in our upper bounds on the computational cost.
His work prompted intensive study of root-finding by researchers in the Theory of Computing in the 1980s and 1990s, culminated at ACM STOC 1995 with our near-optimal divide and conquer root-finder of [78] (also see its journal versions [79, 84] and surveys [80, 81]). Both upper bound in $\tilde{O}(bd^{2})$ and lower bound of $0.25(d+1)bd$ bit operations of [78, 84] decrease by a factor of $d$ for polynomials of Def. 22.1 of Sec. 22, whose all zeros and tiny clusters of zeros are pairwise sufficiently well isolated (see Thms. 22.1 and 22.4 and Cor. 22.4 in Sec. 22).
1.2 Reduction to the case of the unit disc
Hereafter $D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$ and $C(c,\rho):=\{x:~{}|x-c|=\rho\}$ denote the disc of radius $\rho>0$ centered at a complex point $c$ and its boundary circle, respectively, and Problem $\widehat{0}$ denotes Problem 0 where the domain $\mathbb{D}$ is the unit disc $D(0,1)$; we can extend its solution to root-finding in the disc $D(c,\rho)$ for any pair of complex $c$ and positive $\rho$ by means of applying the linear map
$x\mapsto\frac{x-c}{\rho},~{}p(x)\mapsto t(x)=p\Big{(}\frac{x-c}{\rho}\Big{)},~{}D(c,\rho)\mapsto D(0,1),$ (1.4)
which shifts the variable $x$ by $c$ and then scales $x-c$ by $1/\rho$.
Next we reduce Problem 0 for any domain $\mathbb{D}$ to Problem $\widehat{0}$.
(i) Two ways from the roots in the unit disc to all roots. Given a value $R$ of (1.3), we can scale the variable, $x\mapsto y=\frac{x}{R}$, $p(x)\mapsto p(y)=p(\frac{x}{R})$, and then approximate every root $x_{j}$, for $j=1,\dots,d$, within $\epsilon R$ by solving Problem $\widehat{0}$ for $p(y)$.
Here is Schönhage’s alternative reduction in [106] of Problem 0 for all $d$ roots to Problem $\widehat{0}$: Given an algorithm $\mathcal{A}$ that solves Problem $\widehat{0}$, compute also approximations $\tilde{y}_{j}$ within $\epsilon$ to the reciprocals $y_{j}=1/z_{j}$ of all roots lying outside the disc $D(0,1)$ – simply by applying this algorithm to the reverse polynomial
$p_{\rm rev}(x):=x^{d}p\Big{(}\frac{1}{x}\Big{)}=\sum_{i=0}^{d}p_{i}x^{d-i},~{}p_{\rm rev}(x)=p_{0}\prod_{j=1}^{d}\Big{(}x-\frac{1}{z_{j}}\Big{)}~{}{\rm if}~{}p_{0}\neq 0.$ (1.5)
Write $\Delta_{j}:=\tilde{y}_{j}-y_{j}$ and deduce that $\frac{1}{y_{j}}-\frac{1}{\tilde{y}_{j}}=\frac{\Delta_{j}}{\tilde{y}_{j}~{}y_{j}}$, and so $\frac{\tilde{y}_{j}}{y_{j}}-1=\frac{\Delta_{j}}{y_{j}}=\Delta_{j}z_{j}$. Recall that $|\Delta_{j}|\leq\epsilon$ by assumption and obtain that $|\frac{\tilde{y}_{j}}{y_{j}}-1|\leq\epsilon|z_{j}|$, thus arriving at the following
Proposition 1.1.
The reciprocal of an approximation within $\epsilon$ to a zero of $p_{\rm rev}(x)$ in the disc $D(0,1)$ approximates the associated zero of $p(x)$ lying outside that disc within a relative error at most $\epsilon$.
(ii) From roots in the unit disc to roots in a bounded domain. Given a bounded domain $\mathbb{D}$ on the complex plane, we can successively (a) cover it with a minimal disc $D(c,\rho)$, (b) map this disc into the disc $D(0,1)$, (c) solve Problem $\widehat{0}$ for the resulting polynomial $t(x)$ of (1.4), and finally (d) use the map (1.4) again – this time to approximate all roots lying in the intersection of the disc $D(c,\rho)$ with the domain $\mathbb{D}$. In this way we also approximate extraneous roots lying in $D(c,\rho)$ but not in $\mathbb{D}$. In Sec. 2.1 we curtail such an extraneous work (see Prop. 2.1 and Remark 2.1).
1.3 Black box root-finders and approximation of matrix eigenvalues
A nontrivial variant of Newton’s iterations by Louis and Vempala at FOCS 2016 [59] approximates within a fixed positive $\epsilon\leq 1/2$ an absolutely largest zero of a real-rooted polynomial $p$ of a degree $d$ at the cost of its evaluation with a precision in $O((b+\log(R))\log(R~{}d)))$ at $O((b+\log(R))\log(d))$ points for $b:=\log_{2}(R/\epsilon)$ and a fixed $R$ satisfying (1.3). While the previous polynomial root-finders have been devised for polynomials given in the monomial basis, that is, with their coefficients, Louis and Vempala in [59] never involve them, working for a black box polynomial $p$ – given with an oracle (black box subroutine) for its evaluation, and then extend their root-finder to efficient approximation of the eigenvalues of a matrix. Next we state the relevant tasks.
Problem 1 (Black Box Polynomial Root Approximation) is Problem 0 for a black box polynomial $p$.
Problem 2: Approximation of the Eigenvalues of a Matrix. [^3] As in [59], the problem covers or can be extended from a matrix to a polynomial matrix, aka matrix polynomial.
INPUT: $\epsilon=1/2^{\tilde{b}}>0$, an integer $d>0$, a $d\times d$ matrix $M$, and a domain $\mathbb{D}$ on the complex plane.
OUTPUT: Approximations within $\epsilon$ to every eigenvalue of $M$ lying in the domain $\mathbb{D}$.
The output of Problems 0, 1, and 2 may include multiple approximations within $\epsilon$ to the same root or eigenvalue, but the next modifications exclude such a confusion.
Problems $0^{*}$, $1^{*}$, $2^{*}$: Isolation of Clusters of Roots and Eigenvalues. Under the assumptions of Problems 0, 1, and 2 compute at most $d$ disjoint discs of radii at most $\epsilon$ (hereafter $\epsilon$-discs) that cover all roots or eigenvalues lying in the domain $\mathbb{D}$; also compute the number of the roots or eigenvalues lying in each $\epsilon$-disc.
Remark 1.1.
If we collapse all zeros of $p$ lying in every $\epsilon$-disc into a single multiple zero placed in the center of the disc, then these zeros will lie at the distances greater than $2\epsilon$ from each other. If a polynomial $p$ is square-free (having only simple zeros) and has integer coefficients, then Problem 0 turns into the well-known Problem of Root-isolation (see Problem ISOL and Remark 23.4 in Sec. 23.5). These comments can be immediately extended to Problems 1 and 2.
1.4 More on Black Box Polynomial Root-finders
1.4.1 Why we should use them
Namely, the algorithm of Fiduccia [32] and Moenck and Borodin [67] (see [12, Sec. 4.5] for historical account) evaluates a polynomial of degree $m$ at $m$ points in $O(m\log^{2}(m))$ ops if its coefficients are given, versus $2m^{2}$ ops involved in Horner’s algorithms. Numerically stable algorithms of [64, 44] yield such acceleration also in terms of bit operation cost.
One can perform black box evaluation at $q>d$ points $x_{1},\dots,x_{q}$ and then interpolate to $p$. The bit operation cost of interpolation, however, is unbounded unless the $q$ points are well isolated from the roots, but it is a challenge to yield such isolation. Furthermore, even if the isolation is ensured, the cost of interpolation exceeds the cost of the solution of Problems 0, 0, 1, 1, 2, and 2 in the cases where domain $\mathbb{D}$ contains only a small number of roots as well as where a polynomial $p$ can be evaluated fast because root-finding for a polynomial is routinely reduced to its evaluation at sufficiently many points [63, 65].
For example, one only needs $O(\log(d))$ ops for evaluation of a Mandelbrot-like polynomial $p:=M_{k}(x)$, where $M_{0}:=x,~{}M_{i+1}(x):=xM_{i}^{2}(x)+1$, $i=0,1,\dots,k$, $d=2^{k}-1$, or the sum of a bounded number of shifted monomials, such as $p:=\alpha(x-a)^{d}+\beta(x-b)^{d}+\gamma(x-c)^{d}$ for six constants $a$, $b$, $c$, $\alpha$, $\beta$, and $\gamma$. For comparison, if we compute the coefficients of such a polynomial and then use them to evaluate it at a point $x$ by applying Horner’s algorithm, then we would need $2d$ ops, and this bound is optimal (see [72, 53] or [112, Section “Pan’s method"]).
In computational practice numerical stability of polynomial evaluation is highly important, and it is ensured when a polynomial is represented in the Bernstein basis [57]. Transition to it from the monomial basis is costly but is not needed where a root-finder works for a black box polynomial.
One can benefit from black box root-finders even where an input polynomial is given with its coefficients.
(i) Application of a black box root-finder to the characteristic polynomial $p(x)=\det(xI-M)$ of a $d\times d$ matrix $M$ only involves the values of this polynomial but not its coefficients, whose overall size exceeds the size (a norm) of the matrix $M$ by a factor of $d$. Hence extension of such a root-finder to Problem 2 saves a factor of $d$ in bit precision and bit operation count versus the known solutions of Problems 2 and 2, based on [106, Sec. 21].
(ii) Map (1.4) is a valuable root-finding tool, but its application is limited because it destroys sparsity and blows up computational precision and bit operation cost. Similar comments can be applied to maps (3.4) and (3.5) of Sec. 3.4. Black box polynomial root-finders keep all benefits of using those maps but avoid the latter notorious problems by involving no coefficients of $t(x)$ and $p_{h}(x)$ of these maps.
(iii) Deflation of a factor of a polynomial is a well-known cause of the coefficient swell, e.g., for degree 300 factors of $x^{1000}+1$, which blows up precision and bit complexity of computations. For an exception, one can readily deflate a small degree factor $f(x)$ and approximate its zeros, but the coefficients of the high degree factor $g(x)=p(x)/f(x)$ still tend to blow up. We, however, avoid their computation when we apply a black box root-finder to the polynomial $g(x)$ represented by the pair of $p(x)$ and $f(x)$.
1.4.2 Progress at FOCS 2016 and its limitations
The known polynomial root-finders, except [59], extensively use the coefficients of $p$, and we must be careful in adopting their techniques for Problems 1 and 1.
Besides proposing novel root-finder for a black box polynomial with extension to approximation of an eigenvalue of a matrix, Louis and Vempala in [59] presented non-trivial techniques for bounding precision of root-finding computations. Extension of these techniques helped us simplify dramatically the estimation of the bit operation complexity of our root-finders compared to those of [106, 78, 84, 52], which bypass estimation of computational precision.
On the other hand, direct extension of the root-finder of [59] cannot help solve even Problem 1. Indeed, repeatedly deflating $p$, by peeling off large roots, can accumulate errors that grow too quickly, together with the coefficient length of the factors: notice, e.g., the swell of the coefficients of $p(x)=x^{d}+1$ in transition to its factors of degree $d/2$, say. This is true even if one performs deflation in the black box setting implicitly, based on Eqn. (3.3) of Sec. 3. Indeed, the high order derivatives involved in the algorithm of [59] are proportional to the coefficients and together with them swell in deflation.
Even if deflation succeeded, the algorithm of [59] would have only solved the well-known and much easier problem of root-finding for a real-rooted polynomial (cf. [75, 18, 25]). Moreover, the resulting algorithm for that much easier problem would involve more than $b^{2}d^{2}$ bit operations, exceeding by a factor of $b$ the upper bound of [78, 84] for general polynomial root-finding.
Furthermore, we accelerate the solution of Problem 0 by a factor of $m/\log^{2}(m)$ by incorporating fast multipoint polynomial evaluation of [32, 67, 52, 64], but we would lose this chance if we successively approximated the zeros of $p$ or matrix eigenvalues by using deflation.
Thus, instead of applying high order Newton’s iterations of [59] and deflation, we recall and advance subdivision iterations of [120, 39, 100, 82].
1.5 Our complexity estimates
Unlike the root-finders of [106, 78, 84, 52], seeking all $d$ roots, of [59], seeking a single absolutely largest root and of [78, 84], seeking all $d$ roots or all roots in a disc well isolated from the external roots, we approximate the roots in any domain $\mathbb{D}$ in the complex plane. Root approximation within $\epsilon$ in a domain $\mathbb{D}$ involves the roots in its $\epsilon$-neighborhood; our complexity estimates grows with the number of roots in a little larger but very small neighborhood of $\mathbb{D}$.
Definition 1.1.
For a polynomial $p=p(x)$, a domain $\mathbb{D}$ in the complex plane having a diameter $\Delta$, and a positive $\beta$ such that $\beta\Delta>2\epsilon$, write $m=m(p,\mathbb{D},\beta)$ if precisely $m$ roots of $p(x)$ lie in the $\beta\Delta$-neighborhood of $\mathbb{D}$.
We will prove our Main Theorems 1-4 below for $m$ of Def. 1.1, any fixed constant $\beta>2\epsilon/\Delta$, and either $p(x)=\det(xI-M)$ in the case of Problems 2 and 2 or $p(x)$ of (1.1) otherwise. [^4] Hereafter, whenever we say that subdivision iterations are applied to a input domain $\mathbb{D}$ of diameter $\Delta$ that contains $m$ roots, we mean that so does its $\beta$-neighborhood.
Theorem 1.1.
(Main Theorem 1.) One can solve Problems 1 and 1 by querying the polynomial evaluation oracle at $\tilde{O}(m^{2})$ locations with a computational precision $O(b\log(d))$.
Theorem 1.2.
(Main Theorem 2.) One can solve Problems 2 and 2 at a Las Vegas randomized bit operation cost
$\mathbb{B}_{\rm eigen}=\tilde{O}(\tilde{b}d^{\omega}m^{2})$ (1.6)
for $\tilde{b}:=\log_{2}(||M||_{F}/\epsilon)$, the Frobenius norm $||M||_{F}$ of a matrix $M$, and $\omega$ denoting an exponent of feasible or unfeasible matrix multiplication (MM). [^5] The current records are about 2.7734 for exponents of feasible MM [73, 88, 90, 56, 42], unbeaten since 1982, and about 2.372 for exponents of unfeasible MM [122]. The exponents below 2.77 have only been reached at the expense of blowing up the size of MM by immense factors.
Theorem 1.3.
(Main Theorem 3.) One can solve Problems 0 and 0 at a Las Vegas randomized bit operation cost $\tilde{O}(bdm))$ for a polynomial $p$ of (1.1) normalized such that $\sum_{i=1}^{d}|p_{i}|=1$.
Random Root Model: The zeros of $p(x)$, for $p(x)=\det(xI-M)$ in Problems 2 and 2 and for $p(x)$ of (1.1) otherwise, are independent identically distributed (iid) random variables sampled under the uniform probability distribution on the disc $D(0,R)$ for a large real $R$ such that the value $\frac{R^{2}}{d}-1$ exceeds a positive constant.666(i) The assumption of the model is a little stronger than that of our Def. 22.1, supporting such a decrease in [78, 84]. (ii) Random coefficient models fails at the basic step of the shift of the variable $x$, which no known probability distribution withstands.
Theorem 1.4.
(Main Theorem 4.) Under the Random Root Model one can decrease by a factor of $m$ the upper bounds of Thms. 1.1-1.3 on the computational precision and bit operation complexity.
In view of Remarks 1.1 and 23.4 we can readily extend our Main Theorems to Polynomial Root Isolation.
2 Fast subdivision root-finding: overview and soft e/i tests
We support Main Theorems 1 – 4 with our accelerated subdivision root-finders. They are reduced to recursive application of soft exclusion/inclusion (e/i) tests, which decide whether a fixed complex disc or its small neighborhood contains roots. In Sec. 3 we outline our two approaches to devising e/i tests for a black box polynomial and the unit disc $\mathcal{D}=D(0,1)$. We expand these two outlines in Part II, made up of Secs. 610, and Part III, made up of Secs. 1115, respectively. Our root-finders of Parts II and III only approximate (rather than isolate) the roots and support complexity estimates that are inferior to those of our Main Theorems by a factor of $b$. In Part IV, made up of Secs. 1621 and outlined in Sec. 4, we fix both deficiencies, thus proving the Main Theorems. In Sec. 5.5 we further comment about the contents of our paper, including its Part V, made up of Secs. 2225.
2.1 Subdivision root-finders and reduction to e/i tests
We depart from popular subdivision root-finders, traced back to [120, 39, 100, 82]. They extend root-finding in a line segment by means of bisection to polynomial root-finding in a square on the complex plane and have been extensively used in Computational Geometry under the name of Quad-tree Construction. Next we outline them.
(A) First consider root-finding in a square on the complex plane. Call that square suspect, recursively divide it and every other suspect square into four congruent sub-squares, and to each of them apply exclusion/inclusion (e/i) test to decide whether the square contains a root or does not. If it does, call the square suspect and process it in the next iteration; otherwise discard it (see Fig. 1a). In Def. 2.1 we generalize e/i tests to $\ell$-tests, which decide whether a disc in the complex plane contains $\ell$ or fewer roots provided that it contains at most $m$ roots. E/i test is a 1-tests, and whenever we need $\ell$-tests for $\ell>1$ (e.g., in Sec. 16 and Part IV), we reduce them to e/i tests (see Sec. 18 and 13.4).
(a) Four roots marked by asterisks lie in suspect sub-squares; the other (empty) sub-squares are discarded.
(b) Both exclusion and $\sqrt{2}$-soft inclusion criteria hold for e/i test applied to the disc bounded by internal circle.
Figure 1:
Observation 2.1.
A subdivision iteration applied to a square with $m$ roots in its $\beta$-neighborhood (cf. Def. 1.1 and Remark 2.1) (i) decreases by twice the diameter of a suspect square and (ii) processes at most $4m$ suspect squares, (iii) whose centers approximate every root within the distance of one half of their diameter.
Proof.
Readily verify claim (i). Claim (ii) holds because a root can lie in at most four suspect squares, and even in at most two unless it is the shared vertex of the four. Claim (iii) holds because a subdivision iteration never discards a square containing a root. ∎
Corollary 2.1.
Let subdivision iterations be applied to a square $S$ with side length $\Delta$ and diameter $\Delta^{\prime}:=\Delta\sqrt{2}$ containing precisely $m$ roots. Then (i) the $k$-th iteration outputs at most $4m$ suspect squares whose centers approximate within $\epsilon:=\Delta^{\prime}/2^{k+1}$ all the $m$ roots lying in $S$, so that $k+1\leq\lceil\log_{2}(\frac{\Delta^{\prime}}{\epsilon})\rceil$, and (ii) the first $k$ iterations together involve at most $4mk$ e/i tests.
(B) Seeking roots in any fixed complex domain $\mathbb{D}$, we apply subdivision iterations to its minimal covering square $S$ with diameter $\Delta^{\prime}$ and while performing an iteration also discard all squares not intersecting the domain $\mathbb{D}$, even if they contain roots. The following proposition bounds the overall number of e/i tests performed in $k$ iterations.
Proposition 2.1.
Let $k$ subdivision iterations be applied to a complex domain $\mathbb{D}$, let precisely $m_{i}$ roots lie in the $\Delta_{i}^{\prime}$-neighborhood of $\mathbb{D}$ for $\Delta^{\prime}_{i}$ denoting the diameter of a suspect square at the $i$-th subdivision iteration, and let precisely $m_{i}$ e/i tests be performed at the $i$-thiteration, so that
$m_{i}\leq 4^{i+1},~{}i=0,1,\dots,k-1.$ (2.1)
Then
(i) the $(\Delta_{h}^{\prime})$-neighborhood of $\mathbb{D}$ covers the union of all suspect squares processed in the $h$-th subdivision iteration, while $\Delta_{h}^{\prime}=\Delta_{0}^{\prime}/2^{h}$ fast converges to 0 as $h$ increases; in particular, for any fixed $\beta$ of Def. 1.1 we have $\Delta^{\prime}_{h}\leq 0.5\beta\Delta^{\prime}_{0}$ already for $h=O(1)$.
(ii) The first $h$ iterations involve at most $\frac{4^{h+1}}{3}$ e/i tests, that is, $O(1)$ tests if $h=O(1)$.
Proof.
We only prove claim (ii). Eqn. (2.1) implies that $\sum_{i=0}^{h-1}m_{i}\leq\sum_{i=0}^{h-1}4^{i+1}=\frac{4^{h+1}}{3}$. ∎
Remark 2.1.
Every root is defined up to within $\epsilon$, and so at the $h$-th subdivision iteration we decide whether we should discard a square or call it suspect and keep subdividing it depending on the roots that lie in the $(\Delta_{h}^{\prime}+\epsilon)$-neighborhood of $\mathbb{D}$. Def. 1.1 for $\Delta:=\Delta_{0}^{\prime}$ implies that $\Delta^{\prime}_{h}\epsilon\leq\beta\Delta$ already for some $h=O(1)$, and so, the outputs of all e/i tests, except for $h=O(1)$ initial ones, depend only on the roots that lie in the $\beta\Delta$-neighborhood of $\mathbb{D}$.
Remark 2.2.
[Real and near-real root-finding.] In the case of Problem 0 of Example 1.2, generalize bisection root-finding as follows: (i) recursively partition (subdivide) the real line segment defining every suspect rectangle into two halves and accordingly subdivide the rectangle into two congruent rectangles; to each of them apply e/i test, and then discard a rectangle if it contains no roots but call it suspect and subdivide it otherwise, (ii) stop where a rectangle has diameter at most $2\epsilon$, so that its center approximates within $\epsilon$ all roots lying in it.
2.2 Subdivision iterations with soft e/i tests
Instead of testing whether a suspect square $S$ contains some roots, Henrici in [39] tested whether the minimal disc $D(c,\rho)$ covering $S$ does this. Furthermore, in the presence of rounding errors no e/i test can decide whether a root lies in or outside a disc if it actually lies on or very close to its boundary circle. We address both problems by fixing $\sigma>1$ and applying soft e/i tests, which are a special case for $\ell=1$ of the following more general soft $\ell$-tests. [^7] Besides 1-tests (that is, e/i tests) we only use $\ell$-tests for sub-discs of discs containing at most $\ell$ roots and well-isolated from the external roots.
Definition 2.1.
For $p(x)$ of (1.1), $\sigma>1$, and an integer $\ell$, $1\leq\ell\leq d$, a $\sigma$-soft $\ell$-test, or just $\ell$-test for short (in particular an e/i test turns into a 1-test), either outputs 1 and stops if it detects that $r_{d-\ell+1}\leq\sigma$, that is, #$\#(D(0,\sigma))\geq\ell$, or outputs 0 and stops if it detects that $r_{d-\ell+1}>1$, that is, [^8] Bounds $r_{d-\ell+1}\leq\sigma$ and $r_{d-\ell+1}>1$ can hold simultaneously (cf. Fig. 1b), but as soon as an $\ell$-test verifies any of them, it stops without checking if the other bound also holds. #$\#(D(0,1))<\ell$. An $\ell$-test applied to the polynomial $t(y)$ of (1.4) is said to be an $\ell$-test for the disc $D(c,\rho)$ and/or the circle $C(c,\rho)$ and also to be an $\ell$-testc,ρ.
Observation 2.2.
Let $\gamma(\sigma)$ denote the maximal number of $\sigma$-covers at a fixed subdivision step that share a complex point, e.g., a root. Then $\gamma(\sigma)\leq 4$ if $\sigma<\sqrt{2}$ and $\gamma(\sigma)$ is in $O(1)$ if so is $\sigma$.
Corollary 2.2.
Cor. 2.1 and the observation of Remark 2.1 still hold if e/i tests are replaced with $\sigma$-soft e/i tests for $\sigma<\sqrt{2}$, while a subdivision root-finder using $\sigma$-soft e/i tests involves $O(km)$ such tests if $\sigma$ is a constant.
3 Fast soft e/i tests
3.1 Background: Newton’s Inverse Ratio
We reduce e/i tests of subdivision root-finders to multipoint evaluation of the ratio
${\rm NIR}(x)=\frac{p^{\prime}(x)}{p(x)}=\sum_{j=1}^{d}\frac{1}{x-z_{j}},$ (3.1)
said to be Newton’s inverse ratio. To prove the latter well-known equation, differentiate $p(x)=p_{d}\prod_{j=1}^{d}(x-z_{j})$.
The ops involved in our evaluation of NIR$(x)$ are more numerous than the other ops involved in our root-finders, even in the case where $p(x)=x^{d}$.
By following [59] we fix a complex $\delta$, for $|\delta|$ small in context, and approximate $p^{\prime}(x)$ with the first order divided difference $\frac{p(x)-p(x-\delta)}{\delta}$, equal to $p^{\prime}(y)$ for some $y\in[x-\delta,x]$, by virtue of Taylor-Lagrange’s formula. Hence $y\mapsto x$ and $p^{\prime}(y)\mapsto p^{\prime}(x)$ as $\delta\mapsto 0$, and
${\rm NIR}(x)\approx\frac{p(x)-p(x-\delta)}{p(x)\delta}=\frac{p^{\prime}(y)}{p(x)}~{}{\rm for~{}some}~{}y\in[x-\delta,x].$ (3.2)
3.2 Background: error detection and correction
A randomized root-finder can lose some roots, although with a low probability, but we can detect such a loss at the end of root-finding process, simply by observing that among the $m\leq d$ roots lying in an input disc $D$ only $m-w$ tame roots, say, $z_{w+1},z_{w+1},\dots,z_{m}$ have been closely approximated, [^9] We call a complex $c$ a tame root for an error tolerance $\epsilon$ if $c$ lies in an isolated disc $D(c,\epsilon)$, whose index #$\#(D(c,\epsilon))$ we can compute fast by applying Thm. 12.1 or the algorithm in Sec. 17.5. while $w>0$ wild roots, say, $z_{1},z_{2},\dots,z_{w}$ remain at large. We can compute the coefficients of the factor $g(x):=\frac{p(x)}{f(x)}$ for $f(x):=\prod_{j=w+1}^{m}(x-z_{j})$, say, by applying evaluation-interpolation techniques, then apply to $g(x)$ a black box root-finder, and repeat this computation recursively to approximate the $w$ wild roots. [^10] By this recipe we detect any output error at the very end of computations in any root-finder that can only lose roots but cannot wrongly generate them. Our subdivision root-finders belong to this class, but with them we can detect the loss of a root even earlier – whenever in a subdivision process applied to a square with $m$ roots we detect that $m$ exceeds the sum of the indices of all suspect squares, that is, the overall number of roots contained in them. We readily bound the complexity and error probability of this recursive process (although we will not use these estimates in the following):
Proposition 3.1.
Let the above recipe for error detection and correction be applied to the output of a randomized root-finder performing at the cost $\alpha$ (under a fixed model), let it be correct with a probability $1/\beta$ for $\beta>1$ and let the algorithm be applied recursively, each time with error detection at the cost $\delta$. Then in $v$ recursive applications, the output is certified to be correct at the cost of less than $(\alpha+\delta)v$ with a probability at least $1-1/\beta^{v}$.
Explicit deflation can blow up the coefficient length and destroy sparsity unless the degree $w$ is sufficiently small, but we avoid these mishaps by combining a fixed black box polynomial root-finder with implicit deflation (cf. example (iii) in Sec. 1.4.1), that is, by applying that root-finder to the quotient polynomial $q(x)$ without computing its coefficients – based on the following expression, implied by (3.1):
$\frac{q^{\prime}(x)}{q(x)}=\frac{p^{\prime}(x)}{p(x)}-\frac{t^{\prime}(x)}{t(x)}=\frac{p^{\prime}(x)}{p(x)}-\sum_{j=w+1}^{m}\frac{1}{x-z_{j}}.$ (3.3)
Remark 3.1.
One can accelerate low precision numerical approximation of the sum $\sum_{j=w+1}^{m}\frac{1}{x-z_{j}}$ simultaneously at many points $x$ by applying Fast Multipole Method (FMM) by Greengard and Rokhlin [37, 20, 24] (see more in Sec. 24.3).
3.3 Background: root-lifting
Our e/i tests of Part II involve $h$-lifting of roots:
$p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{h}(x):=\prod_{g=0}^{h-1}p_{0}(\zeta_{h}^{g}x)=(-1)^{d}\prod_{j=1}^{d}(x-z_{j}^{h}),~{}\zeta_{h}:=\exp\Big{(}\frac{2\pi\sqrt{-1}}{h}\Big{)},$ (3.4)
for $h=1,2,\dots$ and $t_{2^{k}}(x)=p_{k}(x)$ for all $k$. $h$-lifting maps roots to their $h$th powers; this keeps the unit disc $D(0,1)$ and its boundary circle $C(0,1)$ intact but moves the external roots away from the circle, thus strengthening its isolation from those roots.
Remark 3.2.
For $h$ being a power of 2, the $h$-lifting generalizes the DLG classical root-squaring iterations of Dandelin 1826, Lobachevsky 1834, and Gräffe 1837 (cf. [38]):
$p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{2^{i}}(x^{2}):=(-1)^{d}p_{2^{i-1}}(x)p_{2^{i-1}}(-x)=\prod_{j=1}^{d}(x-z_{j}^{2^{i}}),~{}i=1,2,\dots.$ (3.5)
Polynomial root-finding in the 19-th and well into the 20-th century relied on iterations (3.5) combined with Vieta’s formulas, which link the polynomial coefficients $p_{0},p_{1},\dots,p_{d}$ to the elementary symmetric polynomials of the roots. The combination enabled efficient approximation of the minimal and maximal distance of the roots from the origin and, due to the map (1.4), from any complex point (see [69, 94, 96] for details). The approach became obsolete because of severe problems of numerical stability of the DLG iterations, but they disappeared in the case of root-finders that work for black box polynomials. Extensive tests with standard test polynomials have consistently demonstrated efficiency of the particular implementation of this approach in [96], but even the straightforward implementation of iterations (3.5) promises to be highly efficient as well. The paper [36] explores direct root-squaring for NIR of a black box polynomial:
${\rm NIR}_{p_{h+1}}(x)=\frac{1}{2\sqrt{x}}\left({\rm NIR}_{p_{h}}(\sqrt{x}){\rm NIR}_{p_{h}}(-\sqrt{x})\right),~{}x\neq 0,~{}~{}h=0,1,\dots.$
3.4 Soft e/i test with root-lifting: generic algorithm
In Sec. 7 we elaborate upon the following generic soft e/i test applied to the unit disc $D(0,1)$.
Algorithm 3.1.
A generic soft e/i test.
INPUT: $\sigma>1$, an integer $d>0$, and a $d$-th degree black box polynomial $p(x)$ given by an oracle for its evaluation.
INITIALIZATION: Fix two positive integers $h$ and $q$ and $q$ points $x_{1},\dots,x_{q}$ on the circle $C(0,1)$.
COMPUTATIONS: By extending (3.2) to $p_{h}(x)$ of (3.4) (see (6.3)) approximate the values
$v_{g}:=|{\rm NIR}_{p_{h}}(x_{g})|=\Big{|}\frac{p^{\prime}_{h}(x_{g})}{p_{h}(x_{g})}\Big{|},~{}g=1,\dots,q.$ (3.6)
If $1+d/v_{g}\leq\sigma^{h}$ for some $g\leq q$, then certify that #$(D(0,\sigma))>0$ ($\sigma$-soft inclusion). Otherwise, claim that #$(D(0,1))=0$ (exclusion).
Remark 3.3.
(6.3) and (3.4) together express $v_{g}$ for all $g$ through $2hq$ values of $p(x)$, by-palink to the values of $p_{h}(s)$ helps us o prove correctness of the algorithm.
3.5 Soft e/i test with root-lifting: correctness of inclusion claim
Eqn. (3.1) implies two useful corollaries:
Corollary 3.1.
Map (1.4) scales NIR by $\rho$.
Proof.
Eqn. (3.1) implies that NIR$(x-c)=$NIR$(x)$ and that NIR$(x/a)=a$NIR$(x)$. ∎
Corollary 3.2.
It holds that $r_{d}(c)\leq d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}$, that is, #$\#\Big{(}D\Big{(}c,d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}\Big{)}\Big{)}>0$, for any complex $c$.
Proof.
Let $x=c$ in Eqn. (3.1). If $|c-z_{j}|<\rho$ for all $j$, then
$\Big{|}\frac{p^{\prime}(c)}{p(c)}\Big{|}=\Big{|}\sum_{j=1}^{d}\frac{1}{c-z_{j}}\Big{|}\leq\sum_{j=1}^{d}\frac{1}{|c-z_{j}|}<\frac{d}{\rho},~{}{\rm and~{}so}~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}>\frac{\rho}{d}.$
We use Cor. 3.1 later but next apply Cor. 3.2 to $p_{h}$ for a sufficiently large $h$ of order $\log(d)$ to prove correctness of the inclusion claim of Alg. 3.1.
Lemma 3.1.
Let $\rho:=d\Big{|}\frac{p_{h}(x)}{p^{\prime}_{h}(x)}\Big{|}$ and $\sigma:=(|x|+\rho)^{1/h}$ for a positive integer $h$. Then #$(D(0,\sigma))>0$.
Proof.
Apply Cor. 3.2 to the polynomial $p_{h}(x)$ and obtain that the equation $\rho=d\Big{|}\frac{p_{h}(x_{)}}{p^{\prime}_{h}(x)}\Big{|}$ implies that the disc $D(x,\rho)$ contains a zero of $p_{h}(x)$. Hence the disc $D(0,|x|+\rho)$ contains a zero of $p_{h}(x)$ by virtue of triangle inequality. Therefore, the disc $D(0,\sigma)$ contains a zero of $p(x)$ for $\sigma=(|x|+\rho)^{1/h}$. ∎
Apply this lemma for $x=x_{i}$, $|x_{i}|=1$, $\rho=d/v_{i}$, and $v_{i}=|{\rm NIR}_{p_{h}}(x_{i})|$ (cf. Alg. 3.1). Deduce that #$(D(0,\sigma))>0$ for $\sigma^{h}=1+d/v_{i}$, that is, Alg. 3.1 is correct when it claims $\sigma$-soft inclusion.
Observation 3.1.
$\sigma$-soft inclusion test by means of Alg. 3.1 is correct where
$d\geq 2,~{}x_{1}=1,~{}\frac{1}{|{\rm NIR}(1)|}=4\sqrt{d}-\frac{1}{d},~{}h=40\log_{2}(d),~{}{\rm and}~{}\sigma<1.1.$
Proof.
Indeed, in this case
$1+\rho=1+\frac{d}{|{\rm NIR}(1)|}=4d\sqrt{d}=2^{2+1.5\log_{2}(d)}.$
Apply Lemma 3.1 and obtain that #$(D(0,\sigma))>0$ for $\sigma=(1+\rho)^{1/h}=2^{\nu}~{}{\rm and}~{}\nu=\frac{2+1.5\log_{2}(d)}{40\log_{2}(d)}$.
Deduce that $\nu\leq\frac{7}{80}$ because $\frac{2}{40\log_{2}(d)}\leq 0.05$ for $d\geq 2$ and because $\frac{2+1.5\log_{2}(d)}{40\log_{2}(d)}<\frac{3}{80}$.
Hence $\sigma\leq 2^{\frac{7}{80}}=1.062527\dots<1.1$. ∎
3.6 Soft e/i test with root-lifting: further analysis (outline)
We prove (i) correctness of the exclusion claim of Alg. 3.1 separately for two specifications, both strengthening isolation of the unit circle $C(0,1)$ from the roots by means of root-lifting (3.4).
(a) In Sec. 7.3 we choose any integer $q$ exceeding $d$ and then prove correctness of exclusion claim, provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(2\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity. This bound is compatible with the bound of Observation 3.1 because $4\sqrt{d}-\frac{1}{d}>3\sqrt{d}$ for $d\geq 2$. We also extend this e/i test by allowing any $q>m$ (rather than $q>d$) provided that the unit disc contains $m$ roots and $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(3\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity.
(b) In Sec. 7.4 we sample $q=O(\log(d))$ iid random values of $x_{i}$ under the uniform probability distribution on the circle $C(0,1)$ and then prove that the exclusion claim is correct whp under the Random Root Model if $|{{\rm NIR}_{p_{h}}(x_{i})}|\leq 1/d^{a}$ for a sufficiently large constant $a$.
(ii) Counting query locations. Eqn. (3.4) reduces approximation of $p_{h}(x)$ at $x=c$ to approximation of $p(x)$ at all $h$th roots of $c$. We can fix an integer $h>1$ and apply (3.4) or fix $h>1$ being a power of two and apply (3.4). In both cases Alg. 3.1 evaluates polynomial $p(x)$ either at $2jh$ points for $j\leq q$ when it claims inclusion or at $2qh$ points when it claims exclusion. We always choose $h=O(\log(d))$ and then ignore a factor of $h$ in our $\tilde{O}$ estimates.
In our soft e/i test based on Alg. 3.1 we evaluate polynomial $p(x)$ at $\tilde{O}(m)$ points and then combine these estimates with Cor. 2.2, and bound by $\tilde{O}(bm^{2})$ the overall number of evaluation queries in the resulting root-finders.
Under the Random Root Model of Sec. 1.5 the roots are reasonably well isolated whp from the unit circle $C(0,1)$. We strengthen this isolation by applying $O(\log(d))$ root-lifting steps of (3.4) and then whp bound the overall number of queries by $\tilde{O}(bm)$.
(iii) Computational precision and bit-operation complexity. We deduce the same bound $O(b\log(d))$ of [59] on computational precision of our root-finders; our derivation is much simpler because unlike [59] we only use the first order derivatives and divided differences.
We combine this estimate and our count of evaluation queries in our root-finders with (i) Storjohann’s algorithm for Las Vegas computation of the determinant [109] to obtain record fast solution of Problems 2 and $2^{*}$ of matrix eigenvalue approximation (see alternative extension to eigenproblem in part (C) of Sec. 5.4) and (ii) the complexity estimates of Guillaume Moroz [64] for his fast multipoint evaluation of a polynomial represented in monomial basis to arrive at the near-optimal complexity bounds of our Main Theorems 1 – 4 within a factor of $b$, to be removed in Part IV.
3.7 Cauchy-based soft e/i tests: outline
Next we outline our alternative soft e/i test of Part III. Let #$\#(\mathcal{D})$ denote the number of roots lying in a complex domain $\mathcal{D}$ with the boundary $\Gamma$ and recall from [1] that
[…]
that is, #$\#(\mathcal{D})$ is the average of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ on the boundary $\Gamma$.
Having an approximation of integral (3.7) within less than 1/2, we can round it to a closest integer and obtain $s_{0}=\#(D(0,1))$. This root-counter defines $\ell$-tests for all $\ell\leq d$ and in particular an e/i test, for $\ell=1$.
In Part III we approximate the integral (3.7) with finite sums of the values of $\frac{x}{q}$$(x)$ at the scaled $q$th roots of unity. Namely, let $\zeta:=\zeta_{q}$ denote a primitive $q$th root of unity (cf. (3.4)) and devise randomized $\sigma$-soft e/i tests as follows.
Algorithm 3.2.
[Cauchy e/i test].
(i) Fix a constant $\gamma>1$, say, $\gamma=2$, and choose a real $\rho$ at random in a fixed range in $(1,\sigma)$.
(ii) Fix a sufficiently large integer $q$ of order $m\log(d)$.
[…]
(iv) Unless $\bar{s}_{0}=0$ certify inclusion.
Otherwise, repeat the test $v$ times for a fixed integer $v>0$ (at the cost of the evaluation of $p(x)$ at $O(mv\log(d))$ points). Unless inclusion is certified in these tests, claim exclusion (correctly with a probability at least $1-1/\gamma^{v}$).
4 Faster subdivision with compression: outline
In Part IV we accelerate the algorithms of both Parts II and III by a factor of $b/\log(b)$ to ensure pairwise isolation of clusters of roots and eigenvalues and to prove Main Theorems 1 – 4; we prove the first three of them in two ways – by devising two distinct supporting root-finders based on Parts II and III, respectively. Next we outline that acceleration.
(i) Component tree and compact components. For fixed integer $m$, positive $\epsilon$, and $\sigma$, $1<\sigma<\sqrt{2}$, apply subdivision process to Problem 1 and at every subdivision step partition the union of the suspect squares into connected components.
Figure 2: Roots (asterisks) define compact components. A subdivision step halves their diameters and accordingly strengthens isolation of their minimal covering squares and discs.
Represent all partitions of components with a component tree whose at most $m$ leaves represent components covered by isolated $\epsilon$-discs. The other (at most $m-1$) vertices of the tree represent components-parents, each broken by a subdivision step into at least two components-children. The initial suspect square is the root of the tree. The tree has at most $22m-$ 2 edges, connecting parents with their children and representing partition of the components but not showing modification of components between partitions.
Every component contains $O(m)$ suspect squares whose diameter is halved at every subdivision step. Hence in $O(b)$ subdivision iterations, for $b:=\log_{2}(2\Delta^{\prime}/\epsilon)$ and the diameter $\Delta^{\prime}$ of the initial suspect square, every component has diameter below $\epsilon/2$ and hence lies in an $\epsilon$-disc.
(ii) Faster subdivision with compression. Call a component $\mathcal{C}$ compact if all (at most four) its squares share a vertex; subdivision steps fast decrease the diameter of $\mathcal{C}$ but never decrease its distance from external roots (see Fig. 2).
Subdivision becomes inefficient where a component stays unbroken and compact in many subdivision steps (see Fig. 3), and we replace them with its compression with no loss of roots, of which the component very soon becomes strongly isolated.
The compression outputs either an $\epsilon$-disc, and then we stop, or a rigid disc $D$, that is, minimal up to a fixed constant factor, and then we subdivide a minimal superscribing square $S$ of $D$. Since the disc $D$ is rigid, the iterations very soon output more than one component or a sub-domain of an $\epsilon$-disc.
Figure 3: Five blue marked centers of suspect squares converge to a green-marked root cluster with linear rate; two red marked Schröder’s iterates of (17.1) converge with quadratic rate.
This way we solve Problem 1 by applying at most $22m-$ 2 compression steps and in addition $O(m)$ soft e/i tests, as we deduced in Sec. 19.
(iii) Compression of components. We compress a compact component $\mathcal{C}$ into a disc $D=D(c,\rho)$ by computing:
(a) the number of roots in a component,
(b) a point $c$ in or near the minimal disc covering the root set of $\mathcal{C}$ (hereafter the MCD of $\mathcal{C}$) and
(c) a reasonably close upper bound $\rho$ on the maximal distance from $c$ to the roots in $\mathcal{C}$.
In Secs. 17.4 and 17.5 we readily perform steps (a) and (b) at a low computational cost because subdivision iterations fast strengthen the isolation of the component from external roots.
We begin step (c) with a range $[\rho_{-},\rho_{+}]$ for the bound $\rho$, $0<1/2^{b}=\rho_{-}<\rho_{+}$. We can approximate $\rho$ by applying bisection to this range but proceed faster by applying bisection of the logarithms, said to be BoL, to the range $[\log_{2}(\rho_{-}),\log_{2}(\rho_{+})]=[b,\log_{2}(\rho_{+})]$, for $\log_{2}(\rho_{+})\leq 0$.
We apply fast soft $m$-tests (see part (A) of Sec. 2.1) to define a pivot step of our accelerated bisection, with which we ultimately reduce compression to $O(m\log(b))$ e/i tests. Their computational cost dominates the overall cost of the solution of Problems 0,0, 1, and 1, and this enables us to complete the proof of our Main Theorems.
Remark 4.1.
Many “natural" directions to proving our Main Theorems have led us into dead end. E.g., we tried to use a tree whose vertices represented components made up of non-adjacent suspect squares; then we readily bounded by $12m-$ 1 their overall number at all partition stages but could not count them so readily between these stages until we decided to count all suspect squares. See more examples of this kind in the next remark and Remark 7.1.
Remark 4.2.
By using some of the advanced techniques of Louis and Vempala in [59], we readily bound the precision of computing in our root-finders. This enables us to deduce a priori estimates for the bit operation cost of our subdivision root-finders versus inferior known a posteriori estimates (see, e.g., [97, 17]). In particular, the bit operation cost estimates of the Main Theorem of [17] contain the term
$\log(1/{\rm Discr}(p(x))),$
with “Discr" standing for “Discriminant", such that
$\log(1/{\rm Discr}(p(x/2^{b})))=(d+1)db\log(1/{\rm Discr}(p(x))).$
Subdivision iterations of the known root-finders, in particular, of those of [17], scale the variable $x$ order of $m$ times by factors of $2^{b}$. Hence the bit-complexity specified in the Main Theorem of [17] has order of $mbd^{2}$, which exceeds our bound and that of [78, 84] by a factor of $d$ (see more comments in Sec. 10.4).
5 Our novelties, related works, implementation, and further progress
5.1 Our novelties (a short list)
Root-finding for polynomial given in the monomial basis, with their coefficients, has been studied for four millennia and very intensively in the last decades, but we had to introduce a number of nontrivial novel techniques (i) to handle efficiently black box polynomials and (ii) to accelerate the known subdivision algorithms to a near-optimal level.
Some of our techniques such as randomization, numerical stabilization of root-squaring, compression of a disc without losing roots, bisection of logarithms, approximation of root radii and near-real roots, and even our simple error detection recipe can be of independent interest.
5.2 Related works
From the pioneering paper [59] we adopt the model of computation, extension of a black box polynomial root-finder to approximation of matrix eigenvalues, and the approach to estimation of the computational precision of root-finders, but otherwise we use distinct approach and techniques.
Estimating the bit operation complexity of our root-finders we use [106, Eqn. (9.4)] and [64, Thm. 2] for multipoint polynomial evaluation. [^11] The ingenious algorithm of Moroz supporting his [64, Thm. 2] departs from the simple observation that a polynomial $p(x-c)\mod(x-c)^{\ell-1}$ approximates $p(x)$ in the disc $D(c,\rho)$ within $O(\rho^{\ell+1})$ for small $\rho$. To approximate $p(c)$ for any $c\in D(0,1)$ Moroz used $\tilde{O}((\ell+\tau)d)$ bit operations to compute a set of complex $c_{i}$ and positive $\rho_{i}$ for $i=1,\dots,K$ and $K=O(d/\ell)$ such that $D(0,1)\subseteq\cup_{i=1}^{K}D(c_{i},\rho_{i})$; then he ensured approximation to $p(c)$ for every point $c\in D(0,1)$ within $\epsilon:=3\sum_{{}_{i}=0}^{d}|p_{i}|/2^{\ell}$ by means of approximation within $\epsilon/2$ of the lower degree polynomial $p(x-c_{i})\mod(x-c)^{\ell+1}$ for an appropriate $i$. At this point, application of [52, Alg. 5.3 and Thm. 3.9] yields new record fast algorithm for multipoint approximation of a polynomial given with its coefficients. Ignoring Kirinnis’s paper [52], Moroz cited a rediscovery of his result by Kobel and Sagraloff, made two decades later.
Instead of applying high order Newton’s iterations of [59], we accelerate classical subdivision root-finders. The closest earlier variations in [100, 82, 16, 17] use mach slower e/i tests, involving coefficients of $p$. Two our fast novel e/i tests – based on root-lifting and approximation of Cauchy integrals, respectively, work for a black box polynomial.
The algorithms of [100, 82] dramatically accelerate convergence of subdivision process by means of applying Newton’s or Schröder’s iterations (see Fig. 3). They converge with superlinear rate to the convex hull of all roots lying in $C$. The diameter of the minimal covering disc, however, can be too large, and then the papers [100, 82] apply subdivision iterations again. The resulting root-finders recursively intertwine subdivision iterations with Newton’s or Schröder’s, and this complicates the complexity estimates for a root-finder.
[100, 82] have only estimated the number of ops but not of the bit operations involved. [16, 17] estimated the bit operation complexity in terms of the unknown minimal distance between the unknown roots. Specified in terms of $b$ and $d$, the estimates turned out to be inferior to ours by a factor of $d$ (see Remark 4.2 and Sec. 10.4).
Instead of Newton’s or Schröder’s iterations, one can apply the algorithm of [106]. It first splits $p(x)$ into the product of two polynomials $f(x)$ and $g(x)$ whose zero sets lie in and outside the minimal disc covering the component $C$, respectively. Then it recursively splits the polynomial $p(x)$ into factors $f=f(x)$ and $g=g(x)$. The resulting algorithm of [106] is slower than ours by a factor of $d$. It has been accelerated to a near-optimal level in [78, 84], although that root-finder is much involved and extensively uses the coefficients of $p$, like the root-finders of [106, 100, 82, 16, 17].
[64, 44] presented efficient root-finders for polynomials having only well-conditioned zeros, but not for worst case polynomials $p(x)$. Furthermore, the algorithms of [64, 44] extensively use the coefficients of $p(x)$.
5.3 Implementation: State of the Art (briefly)
The near-optimal but quite involved algorithm of [78, 84] has never been implemented, while the slower algorithm of [106] has been implemented by Xavier Gourdon in [34] for the Magma and PARI/GP computer algebra systems. By now it is superseded by implementations of various non-optimal complex polynomial root-finders such as MATLAB “roots”, Maple “solve”, CCluster, and MPSolve; the latter one relies on Ehrlich’s iterations and is user’s current choice for approximation of all $d$ complex roots (see some competitive implementations of Newton’s iterations in [115, 102, 68]).
Current user’s choice for real root-finding is the algorithm of [54], although it has already been superseded in [47] according to extensive tests for standard test polynomials.
We present algorithms that promise to compete for user’s choice. Already incorporation in 2020 of some of our techniques into the previous best implementation of subdivision root-finder [49] has enabled its noticeable acceleration (see [45]). The more advanced publicly available implementation in [48] “relies on novel ideas and techniques" from [93] (see [48, Sec. 1.2]). According to the extensive tests with standard test polynomials, the algorithm has performed dramatically faster than [49], has at least competed with user’s choice package MPSolve even for approximation of all $d$ zeros of a polynomial $p$ given with its coefficients, [^12] The computational cost of subdivision root-finders in a domain $\mathbb{D}$ as well as the root-finder of [78, 84] decreases at least proportionally to the number of roots in $\mathbb{D}$, while Ehrlich’s, Weierstrass’s, and a number of other functional iterations approximate the zeros of $p$ in a domain $\mathbb{D}$ almost as slow as all its $d$ complex roots. Thus already implementation in [49] of subdivision iterations, slightly outperformed MPSolve for root-finding in a domain $\mathbb{D}$ containing a small number of roots although is by far inferior for the approximation of all $d$ roots. and superseded MPSolve more and more significantly as the degree $d$ of $p$ grew large.
There is still a large room for further improvement because the root-finder of [48] inherits some earlier sub-algorithms from [16, 49], which are slower by order of magnitude than our current ones (see the next subsection).
Application of our near-optimal subdivision root-finder in Remark 2.2 to approximation of real and near-real roots promises to be at least competitive with user’s current choice [54] and its improvement in [47]. It can be applied to a black box polynomial, counters the impact of input and rounding errors, and can be extended to approximation of real and near-real matrix eigenvalues.
5.4 Further progress and promises
(A) Acceleration of e/i tests. E/i tests are basic ingredient of subdivision root-finders. [45, 48] used the Cauchy-based e/i tests of our Part III and their heuristic acceleration. We can devise and test similar heuristic acceleration of the formally supported e/i tests of our Part II, based on root-lifting. Namely, we can merely apply Alg. 7.2 with an integer $q$ decreased dynamically from $m+1$, as long as the exclusion claim remains empirically correct.
(B) Acceleration of the compression of a disc. Assuming that the unit disc $D(0,1)$ contains $m$ roots and is isolated from the $d-m$ external roots, we compress that disc towards an $\epsilon$-disc or a rigid disc, in both cases losing no roots. Algorithmic extras 1 (for compression of a disc with a small number of roots) and 2 of Sec. 21 can help strengthen this approach.
Like us, [48] reduces compression to (a) non-costly root-counting in an isolated disc, (b) computing a point $c$ lying in or near the MCD, that is, the minimal covering disc  [^13] One may use the convex hull instead of the minimal covering disc. of the root set in $D(0,1)$, and (c) approximation of the $m$th smallest root radius from $c$.
At stage (b) [48] computes a close approximation of the sum $s_{1}$ of the $m$ roots lying in the disc $D(0,1)$ and then outputs $c=\frac{s_{1}}{m}$. In Sec. 17 we greatly accelerate this algorithms in the case where the radius of the output disc is small.
At stage (c) [48] applies the algorithm of our Sec. 18, which performs a constant but fairly large number of e/i tests in a disc. One may test this approach versus our alternative of Sec. 13.4 using a single e/i test for an annulus and versus the combination of the Las Vegas randomized algorithm of Sec. 13 with BoL of Sec. 16.
Finally, implementation of compression in [45, 48] is slowed down at the stage of verification based on Pellet’s theorem, but we can drop this verification because we readily detect highly unlikely output errors of our root-finders at a dominated cost (see Sec. 3.2).
(C) Fast approximation of the eigenvalues of a matrix. Our root-finders are essentially reduced to the evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$, and we can extend them to approximation of the eigenvalues of a matrix or a matrix polynomial by following [59] or based on the following alternative expressions:
$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}((xI-M)^{-1})}$
where $M$ is a matrix, while $p(x)=\det(xI-M)$ (cf. (3.1)) and
$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}(M^{-1}(x)M^{\prime}(x))}$
where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [13, Eqn. (5)]. [^14] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and of its determinant.
These expressions reduce approximation of the eigenvalues of a matrix or a matrix polynomial to its inversion. One can use the matrix inversion algorithm and its bit operation cost estimates of [110] towards alternative derivation of our Main Theorem 2 for any input matrix or matrix polynomial, but the inversion can be greatly accelerated in the highly important case where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
Bini and Noferini in [13] proposed and explored such a reduction of approximation of all $d$ eigenvalues to root-finding by means of Ehrlich’s iterations and showed various benefits of this approach, but we can readily combine this reduction with subdivision iterations instead of Ehrlich’s and then obtain additional benefits: our algorithms (a) are much faster for approximation of a small number of eigenvalues that lie in any fixed domain and (b) isolate the eigenvalues instead of just approximating them.
(D) Acceleration of real and near-real root-finding. In Remark 2.2 we adjust our near-optimal subdivision iterations to real and near-real root-finding, which counters the impact of input and rounding errors.
(E) Acceleration by using fast multipoint polynomial evaluation. Mere incorporation of fast algorithms of [64, 44] should greatly accelerate the performance of [45, 48] in the case of root-finding for polynomials given with their coefficients.
5.5 Organization of the rest of the paper
In Parts II and III we devise and analyze our two e/i tests. We use Las Vegas randomization in both parts to bound computational precision and bit-operation cost. In Part II the same randomization helps us devise fast e/i tests. In Part III we support that step with additional randomization.
In Part II we further accelerate by a factor of $d$ e/i tests and overall root-finding under the Random Root Model.
We make our presentation in Part III largely independent of Part II by reusing some definitions, results, and arguments.
In Part IV we devise and analyze our compression algorithm and incorporate it into the algorithms of Parts II and III – to accelerate them by a factor of $b/\log(b)$ and to ensure pairwise isolation of the output roots and their clusters.
In Part V we cover (i) low bounds on the bit-operation complexity of polynomial factorization, root-finding, and root isolation, (ii) acceleration with FMM (cf. Remark 3.1) and initialization of root-finding by means of functional iterations, and (iii) some additional algorithms for root radii approximation.
PART II: Classical subdivision root-finder with novel e/i tests based on root-lifting
Organization of Part II: We devote the next section to background. In Sec. 7 we cover our e/i tests. In Secs. 8 and 10 we estimate computational precision of our root-finders and their bit operation cost, respectively. In Sec. 9 we extend our study of root-finding to approximation of matrix eigenvalues.
6 Background
6.1 Basic definitions
We extend the list of definitions used in the Introduction.
  • ${\rm NR}_{t}(x):=\frac{t(x)}{t^{\prime}(x)}$ and ${\rm NIR}_{t}(x):=\frac{t^{\prime}(x)}{t(x)}$ are Newton’s ratio and Newton’s inverse ratio, respectively, for a polynomial $t(x)$. We write NR$(x):={\rm NR}_{p}(x)$ and NIR$(x):={\rm NIR}_{p}(x)$ (cf. (3.1)).
  • Define squares, discs, circles (circumferences), and annuli on the complex plane: $S(c,\rho):=\{x:~{}|\Re(c-x)|\leq\rho$, $|\Im(c-x)|\leq\rho\}$$D(c,\rho):=\{x:~{}|x-c|\leq\rho\},$ $C(c,\rho):=\{x:~{}|x-c|=\rho\}$,  $A(c,\rho,\rho^{\prime}):=\{x:~{}\rho\leq|x-c|\leq\rho^{\prime}\}.$
  • $\Delta(\mathbb{S})$, $MCD(\mathbb{S})$, $X(\mathbb{S})$, and #$\#(\mathbb{S})$ denote the diameter, root set, minimal covering disc, and index (root set’s cardinality) of a set $\mathbb{S}$ on the complex plane, respectively.
  • A disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is $\theta$-isolated for $\theta\geq 1$ if $X(D(c,\rho))=X(D(c,\theta\rho))$, $X(C(c,\rho))=X(A(c,\rho/\theta,\rho\theta)$ or $X(S(c,\rho))=X(S(c,\theta\rho)),$ respectively.
    The isolation of a disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is the largest upper bound on such a value $\theta$ (see Fig. 4) and is denoted $i(D(c,\rho))$, $i(C(c,\rho))$, or $i(S(c,\rho))$, respectively.
    We say “isolated" for “$\theta$-isolated" if $\theta-1$ exceeds a positive constant.
    $\rho$$\rho\theta$$\rho\theta^{2}$c********
    Figure 4: The roots are marked by asterisks. The red circle $C(c,\rho\theta)$ and the disc $D(c,\rho\theta)$ have isolation $\theta$. The disc $D(c,\rho)$ has isolation $\theta^{2}$.
  • A disc $D$ has rigidity $\eta(\mathbb{S})=\Delta(X(D))/\Delta(D)\leq 1$. A disc is $\eta$-rigid for any $\eta\leq\eta(D)$.
    We say “rigid" for “$\eta$-rigid" if $\eta$ exceeds a positive constant.
  • $|u|:=||{\bf u}||_{1}:=\sum_{i=0}^{q-1}|u_{i}|$, $||{\bf u}||:=||{\bf u}||_{2}:=(\sum_{i=0}^{q-1}|u_{i}|^{2})^{\frac{1}{2}}$, and $|u|_{\rm max}:=||{\bf u}||_{\infty}:=\max_{i=0}^{q-1}|u_{i}|$ denote the 1-, 2-, and $\infty$-norm, respectively, of a polynomial $u(x)=\sum_{i=0}^{q-1}u_{i}x^{i}$ and its coefficient vector ${\bf u}=(u_{i})_{i=0}^{d-1}$, and then (see [35, Eqs. (2.2.5)–(2.2.7)])
    $||{\bf u}||_{\infty}\leq||{\bf u}||_{2}\leq||{\bf u}||_{1}\leq{\sqrt{q}}~{}||{\bf u}||_{2}\leq q||{\bf u}||_{\infty}.$ (6.1)
  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex point $c$ to the roots $y_{1},\dots,y_{d}$ of a $d$th degree polynomial $t(x)$. Write $r^{h}_{j}(c,t):=(r_{j}(c,t))^{h}$, $r_{j}(c):=r_{j}(c,p)$.
Observation 6.1.
Taylor’s shift, or translation, of the variable $x\mapsto y:=x-c$ and its scaling combined (see (1.4)) map the zeros $z_{j}$ of $p(x)$ into the zeros $y_{j}=\frac{z_{j}-c}{\rho}$ of $t(y)$, for $j=1,\dots,d$, and preserve the index and the isolation of the disc $D(c,\rho)$.
6.2 Reverse polynomial
Immediately observe the following properties of the reverse polynomial (1.5).
Proposition 6.1.
(i) The roots of $p_{\rm rev}(x)$ are the reciprocals of the roots of $p(x)$, and hence
$r_{j}(0,p)r_{d+1-j}(0,p_{\rm rev})=1~{}{\rm for}~{}j=1,\dots,d.$ (6.2)
(ii) The unit circle $C(0,1)$ has the same isolation for $p(x)$ and $p_{\rm rev}(x).~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
${\rm(iii)}~{}p^{\prime}_{\rm rev}(x)=dx^{d-1}p\Big{(}\frac{1}{x}\Big{)}-x^{d-2}p^{\prime}\Big{(}\frac{1}{x}\Big{)}~{}{\rm for}~{}x\neq 0,~{}p_{d}=p_{\rm rev}(0),~{}p_{d-1}=p^{\prime}_{\rm rev}(0).$
${\rm(iv)}~{}~{}~{}~{}\frac{p_{\rm rev}(x)}{p^{\prime}_{\rm rev}(x)}=\frac{1}{x}-\frac{{\rm NIR}(x)}{x^{2}},~{}x\neq 0,~{}{\rm NIR}(0)=\frac{p_{d-1}}{p_{d}}=-\sum_{j=1}^{d}z_{j}.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
[…]
6.3 Root-lifting with divided differences
[…]
${\rm NIR}_{p_{h}}(x)\approx\frac{p_{h}(x)-p_{h}(x-\delta)}{p_{h}(x)\delta}=\frac{p_{h}^{\prime}(y)}{p_{h}(x)}.$ (6.3)
Remark 6.1.
[…]
Observation 6.2.
(i) Root-lifting preserves the index #$\#(D(0,1))$ of the unit disc. (ii) $h$-lifting (achieved with root-squaring for $h=2^{k}$ and a positive integer $k$) maps the values $i(D(0,1))$, $\sigma(D(0,1))$, $\eta(D(0,1))$, and $r_{j}(0,p)$ for all $j$, that is, the isolation of the unit disc $D(0,1)$, its softness, its rigidity and root radii, to their $h$-th powers.
6.4 The distance to a closest root and certified inclusion
[63, Sec. 1.10] lists various bounds on $r_{d}$ and $r_{1}$ in terms of the coefficients of $p$, in particular [31]:
$r_{1}\leq 2\max_{i=1}^{d}|p_{d-j}/p_{d}|^{1/j}$ unless $p_{d}=0$ and $r_{d}\geq 2\min_{i=1}^{d}|p_{j}/p_{0}|^{1/j}$ unless $p_{0}=0$.
For a complex $c$ and positive integers $j$ and $h$, Eqn. (3.1) and Observation 6.2 combined imply that
$r_{d}(c)\leq d~{}|{\rm NR}(c)|,~{}{\rm and}~{}r^{h}_{j}(0,p)=r_{j}(0,p_{h}).$ (6.4)
The roots lying far from a point $c$ little affect NIR$(c)$, and we extend bounds (6.4) as follows.
Proposition 6.2.
If #$\#(D(c,1))=\#(D(c,\theta))=m$ for $\theta>1$, then $|{\rm NIR}(c)|-\frac{d-m}{\theta-1}<\frac{m}{r_{d}(c)}.$
Proof.
Assume that $|z_{j}|>\theta$ if and only if $j>m$ and then deduce from Eqn. (3.1) that $|{\rm NIR}(c)|\leq\sum_{j=1}^{m}\frac{1}{|c-z_{j}|}+\sum_{j=m+1}^{d}\frac{1}{|c-z_{j}|}<\frac{m}{r_{d}(c)}+\frac{d-m}{\theta-1}$. ∎
Certify inclusion into the disc $D(0,\sigma)$ if $\rho_{d}(0,p_{h})\leq\sigma^{h}$, $\sigma>1$ and $h>0$, although this bound is extremely poor for worst case polynomials $p$:
Proposition 6.3.
The ratio $|{\rm NR}(0)|$ is infinite for $p(x)=x^{d}-u^{d}$ and $u\neq 0$, while
$r_{d}(c,p)=r_{1}(c,p)=|u|.$
Proof.
Notice that $p^{\prime}(0)=0\neq p(0)=-u^{d}$, $p(x)=\prod_{j=0}^{d-1}(x-u\zeta_{d}^{j})$ for $\zeta_{d}$ of (3.4). ∎
Notice that $\frac{1}{r_{d}(c)}\leq|{\rm NIR}(c)|=\frac{1}{d}\Big{|}\sum_{j=1}^{d}\frac{1}{c-z_{j}}\Big{|}$ by virtue of (3.1), and so approximation to the root radius $r_{d}(c,p)$ is poor if and only if severe cancellation occurs in the summation of the $d$ roots. It seems that such a cancellation only occurs for a very narrow class of polynomials $p$, and if so, then the above non-costly test certifies inclusion for a very large class of polynomials $p$ [^15] Shifts of the variable $p(x)\leftarrow t(x)=p(x-c)$ for random $c$ should further enlarge this class.
6.5 Subdivision root-finding involves exclusion
It is not enough for us to have non-costly certification of inclusion because exclusion occurs in more than 25% of all e/i tests in a subdivision process:
Proposition 6.4.
Let $\sigma_{i}$ suspect squares enter the $i$-th subdivision step for $i=1,\dots,t+1$. Let it stop at $f_{i}$ of them, discard $d_{i}$ and further subdivide $\sigma_{i}-d_{i}-f_{i}$ of them. Write $\sigma_{1}:=4$, $\sigma_{t+1}:=0$, $\Sigma:=\sum_{i=1}^{t}\sigma_{i}$, $D:=\sum_{i=1}^{t}d_{i}$, and $F:=\sum_{i=1}^{t}f_{i}$. Then $3\Sigma=4(D+F-1)$ and $D>\Sigma/4$.
Proof.
Notice that $\sigma_{i+1}=4(\sigma_{i}-d_{i}-f_{i})~{}{\rm for}~{}i=1,2,\dots,t+1$, sum these equations for all $i$, substitute $\sigma_{1}=4$ and $\sigma_{t+1}=0$, and obtain $3\Sigma+4\geq 4D+4F$. Represent subdivision process by a tree with at most $F\leq m$ leaves whose other nodes are components made up of suspect squares, to obtain pessimistic lower bounds $\Sigma+1\geq 2m\geq 2F$ and hence $D\geq\Sigma/4+1/2$. ∎
7 Our e/i tests
7.1 Overview
Our Alg. 7.1 of Sec. 7.2 and Alg. 7.2 of Sec. 7.3 are deterministic specifications of Alg. 3.1. By virtue of Lemma 3.1 all of them are correct when they claim inclusion, but we will prove that they are also correct when output exclusion. Namely, Algs. 7.1 and 7.2 evaluate ${\rm NIR}(x)$ at $hq$ equally spaced points of the unit circle $C(0,1)$ for $h=O(\log(d))$ and certify exclusion if at all these points $|{\rm NIR}_{p_{h}}(x)|\leq\frac{1}{2\sqrt{q}}$ for $q>d$ or if $|{\rm NIR}_{p_{h}}(x)|\leq\frac{1}{3\sqrt{q}}$ for $q>m$, respectively, and otherwise certify inclusion. Our Example 7.1 shows that we cannot devise such deterministic algorithms for $q<m$, but our Las Vegas e/i test evaluates polynomial $p(x)$ at $O(\log(d))$ points $x$ overall under the Random Root Model.
7.2 Basic deterministic e/i test
[…]
Algorithm 7.1.
Basic deterministic e/i test.
INPUT: a black box polynomial $p(x)$ of a degree $d$ given by an oracle for its evaluation and $\sigma$, $1<\sigma<\sqrt{2}$.
INITIALIZATION: Fix two integers $q>d$ and $h:=\lceil\log_{\sigma}(1+2d\sqrt{q})\rceil$, such that $\sigma^{h}\leq 1+2d\sqrt{q}$.
COMPUTATIONS: For $g=0,1,\dots,q-1$ recursively compute the values
$v_{g}:=\Big{|}\frac{p^{\prime}_{h}(\zeta^{g})}{p_{h}(\zeta^{g})}\Big{|}.$ (7.1)
Stop and certify $\sigma$-soft inclusion #$(D(0,\sigma))>0$ if $1+d/v_{g}\leq\sigma^{h}$ for an integer $g<q$. Otherwise stop and certify exclusion, that is, #$(D(0,1))=0$.
Alg. 7.1 outputs $\sigma$-soft inclusion correctly by virtue of Lemma 3.1.
Next we will prove that it also correctly claims exclusion.
Lemma 7.1.
A polynomial $p$ has no roots in the unit disc $D(0,1)$ if $|p(\gamma)|>2\max_{w\in D(0,1)}|p^{\prime}(w)|$ for a complex $\gamma\in C(0,1)$.
Proof.
By virtue of Taylor-Lagrange’s theorem $p(x)=p(\gamma)+(x-\gamma)p^{\prime}(w)$, for any pair of $x\in D(0,1)$ and $\gamma\in C(0,1)$ and for some $w\in D(0,1)$. Hence $|x-\gamma|\leq 2$, and so $|p(x)\geq|p(\gamma)|-2~{}|p^{\prime}(w)|.$ Therefore, $|p(x)|>0$ for $\gamma$ satisfying the assumptions of the lemma and for any $x\in D(0,1)$. ∎
Proposition 7.1.
#$\#(D(0,1))=0$ for a black box polynomial $p(x)$ of a degree $d$ if
[…]
Proof.
Bounds (7.1) and (7.2) together imply that $|p^{\prime}(\zeta^{g})|/v\leq|p(\zeta^{g})|$ for $g=0,\dots,q-1$ and hence $2\sqrt{q}||(p^{\prime}(\zeta^{g}))_{g=0}^{q-1}||_{\infty}<\widehat{P}~{}{\rm for}~{}\widehat{P}:=||(p(\zeta^{g}))_{g=0}^{q-1}||_{\infty}.$ Combine this bound with (6.1) to obtain
$2||(p^{\prime}(\zeta^{g}))_{g=0}^{q-1}||_{2}<\widehat{P}.$ (7.3)
Pad the coefficient vector of the polynomial $p^{\prime}(x)$ with $q-d-1$ initial coordinates 0 to arrive at $q$-dimensional vector ${\bf p^{\prime}}$. Then $(p^{\prime}(\zeta^{g}))_{g=0}^{q-1}=F{\bf p^{\prime}}$ for the $q\times q$ matrix $F:=(\zeta^{ig})_{i,g=0}^{q-1}$ of discrete Fourier transform. Substitute $F{\bf p^{\prime}}$ for $(p^{\prime}(\zeta^{g}))_{g=0}^{q-1}$ in (7.3) and obtain $2||F{\bf p^{\prime}}||_{2}<\widehat{P}$. Hence $2\sqrt{q}||{\bf p^{\prime}}||_{2}<\widehat{P}$ because the matrix $\frac{1}{\sqrt{q}}F$ is unitary (cf. [35, 83]). It follows (cf. (6.1)) that $2||{\bf p^{\prime}}||_{1}<\widehat{P}.$ Since $\max_{x\in D(0,1)}|p^{\prime}(x)|\leq||{\bf p^{\prime}}||_{1}$, deduce that $2\max_{x\in D(0,1)}|p^{\prime}(x)|<\widehat{P}=||(p(\zeta^{g}))_{g=0}^{q-1}||_{\infty}$. Complete the proof by combining this bound with Lemma 7.1. ∎
If Alg. 7.1 outputs exclusion, then $1+\frac{d}{v_{g}}>\sigma^{h}$ for $g=0,1,\dots,q-1$. Hence $\frac{1}{v_{g}}>2\sqrt{q}$ because $\sigma^{h}>1+2d\sqrt{q}$ according to initialization of Alg. 7.1. Therefore, the assumptions of Prop. 7.1 hold for $p_{h}(x)$ replacing $p(x)$, and hence the disc $D(0,1)$ contains no zeros of $p_{h}(x)$, but then it contains no zeros of $p(x)$ as well (see Observation 6.2 (iii)), and so exclusion is correct.
Readily verify the following estimates.
Proposition 7.2.
Alg. 7.1 is a $\sigma$-soft e/i test, certifying inclusion at the cost of evaluation of polynomial $p(x)$ at $2hj$ points $x$ or exclusion at the cost of evaluation at $2hq$ points $x$ for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m$, and $h:=\lceil\log_{\sigma}(1+2d\sqrt{q})\rceil$, $h=O(\log(d))$ for $\sigma-1$ exceeding a positive constant.
7.3 A faster deterministic e/i test
[…]
Next we supply relevant estimates.
Algorithm 7.2.
A faster deterministic e/i test.
INPUT: a black box polynomial $p(x)$ of a degree $d$, two integers $m$ and $q>m$, and real $\sigma$ and $\theta>1$ such that $1<\sigma<\sqrt{2}$,
$|z_{j}|\leq 1~{}{\rm for}~{}j\leq m,~{}|z_{j}|>\theta~{}{\rm for}~{}j>m,$
$\frac{d-m}{\theta^{h}-1}\leq\frac{1}{6\sqrt{q}}~{}{\rm for}~{}h:=\lceil\log_{\sigma}(1+6m\sqrt{q})\rceil~{}{\rm such~{}that}~{}\sigma^{h}\geq 1+6m\sqrt{q}.$
COMPUTATIONS: Recursively compute the values $v_{g}$ of (7.1) for $g=0,1,\dots,q-1$. Certify $\sigma$-soft inclusion, that is, #$(D(0,\sigma))>0$, if $v_{g}\geq\frac{1}{3\sqrt{q}}$ for some integer $g<q$. Otherwise certify exclusion, #$(D(0,1))=0$.
Prop. 6.2, for $p_{h}$ replacing $p$, and the assumed bound $\frac{d-m}{\theta^{h}-1}\leq\frac{1}{6\sqrt{q}}$ together imply
$\frac{m}{r_{d}(\zeta^{g},p_{h})}\geq v_{g}-\frac{d-m}{\theta^{h}-1}\geq v_{g}-\frac{1}{6\sqrt{q}}~{}{\rm for~{}all}~{}g.$
Alg. 7.2 claims inclusion if $v_{g}\geq\frac{1}{3\sqrt{q}}$ for some $g$, but then
$\frac{m}{r_{d}(\zeta^{g},p_{h})}\geq\frac{1}{6\sqrt{q}},~{}r_{d}(\zeta^{g},p_{h})\leq 6m\sqrt{q},$
and hence
$r_{d}(0,p_{h})\leq 1+6m\sqrt{q}~{}\leq\sigma^{h}~{}{\rm for}~{}|\zeta|=1.$
Now Observation 6.2 (iii) implies that Alg. 7.2 claims inclusion correctly.
Next we prove that it also claims exclusion correctly.
Proposition 7.3.
Let a polynomial $p(x)$ of a degree $d$ have exactly $m$ roots in the disc $D(0,\theta)$ for
$\theta^{h}\geq 1+6(d-m)\sqrt{q},~{}h:=\lceil\log_{\sigma}(1+6m\sqrt{q})\rceil,~{}{\rm and~{}an~{}integer}~{}q>m$
and let $v<\frac{1}{3\sqrt{q}}$ for $v$ of (7.2). Then #$\#(D(0,1))=0$.
Proof.
Let $p(x)$ and its factor $f(x)$ of degree $m$ share the sets of their zeros in $D(0,\theta)$. Then (3.1) implies that $\Big{|}\frac{f^{\prime}(x)}{f(x)}\Big{|}<\Big{|}\frac{p^{\prime}(x)}{p(x)}\Big{|}+\frac{d-m}{\theta-1}~{}{\rm for}~{}|x|=1,$ and so $\max_{g=0}^{q-1}\Big{|}\frac{f^{\prime}(\zeta_{q}^{g})}{f(\zeta_{q}^{g})}\Big{|}<v+\frac{d-m}{\theta-1}$. Recall that $\frac{d-m}{\theta-1}\leq\frac{1}{6\sqrt{q}}$ and $v<\frac{1}{3\sqrt{q}}$ by assumptions.
Combine these bounds to obtain $\max_{g=0}^{q-1}\Big{|}\frac{f^{\prime}(\zeta_{q}^{g})}{f(\zeta_{q}^{g})}\Big{|}<\frac{1}{2\sqrt{q}}$. To complete the proof, apply Prop. 7.1 for $f(x)$ replacing $p(x)$ and for $m$ replacing $d$. ∎
Readily extend Prop. 7.2 to Alg. 7.2 by decreasing its bounds on $q$ and $h$ provided that #$\#(D(0,\theta))=m$ for $\theta\geq 1+6(d-m)\sqrt{q}$.
Proposition 7.4.
Alg. 7.2 is a $\sigma$-soft e/i test: it certifies either inclusion or exclusion at the cost of evaluation of $p(x)$ at $2hj$ or $2hq$ points $x$, respectively, for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m:=\#(D(0,\theta)),~{}\sigma^{h}>1+6m\sqrt{q}$, $\theta\geq 1+6(d-m)\sqrt{q}$, and $h:=\lceil\log_{\sigma}(1+6m\sqrt{q})\rceil$, so that $h=O(\log(m))$ if $\sigma-1$ exceeds a positive constant, and in this case we only need $O(m\log(m)$ queries of the oracle to perform an e/i test.
7.4 A randomized e/i test
7.4.1. Introductory comments. Would Algs. 7.1 and 7.2 stay correct if we apply them at fewer evaluation points? They have consistently stayed correct for $q$ of order $\log^{2}(d)$ in extensive numerical tests of [45, 48] applied to $p(x)$ rather than to $p_{i}(x)$ for $i>0$. The tests reported exclusion where the values $|s_{h,q}|$ were reasonably small for $s_{h,q}:=\frac{1}{q}\sum_{g=0}^{q-1}\zeta^{(h+1)g}~{}\frac{p^{\prime}(\zeta^{g})}{p(\zeta^{g})},~{}{\rm for}~{}h=0,1,2$ (cf. Eqn. 12.2 in Part III), and reported inclusion otherwise. Clearly, $|s_{h,q}|\leq\delta$ for all $h$ if $|p(\zeta^{g})|\leq\delta/q$ for $g=0,1,\dots,q-1$, while inclusion is certified unless $|p(\zeta^{g})|$ is small for all $g$.
The test results of [45, 48], however, cannot be extended to worst case input polynomials $p(x)$ already for $q=m-1$.
Example 7.1.
Let $q=m-1$, $p(x)=(1-ux^{d-m})(x^{m}-mx+u)$, $u\approx 0$ but $u\neq 0$. (i) First let $m=d$. Then $p(x)=(1-u)(x^{d}-dx+u)$, $p^{\prime}(x)=(1-u)d(x^{d-1}-1)$ and NIR$(x)=0$ for $x$ being any $(d-1)$-st roots of unity. Then our e/i test for $q=d-1$ would imply that #$\#(D(0,1))=0$, while $p(y)=0$ for $u\approx 0$ and some $y\approx 0$. (ii) For $0<m<d$ notice that $p^{\prime}(x)=uf(x)+m(1-ux^{d-m})(x^{m-1}-1)$ where $f(x)=(d-m)x^{d-m-1}(x^{m}-mx+u)$ and so $uf(x)\approx 0$ while $1-ux^{d-m}\approx 1$ for $|x|\leq 1$ and $u\approx 0$. Therefore, $p^{\prime}(x)\approx 0$ where $x$ is any $(m-1)$-st root of unity, and then again extension of our e/i test to $q=m-1$ would imply that #$\#(D(0,1))=0$, while $p(y)=0$ for $u\approx 0$ and some $y\approx 0$.
Clearly, [45, 48] have never encountered worst case inputs in their extensive tests. Next, in line with this observation, we prove that our e/i tests, which evaluate $p(x)$ at $O(\log(d)$ points, are correct whp under the Random Root Model.
[…]
Lemma 7.2.
Fix any constant $\beta>1$ and a sufficiently large integer $h$ of order $\log(d)$. Then under the Random Root Model, Alg. 3.1(rrm) is correct with a probability at least $1-P$, where
$P\leq\frac{\gamma d}{\pi R^{2}}~{}{\rm for}~{}\gamma:=\pi\max\Big{\{}\frac{\beta^{2}}{\sigma^{2}},O\Big{(}\frac{1}{\beta^{2h}}\Big{)}\Big{\}}.$ (7.4)
Proof.
By virtue of Lemma 3.1 Alg. 3.1(rrm) outputs $\sigma$-soft inclusion correctly and, therefore, can be incorrect only where it outputs exclusion for a disc $D(0,1)$ containing a root.
Let this root be $z_{1}$, say. Then fix the values of the random variables $z_{2},\dots,z_{d}$ and $x_{1}\in C(0,1)$ and let $\gamma$ denote the area (Lebesgue’s measure) of the set $\mathbb{S}$ of the values of the variable $z_{1}$ for which
$v_{1}=|{\rm NIR}_{t_{h}}(x_{1})|<\frac{d}{\sigma^{h}-1}.$ (7.5)
Then Alg. 3.1(rrm) is incorrect with a probability at most
$P=\frac{\gamma d}{\pi R^{2}}$ (7.6)
where the factor of $d$ enables us to cover the cases where the value of any of the $d$ iid random variables $z_{1},\dots,z_{d}$, and not necessarily $z_{1}$, lies in the disc $D(0,1)$.
Fix two values $\bar{z},z\in\mathbb{S}$ such that
$\bar{z}={\rm argmax}_{y\in\mathbb{S}}~{}|y|~{}{\rm and}~{}z={\rm argmax}_{y\in\mathbb{S}}~{}|y-\bar{z}|$
and notice that
$\gamma\leq\pi\min\{|\bar{z}|^{2}|,|\bar{z}-z|^{2}\}.$ (7.7)
To estimate $|\bar{z}-z|$, first write $y:=z^{h}\in D(0,1),~{}\bar{y}:=\bar{z}^{h}\in D(0,1)$ and prove that
$|\Delta|=|\bar{y}-y|<\nu_{h}~{}{\rm for}~{}\nu_{h}:=\frac{8d}{\sigma^{h}-1}.$ (7.8)
[…]
$|\Delta_{t_{h}}|<\frac{2d}{\sigma^{h}-1}~{}{\rm for}~{}\Delta_{t_{h}}:={\rm NIR}^{\prime}-{\rm NIR}.$ (7.9)
Now let $y_{1}:=z_{1}^{h},\dots,y_{d}=z_{d}^{h}$ denote the $d$ zeros of $t_{h}(x)$ and deduce from (3.1) that
$|{\rm NIR}|=\Big{|}\frac{1}{x_{1}-y}+S\Big{|}~{}{\rm and}~{}|{\rm NIR}^{\prime}|=\Big{|}\frac{1}{x_{1}-\bar{y}}+S\Big{|}~{}{\rm for}~{}S:=\Big{|}\sum_{j=2}^{d}\frac{1}{x_{1}-y_{j}}\Big{|}.$
Hence
$|\Delta_{t_{h}}|=|{\rm NIR}-{\rm NIR}^{\prime}|=\Big{|}\frac{1}{x_{1}-y}-\frac{1}{x_{1}-\bar{y}}\Big{|}=\frac{|\Delta|}{|(x_{1}-y)(x_{1}-\bar{y})|}.$ (7.10)
Recall that $x_{1}\in C(0,1)$, while $y,\bar{y}\in D(0,1)$, and so $\max\{|x_{1}-y|,|x_{1}-\bar{y}|\}\leq 2.$
Combine this inequality with (7.10) and deduce that $|\Delta_{t_{h}}|\geq|\Delta|/4.$
Combine the latter bound with (7.10) and obtain (7.8).
Now fix a large integer $g$, substitute $\bar{y}=\bar{z}^{h}$ and $y=z^{h}$ for $h=g$ and $h=g+1$, and obtain
$\lambda_{g+1}:=z^{g+1}-\bar{z}^{g+1},~{}\lambda_{g}:=z^{g}-\bar{z}^{g},~{}{\rm and~{}so}~{}z^{g+1}=z\bar{z}^{g}+z\lambda_{g},~{}|\lambda_{h}|\leq\nu_{h}~{}{\rm for}~{}h=g,g+1.$
Therefore, $\lambda_{g+1}+\bar{z}^{g+1}=\lambda_{g}z+\bar{z}^{g}z$, and so
$z-\bar{z}=\frac{\lambda_{g+1}-z\lambda_{g}}{\bar{z}^{g}},~{}|\bar{z}-z|\leq\frac{|\lambda_{g+1}|+|z|\cdot|\lambda_{g}|}{|\bar{z}|^{g}}<\frac{\nu_{g+1}+|z|\nu_{g}}{|\bar{z}|^{g}}=O\Big{(}\frac{d}{(\sigma|\bar{z}|)^{g}}\Big{)}.$
Hence $|z-\bar{z}|=O(1/\beta^{g})$ for a fixed $\beta>1$. Combine this bound with (7.6) and (7.7), choosing sufficiently large $g$ of order $\log(d)$ unless $\sigma|\bar{z}|<\beta$.
Combine this bound with (7.6) and (7.7), recall that $h=g$ or $h=g+1$, and obtain the lemma. ∎
For a fixed $\sigma>1$, e.g., for $\sigma=1.2$, choose a constant $\beta>1$ and a sufficiently large $g=O(\log(d))$ such that $\gamma<\pi$ in (7.4). Then under the Random Root Model the probability $P$ of the output error for Alg. 3.1(rrm) is less than a constant $\nabla<1$. We can decrease the error probability below $\nabla^{v}$ by applying the algorithm $v$ times, in particular below $d^{a}$ for any fixed positive constant $a$ and for $v=O(\log(d))$. We arrive at the following
Corollary 7.1.
Under Random Root Model one only needs to query an evaluation oracle to perform an e/i test correct;y whp.
Remark 7.1.
It is tempting to combine root-lifting or root-squaring with the algorithms of Part III to obtain an alternative randomized e/i test supporting Cor. 7.1, but for that we would need to raise disc isolation from order of $1+1/m$ to a constant; this would require too costly lifting of order $m$.
7.5 Counting the evaluation queries of our root-finders
Combine Cors. 2.2 and 7.1 and Prop. 7.4 and obtain
Corollary 7.2.
We can reduce root-finding to deterministic evaluation of $p(x)$ at $O(m^{2}b\log(d))$ points as well as to the evaluation at $O(mb\log^{2}(d))$ points under the Random Root Model.
8 Precision of computing
8.1 Doubling computational precision towards its a posteriori estimation
Given an algorithm $\mathbb{A}$ for black box polynomial computations and target tolerance TOL to the output errors, how should one choose computational precision? Here is a simple practical recipe from [6, 9, 15] and MPSolve for obtaining a posteriori estimate within a factor of two from optimal: first apply algorithm $\mathbb{A}$ with a lower precision, then recursively double it and stop where an output error decreases below TOL.
8.2 A priori estimates for computational precision
In the rest of this section we deduce a priori estimate of $O(b\log(d))$ bits for computational precision. Namely, we prove that this precision supports correctness of our subdivision root-finders; our analysis includes some advanced techniques of [59].
Estimating computational precision of our soft e/i tests we assume that they are applied to the unit disc $D(0,1)$ because we can extend the results to any disc $D(c,\rho)$ by applying map (1.4) and increasing the upper bounds on the precision by a factor of $\log_{2}(1/\rho)$. This factor is at most $b$ because we only handle discs of radii $\rho>\epsilon=1/2^{b}$ until we reach $\epsilon$-discs and stop. Therefore, our bound $O(\log(d))$ on the computational precision of our e/i tests for the disc $D(0,1)$ will imply the bound $O(b\log(d))$ for any disc.
We also make Assumption 0 that we can request and obtain from evaluation oracle (black box subroutine) the values of $p$ with a relative error bound within $1/d^{\nu}$ for any fixed constant $\nu>0$ of our choice; we can represent these values with the bit-precision $\log_{2}(d)+\log_{2}(\nu)$.
In our e/i tests we only need to compare the values $|{\rm NIR}_{t_{h}}(x)|$ at some complex points $x=v_{d}\geq 1/d^{\beta}$ for a constant $\beta>0$, and next we estimate that under Assumption 0 we can indeed do this by performing our computations with a precision of $O(\log(d))$ bits.
Multiplication with a precision $s$ contributes a relative error at most $2^{s}$ (cf. [14, Eqn. A.3 of Ch. 3]) and otherwise just sums the relative error bounds of the factors for we ignore the dominated impact of higher order errors. Hence we deduce that $h=O(\log(d))$ multiplications for root-lifting to their $h$th powers increase these bounds by at most a factor of $h$, which is immaterial since we allow relative errors of order $1/d^{\nu}$ for a constant $\nu>0$ of our choice.
We apply this analysis to approximation of NIR$(x)$ for $p_{h}(x)$ based on (3.2) and under the following
Assumption 1: Whp all $q$ evaluation points $x\in C(0,1)$ of an e/i test for the unit disc $D(0,1)$ lie at the distance at least $\phi:=\frac{2\pi}{mqw}$ from all roots for any fixed $w>1$ of our choice.
This assumption holds whp because we apply our e/i tests under random rotation $x\mapsto vx$ for random variable $v$ specified in our next proposition.
Proposition 8.1.
Let $\phi_{g}:=\min_{j=1}^{d}|v\zeta^{g}-z_{j}|$ and let $\phi:=\min_{g=0}^{q-1}\phi_{g}$ for the roots $z_{j}$, $\zeta$ of (3.4) and random variable $v$ sampled from a fixed arc of the circle $C(0,1)$ of length $2\pi/q$ under the uniform probability distribution on that arc. Let $i(D(0,1))=m$. Then $\phi\leq\frac{2\pi}{mqw}$ for $w>1$ with a probability at most $\frac{\pi}{w}$.
Proof.
Let $\psi_{j}$ denote the length of the arc $C(0,1)\cap D(z_{j},\frac{2\pi}{mqw})$, $j=1,2,\dots,d$. Then $\psi_{j}=0$ for at least $d-m$ integers $j$, and $\max_{j=1}^{d}\psi_{j}\leq\frac{\pi^{2}}{mqw}$ for sure. Hence $\sum_{j=1}^{d}\psi_{j}<\frac{\pi^{2}}{qw}$, and so $\phi_{g}\leq\frac{2\pi}{mqw}$ for a fixed $g$ with a probability at most $\frac{\pi}{qw}$, and the lemma follows. ∎
Next, by applying Prop. 8.1 for a sufficiently large $w=d^{O(1)}$, we prove that a precision of $O(\log(d))$ bits supports our e/i tests under Assumptions 0 and 1. We will (i) prove that $|{\rm NIR}(z)|=d^{O(1)}$, (ii) will approximate this value within $d^{\nu}\delta$ for any fixed constant $\nu$ and for $\delta$ of (3.2) chosen such that $\log(\frac{1}{\delta})=O(\log(d))$, and (iii) will estimate that such a bound withstands the impact of rounding errors of computations for (3.2).
Eqn. (3.1) immediately implies
Proposition 8.2.
Under Assumption 1, $\max_{g=1}^{q}|{\rm NIR}(\zeta^{g})|\leq{d^{\eta}}$ for $\eta=O(1)$.
Corollary 8.1.
A precision of $\lceil(\eta+\beta)\log_{2}(d)\rceil$ bits is sufficient to represent ${\rm NIR}(x)$ within $\frac{1}{d^{\beta}}$ for $g=1,2,\dots,q$.
Instead of ${\rm NIR}(x)=\frac{t^{\prime}(x)}{t(x)}$ for $x\in C(0,1)$ we actually approximate $\frac{t^{\prime}(y)}{t(x)}$ where $|y-x|\leq\delta$ and $t^{\prime}(y)$ is equal to the divided difference of (3.2). Thus we shall increase the above error bound by adding upper bounds $\alpha$ on $|\frac{t^{\prime}(x)}{t(x)}-\frac{t^{\prime}(y)}{t(x)}|$ and $\beta$ on the rounding error of computing $\frac{t(x)-t(x-\delta)}{\delta t(x)}$, for $x=\zeta^{g}$ and $g=1,2,\dots,q$.
In the proof of the next theorem we use the following lemma, which is [59, Fact 3.5].
Lemma 8.1.
For $t(x)=\prod_{j=1}^{d}(x-y_{j})$ and a non-negative integer $j\leq d$ it holds that
$t^{(j)}(x)=j!~{}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}$
where the summation $\sum_{S_{j,d}}$ is over all subsets $S_{j,d}$ of the set $\{1,\dots,d\}$ having cardinality $j$.
Proposition 8.3.
Under the assumptions of Prop. 8.2 it holds that
$|t(x)\alpha|=\Big{|}\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)\Big{|}\leq\Big{|}t(x)\frac{(d\xi^{\prime})^{2}\delta}{1-d\xi^{\prime}\delta}\Big{|}~{}{\rm for}~{}\xi^{\prime}=\frac{1}{\rho}.$
Proof.
The claimed bound on $|\xi^{\prime}|$ follows from (3.1). It remains to apply the first five lines of the proof of [59, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, respectively. Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
Substitute the expressions of Lemma 8.1 and obtain
$t(x)-t(x-\delta)=\delta t^{\prime}(x)+\sum_{j=2}^{\infty}\delta^{j}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}.$
Combine this equation with the assumed bounds on $\frac{1}{|x-y_{j}|}$ and deduce that
$\Big{|}\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)\Big{|}\leq\Big{|}\frac{f(x)}{\delta}\sum_{j=2}^{\infty}\delta^{j}w^{j}d^{j}\Big{|}\leq\Big{|}f(x)\frac{(d~{}w)^{2}~{}\delta}{1-\delta~{}d~{}w}\Big{|}.$
Corollary 8.2.
We can ensure that $\alpha<1/d^{\bar{\eta}}$ for any fixed constant $\bar{\eta}$ by choosing $\delta$ of order $d^{O(1)}$, represented with a precision of $O(\log(d))$ bits.
Proposition 8.4.
Suppose that the values $t(y)$ have been computed by an oracle within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor. 8.2. Then one can ensure that $\beta<\frac{1}{d^{\bar{\eta}}}$ for any fixed constant $\bar{\eta}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $O(\log(d))$ bits, which we can do under Assumption 0.
Proof.
Write
$\beta\delta=\frac{t(x-\delta)(1+\nabla(x-\delta))}{t(x)(1+\nabla(x))}-\frac{t(x-\delta)}{t(x)}=\frac{t(x-\delta)}{t(x)}~{}\frac{\nabla(x)-\nabla(x-\delta)}{1+\nabla(x)}.$
Taylor-Lagrange’s formula implies that
$\frac{t(x-\delta)}{t(x)}=1+\frac{t(x-\delta)-t(x)}{t(x)}=1+\delta\frac{t^{\prime}(u)}{t(x)},~{}{\rm for}~{}u\in[x-\delta,x].$
Hence $|\frac{t(x-\delta)}{t(x)}|\leq 1+|\delta\alpha^{\prime}|+|\frac{t^{\prime}(x)}{t(x)}|$ for $\alpha^{\prime}=\frac{t^{\prime}(u)-t(x)}{t(x)}$.
By extending Cor. 8.2 obtain $|\alpha^{\prime}|\leq\frac{1}{d^{O(1)}}$ and deduce from (3.1) and Assumptions 0 and 1 that $|\frac{t^{\prime}(x)}{t(x)}|\leq\frac{1}{d^{O(1)}}$ (cf. the proof of Prop. 14.1). Hence
$w:=\Big{|}\frac{t(x-\delta)}{t(x)}\Big{|}\leq d^{O(1)}~{}{\rm and}~{}|\beta\delta|\leq|w|\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}.$
Now choose $\nabla(y)$ such that $|\nabla(y)|\leq\frac{|\delta|}{20|w|}$ for $y=x$ and $y=x-\delta$. Then verify that
$\log\Big{(}\frac{1}{|\nabla(y)|}\Big{)}=O(\log(d)),~{}1-|\nabla(x)|\geq 1-\frac{|\delta|}{20|w|}>\frac{19}{20},~{}{\rm and}~{}|\delta\beta|<\frac{|\delta|}{8}.$
Combine Cors. 8.1 and 8.2 and Prop. 8.4 to obtain
Corollary 8.3.
Given an oracle for the evaluation of a polynomial $p(x)$ of a degree $d$ with a relative error of any order in $1/d^{O(1)}$ (this holds under Assumption 0), we can perform our black box subdivision root-finders with a precision of $O(\log(d))$ bits.
Remark 8.1.
We derived our precision bound $O(\log(d))$ (in the case of root-finding in the unit disc $D(0,1)$) based on operating entirely with NIR$(x)$. For comparison, if we relied on computing $p(x)$ and on the bound $\min_{x:~{}|x|=1}(|p(x)|/\sum_{i=0}^{d}|p_{i}|)\geq(\frac{\theta-1}{2\theta})^{d}$ of [106, Eqn. (9.4)], we would have only needed a bound of order $m^{2}d$ on the computational precision of our root-finders.
9 Approximation of matrix eigenvalues
Given a $d\times d$ matrix $M$ that has characteristic polynomial $p(x)=\det(xI-M)$, our root-finders can approximate all its $m$ zeros (that is, the eigenvalues of $M$) lying in a fixed disc $D$ on the complex plane isolated from its external zeros (or eigenvalues of $M$).
Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [111, Thm. 1.3.2]:
$|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|~{}{\rm for~{}some}~{}i,~{}1\leq i\leq d.$
Hence, clearly, $|\lambda|\leq||M||_{F}$.
Combine this bound with the a priori precision estimates of the previous section for our root-finders [^16] For a posteriori estimates for the bit operation complexity of Problems 0, $0^{*}$, 2, and 2, one can combine our root-finders with the recipe of doubling the precision of computing of Sec. 8.1. and with the following result of Storjohann [109]:
Theorem 9.1.
One can evaluate the determinant of a $d\times d$ integer matrix $A$ by using an expected number of $O(d^{\omega}\log^{2}(d)\log(|A||_{F}))$ bit operations.
For any $d\times d$ matrix $M$ we arrive at a record bit operation complexity bound $\tilde{O}(m^{2}\tilde{b}bd^{\omega})$ for approximation within $\epsilon$ of all $m$ eigenvalues of any matrix $M$ lying in the disc $D$ provided that $\tilde{b}=\log_{2}(||M||_{F}/\epsilon)$. This bound exceeds (1.6) by a factor of $b$ and can be decreased by a factor of $m$ under the Random Root Model for the roots of the characteristic polynomial of $M$.
10 Bit operation complexity of polynomial root-finding
10.1 Straightforward extension of the precision bound and a direction to saving evaluation queries
Recall the bound
$\mu(s):=O(s\log(s)=\tilde{O}(s)$ (10.1)
on the number of bit operations for integer multiplication modulo $2^{s}$ [40] (cf. [114, 53, 106, 107]), combine it with the estimate $O(bm)$ of Cor. 7.2 for the number of evaluation queries in our root-finders, the upper bound $2d$ ops for the evaluation of a polynomial $p(x)$ given with its coefficients, and the precision bound $O(b\log(d))$ of Sec. 8, and obtain the bit-operation complexity bound $\tilde{O}(b^{2}dm)$ for our root-finders. [^17] By extending precision bound of [59] we greatly simplified derivation of our bit-complexity estimates versus the papers [106, 52], each of 46 pages, which bypass precision bound and had to use long tedious step-by-step analysis of their much involved root-finders.
Since $p(x)$ is given with its coefficients, we apply fast multipoint evaluation of a polynomial in [32, 67] and use fewer ops, roughly by a factor of $d/\log^{2}(d)$. Peter Kirrinnis in [52] and Moroz in [64] have non-trivially extended this progress to saving bit-operations:
Theorem 10.1.
Given a positive $b$, the coefficients of a polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}$ such that $||t(x)||_{1}=\sum_{i=0}^{d}|t_{i}|\leq 2^{\tau}$, and $q$ complex points $x_{1},\dots,x_{q}$ in the unit disc $D(0,1)$, one can approximate the values $p(x_{1}),\dots,p(x_{q})$ within $1/2^{b}$ by using $\tilde{O}((d+q)(b+\tau))$ bit operations.
Proof.
For $q=d$ this is [64, Thm. 2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Moroz in [64]. Trivially extend this theorem to the cases where $q>d$ (by considering $t(x)$ a polynomial of degree $q$ with $q-d$ leading coefficients 0) and where $q<d$ (by removing $d-q$ points $x_{i}$, with no increase of the complexity bound). ∎
We only apply Thm. 10.1 for $\tau=0$, in which case its bit operation cost bound turns into
$\mathbb{B}_{\rm Moroz}=\tilde{O}((d+q)b).$ (10.2)
For comparison, the State of the Art before [64] was the bound
$\mathbb{B}_{\rm Kirr}=\tilde{O}((d+q)(b+q))$ (10.3)
of [52, Thm. 3.9, Alg. 5.3, and Appendix A.3] by Kirrinnis, who combined his and Schönhage’s study of the bit operation complexity of basic polynomial computations (cf. [106]) with the algorithm of [32, 67]. [^18] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by reducing the task to multiplication of Vandermonde matrix by a vector applying FMM (cf. [86, 87]). One can, however, support bounds of Eqns. (10.2) and (10.3) on the overall bit operation complexity by computing with multiple precision for larger values $b$. Moroz obtained [64, Thm. 2] by applying bound (10.3) to his lower degree polynomials approximating $p(x)$.
Bound (10.2) combined with the precision bound $O(\log(d))$ of Sec. 8 implies the bit operation cost bound $\tilde{O}((d+q)b)$, dominated by an estimated bit operation cost $\mathbb{B}$ for the evaluation of $p(x)$ with relative error $1/d^{O(1)}$ at $\tilde{O}(db)$ points. It remains to estimate $\mathbb{B}$, which we do next.
10.2 Reduction of our root-finders to multipoint polynomial approximation
Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ roots, we approximate all these roots within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with e/i tests of Sec. 7. Then every subdivision step is reduced to application of our e/i tests to $\bar{m}=O(m)$ discs $D(c_{\lambda},\rho_{\lambda})$, where $\rho_{\lambda}\geq R/2^{b}$ for all $\lambda$. Such a test amounts to approximation of $p(x)$ at $q$ equally spaced points $x$ on each circle $C(c_{\lambda},\rho_{\lambda})$.
In Thm. 10.1 a polynomial $t(x)$ with $||t(x)||_{1}\leq 1$ is approximated at some points $x\in D(0,1)$. To ensure the assumption about $x$, we scale the variable $x\mapsto Rx$ to map the discs $D(c_{\lambda},\rho_{\lambda}\sigma)\mapsto D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})\subseteq D(0,1)$, for $\bar{c}_{\lambda}=c_{\lambda}/R$, $\bar{\rho}_{\lambda}=\rho_{\lambda}/R\geq 1/2^{b}$, and all $\lambda$.
Then we write $t(x):=p(Rx)/\psi$ for $\psi=O(R^{d})$ such that $||t(x)||_{1}=1$ and approximate $t(x)$ within $1/2^{\bar{b}}$ or equivalently within the relative error bound $1/(|t(x)|2^{\bar{b}})$ at $\bar{m}$ points $x$ in $\bar{m}$ e/i tests. [^19] Under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $\bar{m}/\log(\bar{m})$ and then extend this decrease to the related subsequent estimates. Due to Cor. 8.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we deduce it by choosing $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$ (cf. Assumption 1 of Sec. 8.2). Then $\log(\frac{1}{|t(x)|})=O(d\log(m))$ by virtue of [106, Eqn. (9.4)]). [^20] That bound follows from the estimate $|f|\geq 1/2^{d}$ of [106, Thm. 4.2], which holds provided that $t(x)=f\prod_{j=1}^{m}(x-x_{j})\prod_{j=m+1}^{d}(\frac{x}{x_{j}}-1)$, $||t(x)||_{1}=1$, $|x_{j}|<1$ for $j\leq m$, and $|x_{j}|>1$ for $j>m$.
Substitute $\bar{\rho}_{\lambda}\geq 1/2^{b}$ and obtain $\bar{b}=\log(\frac{1}{|t(x)|})+O(\log(d))=O(d\log(m)+b)$.
Remark 10.1.
For $m=d$, the above bound on $\log(\frac{1}{|t(x)|})$ is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$. For $m<d$, [106, Thm. 4.5] implies a little stronger bounds, namely, $\log(\frac{1}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.
10.3 Bit operation cost of our e/i tests and root-finders
Apply Thm. 10.1 with $q$ and $b$ replaced by $\bar{q}=O(m^{2})$ and $\bar{b}=O(b+d+m\log(m))$, respectively, and obtain that the bit operation cost of our $O(m)$ e/i tests of Sec. 7 at any fixed subdivision step is in $\tilde{O}((b+d+m)m^{2})$ for $1\leq m\leq d$. This implies the upper estimate
$\mathbb{B}_{\rm roots}=\tilde{O}((b+d)(m^{2}+d)b)$ (10.4)
for the overall bit operation cost of our $O(b)$ subdivision steps. Based on [52, Alg. 5.3] and Eqn. (10.3) we obtain a little larger but still reasonably small upper bound
$\mathbb{B}_{\rm roots}^{\prime}=\tilde{O}(((m^{2}+d)(m^{2}+d+b)b).$ (10.5)
10.4 Higher bit operation complexity of general polynomial root-finding based on Pellet’s theorem, map (1.4), and root-lifting
Recall the followings theorem of Pellet [71], which extends the previous results by Cauchy and Rouché:
Theorem 10.2.
Given a monic polynomial $t(x)=\sum_{i=0}^{d}t_{i}x^{i}$ with complex coefficients and $t_{d}=1$, let the polynomial $s(x)=x^{d}+p_{d-1}x^{d-1}+\cdots+|p_{i+1}|x^{i+1}-|p_{i}|x^{i}|+p_{i-1}|x^{i-1}+\cdots+|p_{0}|$ have two distinct positive roots $r$ and $R$, $r<R$. Then the polynomial $t(x)$ has exactly $i$ zeros in or on the circle $C(0,r)=\{z:~{}|z|=r\}$ and no zeros in the annulus $A(0,r,R)=\{z:~{}r<|z|<R\}$.
One must know the coefficients of a polynomial $t(x)$ to apply this theorem and then can perform very fast an e/i test for a disc centered at the origin.
Such soft e/i tests were the basis of the subdivision root-finders of [16, 17], which apply them to polynomials $t(x)$ obtained from $p(x)$ of (1.1) by means of shifts of the variable $x$, scaling it by factors ranging from $R/2^{b}$ to $R=\max_{j-1}^{d}|z_{j}|$, and order of $\log(\log(d))$ root-squaring steps.
To apply Pellet’s theorem to such a polynomial $t(x)$ one must compute the values $|\frac{t_{i}}{t_{d}}|$ for $i=0,1,\dots,d-1$. Because of scaling alone, the coefficient $t_{i}$ must be computed within $1/2^{b-i}$ to approximate the zeros of $p$ within $1/2^{b}$ for every $i$, $i=0,1,\dots,d$.
As in the proof of Observation 22.2, it follows that one must process at least $0.5(d+1)db$ bits of the coefficients of $t(x)$ and must perform at least $0.25(d+1)db$ bit operations to support even a single e/i tests towards approximation of the zeros of $p(x)$ within $1/2^{b}$.
Already a single subdivision iteration for Problem 1m can involve up to $4m$ soft e/i tests and hence up to at least $0.25(d+1)dbm$ bit-operations. This is by a factor of $d$ exceeds the bound of [78, 84] for approximation ofall $m$ roots in an isolated disc.
Such a bit-count and the bit operation cost bounds for the root-finders of [78, 84, 17] decrease by a factor of $d$ where the minimal pairwise distance between the zeros of $p$ exceeds a positive constant; the ratio of the bounds of [17] and [78, 84] remains at least $d$.
Similar argument shows that already for a singe e/i test representation of the leading coefficients $t_{d}$ of $t(x)$ alone under scaling involves order of $bd$ bits. Moreover, this is amplified to $d^{2}b$ for $t_{d}^{d}$ in every e/i test and up to $d^{2}bm$ per a subdivision iteration with $m$ e/i tests.
It may be interesting that a shift by $c$ such that $|c|\leq 2$ is much less costly than scaling: [^21] We can relax the restriction $|c|\leq 2$ above by means of scaling the variable $x$, but as we showed above, scaling can greatly increase the bound on the overall number of bit operations involved.
Theorem 10.3.
Bit operation cost of the shift of the variable, [108, Lemma 2.3] (cf. [52, Lemma 3.6]). Given a positive $b$, the coefficients of a polynomial $p(x):=\sum_{i=0}^{d}p_{i}x^{i}$ such that $||p(x)||_{1}\leq 1$, and a complex $c$ such that $|c|\leq 2$, one can approximate within $1/2^{b}$ the coefficients $t_{0},t_{1},\dots,t_{d}$ of the polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}:=p(x-c)$ by performing $O(\mu((b+d)d))$ bit operations for $\mu(s)$ of (10.1).
PART III: Classical subdivision root-finder with Cauchy-based novel e/i tests
Next we elaborate upon Alg. 3.2 and analyze it.
Organization of the presentation in Part III: We devote the next section to some basic definitions. In Sec. 12 we count roots in a disc on the complex plane by means of approximation of a Cauchy integral over its boundary circle. In Sec. 13 we analyze the resulting root-counters and e/i tests in a disc. In Secs. 14 we estimate the computational precision supporting our root-finders. In Sec. 15 we estimate their bit operation complexity assuming that a polynomial $p(x)$ is represented with its coefficients.
11 Background: basic definitions
  • “Wlog" stands for “without loss of generality".
  • $S(c,\rho):=\{x:~{}|\Re(c-x)|\leq\rho$, $|\Im(c-x)|\leq\rho\},~{}D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$,
    $C(c,\rho):=\{x:~{}|x-c|=\rho\}$, and $A(c,\rho,\rho^{\prime}):=\{x:~{}\rho\leq|x-c|\leq\rho^{\prime}\}$
    denote a square, a disc, a circle (circumference), and an annulus (ring) on the complex plane, respectively.
  • The largest upper bound on such a value $\theta$ is said to be the isolation of the disc $D(c,\rho)$, the circle $C(c,\rho)$, or the square $S(c,\rho)$, respectively (see Fig. 4), and is denoted $i(D(c,\rho))$, $i(C(c,\rho))$, and $i(S(c,\rho))$, respectively.
  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex point $c$ to the zeros $y_{1}\dots,y_{d}$ of a $d$th degree polynomial $t(x)$. $r_{j}(c):=r_{j}(c,p)$, $r_{j}:=r_{j}(0)$ for $j=1,\dots,d$.
Observation 11.1.
Eqn. (1.4) maps the zeros $z_{j}$ of $p(x)$ into the zeros $y_{j}=\frac{z_{j}-c}{\rho}$ of $t(y)$, for $j=1,\dots,d$, and preserves the index #$\#(D(c,\rho))$ and the isolation $i(D(c,\rho))$.
12 Cauchy root-counting in a disc
12.1 The power sums of the roots in a disc
Generalize (3.7) as follows (cf. [1]):
[…]
For $\mathcal{D}=D(0,1)$ and $\zeta$ of (3.4) approximate $s_{h}$ with finite sums (cf. [106, 117])
$s_{h,q}:=\frac{1}{q}\sum_{g=0}^{q-1}\zeta^{(h+1)g}~{}\frac{p^{\prime}(\zeta^{g})}{p(\zeta^{g})}~{}{\rm for}~{}h=0,1,\dots,q-1$ (12.2)
and recall that $s_{0}(D(0,1))=\#(D(0,1))$ (cf. (3.7)), while an e/i test is just the decision whether #$\#(D(0,1))>0$.
12.2 Cauchy sums in the unit disc as weighted power sums of the roots
Cor. 12.1 of the following lemma expresses the $h$th Cauchy sum $s_{h,q}$ as the sum of the $h$th powers of the roots $z_{j}$ with the weights $\frac{1}{1-x^{q}}$.
Lemma 12.1.
For a complex $z$, two integers $h\geq 0$ and $q>1$, and $\zeta$ of (3.4), it holds that
$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=\frac{z^{h}}{1-z^{q}}.$ (12.3)
Proof.
First let $|z|<1$ and obtain
$\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=\frac{\zeta^{hg}}{1-\frac{z}{\zeta^{g}}}=\zeta^{hg}\sum_{u=0}^{\infty}\Big{(}\frac{z}{\zeta^{g}}\Big{)}^{u}=\sum_{u=0}^{\infty}\frac{z^{u}}{\zeta^{(u-h)g}}.$
The equation in the middle follows from Newman’s expansion
$\frac{1}{1-y}=\sum_{u=0}^{\infty}y^{u}~{}{\rm for}~{}y=\frac{z}{\zeta^{g}}.$
We can apply it because $|y|=|z|$, while $|z|<1$ by assumption.
Sum the above expressions in $g$, recall that
$\zeta^{u}\neq 1~{}{\rm and}~{}\sum_{s=0}^{q-1}\zeta^{u}=0~{}{\rm for}~{}0<u<q,~{}\zeta^{q}=1,$ (12.4)
and deduce that
$\frac{1}{q}\sum_{g=0}^{q-1}\frac{z^{u}}{\zeta^{(u-h)g}}=z^{u}$
for $u=h+ql$ and an integer $l$, while $\sum_{g=0}^{q-1}\frac{z^{u}}{\zeta^{(u-h)g}}=0$ for all other values $u$. Therefore
$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=z^{h}\sum_{l=0}^{\infty}z^{ql}.$
Apply Newman’s expansion for $y=z^{q}$ and deduce (12.3) for $|z|<1$. By continuity extend (12.3) to the case of $|z|=1.$
Now let $|z|>1$. Then
$\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=-\frac{\zeta^{(h+1)g}}{z}~{}~{}\frac{1}{1-\frac{\zeta^{g}}{z}}=-\frac{\zeta^{(h+1)g}}{z}\sum_{u=0}^{\infty}\Big{(}\frac{\zeta^{g}}{z}\Big{)}^{u}=-\sum_{u=0}^{\infty}\frac{\zeta^{(u+h+1)g}}{z^{u+1}}.$
Sum these expressions in $g$, write $u:=ql-h-1$, apply (12.4), and obtain
$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=-\sum_{l=1}^{\infty}\frac{1}{z^{ql-h}}=-z^{h-q}\sum_{l=0}^{\infty}\frac{1}{z^{ql}}.$
Apply Newman’s expansion for $y=z^{q}$ again and obtain
$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=-\frac{z^{h-q}}{1-\frac{1}{z^{q}}}=\frac{z^{h}}{1-z^{q}}.$
Hence (12.3) holds in the case where $|z|>1$ as well. ∎
Sum equations (12.3) for $z=z_{j}$ and $j=1,\dots,d$, combine Eqns. 3.1 and (12.2), and obtain
Corollary 12.1.
For the roots $z_{j}$ and all $h$, the Cauchy sums $s_{h,q}$ of (12.2) satisfy $s_{h,q}=\sum_{j=1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}$ unless $z_{j}^{q}=1~{}{\rm for~{}some}~{}j.$
12.3 Approximation of the power sums in the unit disc
Cor. 12.1 implies that the values $|s_{h,q-}s_{h}|$ decrease exponentially in $q-h$ if the unit disc $D(0,1)$ is isolated.
Theorem 12.1.
[106, Eqn. (12.10)]. [^22] Unlike the proof in [106], we rely on Cor. 12.1 of independent interest. Thm. 12.1 does not follow from [117, Thms. 2.1 and 2.2] but can possibly be proved with some techniques of [117]. Let the unit circle $C(0,1)$ be $\theta$-isolated and let $d_{\rm in}$ and $d_{\rm out}=d-d_{\rm in}$ denote the numbers of the roots lying in and outside the disc $D(0,1)$, respectively. Then
$|s_{h,q}-s_{h}|\leq\frac{d_{\rm in}\theta^{-h}+d_{\rm out}\theta^{h}}{\theta^{q}-1}\leq\frac{d\theta^{h}}{\theta^{q}-1}~{}{\rm for}~{}h=0,1,\dots,q-1.$ (12.5)
Proof.
Enumerate the roots $z_{j}$ so that $|z_{j}|<1$ if and only if $j\leq d_{\rm in}$; then Cor. 12.1 implies that
$s_{h,q}-s_{h}=s_{h,q,{\rm in}}+s_{h,q,{\rm out}}$ (12.6)
where
$s_{h,q,{\rm in}}=\sum_{j=1}^{d_{\rm in}}z_{j}^{h}\Big{(}\frac{1}{1-z_{j}^{q}}-1\Big{)}~{}{\rm and}~{}s_{h,q,{\rm out}}=\sum_{j=d_{\rm in}+1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}=\sum_{j=d-d_{\rm out}+1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}.$ (12.7)
Notice that $\frac{1}{1-z_{j}^{q}}-1=\frac{z^{q}}{1-z_{j}^{q}}$, and so $|z_{j}^{h}(\frac{1}{1-z_{j}^{q}}-1)|\leq\frac{\theta^{-h}}{\theta^{q}-1}$ for $z_{j}\in D(0,1)$, while $\Big{|}\frac{z_{j}^{h}}{1-z_{j}^{q}}\Big{|}\leq\frac{\theta^{h}}{\theta^{q}-1}$ for $z_{j}\notin D(0,1)$. Combine these bounds with Eqns. (12.6) and (12.7) and obtain (12.5). ∎
By choosing $q$ that exceeds $\log_{\theta}(2d+1)$ we ensure that $|s_{0}-s_{0,q}|<1/2$.
Example 12.1.
Let the unit circle $C(0,1)$ be 2-isolated. Then Thm. 12.1 implies that $|s_{0,q}-s_{0}|<1/2$ for $d\leq 1,000$ if $q=10$ and for $d\leq 1,000,000$ if $q=20$.
Algorithm 12.1.
INPUT: a black box polynomial $p$ of a degree $d$ and $\theta>1$.
INITIALIZATION: Compute the integer $q=\lfloor\log_{\theta}(4d+2)\rfloor>\log_{\theta}(2d+1)$.
COMPUTATIONS: Compute Cauchy sum $s_{0,q}$ and output an integer $\bar{s}_{0}$ closest to it.
Observation 12.1.
(i) Alg. 12.1 evaluates $p(x)$ at $q=\lfloor\log_{\theta}(4d+2)\rfloor$ points. (ii) It outputs $\bar{s}_{0}=\#(D(0,1))$ if the circle $C(0,1)$ is $\theta$-isolated for $\theta>1$. (iii) If the algorithm outputs $\bar{s}_{0}>0$, then #$\#(D(0,\theta))>0$.
Proof.
Thm. 12.1 immediately implies claim (ii) but also implies that #$\#(D(0,1))>0$ if $\bar{s}_{0}>0$ unless #$\#(A(0,1/\theta,\theta))>0$. In both cases #$\#(D(0,\theta))>0$. ∎
12.4 Extension to any disc
Map (1.4), preserving #$\#(D(0,1))$ (cf. Observation 11.1), enables us to extend definition (12.2) of Cauchy sum to any disc $D(c,\rho)$ as follows:
$s_{0,q}(p,c,\rho):=s_{0,q}(t,0,1):=\frac{\rho}{q}\sum_{g=0}^{q-1}\zeta^{g}~{}\frac{p^{\prime}(c+\rho\zeta^{g})}{p(c+\rho\zeta^{g})},$ (12.8)
for $\zeta=\zeta_{q}$ of (3.4), that is, $s_{h,q}(p,c,\rho)$ is the Cauchy sums $s_{h,q}(t,0,1)$ of the zeros of the polynomial $t(y)=p(\frac{y-c}{\rho})$ in the unit disc $D(0,1)$.
Observation 12.2.
Given a polynomial $p(x)$, a complex $c$, a positive $\rho$, a positive integer $q$, and the $q$-th roots of unity, we can compute $s_{0,q}(p,c,\rho)$ by means of the evaluation of NIR$(\zeta^{g})$ for $g=0,1,\dots,q-1$, reduced to evaluation of $p$ at $2q$ points, and in addition performing $q+1$ divisions, $12q-$ 1 multiplications, and $12q-$ 1 additions.
Refer to Alg. 12.1 applied to the polynomial $t(y)$ of (1.4) for an isolation parameter $\theta>1$ and $q=\lfloor\log_{\theta}(4d+2)\rfloor$ as Alg. 12.1c,ρ,q and also as Alg. 12.1 applied to $\theta$-isolated circle $C(c,\rho)$ for $2q$ evaluation points. Combine Observations 11.1 and 12.2 to extend Observation 12.1 as follows.
Proposition 12.1.
(i) Alg. 12.1c,ρ,q evaluates polynomial $p(x)$ at $q=\lfloor\log_{\theta}(4d+2)\rfloor$ points $x$ and in addition performs $25q-$ 2 ops. (ii) It outputs $\bar{s}_{0}=\#(D(c,\rho))$ if the circle $C(c,\rho)$ is $\theta$-isolated. (iii) If the algorithm outputs $\bar{s}_{0}>0$, then #$\#(D(c,\theta\rho))>0$.
By applying Thm. 12.1 we can estimate the errors $|s_{h,q}(t,0,1)-s_{h}(t,0,1)|$ of the approximation of the $h$th power sums $s_{h}(t,0,1)$ of the roots of $t(x)$ in the unit disc $D(0,1)$. Next we specify this for $h=1$.
Proposition 12.2.
(Cf. [48].) Let for $\theta>1$ a $\theta$-isolated disc $D(c,\rho)$ contain exactly $m$ roots. Fix a positive integer $q$ and define the Cauchy sum $s_{1,q}(p,c,\rho)$ of (12.8). Then
$|mc+\rho s_{1,q}(p,c,\rho)-s_{1}(p,c,\rho)|\leq\frac{\rho d\theta}{\theta^{q}-1}.$
13 Root-counting and $\ell$-tests
13.1 Overview
We seek fast root-counters and $\ell$-tests based on computing Cauchy sum $s_{0,q}$ for a smaller integer $q$, but the output can deceive us if we decrease $q$ below $m$ (see Example 7.1).
Extensive tests for $q$ of order $\log^{2}(d))$ in [45, 48] have never encountered such unlucky inputs, and we only provide some indirect formal support for this empirical behavior in Sec. 7.4.
In the next two subsections we support Las Vegas randomized root-counters and $\ell$-tests that evaluate polynomial $p$ at $O(m\log(d))$ points.
Combine this estimate with Cor. 2.2 for $k$ of order $b$ and obtain Las Vegas randomized solution of Problems 0, 1, 0, and 1 at the cost of evaluation of the polynomial $p$ at $O(m^{2}b)$ points.
13.2 A basic randomized root-counter
Next we fix a parameter $\gamma>1$ and devise a root-counter for a disc $D(c,\rho)$ with a fixed center $c$ (wlog we let $c=0$) and a radius $\rho$ sampled at random in a fixed small range. Then we prove that under such a Las Vegas randomization the test fails with a probability at most $1/\gamma$.
Algorithm 13.1.
Basic randomized root-counter.
INPUT: $\gamma\geq 1$ and a $d$-th degree black box polynomial $p$ having at most $m$ zeros in the disc $D(0,\sqrt{2})$.
INITIALIZATION: Sample a random value $w$ in the range $[0.2,0.4]$ under the uniform probability distribution in that range and write $\rho:=2^{w}$ and $q=\lfloor 10m\gamma\log_{2}(4d+2)\rfloor$.
COMPUTATIONS: Apply Alg. 12.10,ρ,q.
Proposition 13.1.
Alg. 13.1 queries evaluation oracle at $q=\lfloor 10m\gamma\log_{2}(4d+2)\rfloor$ locations and outputs the value $\bar{s}_{0}=\#(D(0,\rho))$ with a probability at least $1-1/\gamma$.
Proof.
Readily verify the bound on the number of query locations claimed in the proposition.
By assumption, the annulus $A(0,1,\sqrt{2})$ contains at most $m$ roots.
Hence at most $m$ root radii $r_{j}=2^{e_{j}}$, for $j=1,\dots,m^{\prime}\leq m$, lie in the range $[1,\sqrt{2}]$, that is, $0\leq e_{j}\leq 0.5$ for at most $m^{\prime}\leq m$ integers $j$.
Fix $m^{\prime}$ intervals, centered at $e_{j}$, for $j=1,\dots,m^{\prime}$, each of a length at most $1/(5m^{\prime}\gamma)$. Their overall length is at most $1/(5\gamma)$. Let $\mathbb{U}$ denote their union.
Sample a random $u$ in the range $[0.2,0.4]$ under the uniform probability distribution in that range and notice that Probability$(u\in\mathbb{U})\leq 1/\gamma$.
Hence with a probability at least $1-1/\gamma$ the circle $C(0,\rho)$ is $\theta$-isolated for $\theta=2^{\frac{1}{10m\gamma}}$, in which case Alg. 13.1 outputs $\bar{s}_{0}=s_{0}=\#(D(0,\rho))$ by virtue of claim (ii) of Prop. 12.1. ∎
13.3 Refining the bound on the error probability
When we apply Alg. 13.1 $v$ times, the number of query locations of the evaluation oracle grows linearly – by a factor of $v$, while the bound on Las Vegas error probability of outputting errors decreases exponentially – from $1/\gamma$ to $1/\gamma^{v}$.
First let $\ell=1$ or $\ell=m$ and then again reduce the root radius approximation problem to the decision problem of $\ell$-test. Namely, narrow Alg. 13.1 to an $\ell$-test for a fixed $\ell$ in the range $1\leq\ell\leq m$, so that with a probability at least $1-1/\gamma$ this $\ell$-test outputs 0 if #$\#(D(0,1))<\ell$ and outputs 1 if #$\#(D(0,2^{0.5}))\geq\ell$. Moreover, output 1 is certified by virtue of claim (iii) of Prop. 12.1 if $\ell=1$; similarly the output 0 is certified if $\ell=m$.
Next extend the $\ell$-test by means of applying Alg. 13.1 for $v$ iid random variables $w$ in the range $0.2\leq w\leq 0.4$. Call this $\ell$-test Alg. $\ell$.
[…]
Recall Prop. 13.1 and then readily verify the following proposition.
Proposition 13.2.
(i) For two integers $1\leq\ell\leq m$ and $v\geq 1$ and a real $\gamma\geq 1$, Alg. $\ell$ evaluates $p(x)$ at $\lfloor 10m\gamma\log_{2}(4d+2)\rfloor v$ points; this evaluation involves more ops than the other stages of the algorithm, even in the cases of the Mandelbrot and sparse polynomials $p$ (cf. Observation 12.2). (ii) Output 1 of Alg. 13.1v,1 and the output 0 of Alg. 13.1v,$m$ are certified. (iii) Output 0 of Alg. 13.1v,1 and output 1 of Alg. 13.1v,$m$ are correct with a probability at least $1-1/\gamma^{v}$.
The proposition bounds the cost of $\ell$-tests for $\ell=1$ (e/i tests) and $\ell=m$. This is sufficient for our root-finding, but our next natural extension covers $\ell$-tests for any integer $\ell$ in the range $1<\ell<m$.
Assign the values 0 or 1 to the output integers $g$ as follows: compute the average of the $v$ outputs in $v$ independent applications of the test and output $g$ denoting the integer closest to that average; if the average is 1/2, write $g=0$ to break ties.
This extension of Alg. $\ell$ is only interesting for $1<\ell<m$; otherwise its cost/error estimates are inferior to those of Prop. 13.2.
Proposition 13.3.
For an integer $\ell$ such that $1<\ell<m\leq d$ and real $v\geq 1$ and $\gamma>1$, Algs. $\ell$ queries evaluation oracle at $\lfloor 10m\gamma\log_{2}(4d+2)\rfloor v$ locations, and its output is correct with a probability at least $1-(4/\gamma)^{v/2}$.
Proof.
Since $D(0,1)\subset D(0,\sqrt{2})$, we only consider the following three cases:
(i) #$\#(D(0,1))<\ell$ and #$\#(D(0,2^{0.5}))\geq\ell$,
(ii) #$\#(D(0,2^{0.5}))<\ell$, and then also #$\#(D(0,1))<\ell$,
(iii) #$\#(D(0,1))\geq\ell$, and then also #$\#(D(0,2^{0.5}))\geq\ell$.
In case (i) the output $g$ is always correct.
In case (ii) the output 0 is correct, while the output 1 is wrong and occurs if and only if more than $v/2$ tests go wrong, by outputting 1.
Likewise, in case (iii) the output 1 is correct, while the output 0 occurs if and only if more than $v/2$ tests go wrong, by outputting 0.
In both cases (ii) and (iii) we deal with the binomial distribution having failure probability $1/\gamma$, and so (cf. [19]) the probability $P$ of at least $v/2$ failures among $v$ tests is at most $\sum_{j=\bar{v}}^{v}\begin{pmatrix}v\\ j\end{pmatrix}T_{j}$ for $T_{j}=\gamma^{-j}(1-1/\gamma)^{v-j}\leq\gamma^{-j}$.
Therefore, $P\leq\sum_{j=0}^{v}\begin{pmatrix}v\\ j\end{pmatrix}=T_{\bar{v}}2^{v}\max_{j}T_{j}$. Substitute $\sum_{j=0}^{v}\begin{pmatrix}v\\ j\end{pmatrix}=2^{v}$ and $\bar{v}=\lceil v/2\rceil$ and obtain $P<2^{v}T_{\bar{v}}\leq 2^{v}/\gamma^{v/2}=(4/\gamma)^{v/2}$. ∎
13.4 E/i test for an annulus
The power sums of the roots lying in an annulus $A(c,\rho,\rho^{\prime})$ can be expressed as the difference of the power sums of the roots lying in the two discs $D(c,\rho^{\prime})$ and $D(c,\rho)$, and our algorithms and their analysis can be readily extended, based on Thm. 12.1 supporting e/i tests for an annulus. We can immediately replace the $m$-test of Sec. 18 with such e/i tests.
14 Estimation of the precision of computing
14.1 Doubling computational precision towards its a posteriori estimation
As in Part II we can obtain near-optimal posteriori estimates for computational precision by recursively doubling it.
Next we obtain its a priori estimate of $O(b\log(d))$ bits. The derivation is a little simpler than in Part II due to observations in the next subsection.
14.2 Reduction of a priori estimates to approximation of $x$NIR$(x)$ within $\frac{3}{8}$
In our e/i tests we approximate NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ with the first order divided difference of (3.2); whp we only need to apply the tests to $\theta$-isolated discs $D(c,\rho)$ for $\theta-1>0$ of order $\frac{1}{m}$ and $\rho\geq\epsilon=1/2^{b}$.
Map (1.4) reduces such a test for $p(x)$ to a test in the disc $D(0,1)$ for the polynomial $t(x)=p(\frac{x-c}{\rho})$ replacing $p(x)$ and decreases the error estimate by a factor of $\rho$. Hence map (1.4) increases the precision bound by at most a factor of $b$, and we obtain the claimed precision bound in $O(b\log(d))$ in case of any disc $D(c,\rho)$ from the bound $O(\log(d))$ for the unit disc $D(0,1)$.
The output of an e/i test is an integer, and we only need to certify that the output error bound of Alg. 16.1 is less than $\frac{1}{2}$. The error can be magnified by a factor of $v$ in $v$ applications of Alg. 16.1 but this factor is immaterial in our notation $\tilde{O}$.
Our e/i tests compute the $q$ values $\rho x\frac{p^{\prime}(c+\rho x)}{p(c+\rho x)}=x\frac{t^{\prime}(x)}{t(x)}$, for $x=\zeta^{g}$, $\zeta$ of (12.2), and $g=0,1,\dots,q-1$, sum these values, and divide the sum by $q$.
By slightly abusing notation write ${\rm NIR}(x):=\frac{t^{\prime}(x)}{t(x)}$ rather than ${\rm NIR}(x)=\frac{p^{\prime}(x)}{p(x)}$. According to straightforward error analysis (cf. [14, Part A of Ch.3] and the references therein) rounding errors introduced by these additions and division are dominated, and we can ignore them; then it is sufficient to compute $\zeta^{g}{\rm NIR}(\zeta^{g})$ within, say, $\frac{3}{8}$ for every $g$. Indeed, in that case the sum of $q$ error bounds is at most $\frac{3q}{8}$ and decreases below $\frac{1}{2}$ in division by $q$.
14.3 Precision of representation of NIR$(x)$ within $\frac{1}{8}$
Recall Eqn. (3.1) and represent $x{\rm NIR}(x)$ for $|x|=1$ within $\frac{1}{8}$ by using a precision of $O(\log(d))$ bits.
Proposition 14.1.
Write $C:=C(c,\rho)$ and assume that $|x|=1$ and the circle $C$ is $\theta$-isolated for $\theta>1$. Then $|x{\rm NIR}(x)|\leq\frac{d}{1-1/\theta}=\frac{\theta d}{\theta-1}$.
Proof.
Eqn. (3.1) implies that $|{\rm NIR}(x)|=|\sum_{j=1}^{d}\frac{1}{x-z_{j}}|\leq\sum_{j=1}^{d}\frac{1}{|x-z_{j}|}$ where $|x-z_{j}|\geq(1-1/\theta)\rho$ for $|z_{j}|<1$, while $|x-z_{j}|\geq(\theta-1)\rho$ for $|z_{j}|>1$ since $i(C)\geq\theta$. Combine these bounds, write $m:=\#(D(c,\rho))$, and obtain $|x{\rm NIR}(x)|\leq\frac{m}{1-1/\theta}+\frac{d-m}{\theta-1}\leq\frac{d}{1-1/\theta}$ for $|x|=1$. ∎
Corollary 14.1.
One can represent $x{\rm NIR}(x)$ within $\frac{1}{8}$ for $|x|=1$ by using a precision of $3+\lceil\log_{2}(\frac{d}{1-1/\theta})\rceil$ bits, which is in $O(\log(d))$ provided that $\frac{1}{\theta-1}=d^{O(1)}$.
14.4 What is left to estimate?
Instead of $x{\rm NIR}(x)=\frac{xt^{\prime}(x)}{t(x)}$ we actually approximate $\frac{xt^{\prime}(y)}{t(x)}$ where $|y-x|\leq\delta$ and $t^{\prime}(y)$ is equal to the divided difference of (3.2). Thus, we shall increase the above error bound $\frac{1}{8}$ by adding to it the upper bounds $\frac{1}{8}$ on $\alpha:=|\frac{xt^{\prime}(x)}{t(x)}-\frac{xt^{\prime}(y)}{t(x)}|$ for $|x|=1$ and on the rounding error $\beta$ of computing $x\frac{t(x)-t(x-\delta)}{\delta t(x)}$. Next we estimate $\alpha$ in terms of $d$, $\theta$, and $\delta$ – by extending the proof of [59, Lemma 3.6], then readily ensure that $\alpha<\frac{1}{8}$ by choosing $\delta$ with a precision $\log_{2}(1/|\delta|)$ of order $\log(d)$, and finally estimate $\beta$. We simplify our estimates by dropping the factor $x$ where $|x|=1$.
14.5 Precision of approximation with divided difference
Recall the following lemma, which is [59, Fact 3.5].
Lemma 14.1.
For $t(x)=\prod_{j=1}^{d}(x-y_{j})$ and a non-negative integer $j\leq d$ it holds that
$t^{(j)}(x)=j!~{}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}$
where the summation $\sum_{S_{j,d}}$ is over all subsets $S_{j,d}$ of the set $\{1,\dots,d\}$ having cardinality $j$.
Proposition 14.2.
Under the assumptions of Prop. 14.1 let $w=w(x):=\max_{j=1}\frac{1}{|x-y_{j}|}$. Then
$|t(x)|\alpha=|\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)|\leq|t(x)\frac{(d~{}w)^{2}\delta}{1-\delta~{}d~{}w}|$.
Proof.
The claimed bound on $|\xi^{\prime}|$ follows from (3.1). It remains to apply the first five lines of the proof of [59, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, s respectively. Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
Substitute the expressions of Lemma 14.1 and obtain
$t(x)-t(x-\delta)=\delta t^{\prime}(x)+\sum_{j=2}^{\infty}\delta^{j}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}.$
Combine this equation with the assumed bounds on $\frac{1}{|x-y_{j}|}$ and deduce that
$\Big{|}\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)\Big{|}\leq\Big{|}\frac{f(x)}{\delta}\sum_{j=2}^{\infty}\delta^{j}w^{j}d^{j}\Big{|}\leq\Big{|}f(x)\frac{(d~{}w)^{2}~{}\delta}{1-\delta~{}d~{}w}\Big{|}.$
In our estimates for $\alpha$ and $\delta$ we apply Prop. 14.1 for $w\leq\frac{\theta}{\theta-1}$ and obtain
Corollary 14.2.
Under the assumptions of Prop. 14.1, let $w\leq\frac{\theta}{\theta-1}$. Then
$\alpha<1/8~{}{\rm for}~{}|\delta|\leq\frac{(\theta-1)^{2}}{10d^{2}\theta^{2}}.$
14.6 The impact of rounding errors
To approximate $|x{\rm NIR}_{c,\rho}(x)|$ within $\frac{3}{8}$ it remains to approximate $\beta$ within $\frac{1}{8}$.
Proposition 14.3.
Suppose that a black box oracle has computed the values $t(y)$ within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor. 14.2. Then we can ensure that $|\beta|<\frac{1}{8}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $\log(\frac{1}{\nabla(y)})=O(\log(d))$ bits.
Proof.
Write $\beta\delta=\frac{t(x-\delta)(1+\nabla(x-\delta))}{t(x)(1+\nabla(x))}-\frac{t(x-\delta)}{t(x)}=\frac{t(x-\delta)}{t(x)}~{}\frac{\nabla(x-\delta)-\nabla(x)}{1+\nabla(x)}$.
Taylor-Lagrange’s formula implies that $\frac{t(x-\delta)}{t(x)}=1+\frac{t(x-\delta)-t(x)}{t(x)}=1+\delta\frac{t^{\prime}(u)}{t(x)}$, for $u\in[x-\delta,x]$. Hence $|\frac{t(x-\delta)}{t(x)}|\leq 1+|(1+\alpha^{\prime})\delta|$ for $\alpha^{\prime}=\frac{t^{\prime}(u)-t(x)}{t(x)}$.
By extending Cor. 14.2 obtain $|\alpha^{\prime}|\leq\frac{1}{8}$.
Hence $|\frac{t(x-\delta)}{t(x)}|\leq w:=1+\frac{9|\delta|}{8}$, while $|\beta\delta|\leq w\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}$.
Now choose $\nabla(y)$ such that $|\nabla(y)|\leq\frac{|\delta|}{17w}$ for $y=x$ and $y=x-\delta$, and so $w(|\nabla(x)|+|\nabla(x-\delta)|)\leq\frac{2}{17\delta}$. Then verify that $\log(\frac{1}{|\nabla(y)|})=O(\log(d))$, $1-|\nabla(x)|\geq 1-\frac{|\delta|}{17w}>\frac{16}{17}$, and $|\delta\beta|<\frac{|\delta|}{8}$. ∎
By combining Cors. 14.1 and 14.2 and Prop. 14.3 obtain
Corollary 14.3.
Suppose that a black box oracle can output approximations of a $d$-th degree polynomial $t(x)$ with any relative error in $1/d^{O(1)}$ for $|x|\leq 1$ and let the unit disc be $\theta$-isolated for $t(x)$ and for $\frac{1}{\theta-1}=d^{O(1)}$. Then we can perform e/i test for $t(x)$ on the unit disc by using a precision of $O(b\log(d))$ bits.
15 Bit operation cost of Cauchy e/i tests and root-finding
We omit extension of the precision bound to the bit complexity estimate for the eigenvalue problem because it can be verbatim the same as in Part II.
15.1 Our goal and a basic theorem
Next we estimate the bit operation complexity of our root-finders by combining them with fast multipoint evaluation of a polynomial by means of the algorithm of [32, 67] non-trivially extended in [52, 64] to support the following theorem:
Theorem 15.1.
Given a positive $b$, the coefficients of a polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}$ such that $||t(x)||_{1}=\sum_{i=0}^{d}|t_{i}|\leq 2^{\tau}$, and $q$ complex points $x_{1},\dots,x_{q}$ in the unit disc $D(0,1)$, one can approximate the values $t(x_{1}),\dots,t(x_{q})$ within $1/2^{b}$ by using $\tilde{O}((d+q)(b+\tau))$ bit operations.
Proof.
For $q=d$ this is [64, Thm. 2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed in [64]. Trivially extend this theorem to the cases where $q>d$ (by considering $t(x)$ a polynomial of degree $q$ with $q-d$ leading coefficients 0) and where $q<d$ (by removing $d-q$ points $x_{i}$, with no increase of the complexity bound). ∎
We only apply this theorem for $\tau=0$, in which case its bit operation bound turns into
$\mathbb{B}_{\rm Moroz}=\tilde{O}((d+q)b)$ (15.1)
For comparison, the previous State of the Art can be represented by the bound
$\mathbb{B}_{\rm Kirr}=\tilde{O}((d+q)(b+q)),$ (15.2)
due to [52, Thm. 3.9, Alg. 5.3, and Appendix A.3] based on combining Kirrinnis’s and Schönhage’s study of the bit operation cost of basic polynomial computations (cf. [106]) with the algorithm of [32, 67]. [^23] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by applying FMM (see Remark 3.1) to the equivalent task of multiplication of Vandermonde matrix by a vector. One can, however, support bounds of Eqns. (15.1) and (15.2) on the overall bit operation complexity by computing with multiple precision for larger values $b$. Moroz obtained [64, Thm. 2] by applying bound (15.2) to his lower degree polynomials approximating $p(x)$.
15.2 Acceleration with fast multipoint polynomial approximation
Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ roots, we approximate all these roots within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with e/i tests of Sec. 13. Then every subdivision step is reduced to application of e/i tests to $\bar{m}=O(m)$ discs $D(c_{\lambda},\rho_{\lambda})$ for $\rho_{\lambda}\geq R/2^{b}$ and all $\lambda$.
Such a test, based on our e/i test of Sec. 13, amounts to approximation of $p(x)$ at $O(m\log(d))$ equally spaced points $x$ on each of $v$ circles sharing the center $c_{\lambda}$ and having radii in the range $[\rho_{\lambda}/\sigma,\rho_{\lambda}\sigma]$ for $1<\sigma<\sqrt{2}$.
We only use Thm. 15.1 for $\tau=0$ to cover the case where a polynomial $t(x)$ with $||t(x)||_{1}\leq 1$ is evaluated at points $x_{j}\in D(0,1)$. To ensure these assumptions we first scale the variable $x\mapsto Rx$ to map the discs $D(c_{\lambda},\rho_{\lambda}\sigma)\mapsto D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})\subseteq D(0,1)$, for $\bar{c}_{\lambda}=c_{\lambda}/R$, $\bar{\rho}_{\lambda}=\rho_{\lambda}/R\geq 1/2^{b}$, and all $\lambda$.
Then we write $t(x):=p(Rx)/\psi$ for $\psi=O(R^{d}||p(x)||_{1})$ such that $||t(x)||_{1}=1$ and approximate $t(x)$ within $1/2^{\bar{b}}$ or equivalently within the relative error bound $1/(|t(x)|2^{\bar{b}})$ at $\bar{q}:=qv$ points $x$ in $\bar{m}$ e/i tests. [^24] Scaling $p$ and $x$ above is done once and for all e/i tests; to compensate for this we only need to change our bound on the output precision $b$ by $\lceil|\log_{2}R|\rceil$ bits.
Due to Cor. 14.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we deduce it by choosing any $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$. Map the disc $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ into the unit disc $D(0,1)$. This does not change isolation of the boundary circle and may increase the norm $||t(x)||_{1}$ by at most factor of $2^{b+d}$ because $|\bar{c}_{\lambda}|\leq 1$ and $\bar{\rho}_{\lambda}\geq 1/2^{b}$. Now substitute $||t(x)||_{1}\leq 2^{b+d}$ and $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d\log(m))$ (cf. [106, Eqn.(9.4)]) [^25] This bound is readily deduced from the following non-trivial result [106, Thm. 4.2]: Let $t(x)=f\prod_{j=1}^{m}(x-z_{j})\prod_{j=m+1}^{d}(\frac{x}{z_{j}}-1)$, $||t(x)||_{1}=1$, $|z_{j}|<1$ for $j\leq m$, and $|z_{j}|>1$ for $j>m$. Then $|f|\geq 1/2^{d}$. and obtain that $\bar{b}=\log(\frac{1}{|t(x)|})+O(\log(d))=O(d\log(m)+b)$.
Remark 15.1.
For $m=d$ the above bound on $\log(\frac{||t(x)||_{1}}{|t(x)|})$ of [106, Eqn.(9.4)] is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$. For $m<d$ [106, Thm. 4.5] implies a little stronger bounds, namely, $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.
15.3 The bit operation complexity estimates
Apply Thm. 15.1 with $q$ and $b$ replaced by $\bar{q}=O(m^{2})$ and $\bar{b}=O(b+d+m\log(m))$, respectively, [^26] Under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $m$ and then extend this decrease to the related subsequent estimates. and obtain that the bit operation cost of our $O(m)$ e/i tests of Sec. 13 at any fixed subdivision step is in $\tilde{O}((b+d+m)m^{2})$ for $1\leq m\leq d$. This implies the upper estimate
$\mathbb{B}_{\rm roots}=\tilde{O}((b+d)m^{2}b)$ (15.3)
for the overall bit operation cost of our $O(b)$ subdivision steps, which includes the time-complexity of the evaluation oracle. Based on [52, Alg. 5.3] and Eqn. (15.2) obtain a little larger upper bound on the bit operation cost of our subdivision root-finders:
$\mathbb{B}_{\rm roots}^{\prime}=\tilde{O}((m^{2}+d+b)(m^{2}+d)b).$ (15.4)
PART IV: Acceleration of Root-Finding With Compression
In this part of the paper we decrease the number of subdivision steps of our root-finders of Parts II and III by a factor of $b/\log(b)$, thus arriving at our Main Theorems 1 – 4.
In Sec. 17 we extend the outline of this acceleration from Sec. 4. It relies on an algorithm for compression of an isolated disc, which we devise and analyze also in Sec. 17. The algorithm involves estimation of the $\ell$th smallest root radius, which is of independent interest and is covered in Sec. 16. It is reduced to recursive application of an $\ell$-test; we present it in Sec. 13 for any positive $\ell\leq d$ and any input disc, but in Sec. 18 we specify simpler tests that are sufficient in application to compression. In Sec. 20 we estimate the overall complexity of the resulting root-finder, which intertwines at most $22m-$ 2 compression steps with e/i tests, whose number we estimate in Sec. 19. In Sec. 21 we further accelerate polynomial root-finding for a large class of inputs.
16 Estimation of the $\ell$-th smallest root radius
In this section we narrow a fixed range $[0,R]$ for the $\ell$-th smallest root radius $r_{d-\ell+1}(c)$ from a fixed complex point $c$. To simplify our exposition we let $R=1$ and $c=0$, wlog due to (1.4).
Problem RR. Compute a range $[\rho_{-},\rho_{+}]$ such that $0\leq\rho_{-}\leq r_{d-\ell+1}\leq\rho_{+}$ and $\rho_{+}\leq\epsilon=1/2^{b}$ or $\rho_{+}\leq\beta\rho_{-}$ for fixed $b>0$ and $\beta=2^{\phi}>1$, provided that we are given a positive integer $\ell\leq d$, initial range $[0,1]$ for $r_{d-\ell+1}$, and a $\sigma$-soft $\ell$-test for a fixed $\sigma>1$.
We first use this test to check whether $r_{d-\ell+1}\leq 1/2^{b}$. If so, we are done. Otherwise we seek $r_{d-\ell+1}$ in the range $[\frac{1}{2^{b}\sigma},1]$. We apply binary search, said to be bisection of logarithms or BoL, to the range $[-\bar{b},0]=[-\log_{2}(\frac{\sigma}{\epsilon}),\log_{2}(1)]$, for $\bar{b}=\log_{2}(\frac{\sigma}{\epsilon})$, to approximate $\log_{2}(r_{d-\ell+1})$.
Since the test is soft, so is BoL as well: for a range $[b_{-},b_{+}]$ with midpoint $\widehat{b}=0.5(b_{+}+b_{-})$ regular rigid bisection continues the search in the range $[b_{-},\widehat{b}]$ or $[\widehat{b},b_{+}]$, but our soft BoL continues it in the range $[b_{-},(1+\alpha)\widehat{b}]$ or $[(1-\alpha)\widehat{b},b_{+}]$ for a fixed small positive $\alpha$. Accordingly we slightly increase the range of search for $r_{d-\ell+1}$ to $[-\bar{b},0]$ for $\bar{b}:=\log_{2}(\frac{\sigma}{\epsilon})$. BoL still converges linearly and just a little slower than the regular rigid bisection: the approximation error decreases by a factor of $\frac{2}{1+\alpha}$ rather than by twice. [^27] If we applied BoL to the range $[0,1]$, we would have to bisect the infinite range $[-\infty,0]$, but we deal with the narrower range $[-b,0]$ since we have narrowed the range $[0,1]$ to $[\epsilon,1]$.
Algorithm 16.1.
Bracketing the $\ell$-th smallest root radius based on $\ell$-test.
INPUT: An integer $\ell$ such that $1\leq\ell\leq d$ and #$\#(D(0,1))\leq\ell$, $\sigma=1+\alpha$ for $20<\alpha<1/$ 2, a $\sigma$-soft $\ell$-test0,ρ for a positive $\rho$ (see Sec. 13.4 for such tests), and two tolerance values $\beta:=2^{\phi}>1$ and $\epsilon=1/2^{b}$, $b\geq 1$.
OUTPUT: A solution $[\rho_{-},\rho_{+}]$ of Problem RR.
INITIALIZATION: Write $\bar{b}:=\log_{2}(\frac{\sigma}{\epsilon})$, $b_{-}:=-\bar{b}$, and $b_{+}:=0$.
[…]
$\rho_{-}=2^{b_{-}},~{}\rho_{+}=2^{b_{+}},~{}\rho:=2^{\widehat{b}},~{}{\rm and}~{}\widehat{b}:=\lceil(b_{-}+b_{+})/2\rceil.$ (16.1)
Write $b_{-}:=(1-\alpha)\widehat{b}$ if the $\ell$-test outputs 0; write $b_{+}:=(1+\alpha)\widehat{b}$ if the test outputs 1. Then update $\widehat{b}$ and $\rho$ according to (16.1).
Continue recursively: stop and output the current range $[\rho_{-},\rho_{+}]=[2^{b_{-}},2^{b_{+}}]$ if $b_{+}-b_{-}\leq\phi$, that is, if $\rho_{+}\leq\beta\rho_{-}$; otherwise update $\widehat{b}$ and $\rho$ and reapply the computations.
The disc $D(0,\epsilon)$ contains at least $\ell$ roots, that is, the range $[0,\epsilon]$ brackets the $\ell$th smallest root radius $r_{d-\ell+1}$ if the $\sigma$-soft $\ell$-test applied to the disc $D(0,\frac{\epsilon}{\sigma})$ outputs 1. Otherwise, the disc $D(0,\frac{\epsilon}{\sigma})$ contains less than $\ell$ roots, and so the root radius $r_{d-\ell+1}$ lies in the range $[\frac{\epsilon}{\sigma},1]$, in which case we continue our search.
Readily verify that our policy of recursively updating the range $[b_{-},b_{+}]$ maintains this property throughout and that updating stops where $b_{+}-b_{-}\leq\phi$, that is, where $\rho_{+}\leq\beta\rho_{-}$. This implies correctness of Alg. 16.1.
Let us estimate the number $\mathbb{I}$ of the invocations of an $\ell$-test.
Proposition 16.1.
Alg. 16.1 invokes an $\ell$-test $\mathbb{I}$ times where $\mathbb{I}\leq\lceil\log_{\nu}(\frac{\bar{b}}{\phi})\rceil+1$ for $\nu=\frac{2}{1+\alpha}$, and so $\mathbb{I}=O(\log(b))$ provided that $\nu-1$, $\sigma-1$, and $\phi$ are positive constants.
Proof.
Alg. 16.1 stops if the $\ell$-test outputs 1. Otherwise the test decreases the length $\bar{b}=b_{+}-b_{-}$ of its input range $[b_{-},b_{+}]$ for the exponent to $0.5(1+\alpha)\bar{b}$. $i$ applications change it into $\frac{(1+\alpha)^{i}}{2^{i}}\bar{b}$ for $i=1,2,\dots$, and the proposition follows. ∎
Remark 16.1.
In the case where we know that #$\#(D(0,\rho))\geq\ell$ for a fixed $\rho$ in the range $[\epsilon,1]$, we only need to refine the initial range $[\rho,1]$ for the root radius $r_{d-\ell+1}$.
Clearly, $\mathbb{I}$ applications of $\ell$-test increase the number of evaluation points for the ratio $\frac{p^{\prime}(x)}{p(x)}$ by at most a factor of $\mathbb{I}$, and if the $\ell$-test is randomized, then the error probability also grows by at most a factor of $\mathbb{I}$. Let us specify these estimates for Alg. 16.1.
Corollary 16.1.
Let precisely $\ell$ zeros of a black box polynomial $p(x)$ of a degree $d$ lie in the unit disc, that is, let #$\#(D(0,1))=\ell$, and let $i(D(0,1))\geq\sigma$, for $\sigma-1$ exceeding a positive constant. Given two positive tolerance values $\epsilon=1/2^{b}$ and $\phi$ and a black box $\sigma$-soft $\ell$-test, which evaluates the polynomial $p(x)$ at $\mathcal{NR}$ points, Alg. 16.1 approximates the $\ell$-th smallest root radius $|z_{j}|$ within the relative error $2^{\phi}$ or determines that $|z_{j}|\leq\epsilon$, at the cost of evaluation of polynomial $p(x)$ at $\mathbb{I}\cdot\mathcal{NR}$ points for $\mathbb{I}$ estimated in Prop. 16.1, $\mathbb{I}=O(\log(b))$ provided that $\nu-1$ and $\phi$ are positive constants.
17 Compression of an isolated disc
17.1 A motivation: compact components in the union of suspect squares
As in Sec. 4, partition the union of suspect squares at every subdivision iteration into components and associate breaking (partition of) components into sub-components with component tree, which has at most $22m-$ 2 edges. Next we will bound the numbers of subdivision steps (and hence of e/i tests) at and between partitions of components.
We call a pair of suspect squares non-adjacent if they share no vertices. Call a component compact if it contains just a single non-adjacent suspect square, that is, consists of at most four suspect squares sharing a vertex (see Fig. 5).
Figure 5: Roots (asterisks) define compact components. A subdivision step halves their diameters and at least doubles isolation of their minimal covering squares and discs.
Figure 6: Two discs superscribing two non-adjacent suspect squares are not minimal but do not overlap.
Observation 17.1.
(See Fig. 6.) The $\sigma$-covers of two non-adjacent suspect squares are separated by at least the distance $a\Delta$ for the side length $\Delta$ and separation coefficient
$a:=2-\sigma\sqrt{2}>0~{}{\rm for}~{}1<\sigma<\sqrt{2}.$
Lemma 17.1.
Let $|z_{g}-z_{h}|\geq\sigma\Delta$, for a pair of roots $z_{g}$ and $z_{h}$, the softness $\sigma$ of e/i tests, and the side length $\Delta$ of a suspect square, and let $i>0.5+\log_{2}(m)$. Then in at most $i$ subdivision steps the roots $x$ and $y$ lie in distinct components.
Proof.
Two roots $z_{g}$ and $z_{h}$ lying in the same component at the $i$th subdivision step are connected by a chain of $\sigma$-covers of at most $m$ suspect squares, each having diameter $\sigma\Delta/2^{i-0.5}$. Hence
$|z_{g}-z_{h}|\leq m\sigma\Delta/2^{i-0.5},~{}i\leq\log_{2}\Big{(}\frac{m\sigma\Delta}{|z_{g}-z_{h}|}\Big{)}+0.5.$
Substitute the bound $|z_{g}-z_{h}|\geq\sigma\Delta$ and deduce that $m\leq 2^{i-0.5}$ or equivalently $i\leq 0.5+\log_{2}(m)$ if the roots $z_{g}$ and $z_{h}$ are connected by a chain of suspect squares, that is, lie in the same component. This implies the lemma. ∎
Corollary 17.1.
A component of suspect squares is broken in at most $1+\lfloor 0.5+\log_{2}(m)\rfloor$ successive subdivision steps unless it is compact.
Observation 17.2.
(i) A compact component is surrounded by a frame of discarded squares (see Fig. 5) and hence is covered by a 2-isolated square. (ii) If it stays compact in $h$ successive subdivision steps, then its minimal covering square and disc become $2^{h+1}$-isolated and $2^{h}\sqrt{2}$-isolated, respectively.
Proof.
Readily verify claim (i). Claim (ii) follows because a subdivision step can only thicken the frame of discarded squares while decreasing at least by twice the diameter of any component that stays compact. ∎
Cor. 17.1 bounds the number of subdivision steps between partitions of non-compact components, but compact components can be processed very slowly (see Fig. 3). For a remedy, we intertwine subdivision process with compression of MCD of a component that stays compact in three subdivision steps, in which case the MCD is $4\sqrt{2}$-isolated by virtue of Observation 17.2.
Next we elaborate upon such compression of MCDs.
17.2 The compression problem and flowchart of its solution
Problem COMPR. INPUT: a positive $\epsilon\leq 1$, a constant $\eta>0$, a black box polynomial $p(x)$ of a degree $d$, and an isolated disc $D$ with a non-empty root set.
OUTPUT: an integer $m$, a complex $c\in D$, and positive $\rho$ and $\eta$ such that #$\#(D(c,\rho))=\#(D)$ and the disc $D(c,\rho)$ is $\eta$-rigid unless $\rho\leq\epsilon$.
Flowchart of a compression algorithm. To solve Problem COMPR we successively compute (a) $m=\#(D)$, (b) complex point $c$ lying in or near the minimal covering disc of the root set $X(D)$, and (c) positive $\rho$ and and $\eta$ such that #$\#(D(c,\rho))=m$ and the disc $D(c,\rho)$ is $\eta$-rigid unless $\rho\leq\epsilon$.
Wlog assume that $D$ is a unit disc $D(0,1)$ and (a) compute $m=\#(D)$ for isolated unit disc $D:=D(0,1)$ by using $O(\log(d))$ queries of an evaluation oracle (see Observations 12.1 and 12.2).
Next we elaborate upon the algorithms for stages (b) and (c).
17.3 A point in or near the minimal covering disc for the $d$ roots
Definition 17.1.
A disc is said to be strongly isolated if it is $d^{\nu}$-isolated for a sufficiently large constant $\nu$ of our choice. In particular every disc containing all $d$ roots is strongly isolated.
Next we successively compress the unit disc $D=D(0,1)$ in the s where it contains $d$ roots, is strongly isolated, and is just isolated.
Schröder’s iterations [104],
$z=y-m~{}{\rm NR}(y),$ (17.1)
fast converge to a root of multiplicity $m$ [63, Sec. 5.4] and for $m=d$ outputs a point in or near the MSD of the the set of the $d$ roots if point $y$ lies far enough from that set:
Theorem 17.1.
Write $D:=D(0,1)$ and $\Delta:=\Delta(X(D))$, let #$\#(D)=d$, fix $y$ such that $|y|>2$, let $z:=y-d~{}{\rm NR(y)}$ and $\bar{x}\in MCD(D)$. Then $|z-\bar{x}|\leq\alpha~{}\Delta$ for $\alpha:=\frac{(|y|-1)^{2}}{(y|-2)^{2}}=(1+\frac{1}{(y|-2})^{2}$, and so $\alpha=2.25$ for $|y|=4$, $\alpha=25/16$ for $|y|=6$, and $\alpha\mapsto 1$ as $|y|\mapsto\infty$.
Proof.
Recall that $\bar{x},z_{1},\dots,z_{d}\in X(D)$, $|z_{j}|\leq 1$ and $|\bar{x}-z_{j}|\leq\Delta$ for $j=1,\dots,d$. Hence (cf. (3.1))
$\Big{|}{\rm NIR}(y)-\frac{d}{y-\bar{x}}\Big{|}\leq\sum_{j=1}^{d}\Big{|}\frac{1}{y-z_{j}}-\frac{1}{y-\bar{x}}\Big{|}\leq\sum_{j=1}^{d}\frac{|\bar{x}-z_{j}|}{|y-\bar{x}|\cdot|y-z_{j}|}\leq\frac{\Delta d}{(|y|-1)^{2}}.$
Therefore,
$|u-w|\leq\beta~{}{\rm for}~{}u:=\frac{{\rm NIR}(y)}{d},~{}w:=\frac{1}{y-\bar{x}},~{}{\rm and}~{}\beta:=\frac{\Delta}{(|y|-1)^{2}},$ (17.2)
and so
$|z-\bar{x}|=\Big{|}\bar{x}-y+\frac{d}{{\rm NIR}(y)}\Big{|}=\Big{|}\frac{1}{u}-\frac{1}{w}\Big{|}=\frac{|u-w|}{|uw|}\leq\frac{\beta d}{|{\rm NIR}(y)|}|y-\bar{x}|.$ (17.3)
Deduce from (3.1) and the bounds $\max_{j=1}^{d}|z_{j}|\leq 1<|y|$ that
$|{\rm NIR}(y)|=\Big{|}\sum_{j=1}^{d}\frac{1}{y-z_{j}}\Big{|}\geq d\max_{j=1}^{d}\frac{1}{|y-z_{j}|}\geq\frac{d}{|y|-1}.$
Substitute this bound into Eqn. (17.3) and obtain that
$|z-\bar{x}|\leq\beta~{}||y|-1|\cdot|y-\bar{x}|$
Substitute $\beta:=\frac{\Delta}{(|y|-\Delta)^{2}}$ of (17.2) and obtain
$|\bar{x}-z|\leq\alpha^{\prime}\Delta~{}{\rm for}~{}\alpha^{\prime}:=\frac{(|y|-1)|y-\bar{x}|}{(|y|-\Delta)^{2}}.$
Substitute the bounds $\Delta\leq 2$ and $|\bar{x}|\leq 1$ and obtain the theorem. ∎
Corollary 17.2.
Under the assumptions of Thm. 17.2 the disc $D(z,r_{1}(z))$ is $\frac{1}{2+2\alpha}$-rigid.
Proof.
Verify the claim in the case where $x_{d}=0$, $x_{1}=\Delta=1$, e.g., $0\leq x_{j}\leq 1$ for all $j$, and $z=1+\alpha$ and notice that this is a worst case input. ∎
17.4 Compression algorithm for a disc containing $d$ roots
Next we compress a disc containing $d$ roots towards their minimal covering disc.
Algorithm 17.1.
$\epsilon$-compression 1.
INPUT: a positive $\epsilon=1/2^{b}\leq 1$ and a black box polynomial $p(x)$ of a degree $d$ such that
#$\#(D(0,1))=d.$ (17.4)
OUTPUT: a complex $z$, and a positive $r$ such that #$\#(D(z,r))=d$ and the disc $D(z,r)$ is $\frac{2}{13}$-rigid unless $r\leq\epsilon$.
INITIALIZATION: Fix $y\geq 4$.
COMPUTATIONS:
(i) Compute and output $z=y-d~{}{\rm NR}~{}(y)$.
(ii) Compute an approximation $\tilde{r}\geq 0$ to $r_{1}(z)$ by applying Alg. 16.1 for $\epsilon$ replaced by $\epsilon/4$.
Output $z$ and choose $r$ to define $\epsilon$-compression into the disc $D(z,r)$.
Namely, output $r=\epsilon$ if $\tilde{r}\leq 3\epsilon/4$.
Otherwise output $r=\tilde{r}+0.25\epsilon$, then the disc $D(z,r)$ is rigid.
Correctness of the algorithm is readily verified if Alg. 16.1 outputs a range $[\rho_{-},\rho_{+}]$ of length at most $\epsilon/4$, for in this case $r_{1}\leq r=\tilde{r}+\epsilon/4\leq\epsilon$, and #$\#(D(z,\epsilon))=d$.
Otherwise $1/\beta\leq\tilde{r}/r_{1}\leq\beta$ for a constant $\beta\geq 1$ of Alg. 16.1, while the disc $D(x,r_{1})$ is $\frac{1}{2\alpha+2}$-rigid by virtue of Cor. 17.2 where $\alpha\leq 2.25$ by virtue of Thm. 17.1 because $y\geq 4$ in our case. It follows that the disc $\bar{D}:=D(z,r_{1}(z))$ is $\frac{2}{13}$-rigid.
Remark 17.1.
If we initialize Alg. 17.1 at a sufficiently large $y$, then the same argument ensures that the disc $D(z,r_{1}(z))$ is $\eta$-isolated for any $\eta>4$ unless $\tilde{r}\leq 3\epsilon/4$.
This compression queries evaluation oracle twice while performing Schröder’s single iteration for computing $z$ and otherwise runs at the cost of performing Alg. 16.1, estimated in Cor. 16.1.
We decrease the latter cost for a large class of polynomials $p$ by choosing a larger tolerance $\epsilon$. Then we yield desired compression if the algorithm outputs $r$ exceeding $\epsilon/4$. Otherwise we recursively repeat the computations for the tolerance recursively decreasing to $\epsilon/4$.
17.5 Compression of an isolated disc
If $m<d$ and $i(D)\geq d^{\nu}$ for a sufficiently large constant $\nu$, then Eqn. (3.1) implies that the $d-m$ external roots, lying outside the disc $D$, little affect the proof of Thm. 17.1 and our whole study in the previous two subsections as long as we apply Schröder’s iteration at a complex point $y$ lying far enough from both disc $D$ and the circle $C(0,d^{\nu})$.
The MCD of a compact component can be isolated but not strongly. Then, however, we apply subdivision steps to that component and in order of $\log(d)$ steps have it broken or strongly isolated by virtue of Observation 17.2. Hence we can extend our study in the previous subsections. $O(\log(d))$ subdivision steps for compact components involve $O(m\log(d))$ e/i tests. The factor $\log(d)$ is immaterial for the $\tilde{O}$ estimates of our Main Theorems, but we remove it next.
17.6 From isolation to strong isolation
Let $s_{1}$ denote the sum of all $m$ roots in $D$. Clearly, the point $s_{1}/m$ lies in MCD and even in the convex hull of the root set $X(D)$. We compute approximation $c=s_{1}(b^{\prime})/m$ of $s_{1}/m$ within $1/2^{b^{\prime}}$ at the cost of $O(b^{\prime})$ queries of an evaluation oracle (see Thm. 12.1). For $b^{\prime}$ of order $\log(d)$ we yield error bound $1/d^{\nu}$ for a fixed positive constant $\nu$ of our choice at the cost of $O(\log(d))$ queries of an evaluation oracle 12.1).
By applying Alg. 16.1 we compute approximation $\rho$ of $r_{d-m+1}(c)$ also within $1/d^{\nu}$, and so
$|c-s_{1}/m|\leq 1/d^{\nu}~{}{\rm and}~{}|\rho-r_{d-m+1}(c)|\leq 1/d^{\nu}.$ (17.5)
The rigidity $\eta=\eta(D(c,\rho))$ of the disc $D(c,\rho)$ is given by the value $\frac{1}{\rho}\max^{*}|x_{i}-x_{j}|$ where maximum is over all pairs of roots lying in the disc $D(c,\rho)$. Now notice that
$\max^{*}|x_{i}-x_{j}|>0.5~{}r_{d-m+1}(s_{1}/m).$ (17.6)
Indeed, because of invariance of the claim under map (1.4), we can assume that $s_{1}=0$ and $x_{i}<0$. Then $\Re(x_{j})>0$ for some $j$, and (17.6) follows.
Furthermore, $|r_{d-m+1}(s_{1}/m)-r_{d-m+1}(c)|\leq|s_{1}/m-c|$; combine this bound with Eqn. (17.5) and obtain that $|r_{d-m+1}(s_{1}/m)-\rho|\leq 2/d^{\nu}$. Together with (17.6) this estimate implies that
$\max^{*}|x_{i}-x_{j}|>0.5(\rho-2/d^{\nu}).$
If $\rho\geq 4/d^{\nu}$, then $\rho-2/d^{\nu}\geq\rho/2$, the disc $D(c,\rho)$ is 0.25-rigid, and compression is completed.
If $\rho\leq 4/d^{\nu}$, then the disc $D(c,\rho)$, sharing with the unit disc $D(0,1))$ all its $m$ roots, is strongly isolated, namely, is $d^{\nu}/4$-isolated; then we can apply to it the compression algorithms of the previous sections.
18 Reduction of $m$-test to e/i tests
To complete our compression algorithms we should supply a $\sigma$-soft $\ell$-test for the input of Alg. 16.1, for which we can apply the low cost Las Vegas $\ell$-test in Sec. 13.
In application to compression, however, we can apply two alternative tests because we only need either (i) a $\sigma$-soft $m$-test for a sub-disc $D(0,\rho)$ of the unit disc $D(0,1)$ containing $m$ roots where $\rho<1$ or alternatively (ii) e/i test for the annulus $A(0,\rho,1)$ of Sec. 13.4.
(i) A $\sigma$-soft $m$-test for a sub-disc. Clearly, we can immediately reduce $d$-test in the unit disc $D(0,1)$ for $p(x)$ to performing 1-test (that is, e/i test) for that disc and for the reverse polynomial $p_{\rm rev}(x)$ of (1.5), but next we reduce $m$-test for any $m$, $1\leq m\leq d$, and for any disc $D$ to bounded number of e/i tests for sub-discs. If at least one of the e/i tests fails, then definitely there is a root in a narrow annulus about the boundary circle of an input disc, and then the internal disc of the annulus would not contain $m$ roots. Otherwise we obtain a root-free concentric annulus about the boundary circle of an input disc and then reduce $m$-test to root-counting in an isolated disc, which we can perform at a dominated cost (see claim (ii) of Prop. 12.1).
Algorithm 18.1.
Reduction of $m$-test to e/i tests.
INPUT: a polynomial $p(x)$ of a degree $d$, two values $\bar{\theta}>1$ and $\bar{\sigma}>1$, an integer $m$ such that $1\leq m\leq d$ and #$\#(D(0,\bar{\sigma}))\leq m$, and two algorithms $\mathbb{ALG}~{}0$ and $\mathbb{ALG}~{}1$, which for a black box polynomial perform $\bar{\sigma}$-soft e/i test for any disc and root-counting in an isolated disc, respectively.
INITIALIZATION (see Fig. 7): Fix a positive $\rho$ such that $\bar{\sigma}=O(\rho)$, say, $\rho=\bar{\sigma}/5$, and compute the integer
$v=\lceil 2\pi/\phi\rceil$ (18.1)
where $\phi$ is the angle defined by an arc of $C(0,1)$ with end points at the distance $\rho$, and so $v=O(1)$ for $\bar{\sigma}=O(\rho)$. Then compute the values
$c_{j}=\exp(\phi_{j}\sqrt{-1}),~{}j=0,1,\dots,v,$ (18.2)
such that $|c_{j}-c_{j-1}|=\rho$ for $j=1,2,\dots,v$ and $\phi_{j}-\phi_{j-1}=\phi$.
$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c_{5}$$c_{6}$$c_{7}$$c_{8}$$c_{9}$$c_{0}$
Figure 7: $v=9$, $|c_{j}-c_{j-1}|=\rho>|c_{9}-c_{0}|$ for $j=1,2,\dots,9$.
OUTPUT: Either an upper bound $r_{d-m+1}<1-\rho\frac{\sqrt{3}}{2}$ or one of the two lower bounds $r_{d-m+1}\geq 1-\bar{\sigma}\rho$ or $r_{d-m+1}>1+(\sqrt{3}-1)\rho$.
COMPUTATIONS: Apply algorithm $\mathbb{ALG}~{}0$ (soft e/i test) to the discs $D(c_{j},\rho)$, $j=0,1,\dots,v$. (a) Stop and certify that $r_{d-m+1}\geq 1-\bar{\sigma}\rho$, that is, #$\#(D(0,1-\bar{\sigma}\rho))<m$ for the open disc $(D(0,1-\bar{\sigma}\rho)$, unless algorithm $\mathbb{ALG}~{}0$ outputs exclusion in all its $v+1$ applications.
(b) If it outputs exclusion $v+1$ times, then apply algorithm $\mathbb{ALG}~{}1$ (root-counter) to the $\theta$-isolated disc $D(0,1)$ for
${}\theta:=\min\Big{\{}1+(\sqrt{3}-1)\rho,\frac{1}{1-0.5\rho\sqrt{3}}\Big{\}}.$ (18.3)
Let $\bar{s}_{0}=\#(D(0,1))$ denotes its output integer. Then conclude that
(i) $r_{d-m+1}\leq 1-\rho\frac{\sqrt{3}}{2}$, and so #$\#(D(0,1-\rho\frac{\sqrt{3}}{2}))=m$ if $\bar{s}_{0}=m$, while
(ii) $r_{d-m+1}\geq 1+(\sqrt{3}-1)\rho$, and so #$\#(D(0,1+(\sqrt{3}-1)\rho))<m$ for the open disc $D(0,1+(\sqrt{3}-1)\rho)$ if $\bar{s}_{0}<m$.
To prove correctness of Alg. 18.1 we need the following result.
Theorem 18.1.
For the above values $c_{0},\dots,c_{v}$, and $\rho$, the domain $\cup_{j=0}^{v-1}D(c_{j},\rho)$ covers the annulus $A(0,1-0.5\rho\sqrt{3},1+(\sqrt{3}-1)\rho)$.
Proof.
Let $c_{j-1}$ and $c_{j}$ be the centers of two neighboring circles $C(c_{j-1},\rho)$ and $C(c_{j},\rho)$, let $z_{j}$ and $z^{\prime}_{j}$ be their intersection points such that $|z_{j}|=|z_{j}^{\prime}|-|z_{j}-z^{\prime}_{j}|$ for $j=1,2,\dots,v$, observe that the values $|z_{j}-z_{j}^{\prime}|$, $|z_{j}|$, $|z^{\prime}_{j}|$, and $|c_{j-1}-c_{j}|$ are invariant in $j$, and hence so is the annulus $A(0,|z_{j}|,|z_{j}^{\prime}|)$ of width $|z_{j}-z_{j}^{\prime}|$ as well; notice that it lies in the domain $\cup_{j=0}^{v-1}D(c_{j},\rho)$.
$c_{0}$$c_{1}$$z$$z^{\prime}$
Figure 8: $|c_{1}-c_{0}|=|c_{0}-z|=|c_{1}-z|=|c_{0}-z^{\prime}|=|c_{1}-z^{\prime}|=\rho$.
Let us prove that $|z_{j}-z_{j}^{\prime}|=\rho\sqrt{3}$ for $j=1$ and hence also for $j=2,\dots,v$.
Write $z:=z_{1}$ and $z^{\prime}:=z_{1}^{\prime}$ and observe (see Fig. 8) that $|c_{1}-c_{0}|=|c_{0}-z|=|c_{1}-z|=|c_{0}-z^{\prime}|=|c_{1}-z^{\prime}|=\rho$, and so the points $c_{0}$, $c_{1}$, and $z$ are the vertices of an equilateral triangle with a side length $\rho$. Hence $|z-z^{\prime}|=\rho\sqrt{3}$. The circle $C(c_{j},\rho)$ intersects the line segment $[z_{j},z_{j}^{\prime}]$ at a point $y_{j}$ such that $|y_{j}-z_{j}|>\rho\sqrt{3}/2$ and $|y_{j}-z_{j}^{\prime}|>(\sqrt{3}-1)\rho$ for $\rho<1$, and this implies the theorem if $|c_{v}-c_{0}|=\rho$. If, however, $|c_{v}-c_{0}|<\rho$ and if $C(c_{0},\rho)\cap C(c_{v},\rho)=\{\bar{z},\bar{z}\}$, then $|\bar{z}-\bar{z}^{\prime}|>|z-z^{\prime}|$, and the theorem follows in this case as well. ∎
Corollary 18.1.
The disc $D(0,1)$ is $\theta$-isolated for $\theta$ of (18.3) provided that the discs $D(c_{j},\rho)$ contain no roots for $j=0,1,\dots,v$.
Correctness of Alg. 18.1. (i) If algorithm $\mathbb{ALG}~{}1$ outputs inclusion in at least one of its applications, then one of the discs $D(c_{j},\bar{\sigma}\rho)$ contains a root, and then Alg. 18.1 correctly reports that $r_{d-m+1}\geq 1-\sigma\bar{\rho}$.
(ii) Otherwise Cor. 18.1 holds, and then correctness of Alg. 18.1 follows from correctness of $\mathbb{ALG}~{}1$.
Softness is at most $\frac{1}{1-\bar{\sigma}\rho}$, $\frac{1}{1-0.5\rho\sqrt{3}}$, and $1+(\sqrt{3}-1)\rho$ in cases (a), (b,i), and (b,ii), respectively.
Computational complexity. Alg. 18.1 invokes algorithm $\mathbb{ALG}~{}0$ at most $v+1$ times for $v=O(1)$ of (18.1). If $\mathbb{ALG}~{}0$ outputs exclusion $v+1$ times, then algorithm $\mathbb{ALG}~{}1$ is invoked and runs at a dominated cost.
Remark 18.1.
One can modify Alg. 18.1 as follows (see [48]): instead of fixing a positive $\rho$ and then computing $v$ of (18.1) fix an integer $v>0$, apply algorithm $\mathbb{ALG}~{}0$ at $v$ equally spaced points $c_{j}=\exp(\frac{2j\pi\sqrt{-1}}{v}),~{}j=0,1,\dots,v-1$, let $\phi=\frac{2\pi}{v}$, and then define $\rho$. We also arrive at this algorithm if we compute or guess this value $\rho$ and then apply Alg. 18.1 but notice that $c_{v}=c_{0}$ for $c_{j}$ of (18.2) and only invoke algorithm $\mathbb{ALG}~{}0$ for $c_{0},\dots,c_{v-1}$.
19 Counting e/i tests
19.1 An upper estimate
The upper estimate $22m-$ 2 of part (ii) of Sec. 4 on the number of compression steps is sharp – reached in the case of the complete binary tree $\mathbb{T}$.
Besides compression steps our subdivision root-finders involve e/i tests. Let us count them. Each compression step is preceded by an e/i test, and so we may need to perform at least $22m-$ 2 e/i tests. According to the main theorem of this section below, this lower bound is sharp up to a constant factor:
Theorem 19.1.
Let subdivision iterations with compression be applied to a suspect square containing at most $m$ roots. Then they involve $O(m)$ e/i tests overall.
19.2 The first lemma
We begin our proof of Thm. 19.1 with
Lemma 19.1.
In every component $C$ made up of $\sigma(C)$ suspect squares we can choose $n(C)\geq\sigma(C)/5$ non-adjacent suspect squares.
Proof.
[…]
19.3 The second lemma
Lemma 19.2.
Fix two integers $i\geq 0$ and $g\geq 3$, fix a value $\sigma$ of softness, $1<\sigma<\sqrt{2}$, and define the separation fraction $a:=2-\sigma\sqrt{2}$ of Observation 17.1. For all $j$ let $j$-th subdivision step process $v_{j}$ components made up of $w_{j}$ suspect squares, $n_{j}$ of which are non-adjacent, let $\Delta_{j}$ denote their side length, and let the $j$-th steps perform no compression for $j=i,i+1,\dots,i+g$. Then
$w_{i+g}\geq(n_{i}-v_{i})a2^{g-0.5}~{}{\rm if}~{}\Delta_{i}\geq 2^{g}\Delta_{i+g}.$ (19.1)
Proof.
At the $i$th step a pair of non-adjacent suspect squares of the same component is separated by at least the gap $a\Delta_{i}$, and there are at least $n_{i}-v_{i}$ such gaps with overall length $(n_{i}-v_{i})a\Delta_{i}$. Neither of these gaps disappears at the $(i+g)$th step, and so they must be covered by $w_{i+g}$ suspect squares, each of diameter $\sqrt{2}~{}\Delta_{i+g}=\Delta_{i}/2^{g-0.5}$ under (19.1). Hence $(n_{i}-v_{i})a\Delta_{i}\leq w_{i+g}\Delta_{i}/2^{g-0.5}$. ∎
Next we extend Lemma 19.2 by allowing compression steps. We only need to extend the equation $2\Delta_{j+1}=\Delta_{j}$ to the case where the $j$th subdivision step for $i\leq j\leq i+g$ performs non-final compression, each followed with subdivision steps applied to a single suspect square with side length at most $4\Delta_{j}$.
The next two subdivision steps perform only five e/i tests, making up at most $105m-$ 10 e/i tests performed at all pairs of subdivision steps following all (at most $m-2$) non-final compression steps.
Now we only need to bound by $O(m)$ the number of e/i tests in the remaining subdivision steps, not interrupted with compression steps.
After a compression step we resume subdivision process with a suspect square whose side length bound decreases in the next two subdivision steps from at most $4\Delta_{j}$ to at most $\Delta_{j}$.
Hence we can apply to it Lemma 19.2, because now it accesses the same number $n_{j}$ of non-adjacent suspect squares and an increased number $\sigma_{j}$ of all suspect squares (cf. (19.1)).
19.4 Counting e/i tests at partition steps
Let $h$ denote the height of our component tree and count suspect squares processed in subdivision steps that partition components, temporarily ignoring the other steps.
Clearly, bounds (19.1) still hold and can only be strengthened. Notice that $v_{i}\geq 1$ for all $i$, sum $v_{i+g}$ as well as bounds (19.1) for a fixed $g>0$ and $i=0,1,\dots,h-g$, and obtain
$\sum_{i=0}^{h-g}n_{i}-\lambda\sum_{i=0}^{h-g}\sigma_{i+g}\leq\sum_{i=0}^{h-g}v_{i+g}~{}{\rm for}~{}\lambda:=\frac{\sqrt{2}}{a2^{g}}.$ (19.2)
Notice that
$\sum_{i=0}^{h-g}v_{i+g}\leq\sum_{i=0}^{h}v_{i}-\sum_{i=0}^{g-1}v_{i}<2m-g,$
Combine the above estimate and the bounds $n_{i}\leq m$ and (cf. Lemma 19.1) $w_{i}\leq 5n_{i},$ for all $i$, write $n^{\prime}:=\sum_{i=0}^{h}n_{i}$, obtain
$\sum_{i=0}^{h-g}n_{i}\leq n^{\prime}-\sum_{i=h-g+1}^{h}n_{i}\leq n^{\prime}-mg,~{}\sum_{i=0}^{h-g}w_{i+g}\leq\sum_{i=0}^{h}w_{i}\leq 5n^{\prime},$
and so
[…]
Obtain that
$0.5n^{\prime}\leq 2m-1-g+mg~{}{\rm for}~{}g=1+\Big{\lceil}\log_{2}\Big{(}\frac{5\sqrt{2}}{a}\Big{)}\Big{\rceil}.$
Hence
$n^{\prime}\leq 4m-2-2g+2mg,~{}{\rm and~{}so}~{}n^{\prime}=O(m)~{}{\rm for}~{}g=O(1).$ (19.3)
19.5 Counting all e/i steps
Next consider a non-compact component, count all suspect squares processed in all subdivision steps applied until it is partitioned, and then prove in Lemma 19.3 that up to a constant factor the overall number of these suspect squares is bounded by their overall number at all partition steps. Such a result will extend to all subdivision steps the bound $n^{\prime}=O(m)$ of (19.3), so far restricted to partition steps. Together with Lemma 19.1 this will complete our proof of Thm. 19.1.
Lemma 19.3.
Let $k+1$ successive subdivision steps with $\sigma$-soft e/i tests be applied to a non-compact component for $\sigma$ of our choice in the range $1<\sigma<\sqrt{2}$. Let the $j$-th step output $n_{j}$ non-adjacent suspect squares for $j=0,1,\dots,k$. Then $\sum_{j=0}^{k-1}n_{j}\leq n_{k}\frac{10\sqrt{2}}{a}$ for the separation fraction $a:=2-\sigma\sqrt{2}$ (cf. Observation 17.1.)
Proof.
Substitute $g=k-i$ into (19.3) and obtain $w_{k}\geq(n_{i}-v_{i})a2^{k-i-0.5}$, for $i=0,1,\dots,k-1$.
Recall that $w_{j}\leq 5n_{j}$ (cf. Lemma 19.1) and that in our case $v_{i}=1$ and $n_{i}\geq 2$ and obtain that $n_{i}-v_{i}\geq n_{i}/2$ for all $i$.
Hence $5n_{k}\geq n_{i}a2^{k-i-1.5}$ for all $i$, and so
$\sum_{i=0}^{k-1}n_{i}\leq n_{k}\frac{10\sqrt{2}}{a}\sum_{j=1}^{k}2^{-j}<n_{k}\frac{10\sqrt{2}}{a}.$
Remark 19.1.
The overhead constants hidden in the "O" notation of our estimates are fairly large but overly pessimistic. E.g., (i) the roots $x$ and $y$ involved in the proof of Lemma 19.2 are linked by a path of suspect squares, but we deduce the estimates of the lemma by using only a part of this path, namely, the part that covers the gaps between non-adjacent suspect squares. (ii) Our proof of that lemma covers the worst case where a long chains between the roots $x$ and $y$ can consist of up to $4m$ suspect squares, but for typical inputs such chains are much shorter and can consist of just a few suspect squares. (iii) In the proof of Lemma 19.3 we assume that $n_{i}-1\geq n_{i}/2$; this holds for any $n_{i}\geq 2$, but the bounds are strengthened as $n_{i}$ increases. (iv) In a component made up of five suspect squares at least two are non-adjacent, and so the bound of Lemma 19.1 can be strengthened at least in this case.
20 Overall complexity of our root-finders
We reduced our root-finding in a square with $m$ roots in its close neighborhood to $O(m)$ e/i tests in addition to at most $22m-$ 2 compression steps. We reduced compression first of an isolated and then of strongly isolated disc to (i) counting roots in an isolated disc ($O(\log(d))$ evaluation queries) (ii) application of Schröder’s single iteration (two evaluation queries), and (iii) approximation of the root radius $r_{d-m+1}(z,p)$, that is, performing $O(\log(b))$ $m$-tests.
In Sec. 18 we reduced an $m$-test to $O(1)$ e/i tests, and so the overall complexity of subdivision iterations is dominated by the complexity of performing $O(m\log(b))$ e/i tests and $O(m\log(d))$ queries of an evaluation oracle. By combining the latter estimates with those of Sec. 7 for an e/i test, we complete the proof of Main Theorems 1 – 4.
21 Algorithmic extras
Next we accelerate our root-finders for a large class of inputs.
1. To compress the isolated unit disc $D:=D(0,1)$ containing a small number $m$ of roots $z_{1},\dots,z_{m}$, compute $s_{1},\dots,s_{m}$ the first $m$ powers sums of the roots lying in $D$, then recover the coefficients of the factor $f(x)=\prod_{j=1}^{m}(x-z_{j})$ from Newton’s identities (cf. [83, Eqn. (2.5.4)]), approximate its zeros, and apply a black box root-finder to the polynomial $p(x)/f(x)$ represented by the pair of polynomials $p(x)$ and $f(x)$.
2. In the case of a large integer $m$ the latter algorithm suffers from swelling the coefficients of the factor $f(x)$. For $m=d$ the algorithm [102] empirically approximates all $m$ roots at a rather low cost, which provides more than just compression of $D$. Can we extend this approach to compression in the case where $m<d$ and the disc $D$ is sufficiently strongly isolated?
This is a research challenge, but we propose a similar low cost heuristic recipe which, as we conjecture, accelerates our root-finding for a large class of inputs.
Let the disc $D=D(0,1)$ be strongly isolated and let $i(D)\geq d^{2\nu}$ for a reasonably large $\nu$, and then instead of Newton’s apply Schröder’s single iteration (17.1) at sufficiently many, say 100, equally spaced points $y$ of the circle $C:=C(0,d^{\nu})$. Let $Z$ denote the set of images $z=y-m{\rm NR}(y)$. By extending our study in Sec. 17.5, deduce that all these images lie in the $\alpha\Delta$-neighborhood of the root set $X(D)$ for $\Delta:=\Delta(D)$ and $\alpha$ fast converging to 1 as $\nu$ grows large and hence that $\Delta(Z)\leq(2\alpha+1)\Delta$.
We may run into worst case inputs where $Z$ is a singleton and $\Delta(Z)=0$, but we conjecture that under a random rotation of the set of points $y$, the ratio $\Delta(Z)/\Delta$ exceeds a positive constant whp, and then we can narrow the initial range of BoL in Sec. 16 from $[-b,0]$ to $[\log_{2}(\Delta(Z)),0]$ at a low cost of performing Schröder’s iteration (17.1) at reasonably many, say 100, points.
Part V: Factorization, Root Isolation, Functional Iterations, and Root Radii Estimation
We organize this part of the paper as follows. In Secs. 22 and 23 we cover some results by Schönhage [106] on polynomial factorization and its links to root-finding and root-isolation, and our lower and upper bit-complexity estimates for these problems. In Sec. 24 we recall root-finders by means of Newton’s, Weierstrass’s [119] (aka Durand-Kerner’s [23, 51]), and Ehrlich’s of 1967 [28], aka Aberth’s of 1973 [2] iterations, accelerate them by means of incorporation of FMM, and initialize Newton’s iterations for a black box input polynomial. In Sec. 25 we recall fast approximation of all $d$ root radii where the coefficients of $p$ are given.
22 Polynomial Factorization, Root-finding and Root Isolation:
Linking three Problems and their bit operation complexity
22.1 Linking Factorization and Root-finding: upper bounds on the bit operation complexity
Hereafter write $|u|=\sum_{i=0}^{d}|u_{i}|$ for a polynomial $u(x)=\sum_{i=0}^{d}u_{i}x^{i}$.
The root-finders of [106, 78, 84] solve Problems 0 and 0 by extending the solution of the following problem of independent importance:
Problem F. Approximate Polynomial Factorization: [^28] The solution of Problem F can be also extended to isolation of the zeros of a square-free polynomial with integer coefficients (see our Remark 23.4) and has various other applications to modern computations, listed in Sec. 1.1. For a polynomial of degree $d$ the known sharp upper and lower bounds on the input precision and number of bit operations for root-finding exceed those for numerical factorization by a factor of $d$. Given a real $b$ and $d+1$ coefficients $p_{0},p_{1},\dots,p_{d}$ of (1.1), compute $2d$ complex numbers $u_{j},v_{j}$ for $j=1,\dots,d$ such that
$|p-\prod_{j=1}^{d}(u_{j}x-v_{j})|\leq\epsilon|p|~{}{\rm for}~{}\epsilon=1/2^{b}.$ (22.1)
Hereafter $\mathbb{B}_{0}(b,d)$ and $\mathbb{B}_{F}(b,d)$ denote the optimal number of bit operations required for the solution of Problems 0 and F, respectively, for an error bound $\epsilon=1/2^{b}$, and next we specify a class of polynomials whose zeros make up no large clusters (cf. [8, 22]).
Definition 22.1.
Polynomials with bounded root cluster sizes. The zero set of $p$ consists of clusters, each made up of a constant number of zeros and having a diameter at most $1/2^{3-b/d}$; furthermore, these clusters are pairwise separated by the distances at least $1/d^{\tilde{O}(1)}$.
In [78, 84] we proved the following near-optimal bound for Problem F.
Theorem 22.1.
$\mathbb{B}_{F}(b,d)=\tilde{O}((b+d)d)$.
In Sec. 22.3 we reduce Problem 0 to Problem F and vice versa by applying Schönhage’s results of [106, Cor. 4.3 and Sec. 19]:
Theorem 22.2.
(i) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}(bd,d))$ for any polynomial $p(x)$ of degree $d$, (ii) $\mathbb{B}_{F}(b,d)=\tilde{O}(\mathbb{B}_{0}(b+d+\log(d),d))$ for any polynomial $p(x)$ of degree $d$, and (iii) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}2(b,d))$ for polynomials of Def. 22.1, with bounded cluster sizes.
By combining Thms. 22.1 and 22.2 obtain the following estimates.
Theorem 22.3.
One can solve Problem 0 by using (i) $\tilde{O}((b+d)d^{2})$ bit operations for any polynomial $p$ of (1.1) and (ii) $\tilde{O}((b+d)d)$ bit operations for a polynomial $p$ with bounded cluster sizes.
Corollary 22.1.
One can solve Problem 0 by using $\tilde{O}((b+d)d+(b+m)m^{2})$ bit operation for $m$ of Def. 1.1 of Sec. 1.5. The bound turns into $\tilde{O}((b+d)d)$ if $m^{2}=\tilde{O}(d)$.
Proof.
Proceed in two steps:
(i) by using $\tilde{O}((b+d)d)$ bit operations solve Problem F and then
(ii) apply the algorithm that supports Thm. 22.3 to the $m$th degree factor of the polynomial $\prod_{j=1}^{d}(u_{j}x-v_{j})$ of (22.1) that shares with this polynomial all its zeros lying in the disc $D(1+\beta)$ for positive $\beta$ of Def. 1.1.292929We can slightly speed up stage (i) because at that stage we actually only need to approximate factorization of the factor $f$ of degree $m$ of $p$ that shares with $p$ its zeros in $D$.
22.2 Linking Factorization and Root-finding: lower bounds on the bit operation complexity
Readily verify the following information lower bounds.
Observation 22.1.
To solve Problem 0 or F one must access $(d+1)b$ bits of input coefficients and perform at least $0.5(d+1)b$ bit operations – at most one operation for two input bits.
Next we deduce a stronger lower bound for Problem 0.
Observation 22.2.
Let $p(x)=x^{d}+p_{d-j}x^{d-j}$ for an integer $j$, $1<j<d$, and complex $p_{d-j}$ such that $|p_{d-j}|=1/2^{bj}$. Then the polynomial $p(x)$ has the $j$ zeros $\zeta^{g}2^{-b}$ for $g=0,1,\dots,j-1$, which vanish in the transition to $x^{d}$.
Observation 22.2 implies that one must access the coefficient $p_{d-j}$ within $2^{-jb}$ to distinguish the above $j$ zeros of $p(x)$ from the value 0 or from the zeros of the cluster of $j$ zeros of the polynomials obtained from $x^{d}$ by means of infinitesimal perturbation of the coefficient $p_{d-j}$. Sum these bounds for $j=1,\dots,d$ and obtain
Corollary 22.2.
One must access at least $0.5(d+1)db$ bits of the coefficients of a monic polynomial $p$ and perform at least $0.25(d+1)db$ bit operations to approximate even its single zero within $2^{-b}$ if no restriction is imposed on the number of roots in an input complex domain $\mathbb{D}$ and its $\beta$-neighborhood.
The lower bounds of this subsection imply that the algorithm of [78, 84] and its extension supporting Thm. 22.1 solve Problems 0 for $m=d$ and F by using optimal number of bit operations up to polylogarithmic factors.
Let us extend this study to Problem 0 for $m<d$. In this case isolate from one another the two sets made up of $m$ and $d-m$ roots.
Observation 22.3.
Let $p(x)=(x^{m}+p_{d-j}x^{m-j})(x-z_{1})^{d-m}$ for an integer $j$, $1<j<m$, complex $p_{d-j}$ such that $|p_{d-j}|=1/2^{bj}$, and sufficiently large $|z_{1}|$. Then the polynomial $p(x)$ has the $j$ zeros $\zeta^{g}2^{-b}$ for $g=0,1,\dots,j-1$, which vanish in the transition to $p(x)=x^{m}(x-z_{1})^{d-m}$.
Now we extend Cor. 22.2 as follows.
Corollary 22.3.
For an integer $m$, $1\leq m\leq d$, one must access at least $0.5(m+1)mb+db$ bits of the coefficients of a monic polynomial $p$ and perform at least $0.25(m+1)mb+0.5db$ bit operations to approximate within $2^{-b}$ even its single zero lying in the $\beta$-neighborhood of a complex domain $\mathbb{D}$ for a fixed constant $\beta>0$ if this neighborhood contains at most $m$ roots.
Proof.
Obtain the lower bounds $0.5(m+1)mb$ and $0.25(m+1)mb$ as for Cor. 22.2. Then notice that a $\delta$-perturbation of any trailing coefficient $p_{i}$ of $p(x)$ for $i\leq d-m$ can move the roots lying in the $\beta$-neighborhood of $\mathbb{D}$ by order of $\delta$. ∎
Our Main Theorem 3 implies that these lower bounds are near-optimal for $m$ of order $d$.
22.3 Linking Factorization and Root-finding: proof of Thm. 22.2
Next we prove Thm. 22.2, by applying the results of [106, Cor. 4.3 and Sec. 19].
Recall that $\mathbb{B}_{0}(b,d)$ and $\mathbb{B}_{F}(b,d)$ denote the optimal number of bit operations for the solution of Problems 0 and F, respectively, for an error bound $\epsilon=1/2^{b}$, and furthermore, recall the lower bounds
$\mathbb{B}_{0}(b,d)\geq 0.5(d+1)b~{}{\rm and}~{}\mathbb{B}_{F}(b,d)\geq 0.25(d+1)db$ (22.2)
of [78, 84] and the following result (cf. [106, Thm. 19.1]).
Theorem 22.4.
Let $r_{1}=\max_{j=1}^{d}|z_{j}|\leq 1$. Then a solution of Problem 0 within an error bound $1/2^{b}$ can be recovered at a dominated bit operation cost from the solution of Problem F within an error bound $1/2^{bd+2}$, and so $\mathbb{B}_{0}(b,d)\leq\mathbb{B}_{F}(bd+2,d)$.
Based on [106, Cor. 4.3] we can also readily relax the assumption that $p(x)$ is monic.
Furthermore, in view of (22.2), relaxing the bound $r_{1}\leq 1$ would little affect the latter estimate because we can apply map (1.4), which changes the error bound of Problem 0 by at most $\log_{2}(\max\{\rho,1/\rho\})$, for $\rho\geq 1/2^{b}$.
[108, Thm. 2.7], which we reproduce in Sec. 23.4 as Thm. 23.3, shows an alternative way to relaxing both assumptions that $r_{1}\leq 1$ and $p(x)$ is monic.
Schönhage in [106, Sec. 19] has greatly strengthened the estimate of Thm. 22.4 unless “the zeros of $p$ are clustered too much". Namely, he proved (see [106, Eqn. (19.8)]) the following result.
Lemma 22.1.
Assume that the polynomial $p(x)$ is monic, the values $u_{j}=1$ for all $j$ and complex $v_{1},\dots,v_{d}$ satisfy bound (22.1), and $|z_{j}-v_{j}|\leq r:=2^{2-b/d}$ for all $j$. Furthermore, write $\Delta_{j,1}:=|v_{j}-v_{1}|$ for $j=2,3,\dots,d$ and suppose that $1.25\Delta_{j,1}\leq r$ for $j=2,3,\dots,d$ and $\Delta_{1}:=\min_{j}\Delta_{j,1}>r$. Then $|z_{1}-v_{1}|\leq 2^{2-b/m}/|\Delta_{1}-r|^{\frac{d-m}{m}}$.
If $1/|\Delta-r|^{\frac{d-m}{m}}=2^{\tilde{O}(d)}$, then Lemma 22.1 implies that $|z_{1}-v_{1}|\leq 1/2^{b/m-\tilde{O}(d)}$, and we can similarly estimate $|v_{j}-z_{j}|$ for all $j=1,2,\dots,d$.
By extending these estimates we strengthen Thm. 22.4 as follows.
Corollary 22.4.
Under bounding cluster sizes of Def. 22.1, it holds that $|z_{j}-v_{j}|\leq 1/2^{b/m-\tilde{O}(d)},~{}j=1,2,\dots,d$. Hence if $m=O(1)$ is a constant, then a solution of Problem 0 within an error bound $1/2^{b}$ can be recovered at a dominated bit operation cost from the solution of Problem F with a precision $1/2^{\bar{b}}$ for a sufficiently large $\bar{b}=\tilde{O}(b+d)$, and so $\mathbb{B}_{0}(b,d)\leq\mathbb{B}_{F}(\tilde{O}(b+d),d)$.
For converse reduction of Problem F to Problem 0 recall [106, Cor. 4.3] in the case of $k=2$, covering the product of two polynomials:
Theorem 22.5.
Let $p=gh$ for three polynomials $p$ of (1.1), $g$, and $h$. Then $||g||_{1}||h||_{1}\leq 2^{d-1}||p||_{1}$.
Now consider the products $p=gh$ for $g=x-z_{j}$, for $j=1,\dots,d$. For a fixed $j$, an error of at most $\epsilon/2^{d}$ for the root $z_{j}$ contributes at most $\epsilon||h||_{1}$ to overall factorization error of Problem F. Therefore, $||h||_{1}\leq 2^{d-1}||p||_{1}/||x-z_{j}||_{1}\leq 2^{d-1}||p||_{1}$ by virtue of Thm. 22.5. Hence the approximation errors of all the $d$ roots together contribute at most $2^{d-1}d\epsilon$ to the relative error of the solution of Problem F (ignoring dominated impact of higher order terms), and we arrive at
Corollary 22.5.
If $b=O(b-d-\log(d))$, then Problem F for an error bound $1/2^{b}$ can be reduced at a dominated bit operation cost to Problem 0 for an error bound $1/2^{\widehat{b}}$ and a sufficiently large $\widehat{b}=O(b+d+\log(d))$; hence and so $\mathbb{B}_{F}(b,d)=O(\mathbb{B}_{0}(b,d))$.
It follows that the complexity estimate of Thm. 1.3 can be also applied to the solution of Problem F of factorization provided that $b=O(b-d-\log(d))$. Then again, the bit operation complexity of this solution of Problem F is optimal (up to polylogarithmic factors) under the random root model provided that $d=\tilde{O}(b)$ (cf. Observation 22.1).
23 Recursive Factorization, its bit complexity, and Root Isolation
In this section we briefly recall some algorithms and complexity estimates from [106, 108]. We first recursively decompose a polynomial into the product of linear factors and then extend the factorization to approximation and isolation of the roots.
23.1 Auxiliary norm bounds
Recall two simple lemmas.
Lemma 23.1.
It holds that $|u+v|\leq|u|+|v|$ and $|uv|\leq|u|~{}|v|$ for any pair of polynomials $u=u(x)$ and $v=v(x)$.
Lemma 23.2.
Let $u=\sum_{i=0}^{d}u_{i}x^{i}$ and $|u|_{2}=(\sum_{i=0}^{d}|u_{i}|^{2})^{\frac{1}{2}}$. Then $|u|_{2}\leq|u|$.
The following lemma relates the norms of a polynomial and its factors.
Lemma 23.3.
If $p=p(x)=\prod_{j=1}^{k}f_{j}$ for polynomials $f_{1},\dots,f_{k}$ and $\deg p\leq d$, then $\prod_{j=1}^{k}|f_{j}|\leq 2^{d}|p|_{2}\leq 2^{d}|p|.$
Proof.
The leftmost bound was proved by Mignotte in [61]. The rightmost bound follows from Lemma 23.2. ∎
Remark 23.1.
[108, Lemma 2.6] states with no proof a little stronger bound $\prod_{j=1}^{k}|f_{j}|\leq 2^{d-1}|p|$ under the assumptions of Lemma 23.3. From various factors of the polynomial $p(x)=x^{d}-1$ such as $\prod_{j=1}^{\frac{d}{2}}(x-\zeta_{d}^{j})$ for even $d$ and $\zeta_{q}$ of (3.4), one can see some limitations on strengthening this bound.
23.2 Errors and complexity
Suppose that we split a polynomial $p$ into a pair of factors over an isolated circle and recursively apply this algorithm to the factors until they become linear of the form $ux+v$; some or all of them can be repeated. Finally, we arrive at complete approximate factorization
$p\approx p^{\ast}=p^{\ast}(x)=\prod_{j=1}^{d}(u_{j}x+v_{j}).$ (23.1)
Next, by following [106, Sec. 5], we estimate the norm of the residual polynomial
$\Delta^{\ast}=p^{\ast}-p.$ (23.2)
We begin with an auxiliary result.
Theorem 23.1.
Let
$\Delta_{k}=|p-f_{1}\cdots f_{k}|\leq\frac{k\epsilon|p|}{d},$ (23.3)
$\Delta=|f_{1}-fg|\leq\epsilon_{k}|f_{1}|,$ (23.4)
for some non-constant polynomials $f_{1},\ldots,f_{k},f$ and $g$ such that
$\epsilon_{k}d2^{d}\prod_{j=1}^{k}|f_{j}|\leq\epsilon.$ (23.5)
Then
$|\Delta_{k+1}|=|p-fgf_{2}\cdots f_{k}|\leq\frac{(k+1)\epsilon|p|}{d}.$ (23.6)
Proof.
$\Delta_{k+1}=|p-f_{1}\cdots f_{k}+(f_{1}-fg)f_{2}\cdots f_{k}|.$ Apply Lemma 23.1 and deduce that $\Delta_{k+1}\leq\Delta_{k}+|(f_{1}-fg)f_{2}\cdots f_{k}|$ and, furthermore, that
$|(f_{1}-fg)f_{2}\cdots f_{k}|\leq|f_{1}-fg|~{}|f_{2}\cdots f_{k}|=\Delta|f_{2}\cdots f_{k}|.$
Combine the latter inequalities and obtain $\Delta_{k+1}\leq\Delta_{k}+\Delta|f_{2}\cdots f_{k}|$. Combine this bound with (23.3)-(23.5) and Lemmas 23.1 and 23.3 and obtain (23.6). ∎
Write $f_{1}:=f$ and $f_{k+1}=g.$ Then (23.6) turns into (23.3) for $k$ replaced by $k+1.$ Now compute one of the factors $f_{j}$ as in (23.4), apply Thm. 23.1, then recursively continue splitting the polynomial $p$ into factors of smaller degrees, and finally arrive at factorization (23.1) with
$|\Delta^{\ast}|\leq\epsilon|p|$
for $\Delta^{*}$ of (23.2). Let us call this computation Recursive Splitting Process provided that it begins with $k=1$ and $f_{1}=p$ and ends with $k=d$.
Theorem 23.2.
In order to support (23.3) for $0<\epsilon\leq 1$ and $j=1,2,\ldots,d$ in the Recursive Splitting Process it is sufficient to choose $\epsilon_{k}$ in (23.4) satisfying
$\epsilon_{k}~{}d~{}2^{2d+1}\leq\epsilon~{}\textrm{for~{}all}~{}k.$ (23.7)
Proof.
Prove bound (23.3) by induction on $j$. Clearly, the bound holds for $k=1$. It remains to deduce (23.6) from (23.3) and (23.7) for any $k$. By first applying Lemma 23.3 and then bound (23.3), obtain
$\prod_{i=1}^{k}|f_{i}|\leq 2^{d}|\prod_{i=1}^{k}f_{i}|\leq 2^{d}~{}\Big{(}1+\frac{k\epsilon}{d}\Big{)}~{}|p|.$
The latter bound cannot exceed $2^{d+1}|{p}|$ for $k\leq d$, $\epsilon\leq 1$. Consequently (23.7) ensures (23.5), and then (23.6) follows by virtue of Thm. 23.1.
Remark 23.2.
The theorem implies that we can ensure the output precision bound $b^{\prime}$ by working with the precision of
$\bar{b}\geq\bar{b}_{\rm inf}=2d+1+\log_{2}d+b^{\prime}$
bits throughout Recursive Splitting Process.
23.3 Overall complexity of Recursive Factorization
The overall complexity of recursive factorization is dominated by the sum of the bounds on the complexity of all splittings into pairs of factors. The algorithm of [84] adopts recursive factorization of [106] but decreases the overall complexity bound by roughly a factor of $d$. This is due to additional techniques, which ensure that in every recursive splitting of a polynomial into the product of two factors the ratio of their degrees is at most 11. Consequently, $O(\log(d)$ recursive splitting steps are sufficient in [84], and so the overall complexity of root-finding is in $\tilde{O}(db)$ – proportional to that of the first splitting. In contrast, the degrees of a factor in any splitting in [106] can be low, e.g., 1, and so there can be up to $d-1$ splittings overall, each of the first $0.5d$ of them being almost as expensive as the very first one.
23.4 From factors to roots
Theorem 23.3.
[108, Thm. 2.7]. Suppose that
$p=p_{d}\prod_{j=1}^{d}(x-z_{j}),~{}\tilde{p}=\tilde{p}_{d}\prod_{j=1}^{d}(x-y_{j}),~{}|\tilde{p}-p|\leq\epsilon|p|,~{}\epsilon\leq 2^{-7d}$
and
$|z_{j}|\leq 1~{}{\rm for}~{}j=1,\dots,d^{\prime}\leq d,~{}|z_{j}|\geq 1~{}{\rm for}~{}j=d^{\prime}+1,\dots,d.$
Then up to reordering the roots it holds that
$|y_{j}-z_{j}|<9\epsilon^{\frac{1}{d}}~{}{\rm for}~{}j=1,\dots,d^{\prime};~{}\Big{|}\frac{1}{y_{j}}-\frac{1}{z_{j}}\Big{|}<9\epsilon^{\frac{1}{d}}~{}{\rm for}~{}j=d^{\prime}+1,\dots,d.$
By virtue of Thm. 23.3 for $b^{\prime}=O(bd)$ we can bound the bit operation complexity of the solution of Problem 0 by increasing the estimate for the complexity of factorization in Sec. 23.3 by a factor of $d$.
Corollary 23.1.
The bit operation complexity of the solution of Problem 0. Given a polynomial $p$ of degree $d$ and a real $b^{\prime}\geq 7d$, one can approximate all its $d$ zeros within the error bound $2^{-b^{\prime}}$ by using $O(\mu(b^{\prime}+d)d^{2}\log(d)(\log^{2}d+\log(b^{\prime})))=\tilde{O}((b^{\prime}+d)d^{2})$ bit operations.
Remark 23.3.
The study in this section covers the solution of Problem 1 for the worst case input polynomial $p$. If its roots are isolated from each other, then the upper estimates of Thm. 23.3 and Cor. 23.2 can be decreased by a factor of $d$ (cf. [84]).
23.5 Polynomial Root Isolation
Based on his results recalled in this section, Schönhage in [106, Sec. 20] estimates the bit operation complexity of the following well-known problem by reducing it to Problem F.
Problem ISOL: Polynomial Root Isolation. Given a polynomial $p$ of (1.1) that has integer coefficients and is square-free, that is, has only simple roots, compute $d$ disjoint discs on the complex plane, each containing exactly one root.
Remark 23.4.
Problem 0 can be viewed as the Problem of Polynomial Root Cluster Isolation. It turns into Problem ISOL if all the coefficients of $p$ are integers and the roots are simple, and so our algorithms for Problem 0 as well as the Algorithms of [78, 84] solve Problem ISOL within a record complexity bounds.
In [106, Sec. 20] Schönhage reduces Problem ISOL to factorization Problem F. Problem ISOL is readily solved if the zeros are approximated within $\epsilon=1/2^{b}<0.5\delta$ for $\delta$ denoting the minimal pairwise distance between the zeros of $p$ (see the known lower bounds on $\delta$ in Kurt Mahler and Maurice Mignotte [60, 61, 62]). Schönhage has estimated that the solution of Problem F (e.g., presented in [106, 52, 76, 77, 84]) implies isolation provided that
$b=O((\tau+\log(d))+d)~{}{\rm and}~{}\tau=\log_{2}\max_{i=0}^{d}|p_{i}|.$ (23.8)
Corollary 23.2.
The bit operation complexity of polynomial root isolation. Suppose that a polynomial $p$ of (1.1) has integer coefficients and only simple roots. Let $\sigma_{p}$ denotes its root separation, that is, the minimal distance between a pair of its roots. Write $\epsilon:=0.4\sigma_{p}$ and $b^{\prime}:=\log_{2}(\frac{1}{\epsilon})$. Let $\epsilon<2^{-d}$ and let $m=m_{p,\epsilon}$ denote the maximal number of the roots of the polynomial $p(x)$ in $\epsilon$-clusters of its roots. Then one can solve Problem ISOL by using $\tilde{O}(bdm)$ bit operations for $b=\frac{b^{\prime}}{m}$.
24 Newton’s, Weierstrass’s, and Ehrlich’s iterations
24.1 Contents
In the next subsection we recall Newton’s, Weierstrass’s, and Ehrlich’s iterations and comment on their convergence and acceleration by means of fast multipoint evaluation, which is dramatic but involves the coefficients of $p$. In Sec. 24.3 we cover more limited acceleration by means of incorporation of FMM, which can be applied to a black box polynomial. In Secs. 24.5 and 25 we cover initialization of Newton’s, Weierstrass’s, and Ehrlich’s iterations, required for their fast global empirical convergence. For Newton’s iterations we only need to compute a disc that contains all zeros of a polynomial $p$ (cf. [55, 43, 115, 102, 113]). This is given by the well-known estimates via the coefficients, but we present simple novel initialization recipes in the case of a black box polynomial $p$. The customary initialization is more demanding for Weierstrass’s and Ehrlich’s iterations but is still fast for polynomials represented with their coefficients, as we recall in Sec. 25, where we also discuss applications of this initialization in particular to fast real and nearly real root-finding. It is still a challenge to extend such initialization to the case of a black box polynomial that would be faster than our radii algorithm of Sec. 16.
24.2 Definitions
Consider Newton’s, Weierstrass’s, and Ehrlich’s iterations [55, 43, 102, 115, 119, 23, 51, 28, 2, 15]):
$z\leftarrow z-N(z),~{}z_{j}\leftarrow z_{j}-W(z_{j}),~{}z_{j}\leftarrow z_{j}-E_{j}(z_{j}),~{}j=1,\dots,d,~{}{\rm where}$ (24.1)
$N(x)=W(x)=E_{j}(x)=0~{}{\rm for~{}all}~{}j~{}{\rm if}~{}p(x)=0$ (24.2)
and otherwise
$N(z):=\frac{p(z)}{p^{\prime}(z)},~{}W(x):=\frac{p(x)}{l^{\prime}(x)},~{}\frac{1}{E_{j}(x)}:=\frac{1}{N(x)}-\sum_{i=1,i\neq j}^{d}\frac{1}{x-z_{i}}$ (24.3)
for $l(x):=\prod_{j=1}^{d}(x-z_{j})$.
Newton’s iterations can be initiated at a single point $z$ or at many points and can approximate a single root, many or all $d$ roots at the cost decreasing as the number of roots decreases [55, 43, 115, 113, 68, 102].
Weierstrass’s and Ehrlich’s iterations begin at $d$ points for a $d$th degree polynomial and approximate a subset of its $d$ roots about as fast and almost as slow as all its $d$ roots.
Near the roots Newton’s and Weierstrass’s iterations converge with quadratic rate, Ehrlich’s iterations with cubic rate [116]. Empirically, however, all of them consistently converge very fast globally, right from the start, if they are initialized properly. [^30] Formal support for fast global convergence is only known for for Newton’s iterations empirically they still run faster than according to the known estimates (see [102]). The known fast initialization algorithms involve the coefficients of $p$ (see [9, 15]).
All expressions in (24.3) for the functions $N(x)$, $W(x)$, and $E_{j}(x)$ can be rewritten as ratios $\frac{u(x)}{v(x)}$ for pairs of polynomials $u(x)$ and $v(x)$, whose coefficients are given or readily available at a low computational cost if the polynomial $p(x)$ is given with its coefficients (see [83, Problems 3.6.1 and 3.6.3]). Then multipoint evaluation of such a ratio amounts to multipoint evaluation of $u(x)$ and $v(x)$ and division of the computed values, and one can accelerate Newton’s, Weierstrass’s, and Ehrlich’s iterations by a factor $d/\log^{2}(d)$ by means of incorporation of the algorithms of [64, 44].
24.3 Acceleration with Fast Multipole Method (FMM)
The algorithms of [64, 44] deeply involve the coefficients of $p(x)$. Next we cover acceleration of Newton’s, Weierstrass’s, and Ehrlich’s iterations applied to black box polynomial root-finding. Given $3d$ complex values $w_{j}$, $z_{i}$, $v_{j}$ such that $z_{i}\neq v_{j}$ for all $i,j=1,\dots,d$, FMM computes the sums $\sum_{j=1}^{d}\frac{w_{j}}{x_{i}-z_{j}}$ as well as $\sum_{j=1,j\neq i}^{d}\frac{w_{j}}{z_{i}-z_{j}}$ for all $i$ by using $O(db)$ arithmetic operations, with the precision of order $b$, versus $2d^{2}$ arithmetic operations in the straightforward algorithm. The papers [86, 88] list the recent bibliography on FMM and related computations with Hierarchical Semiseparable (HSS) matrices. Relevant software libraries developed at the Max Planck Institute for Mathematics in the Sciences can be found in HLib,
The acceleration based on FMM is practically valuable in spite of the limitation of the working and output precision to the order of $\log(d)$ and the resulting limitation on the decrease of the bit operation complexity.
Next we incorporate FMM into Newton’s, Weierstrass’s, and Ehrlich’s iterations. First relate the roots of a polynomial $p(x)$ to the eigenvalues of a diagonal-plus-rank-one matrix [^31] Hereafter T denotes transposition of a vector.
$A=D+{\bf v}{\bf e}^{T}~{}{\rm for}~{}D=\operatorname{diag}(z_{j})_{j=1}^{d},~{}{\bf e}=(1)_{j=1}^{d},~{}~{}{\bf w}=(w_{j})_{j=1}^{d},~{}w_{j}=W(z_{j}),$
$W(x)$ of (24.3) and all $j$ (cf. [10]). Let $I_{d}$ denote the $d\times d$ identity matrix, apply the Sherman-Morrison-Woodbury formula, and obtain
$xI_{d}-A=(xI_{d}-D)(I_{d}-(xI_{d}-D)^{-1}{\bf w}{\bf e}^{T}).$
Hence
$\frac{1}{p_{d}}p(x)=\det(xI_{d}-A)=-\sum_{i=1}^{d}w_{i}\prod_{j\neq i}(x-z_{j})+\prod_{j=1}^{d}(x-z_{j}).$ (24.4)
Indeed, the monic polynomials of degree $d$ on both sides have the same values for $x=z_{j}$, $j=1,\dots,d$ (cf. [10])). [^32] Gerschgorin’s theorem (cf. [35]) for the matrix $xI_{d}-A$ defines neighborhoods of its eigenvalues (the roots of $p$), which helps to analyze and to modify the iterations (cf. [15]). Equations (24.4) imply that
$p(z_{i})=-w_{i}\prod_{j\neq i}(z_{i}-z_{j}),~{}p^{\prime}(z_{i})=\prod_{j\neq i}(z_{i}-z_{j})\Big{(}1-\sum_{j\neq i}\frac{w_{i}+w_{j}}{z_{i}-z_{j}}\Big{)},$
and therefore
$\frac{w_{i}}{N(z_{i})}=\sum_{j\neq i}\frac{w_{i}+w_{j}}{z_{i}-z_{j}}-1.$ (24.5)
Hence, given the $2d$ values of $z_{j}$ and $w_{j}=W(z_{j})$ for all $j$, one can fast approximate $N(z_{j})$ and consequently $E_{j}(z_{j})$ for all $j$ (cf. (24.3)) by applying FMM.
Every iteration but the first one begins with some values $z_{j}$, $w_{j}$ and the updated values $z_{j}^{\rm(new)}$ of (24.1), for $j=1,\dots,d$, and outputs the updated values $w_{j}^{\rm(new)}=W(z_{j}^{\rm(new)})$ for all $j$. Then equations (24.3) and (24.5) define $N(z_{j}^{\rm(new)})$ and $E_{j}(z_{j}^{\rm(new)})$ for all $j$. To update the values $w_{j}$, apply [15, equation (17)]:
$w_{i}^{\rm new}=(z_{i}^{\rm new}-z_{i})\prod_{j\neq i}\frac{z_{i}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}\Big{(}\sum_{k=1}^{d}\frac{w_{k}}{z_{i}^{\rm new}-z_{k}}-1\Big{)},~{}i=1,\dots,d.$
In the case where $z_{i}^{\rm new}=z_{i}$ if and only if $i>k$ this updating is still used for $i=1,\dots,k$ but is simplified to
$w_{i}^{\rm new}=w_{i}\prod_{j=1}^{k}\frac{z_{i}-z_{j}}{z_{i}^{\rm new}-z_{j}}~{}{\rm for}~{}i=k+1,\dots,d.$
Now fast compute the values $\sum_{k=1}^{d}\frac{w_{k}}{z_{i}^{\rm new}-z_{k}}$ for all $i$ by applying FMM.
Also compute the values
$R_{i}=(z_{i}^{\rm new}-z_{i})\prod_{j\neq i}\frac{z_{i}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}$
fast in two ways.
(i) Use the following customary trick (see [24, Remark 3]): approximate the values $\ln(R_{i})$ for $i=1,\dots,d$ fast by applying FMM and then readily approximate $R_{i}=\exp(\ln(R_{i}))$ for all $i$. [24] proves numerical stability of the reduction to FMM of a similar although a little more involved rational expressions.
(ii) Observe that
$\ln\Big{(}\frac{R_{i}}{z_{i}^{\rm new}-z_{j}}\Big{)}=\sum_{j\neq i}\ln\Big{(}1-\frac{z_{j}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}\Big{)}~{}{\rm for}~{}i=1,\dots,d$
and that
$\ln\Big{(}1-\frac{z_{j}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}\Big{)}\approx\frac{z_{j}-z_{j}^{\rm new}}{z_{i}^{\rm new}-z_{j}^{\rm new}}~{}{\rm if}~{}|z_{j}^{\rm new}-z_{j}|\ll|z_{i}^{\rm new}-z_{j}^{\rm new}|.$
Let the latter bound hold for $j\in\mathbb{J}_{i}$ and let $\mathbb{J}^{\prime}_{i}$ denote the set of the remaining indices $j\neq i$. Then
$R_{i}\approx\exp(\sigma_{i})\prod_{j\in\mathbb{J}^{\prime}_{i}}\Big{(}1-\frac{z_{j}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}\Big{)},~{}\sigma_{i}=\sum_{j\in\mathbb{J}_{i}}\frac{z_{j}-z_{j}^{\rm new}}{z_{i}^{\rm new}-z_{j}^{\rm new}},~{}i=1,\dots,d.$
By applying FMM approximate $\sigma_{i}$ for all $i$ and compute the values $\prod_{j\in\mathbb{J}^{\prime}}\frac{z_{i}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}$ for all $i$ in at most $4\sum_{i=1}^{d}~{}|\mathbb{J}_{i}^{\prime}|-d$ arithmetic operations for $|\mathbb{J}^{\prime}_{i}|$ denoting the cardinality of the set $\mathbb{J}^{\prime}_{i}$. Compare the latter cost bound with order of $d^{2}$ arithmetic operations in the straightforward algorithm.
Clearly the above updating computation is simpler and more stable numerically for $W(z_{j})$ versus $N(z_{j})$ and even more so versus $E_{j}(z_{j})$, for $j=1,\dots,d$.
Let us comment on the changes where one incorporates FMM into Newton’s, Weierstrass’s, and Ehrlich’s iterations implicitly applied to the wild factor $f(x)$ of $p(x)$. The same expression (3.3) defines Ehrlich’s iterations for both $f(x)$ and $p(x)$ by virtue of Thm. 6.1. Applying Newton’s iterations to a wild factor one additionally subtract the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ (cf. (3.3)) and can apply FMM to accelerate numerical computation of this sum. Applying Weierstrass’s iterations to a wild factor one performs additional division by $t(x)$ and decreases the degree of the divisor $l^{\prime}(x)$ from $d-1$ to $w-1$.
24.4 MPSolve and its competition with Newton’s iterations
Since its creation in 2000 the package MPSolve by Bini and Fiorentino [9], revised in 2014 by Bini and Robol [15], has been remaining user’s choice for approximation of all $d$ complex roots of a polynomial $p$. It implements Ehrlich’s iterations, which empirically converge very fast globally, right from the start, under proper initialization involving the coefficients of $p$. Such convergence has been consistently observed in decades-long worldwide applications of these iterations as well as in the case of some other root-finders by means of functional iterations, notably Weierstrass’s iterations.
Unsuccessful attempts of supporting this empirical behavior formally (e.g., see [3, 41]), however, have recently ended with surprising proofs that Weierstrass’s and Ehrlich’s iterations diverge in the case of inputs that lie on some complex manifolds, being an open domain in the case of Weierstrass’s iterations (see Reinke, Schleicher, and M. Stoll [103, 101]).
Serious competition to MPSolve came in 2001 from the package EigenSolve by Fortune [33], but the later version of MPSolve [15] combines the benefits of both packages. An implementations of Newton’s iterations for approximation of all roots by Schleicher and R. Stoll in [115] competed with MPSolve in 2017, but then MPSolve has regained upper hand by means of incorporation of FMM. As we showed in the previous subsection, however, such implementation is simpler and can be more efficient for Newton’s than Ehrlich’s iterations. The Ehrlich-Newton competition seems to resume in 2024 (see [7, 102, 68]).
24.5 Computation of a disc covering all zeros of a black box polynomial
Customary initialization of Ehrlich’s and Weierstrass’s iterations involves approximation of all root radii, for which we can $d$ times apply our algorithm of Sec. 16, greatly accelerated in the next section if $p(x)$ is given with its coefficients. In that case we immediately estimate the largest root radius $r_{1}$ to initialize Newton’s iterations of [115, 113, 102]. In this subsection we obtain such estimates for a black box polynomial.
24.5.1 Introductory comments
Given the coefficients of a polynomial $p$ of a degree $d$, we can readily compute its largest and smallest root radii $r_{1}$ and $r_{d}$, respectively (cf. [31, 39]). Next we compute a range $[\rho_{-},\rho_{+}]$ that brackets all root radii or, equivalently, compute an upper estimate for the largest root radius $r_{1}$ and a low estimate for the smallest root radius $r_{d}$ of a black box polynomial $p(x)$ of a degree $d$.
Its solution is required, e.g., in order to initialize algorithms for all $d$ roots by means of functional iterations.
Suppose that we have an algorithm that computes a disc $D(0,\rho_{-})$ containing no zeros of a black box polynomial $p(x)$; clearly, $\rho_{-}\leq r_{d}$. By applying this algorithm to the reverse polynomial $p_{\rm rev}(x)$ we obtain the upper bound $1/\rho_{-,{\rm rev}}$ on $r_{1}$.
It remains to describe computation of the disc $D(0,\rho_{-})$. Empirically we can do this by applying the algorithms of [94, 36], but next we study formally some alternative algorithms.
To obtain a range $[\rho_{-},\rho_{+}]$ that brackets all root radii, we need upper estimate for the root radii $r_{1}$ and $r_{d}$, but next we only estimate $r_{d}$ because we can extend them to $r_{1}$ by estimating $r_{d}$ for the reverse polynomial $p_{\rm rev}(x)$. We obtain a low bound on $r_{d}$ by computing a disc that contains no zeros of a black box polynomial $p(x)$.
In the rest of this section $p(x)$ denotes a black box polynomial.
24.5.2 Probabilistic solution
Seek a disc $D(c,\rho)$ containing no zeros of $p(x)$ in its interior, where we can choose any positive $\rho$ Sample a random $c$ in the unit disc $D(0,1)$ under the uniform probability distribution in $D(0,1)$.
The disc has area $\pi$, while the union of the $d$ discs $D(x_{j},\rho)$, for $j=1,2,\dots,d$, has area at most $\pi d\rho^{2}$. Therefore, $c$ lies in that union with a probability at most $d\rho^{2}$, which we can decrease at will by decreasing $\rho$. Unless $c$ lies in such a union, we obtain a desired disc $D(c,\rho)$.
By applying our e/i tests of Secs. 7.47.3 to this disc, we can verify correctness of its choice.
24.5.3 Deterministic solution
We can deterministically compute a desired disc $D(0,\rho_{-})$ at a higher computational cost, based on a soft e/i test.
Namely, apply 1-test to $d+1$ triples of parameters $c_{i}$, $\theta_{i}$, and $\rho_{i}$, for $i=0,1,\dots,d$, such that the discs $D(c_{i},\theta_{i}\rho_{i})$ have no overlap pairwise. Then, clearly, at most $d$ discs contain roots because $p(x)$ has only $d$ zeros, and so at least one of the $d+1$ discs contains no roots. Deterministic e/i tests of Secs. 7.2 and 7.3 involve evaluation of $p(x)$ at $22d+$ 2 points. We obtain a desired disc $D$ by applying these tests to at most $d+1$ discs and thus computing $p(x)$ at at most $2(d+1)^{2}$ points.
25 Fast approximation of the $d$ root radii with applications
The customary initialization recipes of Weierstrass’s and Ehrlich’s iterations involve the $d$ root radii at the origin and consequently require to use the coefficients of $p$. To initialize Ehrlich’s iterations MPSolve first applies Bini’s heuristic recipe of [6], reused in [9, 15], for the approximation of the $d$ root radii. Given such approximations MPSolve defines $k$ concentric annuli $A(c,\rho_{i},\rho_{i}^{\prime})$ and $k$ integers $m_{i}$ such that #$\#(A(c,\rho_{i},\rho_{i}^{\prime}))=m_{i}$ for $i=1,\dots,k$, $k\leq d$, and $\sum_{i=1}^{k}m_{i}=d$. Then MPSolve initializes Ehrlich’s iterations at $d$ points $z_{j}$, $j=1,2,\dots,d$, exactly $m_{i}$ of them being equally spaced on the circle $C(c,(\rho_{i}+\rho_{i}^{\prime})/2)$. This heuristic initialization recipe turns out to be highly efficient.
The algorithm supporting the next theorem also approximates the $d$ root radii. It appeared much earlier than Bini’s, is deterministic, formally supported, and faster than Bini’s. The algorithm has been devised and briefly analyzed in [106] for fast approximation of a single root radius and then readily extended in [74, Sec. 4] (cf. also [75, Sec. 5]) to approximation of all root radii of a polynomial. As in [6, 9, 15] it uses all $d$ coefficients of $p$. A variation of this algorithm in [47] supports the following theorem.
Theorem 25.1.
[47, Prop. 3]. Given the coefficients of a polynomial $p=p(x)$, one can approximate all the $d$ root radii $|x_{1}|,\dots,|x_{d}|$ within the relative error bound $4d$ at a Boolean cost in $O(d\log(||p||_{\infty}))$.
[47] extends the supporting algorithm to approximation of all root radii within a positive relative error bound $\Delta$ by performing $\lceil\log_{2}\log_{1+\Delta}(2d)\rceil$ DLG root-squaring iterations (3.5).
Combine Thm. 25.1 with estimates for the Boolean cost of performing these iterations implicit in the proof of [106, Cor. 14.3] and obtain
Corollary 25.1.
Given the coefficients of a polynomial $p=p(x)$ and a positive $\Delta$, one can approximate all the $d$ root radii $|x_{1}|,\dots,|x_{d}|$ within a relative error bound $\Delta$ in $\frac{1}{d^{O(1)}}$ at a Boolean cost in $O(d\log(||p||_{\infty})+d^{2}\log^{2}(d))$.
The algorithm supporting the corollary and said to be Alg. 25.1a computes at a low cost $d$ suspect annuli $A(0,r_{j,-},r_{j,+})$ for $j=1,\dots,d$ (some may overlap), whose union $\mathbb{U}$ contains all $d$ roots. The algorithm was the point of departure for amazing progress in [44], which extends the ingenious algorithm of [64] for multipoint polynomial evaluation.
We conclude this subsection with a list of some other benefits of fast approximation of the $d$ root radii.
(i) In a subdivision step one can skip testing exclusion and discard a square unless it intersects the set $\mathbb{U}$.
(ii) The search for real roots can be narrowed to the intersection $\mathbb{I}$ of $\mathbb{U}$ with the real axis, which consists of $2m$ segments of small length (see Fig. 9). [98] has tested a real root-finder that approximates the roots in $\mathbb{I}$ by initiating Newton’s iterations at the centers of the segments. According to these tests the algorithm is highly efficient for the approximation of all well-isolated real roots. One can extend Newton’s iterations at the centers of the segments by applying also our root-radii algorithm of Sec. 16. By complementing the root radii approach with some advanced novel techniques, the paper [47] has noticeably accelerated the real root-finder [54], which is currently user’s choice.
(iii) Initialization based on Alg. 25.1a should improve global convergence of Ehrlich’s, Weierstrass’s, and other functional iterations for root-finding as well as of Kirrinnis’s factorization algorithm of [52].
(iv) Suppose that Alg. 25.1a has computed an annulus $A(0,\rho_{-},\rho_{+})$ containing a small number $u$ of the zeros of $p$ and isolated from its external zeros of $p$. Then we can extend the root-finding recipes of [106] from a disc to an annulus as follows. (a) Readily compute the first $u$ power sums of the zeros of $p$ lying in the discs $D(0,\rho_{+})$ and $D(0,\rho_{-})$ and then (b) their differences equal to the power sums of the zeros of $p$ lying in the annuls $A(0,\rho_{-},\rho_{+})$. (c) By solving the system of Newton’s identities, compute the coefficients of the factor $f$ of the $u$th degree that shares with $p$ its root set in the annulus (cf. [106, Sec. 12]). (d) Approximate the zeros of $f$, at a lower cost since $u$ is small.
We can apply this recipe to every annulus containing a small number of zeros of $p$ and isolated from its external zeros and finally approximate the remaining wild zeros of $p$ by applying implicit deflation (cf. (3.3)).
$x$$y$may contain rootsmay contain roots
Figure 9: Two approximate root radii define two narrow suspect annuli and four suspect segments of the real axis.